A Complete Bibliography of Publications in the IBM Systems Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 September 2024 Version 2.65

Title word cross-reference

+ [CEH^{+}94]. e^{x} [Spi66].

0-07-911795-3 [Ano96a]. 0-201-62878-3 [Rac96]. 0-471-12993-3 [Ano96a]. 0-8129-91180-X [Ano02a].

1-884777-09-0 [Ano96a]. 1-Based [AJM^{+}83]. 1.0 [Cas06]. 1980s [Fra79, WGB99]. 1990s [Cor89, Got92]. 1997 [Ano97c]. 1998 [Ano99b].

2 [CM88, CRTW88, Cru84, DO84, Fer91, Fer92, HD88, HZ84, KR88, Miz88, Sal88, Sch73, Sel06, Syk88, TG84, Wat88]. 2.0 [Cas06]. 2.6 [BS05]. 2000 [Abo99]. 20th [Cle95]. 21st [ÖLL07].

3.0 [GHM^{+}05]. 3090 [AGMS89, ALPC88, HMS92, SPS^{+}91]. 3090/VF [ALPC88]. 3270 [Ghi85]. 3270-PC [Ghi85]. 36 [SKG^{+}87]. 360 [Amd64, BFS67, BB64, Bla64, Cen67, Cla66, FIS64, Fre66, Gou69, LGM69, Mea66, Pad64, RM67, Ste64, Tuc67, Wit66]. 370 [AMP76, Att79, Buc86, Cal75, CW81, CKB89, Fla85, HL76, HPSS79, KLS84, Pla89, SGS89, SM79, SS78, Tet79, TB84]. 38 [Bot82]. 390 [DDMS97, DBC^{+}00, KBB^{+}97, KDY97, NMCB97]. 3D [EDG^{+}96].

400 [Cha89, CC89, Pin89, ST89, ST99].

6000 [ABDD^{+}91, CBH^{+}92].

7 [Col74, RGH^{+}76]. 7040/44 [WT64].

801/Reduced [Hop87]. 88 [HS87].

9370 [AG90, BPZ90, GL90].

A.I [Ano02a]. ABC [AOK^{+}95]. ABLE [BSP^{+}02]. Abstract [ALP83, CJ04]. Abstracts [Ano66a, Ano66b, Ano66c, Ano66d, Ano67a, Ano67b, Ano67c, Ano67d, Ano68a, Ano68b, Ano69a, Ano69b, Ano69c, Ano69d, Ano70a, Ano70b, Ano70c, Ano70d, Ano71a, Ano71b, Ano71c, Ano71d, Ano72a, Ano72b, Ano72c, Ano72d, Ano73a, Ano73b, Ano73c, Ano73d, Ano74a, Ano74b, Ano74c, Ano74d, Ano75a, Ano75b, Ano75c, Ano75d, Ano75e, Ano76a, Ano76b, Ano76c, Ano76d, Ano77a, Ano77b, Ano77c, Ano77d, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano80a, Ano88e, Ano93a, Ano94a]. Access [AR76, BGT96, BR92, BE77, GWM75, IKS^{+}95, Nah74, BBFK05, BE01, BET72, BGG^{+}95b, CHNS95, DCB^{+}01, Dei92, ES05, GSN^{+}01, GGW02, HSK^{+}01, SGS89, SS94, WRS04, Ber88, SCO03]. Accessibility [BBFK05, BFH^{+}05, EPTD05, HBC^{+}05, HB05, MFJ05, THR^{+}05]. accessible [HGB05, MDC^{+}05, WSKW^{+}03]. Accessing [PCC^{+}98, Sen77, SAAF73a, SAAF73b, SAAF73c, TLRH98]. accessors [TJ00]. accommodate [BFH^{+}05]. Accounting [GJS75, RS72, Sel08]. Achieving [MBC08]. Acknowledgment [Ano70e]. acoustic [SR96]. Across [AS82]. action [VW06]. actionable [HMP02]. Active [SBE^{+}00, JG07]. activities [Kra03]. Activity [CFL^{+}06, GMM^{+}06, MGM^{+}06, RS98, SLW66, GMM^{+}06]. Activity-centric [GMM^{+}06]. actuation [Par96]. actuator [CCC^{+}08]. AD [AHS90, Hem90, MMNR90]. AD/Cycle [AHS90, Hem90, MMNR90]. adapt [LS00]. adaptable [CNL08]. adapter [LS05, Syk88]. Adaptive [AEE^{+}97, Sha65a]. ADD [Ger62]. Addendum [Ano73e].

Adding [EN07]. addition [Swe65]. Address [AJ73, Ryd83, AJ99]. addressed [Gri63]. Addressing [Buc63, DF07, Hei63]. Adele [Rac96]. administration [TB05]. Administrative [Gla78, Wim71]. Adopting [GHM^{+}05, HLT94]. adoption [JEF04, VCS^{+}08]. Advanced [Alf91, AFM03, Ben86, GHH^{+}83, Kas86, Mac74, PSLB97, SBL^{+}94, SKG^{+}87, BBGB92, DL89, EOO^{+}06, CMW87, DL89, GCF^{+}87, HP93, LS87, dPBM88, dP88, dM93]. Advancements [MLDF08]. Advances [BBC^{+}95]. advantages [Inm98]. Advisor [Rof01]. Aeronautical [Lue70]. Affect [NOR03]. Affecting [Tha84]. agenda [JMP02]. agents [CFL05, DHPB03, Elo95, LM96, RM00]. aggregator [Zha98]. Aid [Sme78]. Aided [Mus78, BCFH65]. AIM [DL89]. air [DDMS95, HR81]. AIX [LS87, CBH^{+}92, WBSK87]. Alan [Lor94b]. alert [Moo88]. Algorithm [ApS64, Bre65, Ahu68, DeM68, Men68, SL67, WB69]. Algorithms [BBDT84, Yar65, AEE^{+}97, Bel66, HK65, MPS71, VAH^{+}02, YFS94, Yel03]. Aligning [Hes05, IJM93]. alignment [BW93, HV93, HV99, LLO93, Nor93]. Alistair [Cra95]. alive [Pen00]. all-atom [DK01]. alliances [PH01]. allocation [LPR^{+}07, WB69, Wit63]. alphabet [LC03]. Alternate [Yuv80]. ambient [ES05, HR81]. Among [Moo83]. Analysis [Bar75, BBDT84, GS04, KKLR05, Man74, MPS71, MPH68, Mis83, Par82, RS73, Yuv80, dL78, ABLU00, ALPC88, BCFH65, BD66, CW65, EDG^{+}96, Fer71, HH66, IFM^{+}01, Kos63a, LR67, LS91, Lew66, McD96, Men68, Mis81, NN01, OT94, Oka68, PCFY07, Rad66, SLW66, STTA08, Ste79, Swe65, TRL^{+}08, VPH^{+}08, WYB98, ZLK^{+}07]. analyst [Ars05]. Analytic [Dei82, AW72]. analytical [GKB^{+}03]. Analytics [LPR^{+}07, ANPT02, GCG^{+}04, MMS^{+}04].

Analytics-driven [LPR^{+}07]. Analyzer [ACP84, Pow83, OT94]. analyzing [CN08]. Anniversary [Cle95]. annotation [LI01]. API [AKR^{+}95]. APL [ATC77, Alf91, ASW91, Ano73e, Fal91, HL76, Ive91, Kol69, McI91, NJ80, TAR85, Tri91, Wah72, Whe91]. APL2 [BC91, JB91, JF91, Tho91, Wil91]. Apollo [Qua70, SH68]. APPC [Vos92]. APPC/MVS [Vos92]. appliance [ELM^{+}99]. Application [Ahu88, ACP84, AJ05, Ano68g, Bay98, Bie97, BCHM80, FV01, Gar83, Gor80, HHM04, HBBWS01, JDK^{+}91, LR02a, MV80, Rog83, SR93, Ste82, Tha84, WG88, WBC04, Wol88, All71, AP90, BFSLT08, BCF^{+}04, BCE^{+}94, BBGB92, BCE67, Blu64, BKP64, BCG^{+}89, CGG90, Coc93, Coc99, CRTW88, DGK96, Dev67, Dew90, EG64, FL04, FP90, GCS97, HDH88, Hen98, Hou93, Kee67, Kre01, LR67, Min92, NNH^{+}05, ORBdSB02, PTE^{+}95, Rym90, SB93, Sew67, Ste93, Sud67, Vel64, Vos92, WM93, WA02b, van00, Cha89, CC89, Pin89, ST89, ST99]. application-oriented [BCE67, Dev67, Kee67, LR67, Sew67, Sud67]. Applications [Bak80, BBG^{+}00, BKK^{+}02, Bjo75, HG86, Mau83, RH77, SS78, Wil86, Alf91, AKR^{+}95, ABW^{+}08, BBE^{+}97, Bev00, BDS88, CdVF^{+}06, CDM^{+}04, CD68, CCC^{+}08, CMCD94, CHL^{+}00, CEH^{+}94, DL89, DBC^{+}00, DMRS88, Fli97, GMTW08, HHSS01, HZ92, Haz96, HGB05, JR04, KKL06, LR97, LH97, MOB^{+}08, MPF97, MGM^{+}06, NNH^{+}05, PBS97, PBv97, Rof92, Sal88, SZ88, SC97, THR^{+}05, WA02a, WRS04, WLE^{+}97, Yeo05, dLMSC89]. applied [HR81, SBL^{+}94]. Applying [Hes05, ZMF02]. APPN [AB92, BBC^{+}95]. Approach [BBN96, Bro85, DM75, Jen79, Jon85, LC80, NJ80, Ryd70, Sch80a, Yuv80, ABLU00, ATJ92, BGP62, BM99, BGL^{+}05, BRM02, BIJ06, CN08, Col95, GKB^{+}03, HLT94, Hen98, Hou93, IG04, KOH^{+}07, KHC02, KBC^{+}07, MQR05, MBBD98, Min92, MKGR07, NC03, Sch93, SKA86, WCV69, Wei79, WBC04]. approaches [CH06, FK02, Shu89, Shu99]. approximating [EW66]. Arabic [ESH89]. Arch [Nae94]. archaeology [Rei89]. Architect [ABK^{+}06, LNH06]. Architectural [ABK^{+}06, JMZ92, McC78, BS04, BAC^{+}81, BAC^{+}99]. Architecture [Ber84, Bir73, Boh98, Buc86, CEH^{+}94, GY82, Hei85, Hen86, JDK^{+}91, Len78, Lor86, ML84, Mat86, MS80, MV80, MS85, PMM83, RROC85, Rog83, SR06, WCG^{+}80, AMM^{+}95, AMM^{+}99, ABLM95, AB92, ABW^{+}08, BL95, BCE^{+}94, BFH^{+}94, Bay98, BBWL05, BBC^{+}95, BCD^{+}07, BET72, BGG^{+}95a, CGHK98, Che01, CK05, CBC^{+}05, Cyp92, DY92, DFGS92, DM88, EOO^{+}06, EPTD05, FS05, FNS04, GS99, GFS03, GGKS02, GCG^{+}04, IKS^{+}95, IG04, JEF04, KKLR05, KBC^{+}07, LMPR07, MSB^{+}07, MICG95, MMS^{+}02, McD99, MMNR90, NMB04, NLN^{+}07, NG05, OCK^{+}92, OJG91, Pad68, RW04, Rym90, SGS89, SHLP05, SL95, SH87, SZ92, THR^{+}05, Wal07, WSP^{+}02, YRPSK99, Zac87, Zac99, van00, Ahu79, Ahu88, Ake89, AR76, Ano76j, CKB89, Cul76, EHJ77, FFK92, FL04, Hak90, Hob76, IG04, JD91, LMJW93, MMS^{+}04, McF76]. Architecture [Pau07, Pla89, SGS89, SB82b, Ste93, WG88]. Architecture/370 [CKB89, Pla89, SGS89]. Architectures [Gre79, SB82b, Sow02, dPBM88, AT06, Ber92a, HRL99, HGB05, KDY97, LG99, VL05]. Archival [CM77]. Archiving [Gra78]. Area [Bux84, Zim96, CHNS95, SVM^{+}03, Sum89]. areas [Ano68g]. Arguing [Ano02a]. Ariadne [LHT^{+}07]. Arithmetic [Swa90, BHS^{+}65]. array [MMS98, RC69]. art [MLDF08]. artful [HYJK06]. Artifact [BCK^{+}07]. Artifact-centered [BCK^{+}07]. artifacts [NC03]. Artificial [HG86]. arts [Mac00]. AS/400 [AM90, Bot89, KDA^{+}94, SLD89]. Asparagus [Neg96]. Aspect [CC05, HOST05]. Aspect-oriented [CC05, HOST05]. AspectJ [CC05]. Aspects [ADD68, Mue67, CGP68, DF07, FS72, Lip68, Oka68, Pad68]. assembly [Ano65f, Ano65g, Bak65, Cal65a, Cal65b, Dzi65, Gor65, Loe65, Sha65a, Tal63]. assertions [CJ04]. Assessing [Pea82]. assessment [Lor79]. Asset [Wei05]. Assets [Lar06, Hua98]. Assignment [Day70]. assistant [PLGB90]. assisted [GH79]. associated [ACO^{+}91]. associative [Sym68]. Asymmetric [HPSS79]. Asynchronous [Cav95]. Athena [AR92]. Atlantic [AOFH92]. ATM [CA95, OSK^{+}95, SB95]. atom [DK01]. Attach [HZ84]. Attached [TB84, GM03]. Attachment [MS80]. attacks [LM93]. audio [SR96, Moo90]. Audit [Bjo75]. auditing [JG07, RCPR07]. Augmentation [DP86, KB00]. augmented [BMMR96]. Author [Ano69e, Ano81a, Ano84a, Ano94b, DFS94, DK79a, DM76, Mor79, WF77a, Ano70l, Ano71j, Ano72j, Ano73k, Ano74m, Ano67e, Ano68c]. Authorization [Fri70, SGS89]. Authors [Ano63a, Ano64a, Ano64b, Ano64c, Ano65a, Ano65b, Ano65c, Ano66e, Ano66f, Ano66g, Ano66h, Ano67f, Ano67g, Ano67h, Ano67i, Ano68d, Ano68e, Ano68f, Ano69f, Ano69g, Ano69h, Ano69i, Ano70f, Ano70g, Ano70h, Ano70i, Ano71e, Ano71f, Ano71g, Ano71h, Ano72e, Ano72f, Ano72g, Ano72h, Ano73h, Ano73f, Ano73g, Ano73i, Ano74e, Ano74f, Ano74g, Ano74h, Ano75f, Ano75g, Ano75h, Ano76e, Ano76f, Ano76g, Ano76h, Ano77e, Ano77f, Ano77g, Ano77h, Ano78e, Ano78f, Ano78g, Ano78h, Ano79d, Ano79e, Ano79f, Ano79g, Ano80b, Ano80c, Ano80d, Ano80e, Ano81b, Ano81c, Ano81d, Ano81e, Ano82a, Ano82b, Ano82c, Ano82d, Ano88e, DFHS90]. Authorship [Ano70e]. Automated [Gar81, Orw00, Pra83, RH77, WA02a, ACGL04, DDK^{+}04, GCS97, ZMF02]. Automatic [BM83, BFVY96, Han70, Sha79, GY67, Wot93, BARM97]. Automating [HB85]. Automation [Col74, LCS82, MCE^{+}86, Ran02, CG77, DDMS95, HFB^{+}08, Irl92, SB93]. Automatist [DM97]. Autonomic [BBC^{+}03, HDS03, AHS^{+}03, BSP^{+}02, BCD^{+}07, CPW03, GC03, JBB03, LPPC03, MLR03, RMC03, RMDN03, SSM^{+}04, THL^{+}03, VSC^{+}03, WPT03]. autonomous [NOR03]. Autopilot [MBB^{+}07]. AutoTune [DHPB03]. Auxiliary [Sym68, RC69, SLW66]. Auxiliary-storage [Sym68, SLW66]. Availability [Aga85, BARM97, Bro85, KW86, STTA08, AFM03, BS83, BFSLT08, CBCS08, DCH^{+}03, HFB^{+}08, HH08, HPPS75, JAS08, LDL^{+}08, LSH^{+}08, MLDF08, MBC08, RMP08]. available [SBC^{+}08]. Avoiding [Hav68]. aware [LLS^{+}06, PY08, SB00a].

backup [KJP03]. bad [HT06]. balanced [SKA86]. balancing [SM01]. balloon [Par96]. bandwidth [Luc96]. Bank [Gra78]. banking [BW93, BL73, GB07]. Banks [LK96]. Base [AA76, Bjo75, Jon77, McG77b, PMM83, RH77, Sen77, Wat79, WFS80, FC92, GM91, Kah84, San81, SAAF73a, SAAF73b, SAAF73c, Zlo77, Yeo05]. Based [AJM^{+}83, CBG^{+}82, HBR86, Kra78, VR86, ACGL04, ABB^{+}03, BEH^{+}06, BBS02, Bov96, BCL^{+}04, CHPW07, CH06, DP95, ETD07, Elo95, FHP02, GSBH^{+}89, GOS05, Hem90, HBM^{+}04, IMMB01, JOZG02, KHC02, LS91, LR97, LHJ^{+}07, LHT^{+}07, MLA91, PORG00, SSR87, Sel08, Sha79, SWS06, SJL^{+}04, TVPL94, THR^{+}05, Wot93]. Bases [BE77]. Basic [JBB03]. basis [CJ04]. Batch [McG77c, Hob72]. Battle [Ano02a]. BEAM [TRL^{+}08]. beginning [TLRH98]. Behavior [Boi74, AHS^{+}03, Boi99, Par96]. being [Neg00]. beings [Neg00]. believing [BEF^{+}07]. Bell [Cra95]. Benchmark [Nai95]. benchmarks [AAC^{+}95, BDF^{+}00, HSY01]. Benefit [Par82, Str72]. Best [KAN^{+}07]. better [LM96]. Between [GHMWB84, Jen79, Gor71]. Beyond [Dav93, Dav99, HYJK06, KJP03, DCH^{+}03]. biased [ABDD^{+}91]. bibliography [Lor71]. big [Pap00]. bills [Chu63]. BIMS [ÖLL07]. biobanking [ÖLL07]. bioinformatics [GSN^{+}01, STF^{+}01]. biomarkers [HCR^{+}07]. biomedical [LRHH04, SRCV06, UMN^{+}04]. biopharmaceutical [HCR^{+}07]. BiProcessor [Ber92a]. bit [Orw96]. blade [STTA08]. blink [BMMR96]. block [BBET71]. blocks [JBB03]. Blue [IBM01]. blueprint [HLM95, Col95]. bodies [Sma96]. Book [Bon84, JCT^{+}97]. Books [Ano77i, Ano77j, Ano78i, Ano78j, Ano79h, Ano80f, Ano82e, Ano82f, Ano87a, Ano87b, Ano88a, Ano88b, Ano88c, Ano89a, Ano89b, Ano89c, Ano90a, Ano90b, Ano90c, Ano90d, Ano91a, Ano91b, Ano91c, Ano92a, Ano92b, Ano92c, Ano93b, Ano93c, Ano93d, Ano93e, Ano96a, Ano97a, Ano97b, Ano98a, Ano99a, Ano00, Ano01a, Ano02a, Ano03a, Bec95, Cav95, Cle95, Cra95, Den03, Hai89, Hed03, Joh95, Lor94a, Lor94b, McQ94, Nae94, Rac96, Vet95, Wat94b]. bookshelf [FHK^{+}97]. both [Sam88]. Box [HM93, MLH87, LTL^{+}08, Odo90, SDE^{+}06, Tag90]. Box-structured [HM93]. BPEL [KKL^{+}04]. brick [vdS00]. Bricks [RMSS96]. BRICO [Haa00a]. Bridging [Bun00, FKS^{+}02, FKS^{+}03, LH97]. bring [VMKB05]. Bringing [SCK02]. BroadBand [IKS^{+}95, Cav95, EOO^{+}06, OIS^{+}06]. broidery [PORG00]. Broker [CDF^{+}98]. Brooks [Cle95]. browser [HBC^{+}05]. BT [Nai95]. buffer [BD66]. buffering [Woo70]. Buffers [TG84]. build [GCG^{+}04]. Building [AAB^{+}05a, Ben86, Buc85, Cav00, DBC^{+}00, FL04, HST06, Ste93, Vre03, Yeo05, BSP^{+}02, DAK00, JBB03, KHC02, Pau07]. bulkload [KM06a]. bullet [May94]. bundling [WLB^{+}07]. bus [CNL08, SHLP05]. Business [Ano96a, ABD^{+}97, Bev00, BS04, CBC^{+}06, CN08, DDK^{+}01, EN07, Hel81, Kat90, KKL06, KKL^{+}04, LTL^{+}08, McD96, MGM^{+}06, NNH^{+}05, NC03, RS73, Rom08, SB82a, AAB^{+}07, Ano73e, Ars05, BKK^{+}02, BCC^{+}01, BW93, BCL^{+}04, CdVF^{+}06, CGH^{+}05, CDK08, CKKS02, Dav93, Dav99, DB08, DM88, FSSS01, FK01, FV01, GGW02, GKB^{+}03, Hae03, HHSS01, Hes05, HYJK06, IJM93, JL00, KKLR05, KBB^{+}05, KHSW05, KNN^{+}01, LA94, LRS02, LMX07, LLO93, LSH^{+}08, MBB^{+}07, McD99, NLN^{+}07, PCC^{+}98, PV97, PCH93, PBv97, Rap04, SMK^{+}02, Sch93, SR06, SD03, Smi01, Ste93, TRL^{+}08, VCS^{+}08, Wah72, Wal07, ZLK^{+}07, Zlo82, SFJ^{+}94, Zac82, van00]. Business-driven [NNH^{+}05]. Business-process-driven [LTL^{+}08]. Business-to-business [DDK^{+}01, BKK^{+}02, GGW02, SMK^{+}02]. Business/enterprise [Kat90]. Businesses [LK05]. bytecode [CFL05].

C [Cra95, Nae94, AOK^{+}95, AKR^{+}95, Fer91, RS85]. Cache [Gro85, Boz89, Lip68, BCF^{+}04]. Cache-DASD [Gro85]. CAD [dLMSC89]. CAFASP [SF01]. calculation [Mon67]. calculations [RC69]. call [LLS^{+}06, VW06]. calls [Fah02]. Cam [Laf68]. Can [Ste82]. Capability [For78, KMT85, KLS84, GTC68, KMT86]. Capacity [Dei82, Mis81, NORS80, Bro80, Coo80, Maj81, Sch80b]. capital [Sin03]. Capitalizing [Hua98]. car [WB69]. cards [HHSS01]. Care [MV80, BBC^{+}07, EBKT07, ETD07]. CAS [BDG^{+}94]. Case [BBC82, FW75, TMF06, BBS02, Bri90, BMK02, CCHJ94, CNS^{+}08, Cha66, CHL^{+}00, DAK00, HS08, JMZ92, Kra03, PB90, PCH93]. Cases [BM83, Han70, RMS06, WA02a]. CASP [SF01]. Catalog [LHJ^{+}07]. Catalog-based [LHJ^{+}07]. Catalyst [DB08]. Catapults [Bou98]. categorization [JOZG02]. Causal [HAB^{+}92]. CCITT [SSR87]. CDI [ITIW07]. celebrates [Bla98]. Cell [EOO^{+}06, OIS^{+}06]. Center [Gha75, Hir81, RMH08, Bau93, CDK08, HKS71, Lew76, LLS^{+}06, Ham82]. centered [BCK^{+}07, KK03]. Centers [KM89]. Centralization [Str73]. Centralized [Lor94a, Fis91, Wei79]. Centre [PSLB97, SBL^{+}94]. centric [BGT96, GMM^{+}06, HST06, PVK08, RHC^{+}06]. Century [Ano02a, Ohm86, ÖLL07]. Cerutti [McQ94]. chain [Alt08, KBB^{+}07, Sin03]. chains [GOS05]. Challenge [Cor89, BT84, Caf08]. Challenges [Sch03, SCK02, BVP93, ES05, GM03, HGW01, KNN^{+}01]. Change [MS85, Orl73, DB08, MQR05, NS05, WAB^{+}07]. Changes [FW75, BFH89]. Changing [CBG^{+}07, Mac79]. Channel [BET72, Pad64, CW65, Gay66, HZ84]. character [Meg64]. Characteristic [MW66]. Characteristics [HSY01, HS03, SL75, BCE67, CC89, HD88]. characterization [AJ05]. Charlie [Joh95]. Charting [Lue70]. charts [Mon62]. Cheap [Orw96]. check [BM72]. check-clearing [BM72]. Checking [BM83, LMX07, MMS98]. Checkout [Mar73]. chess [SF01]. Chief [Bak72]. Children [Eva96]. choice [VPH^{+}08]. Choices [GG87, Tho88]. choreography [KKL^{+}04]. chronic [HAB^{+}92]. chronically [KMSM^{+}07]. CICS [AS82, BCCN01, BDM97, Dei92, Dei82, EHJ77, Sea80]. CICS/ESA [BDM97]. CICS/VS [AS82, EHJ77, Sea80]. cinematic [DAB^{+}00]. CIO [NA07]. CIO-led [NA07]. circus [SDP00]. Clark [Ano96a]. classification [Kos63b, LC01]. Classroom [Abo99]. Cleanroom [HLT94]. clearing [BM72]. Client [HN98, KJBH^{+}00, MPF97, PHG^{+}04]. client/server [KJBH^{+}00, MPF97]. clinical [ITIW07, SD07, WND07]. clinical-decision-intelligence [WND07]. clinical-genomics [SD07]. Clockwork [RMC03]. Clubhouse [RR96]. Cluster [KDY97, NMCB97]. clustering [JAS08, KM06a, LC01]. clusters [SPS^{+}91]. CMS [SNC91, DK79b, Rei79]. Co [Syk88]. Co-Processor [Syk88]. COBOL [Dor63, Fli97, Sal88]. COBOL/2 [Sal88]. Code [ER70, Fag76, BFVY96, Fag99, Hal77, LHT^{+}07, OT94, SRM^{+}98, SL75]. Code-Generation [ER70]. Coding [TC69]. coherency [HKS08]. collaboration [ABLM95, GMM^{+}06, NS05]. collaborative [KM06b, LI01, MGM^{+}06]. Collecting [DP95]. Collection [EGH^{+}05]. collective [CK98]. Colombo [CDK^{+}05]. Color [JB96]. Combining [CBCS08, KKL^{+}04]. combustion [PGM^{+}92]. come [Pru01]. Commerce [Rof01, BKK^{+}02, Con97, DPR^{+}02, GGW02, DPR^{+}02]. CommercePOINT [Con97]. Commercial [HBR86, Som84, Ano85b, Bev00]. commitment [RS98]. Common [GS04, BBD^{+}92, Par96, Ahu88, Ber88, BR92, JDK^{+}91, JD91, LMJW93]. CommonPoint [DGK96]. commonsense [MMS^{+}02]. Communication [GJ87, GHH^{+}83, Jen79, Joh95, LC80, McG77d, Moo79, SB97, VW06, ABLM95, CN08, Eis67, GB00a, HZ84, JB96, SHFG95, SKN06, TC81, Zim96, Zim99]. Communication-minded [VW06]. Communications [Ahu88, DSM83, FG83, GJ87, Har80, Sul83, Vet95, Bra79, Cyp92, EGLT79, GM91, MO90, Omi81, PTE^{+}95, Sch83, Ste79, Wei79]. Communities [LS01, Mar01, BCE^{+}96, GR01]. Community [Ben96, AAB^{+}05a, CGRW07]. Comparative [Lam84]. Comparing [WPT03, Woo69, Ano70k]. comparison [Rob01, Zac82]. compatible [AAB^{+}05b]. compensation [CGV^{+}02]. Competitive [Yel03, BVP93, GSBH^{+}89]. compilability [Jac77]. compilation [Hai65, PMK^{+}06, PVW91]. Compiler [Day70, EGH^{+}05, Mar73, RG85, RS85, BBD^{+}92, Dor63, EOO^{+}06, Hai65, Lar63, SOT^{+}00]. Compilers [ER70]. Complementarity [LM93]. Complex [Duk75, Alm72, BHR89, Far87, KS03, MOB^{+}08, Ste79]. Complexes [MCE^{+}86]. Complexity [BR08a, Bur96]. Compliance [JG07, BEF^{+}07, DF07, KAN^{+}07, LMX07, RCPR07]. compliance-checking [LMX07]. complicated [KS03]. Component [van00, Yel03, CDF^{+}98, EN07]. components [STF^{+}01, TJ00]. Composite [Mye76, Jac77]. Comprehensive [GDD^{+}01]. compressed [Rob01]. Compressing [Cra95]. compression [Luc96]. Computation [Spi66, FS72, Ger96b]. Computational [DK01, HGW01, Luc96, MMS00]. Computationally [BMMR96]. computations [HD88, KKS88, SSH88]. Computer [ATC77, Boi74, CW65, Dea82, Dim63, DM75, Fre73, GJS75, Han80, JMV69, KLS84, LT89, LS00, Mat86, MK73, MCE^{+}86, Mus78, Ohm86, Rei76, RG85, Str75, Sum84, Swa90, ABB^{+}03, BGP62, Bel66, BBET71, Ber88, BCFH65, Boi99, Bre65, BRM02, CC87, COC63, GH79, HS87, HS03, KK03, Koi89, Lew66, Lew76, Meg64, SB00a, SF01, VAH^{+}02, Cla85, Hop87, RR96, TAR85]. Computer-Aided [Mus78, BCFH65]. computer-assisted [GH79]. Computer-Dependent [Str75]. computer-operated [COC63]. Computers [Bra81, KLRP86, Mur84, All99, AS99, BHW99, BG64, CFP^{+}95, Fra79, Pic00, Sha79, d’A89]. Computing [Ano96a, Bra79, Cha70, DP86, Gar83, Gha75, GWM75, Gla78, KMT85, McQ94, Str73, AR92, Ars05, BBC^{+}03, BGT96, BCD^{+}07, BCL^{+}04, CBG^{+}07, CPW03, Col95, CDK^{+}05, DEA99, EAB^{+}04, Fis91, GC03, GW89, HKS71, HHM04, HBBWS01, JBB03, JEF04, KMT86, LD99, LPPC03, Mar99, McG72, MMG^{+}00, Pal04, PORG00, PW88, Rap04, Res00, RW04, RMDN03, Sch83, Sta96, SSM^{+}04, Str72, Swo01, WPT03, WGB99, WB95, Zim99, Lor94b]. concept [CGV^{+}02, Lew76]. Concepts [Bjo75, Kin73, Pla89, RH77, Sch73, Whe74, Ano68h, Cen67, PPTH72, San81]. conceptual [GB07, LVW08]. Conceptualization [BR08a]. Concern [HOST05]. concordance [SSV64]. concurrency [AOK^{+}95]. Concurrent [GW89]. conferencing [CC87]. configuration [Loe65, MSB^{+}07, Maj81, WAB^{+}07]. configurations [LDL^{+}08, McN93]. conformance [FHP02]. Connect [BBC^{+}07]. connected [GM03]. connection [BGG^{+}95a, Gro92, Moo90]. connections [RHC^{+}06]. connector [FNS04]. consequences [ACMS06]. Considerations [Aga85, Mur84, RG85, Sea80, Wag73, Amd64, Ano74k, BDH75, Cha89, Ham82, HPPS75, Pad64, Rof92, Ste63, THL^{+}03]. considering [Sam88]. Consist [vdS00]. Consistent [Rog83, Ber88]. console [Laf68]. constraint [BESZ02]. Constructing [KRR80]. construction [AC68, Dim63, HK65]. constructionist [DAB^{+}00]. Consulting [LeB08]. Content [Hak90, MN06, BCvE^{+}05, GNP^{+}04, Lip00, SCK02, WRS04]. Contention [SA69, Yuv80]. Contents [Ano64d, Ano65d, Ano66i, Ano67j, Ano69j, Ano70j, Ano71i, Ano72i, Ano73j, Ano74i, Ano75i, Ano76i, Ano77k, Ano78k, Ano79i, Ano80g, Ano81f, Ano82g, Ano83a, Ano84b, Ano85a, Ano86a, Ano87c, Ano88d, Ano89d, Ano90e, Ano91d, Ano92d, Ano93f, Ano95a, Ano96b, Ano97c, Ano98b, Ano99c, Ano99b, Ano02b, Ano03b, Ano67n]. Context [SB00a, Sin03, ACMS06, LS00, LLS^{+}06, SL95]. Context-aware [SB00a, LLS^{+}06]. Context-specific [Sin03]. continuing [HP93]. continuity [LSH^{+}08]. continuous [AFM03, BS67, FG90, JAS08, MBC08]. continuously [SBC^{+}08]. Contrasting [BHW99]. Contrasts [Mil81]. Contributions [EGH^{+}05, LVW08, KM89]. Control [Att79, Dav78, Dev67, GY82, GWM75, Gla78, GBK83, Hob76, Qua70, SS78, Yel85, Zac82, Ahu79, Alm72, ABW^{+}08, Bak65, BGT96, BE01, Bre65, Bro68, DDMS95, DHP02, DK74, Fah02, FC92, GFS03, GGW02, Gor65, Gro89, Gro92, Kee68, Kon93, LM93, Mat91, MLA91, Mug62, PBCK95, RS72, Sha65a, Sha65b, TC69, Tuc67, Yar65]. Controlled [HD77]. Controller [DW85, FG83]. Conventions [Eis67]. Convergent [SF01]. conversation [SKN06]. Conversational [Bro68, GTC68]. conversion [DY92]. Cooperative [GHMWB84]. coordinated [BCSW91, BBMA91, MSS91]. Coordination [CK98, NMB04]. coprocessors [SS01]. copy [GMTW08]. CORDS [ABC^{+}95]. Core [NLN^{+}07, JL00]. corporate [CBG^{+}07, CGRW07, YLG05, Woo02]. Correction [Ano74j]. Cost [BEH^{+}06, Par82, Str72, BRM02, DEA99]. Cost-based [BEH^{+}06]. Cost-Benefit [Par82, Str72]. cost-effective [BRM02]. costs [Mar63]. Coupled [ACM74, Mac74, HKS08, Smi63]. Coupler [CG77]. Coupling [Che84, JMNT97]. CPI [AKR^{+}95]. CPI-C [AKR^{+}95]. Craig [Ano96a]. crash [ALPC88, ALPC88]. Creating [DGK96, HKS08, LH97]. Credentials [dMR03]. crew [ATJ92]. crew-pairing [ATJ92]. criteria [Bar71]. critical [ABW^{+}08, CBCS08, CDK08, Mon67, Was93]. Critiquing [HJM^{+}82]. Cross [ORBdSB02, Cre68, Dew90, HDH88]. cross-section [Cre68]. Cryptographic [EMMT78, JDK^{+}91, JD91, KMM^{+}80, LMJW93, MM78, YS91]. Cryptography [Len78]. cryptosystems [Com90]. crystal [JRG^{+}01, Oka68]. crystals [Koi89]. CS [MPF97]. CUA [Ber92b]. Cues [Wat94b]. Cultural [NS05, ACMS06]. Cumulative [Ano88e, Ano67k]. Current [SSS^{+}87, GDD^{+}01]. curricula [DB08]. curves [Ahu68]. Customer [Yel85, BCK^{+}07, HST06, LPR^{+}07, LLS^{+}06, VPH^{+}08]. customer-centric [HST06]. customization [Lei99]. Customized [HZ92]. Customizing [Gor80]. cycle [Alt08, CK05, HBC^{+}94, AHS90, Hem90, MMNR90].

Dan [Ano96a]. dance [SDP00]. Daniel [McQ94]. DASD [Gro85, Gro89, Gro92, Maj81]. Data [Aga85, AC68, Ahu68, Ano68h, Ano68i, Ano68g, ADD68, App68, AA76, Bak80, BM68, Bjo75, BE77, Bro68, Cen67, CD68, Cla66, CDK08, Com68, Cre68, Cru84, Dav78, DZ68, Day70, DY92, DFGS92, DSM83, FG79, Fre73, GTC68, Gay95, GJ87, HLR02, HHP79, Jac68, Joh68, Jon77, Kas86, KL74, Laf68, Lev79, MM90, MPH68, McG77b, McG77d, Mor78, Oka68, Par82, PMM83, Per02, RH77, Rei89, Rul68, SB82a, Sch78, SAAF73a, SAAF73b, SAAF73c, Sen77, SS94, Ste82, Sym68, Wat94b, WFS80, Aik98, AEE^{+}97, Ano74j, Ano74l, BGML96, BBG^{+}00, BCH^{+}06, BHC^{+}94, Bro80, BCG^{+}89, CdVF^{+}06, COC63, DCB^{+}01, DF07, DD64, DK74, ETD07, Eis67, Far87, FG90, GH91, HSK^{+}01, HMP02, Hen69, JG07, JMNT97, Kah84, Lan73, LI01, LRHH04, Loe65, Mil99a, Omi81, Paz89, PCC^{+}98]. data [RS72, RM67, Rub89, San81, Sch99, SCK02, Str97, Tur63, WT64, WSKW^{+}03, Yag63, Zlo77, Zlo82, EMMT78, Uhr73, Wat79]. Data-Base [Jon77, SAAF73a, SAAF73b, SAAF73c]. Data-Managed [Par82]. data-taking [COC63]. Database [CLSW84, Cru84, DO84, HJ84, HZ84, Sel87, TG84, Alf94, CM88, CMCD94, DL89, HLR02, Hir90, HSY01, IFM^{+}01, LRHH04, MSB^{+}07, MSZ^{+}01, MPTW94, Rei88, RLPG96, SBC97, THL^{+}03, JG07]. Dates [Ohm86]. dawning [GC03]. DB [HD77, SB82a, Siw77]. DB/DC [HD77, SB82a, Siw77]. DB2 [BEH^{+}06, BFG^{+}95, BCH^{+}06, JMNT97, MNQ^{+}02, MLR03, MPF97, ÖCKM06, SRW04, SSB^{+}00]. DB2/CS [MPF97]. DBMSs [FCM^{+}98]. DC [HD77, SB82a, Siw77]. deadlock [Hav68]. Dealing [RMDN03]. Debugging [MPF97, HS02]. decades [ABB^{+}03]. Decision [BS94, Rac96, GKT63, ITIW07, JOZG02, JRG^{+}01, WND07]. decision-tree-based [JOZG02]. Decisions [RS73]. Declarative [CBG^{+}82, KHSW05]. Decodability [SL75]. decomposition [JK92a]. Deep [Swo01]. Defect [Jon85, BHC^{+}94, GTB90, MJHS90]. definable [HR81]. Defining [Mar83]. definition [Odo90]. Delivering [BHT03]. delivery [HS08, KPF^{+}04, KBC^{+}07, MBB^{+}07, RAK08, Hou93]. Demand [Ken03, LK05, San07, Ars05, CDM^{+}04, CBC^{+}05, DDK^{+}04, HHM04, PHG^{+}04, SSR^{+}06, TLRH98]. demonstration [HD88]. dependencies [PTE^{+}95]. Dependent [Str75]. deploying [CCC^{+}08]. deployment [HDS03]. Deriving [YFS94]. Describing [Fre73]. Description [EG64, DY92, DM64, DM99, FIS64, McD99, Neu71, YRPSK99]. Design [All74, All75, Ano63b, ALP83, ACM74, BM72, BM96, BDH75, Bir73, BCE67, Bot82, Bot89, CDM^{+}04, CLSW84, Cla81, Cow84, Dor63, Fag76, Fag99, GJ87, GS04, GG87, Gro85, Hed63, Hen86, HD77, JMV69, KL74, Lar63, Mar73, ML84, MPB^{+}04, Mye76, PHBT00, Pea82, PJ86, Pin89, PORG00, Pow83, RH77, RG85, Rof92, Rog83, Sta76, SMC74, Tal63, Tho88, Van76, Wag73, Wat79, WT64, Amd64, Ano74j, Ano74k, Ano74l, Ano75j, BM68, BL95, BFS67, BT84, BD66, BFH^{+}05, BFVY96, Cav00, Cha89, CNL08, Dew90, Eis67, EW66, Fre66, Gay66, HPPS75, Ive63, Ive99, IG04, Jac77, Kes03, KR88, Laf68, Lin80, MW66, McC78, MC96, NOR03, Pad64, PGM^{+}92, Pau07, Sch03, SLW66, Sea66, Sel08, SB00a, SMC99, THL^{+}03, Tur63, ZZC^{+}08, Bec95]. Designed [Aga85]. designer [Ber92b, Har03, MMS00]. Designing [AASW98, DMRS88, Glu08, HGB05, Koi89, Nae94, WBC04, MDC^{+}05, Whi01, BEF^{+}07]. designs [SR93]. desk [WA02a]. Desktop [KLS84]. Detailed [NORS80]. Detecting [SKN06]. Detection [KW86]. Determination [Dav82, For78, WI83]. determined [JB96]. Determining [SHP69]. develop [EAB^{+}04, HKOS96]. developed [BBS02]. Developer [Lam84, LR02a]. Developing [LR02a, LC03, WSKW^{+}03, AW72, AGA^{+}08, CCC^{+}08]. DevelopMate [Hei90]. Development [BL76, BDE^{+}05, CBC^{+}06, Fag76, HB85, MV80, New82, Sme78, Ste82, Tha84, Van76, AT06, AP90, BHW99, BBCM06, BBGB92, Bie97, BIJ06, CBG^{+}07, CGG90, Coc93, Coc99, CMW87, CK98, DMMR92, Dye80, Fag99, FP90, Gol99, GJMT03, GCS97, HT06, HOST05, HP95, HCR^{+}07, Hen98, HM93, Hou93, Hum85, JR04, Kan91, Lar06, MM90, McC78, Moo90, NOR03, Omi81, PV97, Raz93, SB93, SG98, SB97, Sel06, SBL^{+}94, SLD89, TMF06, VMKB05, WM93, dRW04, van00, van98]. developments [JN92]. Device [CG77, Nah74, BET72, Fer91, Fer92, GB00a, Miz88, SR96]. device-to-device [GB00a]. device-to-human [GB00a]. devices [MMS00, Ver96, Zim99]. diagnosis [BRM02]. diagnostic [Sud67]. diagram [BBET71]. diagrams [Mil99a]. Dictionary [SB82a, GY67, Uhr73]. Dictionary/Directories [Uhr73]. did [Pru01]. difference [Kee93]. differential [HH66]. differing [Hau68]. digital [BG64, Blu64, BKP64, Bre65, EG64, Eis67, GC88, HKLS03, MRB01, MC96, Neg00, RR96, Sea66, Vel64, Zim99]. dimensional [Com68, Ric00]. dimensions [App68].

Direct [Nah74, BET72, PMK^{+}06]. Direct-Access [Nah74]. directed [SRM^{+}98]. Direction [SSS^{+}87, Hum85]. Directions [GHMWB84, JMZ92, JN92, PW88, RM88]. directly [Smi63]. Director [Woo02]. Directories [Uhr73]. directory [SSB^{+}00]. disabilities [ES05, KTBE05]. disabled [PWF05]. disaster [CBCS08, LSH^{+}08]. Discipline [Gol86, Bie97, Glu08, Raz93]. Discovering [HMP02, WPSL01]. Discovery [PPDW01, Sch03, BGL^{+}05, ITIW07, LC01, MKGR07, UMN^{+}04]. DiscoveryLink [HSK^{+}01]. discrete [FG90, HK65, VPH^{+}08]. diseases [HAB^{+}92]. disk [GH03]. Dispatching [Ryd70]. Dispersion [Str73]. Display [Ano68h, MOK78, AC68, ADD68, FS72, GTC68, HD88, MPH68]. Displays [MDHM86]. Distributed [AS82, Aga85, Ano96a, Dem88, Har80, Lor79, Lor94a, McQ94, PS83, RGH^{+}76, Rei88, Sch78, Sch99, AEE^{+}97, AR92, AKR^{+}95, BCE^{+}94, BFH^{+}94, BBE^{+}97, CCC^{+}08, Col95, Dei92, Fis91, IFM^{+}01, Kie79, Nai95, PS99, PBS97, Rof92, SRW04, Sch88, SR93, Sum89, Vos92, WLE^{+}97, ZABM92, Zie79, AT79, BBN96, DFGS92, FG79]. Distribution [Bak80, HS83, MM78, AC89, PBCK95, Unl08]. diversity [MMS^{+}02]. divide [HKLS03]. Division [Ken03, Spe02]. DL [MV80]. DNA [IFM^{+}01]. Document [CBG^{+}82, FFK92, Moo83, KM06a, PRZ02, ZS99, SB82b]. Documentation [NS84]. Documents [Cra95, HN98, PWF05, UMN^{+}04, WAV^{+}04]. Doing [Ano96a]. domain [JRST07, LNH06, Whe91, IG04]. domain-specific [JRST07, LNH06]. domains [TVX^{+}04]. dominant [LVW08]. Donna [McQ94]. DOS/360 [BFS67, Cen67, RM67]. DOS/VS [Bir73, Kra78]. DOS/VS-Based [Kra78]. Down [Van76, LC01]. DPPX [AT79, FG79, AG90, BPZ90, GL90, Kie79]. dream [vdS00]. DRIVE [CCC^{+}08]. Driven [CBC^{+}06, BBCM06, Bat06, BGL^{+}05, BIJ06, CT96, Che69, DDK^{+}04, DHK^{+}08, HT06, HRL87, KBB^{+}07, KHSW05, KBC^{+}07, Lar06, LPR^{+}07, LTL^{+}08, NNH^{+}05, Sel06, SNK^{+}08, ZZC^{+}08]. drivers [Fer91, Fer92, IJM93, Miz88]. drives [GH03]. DS [Paz89]. DS-Viewer [Paz89]. DTMS [Wat79]. dump [LS91]. During [Tha84]. duties [BE01]. Dynamic [AJ73, BCF^{+}04, JBB03, Wit63, ABDD^{+}91, AJ99, FCM^{+}98, HKOS96, HK65, McG65, Yel03]. dynamically [WKI03]. dynamics [KS03, SM01].

E-broidery [PORG00]. e-business

[BCC^{+}01, FSSS01, FK01, FV01, GKB^{+}03, Hae03, HHSS01, KNN^{+}01, SR06, SD03, Smi01]. E-commerce [DPR^{+}02]. e-mail [Fah02]. e-marketplaces [GOS05]. early [Pin89]. EAS [PMM83]. EAS-E [PMM83]. Ease [BT84, Vre03]. Eclipse [GHM^{+}05, LHT^{+}07, PG05, dRW04]. Eclipse-based [LHT^{+}07]. econometric [Sar77]. Economic [BK63, DeM68, Gor71, Men68, Pie68, SHP69]. ecosystem [TRL^{+}08]. editing [GB00b]. Edition [Cle95, BFG^{+}95, Cav95]. Editor [Vet95]. Education [HG86, JRD81, RC91]. educational [Abo99]. Edwards [Ano96a]. EE [CM88]. Effective [FP90, SSH88, BRM02, GCS97, VCS^{+}08]. Effectiveness [DK79b]. Effects [JRD81, SA69, PV97, Sel08]. Efficient [PBCK95, DAK00, KM06a, PMK^{+}06]. EGGG [Orw00]. electric [Smi96]. Electronic [DeS81, Moo83, BM96, BKK^{+}02, Con97, GGW02, Zlo82]. Elements [All74, All75, Bec95, Ano75j]. Eliminating [PTE^{+}95]. elliptic [HH66]. embrace [CGRW07]. Emergent [Cav00]. Emerging [CN79, Gol86, HBC^{+}94, HBZ06, LK05, UI00, MRB01]. emotion [Pic00]. Emphasis [CH85, AOFH92]. empirical [BR08a, SB97]. Empowering [Ars05]. empty [WB69]. Emulator [HL76, HS00, HS00]. enabled [CH07, Wal07]. enablement [CDM^{+}04]. Enabling [AHS^{+}03, OSK^{+}95, SRW04, Uhl88, WLE^{+}97, Hei90, Wol88]. Encoding [HLLW72]. encounters [CN08]. Encryption [EMMT78]. End [BS94]. energy [Sha79]. eNetwork [TLRH98]. enforcement [BE01, JG07]. engagements [BCK^{+}07]. Engaging [CGRW07]. engine [PGM^{+}92, BAI^{+}86]. Engineering [AGMS89, Ber84, CH85, GRW86, Gol86, Pec85, Aik98, BFSLT08, Bat06, Bie97, Bri90, BDG^{+}94, CLS88, Dav93, Dav99, Dye80, HLT94, HZ92, Lar08, Lin80, McG72, Mil80, Mil99b, O’N80, Qui80, Smi62, BHT03, Cle95, MS88]. Engine^{T M} [EOO^{+}06, OIS^{+}06]. Enhanced [For78, HBC^{+}05]. Enhancements [Sul83, MOCP85, WBSK87]. Enhancing [Cas06, HFB^{+}08, HYJK06]. enough [HO07, MDC^{+}05]. Enriching [BCE^{+}96]. Enterprise [DCH^{+}03, Har03, Inm98, Lor94b, MST^{+}03, Par82, PH99, RS98, San07, AvKM^{+}07, BBWL05, BC04, IG04, Kat90, Kon93, LLO93, MPB^{+}04, NLN^{+}07, SRW04, SSB^{+}00, Lor94b, Ake89, BCCN01, BFGS98, CDF^{+}98, CKB89, Gro92, KNN^{+}01, Pla89, SGS89, SHLP05, TJ00]. Entertainment [LK96]. entity [SJL^{+}04]. entity-based [SJL^{+}04]. entropy [Ger96b]. Entry [BHS74, SP78]. Environment [Ber84, Duk75, Gar81, HBR86, Mac79, New82, SB82b, Abo99, ABLU00, ALPC88, BDM97, BCE^{+}94, Bra79, CCT72, CPW03, Cle89, CMW87, DO84, DeS81, Dei92, Gro92, Hem90, KNN^{+}01, LPPC03, McG72, SBC^{+}08, SHFG95, TLRH98, dLMSC89, van98, HOST05]. environments [Ber88, BE01, BCC^{+}01, Cav00, CBCS08, ES05, PDI^{+}00, PLGB90, Sim05, McQ94]. epidemiological [MKGR07]. EPISTLE [HJM^{+}82]. equations [HH66]. equipment [WCV69]. era [GC03, TLRH98, WPSL01]. Erich [Bec95]. Errata [Ano65e, Ano67l, Ano70k, Ano74k, Ano76j, Ano04, Ano08a]. Erratum [Ano74l, Ano75j, Ano78l, Ano85b, Ano90f, Ano91e, Ano98c, Ano08b, KMT86]. Error [KW86, LR67, TC69]. Errors [Fag76, Fag99]. ES/3090 [AGMS89, HMS92, SPS^{+}91]. ES/9370 [AG90, BPZ90, GL90]. ESA [BDM97, BCSW91, Cla89, Fis91, GH91, MSS91, McN93, OJG91, SNC91, YS91]. ESA/390 [OJG91, YS91]. ESA/XC [GH91]. Essays [Cle95]. essence [May94]. Essential [Ano96a]. Estimating [CNS^{+}08, Jen92, Gay66]. Estimation [WF77b, ANPT02]. eSuite [Bri98]. Ethical [Pal01]. Ethnographic [KM06b]. European [DMMR92]. evaluate [BBS02]. Evaluating [FW75, MFJ05, NS84, BFSLT08, SWS06]. Evaluation [Bar73, Dru69, GL74, Hob72, MGST70, SL67, TVPL94, BK63, Cre68, DF63, Kee68, Sch80b, Str72]. Event [CT96, SE08, DHK^{+}08, HMP02, MOB^{+}08, ZLK^{+}07]. Event-driven [CT96, DHK^{+}08]. Event-processing [SE08, MOB^{+}08]. Events [NG05, Lew66]. Everything [Ger00]. Everywhere [Res00, SMM^{+}00, BBFK05]. Evolution [Cyp92, Gro89, HHP79, HH79, JEF04, MDHM86, Moo79, SAAF73a, BDM97, BR92, Bie97, BS94, Cla89, FNS04, GBCW00, HDH88, HP93, KK03, KNNL98, MT03, SF01]. evolutionary [Hen98, Hou93, WBC04]. Evolving [CHNS95, GR01, JHMO07, SUW82, Yel85, BBGB92, Bra79, dLMSC89]. example [FL04, Zlo77, Zlo82]. examples [Blu64, BKP64, EG64, Vel64]. Exchange [Che01, Moo83, DD64, PH01]. exchanges [KHC02, SMK^{+}02]. Execution [Pow83, BBS02, OJG91, DLP^{+}05]. Executive [AT79, FG79, LS87, Mil02, CMW87]. exhibits [SDP00]. existing [MBBD98, PCC^{+}98]. Expanded [KMT85, Hed63, HS65, KMT86]. expected [BHT03]. Experience [PWBK90, GR01, KBK^{+}03, RMDN03, Vre03]. Experiences [AAB^{+}05b, GL95, HRL99, HP95, MJHS90, MFJ05, PPDW01]. experiment [Abo99, van00]. Experimental [Bar73, Kes03, Ryd83, She80, ZS99]. Experiments [BGH76, DCB^{+}01, Mus78, MC96]. Expert [VR86, WAV^{+}04]. expertise [BAI^{+}86]. explained [AYBDS06]. explanation [SB00b]. exploit [EOO^{+}06]. exploration [WSG^{+}01]. Explorer [GMM^{+}06]. explosions [Chu63]. Exponentiation [Com90]. Express [HS00]. expressing [Ben00]. Expression [Mus78, MSZ^{+}01, Wil91]. expressive [PHBT00]. extend [Elo95]. Extended [Mac74, Kon93, KA71, SS78]. extended-precision [KA71]. Extending [BCC^{+}01, CMCD94, CGV^{+}02, SZ92, VM03, Whe91]. Extension [Hir90, LMJW93]. Extensions [Pad68, AYBDS06, CMR02, JD91, Kre01]. External [Pea82]. extra [CJ04]. extra-grid [CJ04]. extracting [WMSW01, Smi96]. extraction [KPF^{+}04].

fabric [RS98]. Fabrication [Ano65f, Ano65g, Bak65, Cal65a, Cal65b, Dzi65, Gor65, Loe65, Sha65a, PORG00]. face [ACO^{+}91, JB91]. faces [KSWC93]. facilitate [HGB05]. Facilities [McG77b, McG77c, McG77d, McG77e, Mye76, Str73, AT79, Cla89, GH91]. Facility [MLA83, SP78, Yel85, BEIW97, Cha89, JMNT97, Paz89, YS91, Got92, Jos81, MS88, Sam88, SZ88, Sor84]. Factors [Hir81, Tha84, Omi81, TC81, TKE01, WBSK87, Was93]. factory [Gri93]. failure [Caf08]. Family [PBv97, SB82b, Fal91, Omi81]. Fastfinger [Rob01]. fault [BRM02, HS87, Jen92]. fault-tolerant [HS87]. Feasibility [AG90, Elo95]. Feature [CH06]. Feature-based [CH06]. features [BM72, Dev67, WMSW01]. federation [HLR02]. feeds [DW00]. Field [Smi96, GJMT03, KK03, Zim96]. File [Buc63, SNC91, AC89, CFP^{+}95, Dei92, KBB^{+}97, MPR^{+}03]. Files [Fri70, ALE^{+}04, BCSW91, Dem88]. Filing [GB84]. Financial [Gor73, Dzi73, Kin73]. finite [CCO02]. First [Ano02a]. Fitting [BAI^{+}86, Fli97]. Five [Ano67j, Ano67l, Ano70l, Ano70m, Ano71j, Ano71k, Ano72j, Ano72k, Ano73k, Ano73l, Ano74m, Ano74n]. Five-year [Ano70l, Ano70m, Ano71j, Ano71k, Ano72j, Ano72k, Ano73k, Ano73l, Ano74m, Ano74n]. FLAVERS [CCO02]. flexibility [Hae03, Sam88, SD03, VCS^{+}08]. flexible [BKK^{+}02, GDD^{+}01, MBB^{+}07]. Flight [JD70]. floating [Pad68, Swe65, Woo70]. floating-point [Pad68, Swe65]. Flow [GY82, GBK83, Ste82, Ahu79, BG71, HMS92, Mil99a, Mon62]. flows [LR02b]. focus [KK03]. folding [DK01]. Following [TMF06]. font [GC88]. fonts [GC88]. footwear [PHBT00]. Force [Fle96, LPR^{+}07]. Forecasting [Ais73]. Foreword [Bra81]. Forging [May94]. Formal [Jen79, All71, FIS64, Neu71]. formalizing [SZ92]. Formatter [CBG^{+}82]. Forms [CF90]. formulation [BESZ02, Kol69]. FORTRAN [Hai65, KA71, Lar63, Rul68, SPS^{+}91, TPS^{+}88, RG85]. Forum [CS76, DK79a, DM76, Enc81, Hal77, Jac77, Lev79, Mor79, Rei79, WF77a, ABD^{+}97, Bau93, Ben94, Col95, Con97, Fah02, TG93, VW06, Ygl93]. foundation [HHM04, LVW08, OJG91]. foundations [JF91]. Four [Ano66i, IG04]. Four-Domain [IG04]. FramerD [Haa96]. Frames [KB00]. framework [ANPT02, Bun00, DDK^{+}01, EAB^{+}04, IMMB01, JK92b, KBB^{+}05, KHC02, LMX07, PRZ02, Ric00, SL95, SS94, SWS06, SZ92, TOV^{+}95, TRL^{+}08, Zac87, Zac99, Eve96, FV01, Ran02]. framework-based [KHC02]. Frameworks [ABD^{+}97, Rac96, vGLS98, Boh98, MBBD98, Pol98, UI00, van98]. Francisco [ABD^{+}97, Boh98, CHL^{+}00, Inm98, Pol98, RCB98]. Frederick [Cle95]. Free [BBDT84, MPS71]. Free-Storage [BBDT84, MPS71]. freight [WB69]. Frontier [JK92b]. Full [Cas06, Mau83, AYBDS06]. Full-Screen [Mau83]. Full-Text [Cas06, AYBDS06]. Function [AJ73, BFS67, HP93, Mac74, dPBM88, dP88, dM93, Bev00, CT96, Hed63, Jen92, WSG^{+}01]. Functional [AJ99, Bir73, LC01, Sch73, Cla66, HGW01, Mea66, Wit66]. functions [AFM03, GKT63]. fundamentals [Alt08]. furniture [OPH^{+}00]. fusion [Kee93]. Future [Hen86, JMW83, Mac00, Sen77, BCE^{+}96, CF05, Gay95, JN92, Rap04]. fuzzy [DHP02].

gained [PWBK90]. game [Orw00]. games [DMB^{+}06, Kes03, LLS^{+}06, MN06, SSR^{+}06, SR06, YC06]. Gamma [Bec95]. gap [Bun00]. GARDEN [dLMSC89]. Gary [Lor94a]. gathering [SH68]. gear [ApS64]. gear-train [ApS64]. Gemini [Mue67]. gene [MSZ^{+}01, IBM01]. GeneMine [LI01]. General [CGP68, McG77a, Per02, Ano73e, Blu64, BKP64, DCH^{+}03, EG64, Gor62, Gou69, HS65, Inm98, Vel64, Wah72, LeB08]. Generalized [Bjo75]. Generating [MOB^{+}08, Ahu68]. Generation [BM83, ER70, Han70, MM78, PMK^{+}06, Yag63, Baa91, BESZ02, BTB95, BC04, BFVY96, Chu63, DMB^{+}06, DDMS95, EFN^{+}02, Orw00, RWKN02, Sal88, SWS06, WA02a, ZMF02]. Generator [HLM^{+}75, Hei85, ABDD^{+}91, Dew90, FHP02, HDH88, Haz96, LGM69, Rec72, SSV64, ASW91]. generic [AASW98, YFS94]. Genesis [VR86]. GeneX [MSZ^{+}01]. genome [LI01]. genome/proteome [LI01]. genomic [DCB^{+}01]. genomics [HGW01, SD07]. geographic [Jac68]. Geometric [Jac68]. Geometry [AC68, Com68, Smi96]. Getting [BC91, LM96]. ghosts [RMDN03]. Gigabytes [Cra95]. Global [ALE^{+}04, GR01, IJM93, LeB08, Rom08, Unl08, ATJ92, DGK96, TMF06]. globally [MQR05]. Glossary [KPF^{+}04]. GNU [EGH^{+}05]. goes [BCH^{+}06]. Goldberg [Rac96]. good [HT06]. governance [VCS^{+}08]. government [Sim05]. GPSS [Gou69, HS65]. GPSS/360 [Gou69]. grammars [PMK^{+}06]. graph [MKGR07]. graph-theoretical [MKGR07]. Graphic [BEP80, BCHM80, Got92, Mus78, BM68, MC96]. Graphical [Enc81, WCG^{+}80, JK92b, Paz89]. Graphics [AC68, Ahu68, Ano68h, Ano68i, Ano68g, ADD68, App68, BM68, Bro68, CD68, Cla85, Com68, Cre68, DZ68, DW85, GTC68, Han80, Jac68, Joh68, Laf68, MPH68, MS80, NJ80, Oka68, RS73, Rul68, She80, Sym68, BBET71, Bur80, Got92, ZMF02]. grappling [Bou98]. grassroots [CBG^{+}07]. gray [LTL^{+}08]. gray-box [LTL^{+}08]. Greenland [AOFH92]. Greenland-Iceland-Norwegian [AOFH92]. GREENPRINT [BEP80]. Grid [Buc85, BGH^{+}04, CJ04, DHB^{+}06, JEF04, MPB^{+}04]. Group [Mil02, Per02, HP95, Klo08, Mil02, Wei05]. Growing [KRS^{+}98, WA02b]. growth [JRG^{+}01]. Guide [Ano96a, Dzi73, Pom72]. guided [Lor71]. Guidelines [DFHS90, DFS94, MDC^{+}05]. GUS [DCB^{+}01].

H [Ano96a, Cra95]. hacking [Pal01]. hand [LS05, SR96, Smi96]. hand-held [SR96]. handling [DGK96, Mat91, SSR87]. Harbingers [BCE^{+}96]. hard [GH03]. Hardware [Mac79, VMKB05]. Harkey [Ano96a]. Harmony [DHK^{+}08]. Harold [Ano96a]. Health [MV80, BBC^{+}07, EBKT07, ETD07, GB07]. health-care [BBC^{+}07, EBKT07, ETD07]. heart [WSG^{+}01]. held [SR96]. Helm [Bec95]. help [Cha89, WA02a]. help-desk [WA02a]. heterogeneous [MPR^{+}03, STF^{+}01]. Heuristic [Ryd70]. hidden [SDE^{+}06]. hiding [BGML96, BBG^{+}00]. Hierarchical [Col74, DM75, Hen69, Kee68]. Hierarchies [GL74, MGST70, CKB89]. hierarchy [HCGT90]. hieroglyphics [McI91]. High [AAC^{+}95, Aga85, Bro85, DMB^{+}06, DPR90, Han80, Mon67, PS83, Siw77, CBCS08, GBCW00, IFM^{+}01, LGCB95, LDL^{+}08, LSH^{+}08, Meg64, MMG^{+}00, PS99, RMP08, SSA^{+}95]. high-availability [LDL^{+}08]. High-Level [PS83, Meg64, PS99]. High-performance [AAC^{+}95, DMB^{+}06, LGCB95, MMG^{+}00]. high-performing [GBCW00]. High-Resolution [Han80]. High-speed [Mon67]. high-throughput [IFM^{+}01]. High-Transaction-Rate [Bro85]. highlights [CM88]. HIPO [Sta76]. Hippocratic [JG07]. histories [CCHJ94]. History [KM89, CFGS05]. HL7 [ETD07]. Holistic [DHK^{+}08]. holographic [Luc96]. HONE [Boo85]. hooks [Bou98]. Host [TLRH98, BGT96]. host-centric [BGT96]. hosting [HBBWS01, SBC^{+}08]. Hosts [GHMWB84]. hot [AHS^{+}03]. House [Ano02a]. hows [Bov96]. HPO [TBB^{+}87]. Human [Bra81, DP86, Omi81, Sta96, TC81, TKE01, AS99, Ber88, Due01, Fle96, GB00a, JL00, KSWC93, KK03, WBSK87, Hir81]. human-computer [Ber88, KK03]. Human-powered [Sta96]. human-technology [Fle96]. Hy [CEH^{+}94]. hybrid [BCH^{+}06]. Hyperlinked [BDCA00]. hypertext [Alf91, Bro91].

I-Patient [MV80]. I/O [AT79, HS03, Maj81, McN93]. Ian [Cra95].

IBCN [GFG^{+}92]. IBM [Ano96a, Klo08, Mil02, Unl08, Wei05, Woo02, AG90, AGMS89, AAC^{+}95, ABDD^{+}91, AJ05, ALPC88, ABD^{+}97, ABB^{+}03, AJ73, AJ99, BL95, Ben94, BBGB92, Bir73, Bla98, BPZ90, Boo85, BFGS98, BDE^{+}05, Buc86, CFGS05, CH85, CLR98, CLSW84, CW86, Cla85, CA95, CG77, Col95, Com90, Con97, CFP^{+}95, Cru84, DO84, DFHS90, DFS94, DH83, DFGS92, DPR^{+}02, DCH^{+}03, Fal91, FV01, GK07, Ghi85, GRW86, GR01, GL90, Got98, GL95, HJ84, HRL87, HZ84, Har03, HMS92, Hen86, Hum85, JBB03, JMZ92, JL00, JAS08, JA00, JRST07, KKS88, KM89, KPF^{+}04, LMPR07, LMPS02, MSB^{+}07, MS88, McC78, MLDF08, MS80, MPR^{+}03, MST^{+}03, MV80, MBC08, ÖCKM06, PGM^{+}92, Per02, PW88, RAK08, RG85, Rof01, RMH08, RCB98, SPS^{+}91, SSH88, Sam88, Sch73, SH87, SKA86, SHFG95]. IBM [SZ88, SOT^{+}00, Sym86, SM92, TLRH98, TAR85, TG84, TG93, TPS^{+}88, Tri91, Tuc86, VGW^{+}07, Vre03, Wal07, Wat79, WR92, YLG05, ZABM92, BS83]. ibm.com [KBK^{+}03]. IBS [vdS00]. ICAP [CLS88]. ICAP/3090 [CLS88]. Iceland [AOFH92]. idea [Pap00]. identification [GKT63]. identifying [GSBH^{+}89, PCFY07]. II [BCE67, Cal65a, DHB^{+}06, EW66, Lip68, McG77b, Men68, O’N80, SAAF73b, Ste64, Tal63, Vel64, Wit66]. III [Amd64, Blu64, BD66, Cla66, Dev67, Hed63, HS65, Lin80, Loe65, McG77c, Pad68, SAAF73c]. ill [KMSM^{+}07]. illustration [HAB^{+}92]. Image [MO90, EKS92, EDG^{+}96, Gla93a, Hel90, KLS90, Pic96, PB90, Rym90, Hak90]. ImagEdit [Rym90]. ImagePlus [AM90, AP90, DPR90, PWBK90, AGLR90]. imagery [SB00b]. Images [Som84, Ano85b, Cra95]. imaging [HCR^{+}07]. Impact [BBWL05, CGH^{+}05, Mat86, CCHJ94, Coc93, Coc99, GH03, SRCV06]. impairments [BFH^{+}05, MFJ05]. imperative [GB07]. Implementation [CLSW84, DZ68, GY82, PJ86, WT64, Bun00, FtCE87, GS04, MPB^{+}04, Nai95, PHBT00, SE08]. implementations [AAC^{+}95, Ste64, Yel03]. Implementing [EMMT78, FT92a, FT92b, GTB90, Lor94b, MQR05, SKG^{+}87, Tag90, Was93, CD68, PWBK90]. Implications [GJ87, Hae03, ML84, SD03, Fra79]. importance [Joh97, KM06a]. Improve [Bon69, Ste82]. improved [Gou69]. improvement [Ben94, BHC^{+}94]. Improving [BCF^{+}04, BDK^{+}81, BW93, BMK02, Gro85, HBC^{+}05, Hel81, HS08, KW86, LLS^{+}06, MBBD98]. IMS [Bla98, McG77a, McG77b, McG77c, McG77d, McG77e, SUW82]. IMS/VS [McG77a, McG77b, McG77c, McG77d, McG77e, SUW82]. in-box [SDE^{+}06]. In-process [BHC^{+}94, KPM01]. incidence [HRL87]. incidence-matrix-driven [HRL87]. incidents [Caf08]. increased [HB05]. Increasing [VCS^{+}08]. independent [AC89, SAAF73c]. Index [Ano93g, Ano69e, Ano69k, Ano70l, Ano70m, Ano71j, Ano71k, Ano72j, Ano72k, Ano73k, Ano73l, Ano74m, Ano74n, Ano81a, Ano81h, Ano84a, Ano84c, Ano88e, Ano94b, Ano94c, Ano94d, Ano67e, Ano67m, Ano68c, Ano68j]. Indexing [Cra95, Wag73, Ano74k]. Indices [Ano67k]. indigenous [Cav00]. individuals [BBWL05]. induction [JOZG02]. industrializing [BGL^{+}05]. industries [CF05, HBZ06, Mil08]. Industry [LK05, BW93, CF05, DF07]. Inertial [Ver96]. Inference [BAI^{+}86]. Influences [AJ73, AJ99]. informal [SKN06]. Information [Ano75k, BS83, Buc85, DL89, Dav82, ETD07, FL04, Hei85, Kee93, Len78, MMS^{+}04, JMP02, McG77a, McG77b, McG77c, McG77d, McG77e, MS85, Par82, RWKN02, San07, SRW04, SAAF73b, Spo99, Yel85, Ygl93, Zac82, BM99, BM96, Bou98, BGH^{+}04, BK63, BVP93, BW93, DeS81, DM88, EBKT07, ELM^{+}99, FC92, Gla93b, GSN^{+}01, GSBH^{+}89, HLLW72, Hei90, HV93, HV99, HBZ06, IJM93, Jac68, KPF^{+}04, KAN^{+}07, LA94, LR02b, LLO93, McD96, MLH87, ÖLL07, PBCK95, Res00, RM00, Ric00, SCO03, SL95, SAAF73a, SRCV06, SD07, SBE^{+}00, SZ92, SB95, TLRH98, VGW^{+}07, WB99, Zac87, Zac99, ZMF02, Zie79, DO84, Eve96, Ham82, SS94]. information-intensive [Gla93b]. Infrastructure [WND07, Ben01, BGH^{+}04, EBKT07, SD07, AM04]. infrastructures [MBC08]. inheritance [AOK^{+}95]. initiatives [Got98, GKB^{+}03, HS08]. initio} [SSH88]. Injecting [CFL05]. Innovation [Ken03, LK05, CGRW07, Mil08, NA07]. Innovations [Ano96a]. innovative [MQR05, WLB^{+}07]. input [FG90, Sar77, Yag63, ZS99]. input-output [Sar77]. inquiry [Dar70]. Inspections [Fag76, Fag99]. Installation [BPZ90, CCJ78, GJS75, Lai75, MM78, OPH^{+}00]. Instance [GGW02]. Instance-level [GGW02]. Institute [CH85]. Instructible [LM96]. Instruction [Hop87]. instrumentation [Abo99]. Integer [CCJ78]. Integrated [Bro91, Buc85, GG87, Jon85, New82, PMM83, Sta76, WAB^{+}07, ALPC88, BKP64, DL89, DCB^{+}01, Dor63, HSK^{+}01, Hed63, HS08, Kan88, Lar63, MSZ^{+}01, Tal63, Wei79, WT64, YS91, dLMSC89]. Integrating [BCH^{+}06, BDS88, PG05, EPTD05, SR06, dRW04]. Integration [BS04, JRST07, ÖCKM06, Wei05, BDL^{+}07, CKKS02, DDK^{+}01, HLR02, KHSW05, LR02b, JMP02, RWKN02, SRW04, STF^{+}01]. Integrator [SRW04]. Integrity [AMP76, DM75, McP74, CS76]. intellectual [Hua98, McI91, Sin03]. Intelligence [HG86, SJL^{+}04, BR02, CKKS02, ES05, ITIW07, WND07]. Intelligent [BRM02, Elo95, JRG^{+}01, PLGB90, Sow02, CF90]. intended [Sew67]. intensive [CF05, Gla93b, SSH88]. Interaction [WCG^{+}80, All99, AS99, KK03, PWF05, SB00a]. Interactive [AC68, Ahu68, Ano68h, Ano68i, Ano68g, ADD68, App68, BM68, BCHM80, Boi74, Bre71, Bro68, BL73, Bur80, CBG^{+}82, CD68, Com68, CMW87, Cre68, DZ68, EKS92, GTC68, GTB90, HD88, ITIW07, JD70, Jac68, Joh68, Laf68, LS87, Lue70, Mau83, MPH68, MLA83, Nae94, Oka68, RS73, Rei76, Rul68, SDP00, Sym68, Tet79, Tha81, Boi99, Hob72, OPH^{+}00, Par96, PHS^{+}00, Paz89, PDI^{+}00, PLGB90, Wim71]. interactive-batch [Hob72]. intercept [Vel64]. Interchange [SB82b, DeS81]. Interconnecting [BHWW83]. Interconnection [Asc86]. interconnections [SSY^{+}87]. interdomain [CJ04]. interests [Eva96]. Interface [Bot82, Ghi85, Syk88, AHS90, Ber88, Bot89, JK92b, Uhl88, Wat88, JDK^{+}91, VL00]. Interfaces [Ben86, HHP79, Mac79, DMRS88, Fle96, UI00]. interference [CW65, Fer71, Gay66]. interior [FT92a]. Interlingual [Haa00a]. Intermediaries [BM99]. Intermediary [IMMB01]. Intermediary-based [IMMB01]. Internal [PGM^{+}92, RM67, Woo69, VGW^{+}07, Ano70k]. International [CN79]. Internet [Ben01, CGHK98, Che01, HKLS03, KRS^{+}98, LK96, SVM^{+}03, vGLS98]. interoperability [DPR^{+}02, EBKT07, SRCV06, VL05]. interpretation [BHC^{+}94, Far87, Meg64]. Interpreter [ATC77, ASW91]. interpretive [BHS^{+}65, OJG91]. interpretive-execution [OJG91]. interrupt [Fah02]. intersection [Lar08]. intrabody [Zim96]. intrinsically [Gri63]. Introducing [van98]. Introduction [GGKS02, Gre79, GR06, Hel90, HW05, HM06, Kee67, LS07, Lic07, PS04, Raz93, SF05, Sab06, SSS04, Sym86, WG88, WR92, Zol07, BFSLT08, BMP08, GC88, MDC^{+}05, NS05]. Introductory [Mea66].

Investigating [BDG^{+}94]. investigation [BR08a]. Investigations [Kra78]. involvement [CFGS05, Pin89]. IP [BTB95]. IPS [KMM^{+}80]. ISBN [Ano96a, Ano02a, Rac96]. ISDN [Cav95]. ISO [FtCE87]. Isolated [Jen79]. Isolation [Ryd83]. Issues [Bux84, JR04, TB05]. item [Pie68]. Items [Day70, Gor71]. Iterative [GJMT03]. ITSM [CHPW07]. ITSM-based [CHPW07]. IV [BKP64, Dye80, Dzi65, LR67, Lar63, Pad64, SLW66, McG77d].

J [Lor94a]. J2EE [BCF^{+}04, KKL^{+}04]. Jalapeño [AAB^{+}00]. JANUS [CBG^{+}82]. Java [SOT^{+}00, ABLU00, BDF^{+}00, BFGS98, CFL05, CLR98, CHL^{+}00, DBC^{+}00, DAK00, EFN^{+}02, Got98, GBCW00, KJBH^{+}00, KNNL98, KNN^{+}01, Kre01, MMS98, MMG^{+}00, Pol98, RCB98, VL00]. JavaBeans [BCCN01, CDF^{+}98, TJ00]. JaViz [KJBH^{+}00]. Jeri [Ano96a]. JES2 [SP78]. jigsaw [Fli97]. Jikes [AAB^{+}05a]. Job [BHS74, Buc69, Cra78, SP78, Wit66, Bro68]. John [Bec95]. Johnson [Bec95]. Jose [Hir81]. Journal [DFHS90, DFS94]. Journalists [SBE^{+}00]. Journey [BCK^{+}94]. Jr. [Cle95]. Junior [SBE^{+}00]. Just [RM00, SOT^{+}00, LM96, Res00]. Just-in-Time [SOT^{+}00, RM00]. Justifying [SB00b]. JVM [DBC^{+}00]. JVMs [DAK00].

K2 [DCB^{+}01]. K2/Kleisli [DCB^{+}01]. K42 [AAB^{+}05b]. Kan [Ano96a]. Kaplan [Ano96a]. Kasparov [SF01]. Kaufman [Joh95]. keeps [LM96]. Keller [Wat94b]. Kenneth [Rac96]. Kernel [Enc81, HH66, AAC^{+}95, AAB^{+}05b]. Key [Che01, EMMT78, Mat91, Ben01, LMJW93, MLA91, WM93]. key-management [MLA91]. Keys [MM78]. kit [MMS00]. Kleisli [DCB^{+}01]. Knowledge [HG86, Hem90, HBR86, MRB01, PH01, PPDW01, Sym86, VR86, BSC^{+}01, Cav00, CKKS02, Due01, Elo95, FSSS01, GSBH^{+}89, Haa96, ITIW07, KM06b, LS91, NN01, Pru01, RS98, Sch03, Sin03, TKE01, UMN^{+}04]. Knowledge-Based [HBR86, VR86, Hem90, Elo95, GSBH^{+}89, LS91]. Knowledge-Systems [Sym86].

Lab [Haa00b]. Laboratory [Col74, Moo79, COC63, CG77, BDCA00, McC78]. Labs [RGH^{+}76]. Labs/7 [RGH^{+}76]. LAN [DDMS97, KBB^{+}97, BL95]. Land [BS94]. landscape [LPPC03]. Language [Ano68i, Cow84, ER70, McI91, Mil81, PS83, PMM83, RS85, Van76, Cha02, Com68, DM64, DM99, Fre66, McD96, Meg64, PS99, Tal63, Tho91, Wil91, Zlo77, Zlo82, SS78]. Languages [Mye76, Neu71]. Large [Ake89, BL76, ER70, Gar83, MCE^{+}86, PB90, WPSL01, BGP62, CC87, CHL^{+}00, CLS88, FS72, GCG^{+}04, Haa96, Hum85, KTBE05, Mon67, SSH88, Sma96, Spi66, Wim71]. Large-Language [ER70]. large-problem [FS72]. Large-scale [PB90, WPSL01, CC87, CHL^{+}00, CLS88, GCG^{+}04, Wim71]. large-systems [Hum85]. last [JCT^{+}97]. late [WGB99]. later [vdS00]. lattice [DK01]. laws [JG07]. layer [FtCE87, SD07]. leads [WPSL01]. learn [LS00]. learned [Rym90]. Learning [Ben00, BW93, Cav00, MQR05, PRZ02, SB00b, ZMF02]. led [NA07]. ledger [Inm98]. legacy [CdVF^{+}06, Fli97, Hes05]. Legal [PVK08]. LEGO [vdS00]. LEO [MLR03]. Lessons [CF05, NOR03, Rym90, BCK^{+}07]. Level [PS83, CGH^{+}05, DHP02, GGW02, Meg64, PS99, Str97]. levels [Lew76]. leverage [Alf94]. Leveraging [HV93, HV99, LDL^{+}08]. Librarian [Pra83]. libraries [Pic96]. library [AGMS89, Gri93, KA71, FT92a, FT92b, MS88, WR92]. Life [Kov01, Alt08, CK05, Ger00, HSK^{+}01, HBC^{+}94, MMS^{+}04, RR96, Swo01]. Lightweight [CDK^{+}05]. like [Ahu68].

Limiter [CW81]. Line

[BGH76, KW86, Dar70, Eva96, SBC97]. Linear [CCJ78, Dzi65, KRR80, FT92a, GKT63, JK92b]. linear-quadratic [JK92b]. lines [Bat06, Hal77]. Link [Moo83, DK74, Eis67, Gay95, Sin03]. Linkage [Mor78, BCFH65, FCM^{+}98, Lev79]. Linking [FSSS01]. Links [GBK83]. Linux [AJ05, AAB^{+}05b, BS05, GMTW08, HST08, TB05, VMKB05, Yeo05]. Linux-compatible [AAB^{+}05b]. living [Abo99]. Load [SM01]. loader [Hed63]. Local [Bux84, CHNS95, DSM83, Sum89]. Local-Area [Bux84, Sum89]. locality [SZ88]. localization [DGK96]. location [Mar63]. locking [BEIW97]. Logger [HBM^{+}04]. Logic [HG86, Alf91, All71, LVW08, Mon62]. Logical [Bak80, RH77, WI83, BB64, CdVF^{+}06]. Logically [Jen79]. Long [Cal65a]. Long-range [Cal65a]. look [AS99]. loosely [HKS08]. Lorin [Ano96a]. lost [Orw96]. lot [DeM68, Gor71, Men68, Pie68]. lot-sizes [Gor71]. lot-sizing [DeM68, Men68, Pie68]. LOTOS [BBD^{+}92]. Lotus [Bri98, PPDW01, Sch03]. Luther [Nae94].

M [Ano96a, Wat94b]. Machine

[BR02, HH79, Mac79, MS70, PRZ02, VL00, AAB^{+}00, CS76, DBC^{+}00, Fer71, GY67, GBCW00, JRD81, McG72, NOR03, PPTH72, ZMF02, AAB^{+}05a]. machine-readable [GY67]. Machines [Jen79, Mac79, SKN06, NOR03]. machining [Alm72]. Macro [Fre66]. magnetic [GH03]. mail [Fah02, Zlo82]. Mainframe [KLRP86, ABB^{+}03]. maintaining [HKS08]. maintenance [JRD81]. Makeup [She80]. Making [LD99, RHC^{+}06, SHLP05, KS03, MMS00]. Man [Cle95]. Man-Month [Cle95]. manage [CGRW07]. Managed [LD08, Par82, Gel89]. Management [AM04, AB92, Bak72, CBC^{+}06, CK05, DO84, EMMT78, FL04, FG79, Gar83, Got86, Ham82, Kre01, Lai75, MMS^{+}04, McG77a, McG77b, McG77c, McG77d, McG77e, PJ86, Per02, SFJ^{+}94, Sor84, Ste92, Sul83, TG93, Wat79, AvKM^{+}07, ACG^{+}95, Ano73e, AC89, ABB^{+}03, BFH^{+}94, BBE^{+}97, BCvE^{+}05, BBMA91, BH80, BDL^{+}07, BN89, BK63, BCD^{+}07, BCL^{+}04, Cen67, CHPW07, CGG90, Cla66, CKKS02, CFL^{+}06, DDMS97, DDK^{+}04, DHP02, Dye80, FSSS01, FK02, GFG^{+}92, Gla93a, Got92, HCGT90, Hen69, HAB^{+}92, Joh97, JHMO07, KJP03, KDA^{+}94, KBS94, Kan88, KBB^{+}05, KOH^{+}07, Kee93, Kes03, KAN^{+}07, KBC^{+}07, LPPC03, Lar08, LRS02, Lin80, LHJ^{+}07, MSB^{+}07, MQR05, MBB^{+}07, MLA91, Mil80, Mil99b, MPTW94, Moo88, NMB04, O’N80, ÖLL07, Owe95, Pru01, Qui80, RMP08, RCPR07, Raz93, RS98, RM67, RM88, RW04, Sch03]. management [Sha79, SH87, SCK02, SM92, THL^{+}03, TKE01, Tri91, VAH^{+}02, Wah72, WAB^{+}07, Wei79, Wit66, d’A89, DL89, DFGS92, GK07, JRST07, LMPR07, MSB^{+}07, Rac96]. Manager [LeB08, MOK78, Per02, CM88, Sag90, Wat88, Ano90f, FP90, LP74, Pau07, Sag90]. Managing [BFH89, DHPB03, DK79b, KTBE05, LC80, LA94, Rei79, RMDN03, Tem92, TG84, AEE^{+}97, BEF^{+}07, CCC^{+}08, CHPW07, NNH^{+}05, SJL^{+}04, Cra95]. manipulating [BM99]. Manipulation [HOST05, CFL05]. manual [JRD81]. manufacturing [Hae03, Haz96, SD03, VMKB05, Don03]. map [Kee93]. Mapping [CF05, WSG^{+}01, RHC^{+}06]. MARC [CM77]. market [FFK92, SJL^{+}04, SJL^{+}04]. Marketing [Boo85, HST06]. marketplace [JA00]. marketplaces [GOS05, SMK^{+}02]. Martha [Vet95]. Martin [Cav95]. Mary [Wat94b]. Mass [Gra78, Mis81]. massively [YC06]. Material [Orl73, Chu63]. materials [Fle96]. Mathematical [Men68, JN92]. matrices [Rob01]. Matrix [Loe65, BHS^{+}65, HRL87, Rad66]. mature [BCK^{+}94]. maturity [HST06]. Maximize [TG84, DHP02]. Maximizing [Alf94]. ME [ZZC^{+}08]. Measurement [Cal75, Fla85, Sta69, WM93, WF77b, BS83, Bar71, MBBD98, SWS06]. measurements [Smi96]. Measures [Att79]. Measuring [Buc69, Gla93b, Jon78]. Mechanism [GWM75, Mor78, GMTW08]. Media [BDCA00, LK96, Lip00, SDP00, Ger96a, Haa00b]. mediated [Mar99]. Medical [HG86, EDG^{+}96]. medium [MBB^{+}07]. MEDLINE [LRHH04]. MedTAKMI [ITIW07]. MedTAKMI-CDI [ITIW07]. Member [SB82b]. Memory [Bar75, Mat86, GMTW08, HG71, Nai95, PBS97, SH87]. Merge [Rad66]. Merge-sort [Rad66]. merger [Ber92a]. mesh [VSC^{+}03]. Message [Don03, Ken03, Klo08, Kov01, LK05, LeB08, Mil02, Per02, Rom08, San07, SSR87, Spe02, Unl08, Wei05, Woo02, d’A89]. messages [Hau68]. Messaging [vGLS98, BGERS08, DHK^{+}08, SMM^{+}00]. MetaCricket [MMS00]. metaprogramming [Bat06]. metering [AM04]. Method [NS84, WF77b, dL78, AGA^{+}08, EW66, FT92b, JK92a, SBC97, Whi01, AR76]. Methodology [Kra03, Coo80, MOCP85, Pal04]. methods [Bur96, FT92a, HM93, Loe65, NS05, PCFY07, Sel08, ZS99]. metric [WA02b]. Metrics [BBS02, Bur96, KPM01]. mice [Smi96]. Microprocessors [ML84]. Microprogram [Tuc67]. microprogramming [Dav72]. microtask [OIS^{+}06]. middleware [AJ05, CDK^{+}05, DHK^{+}08]. Migration [GL90]. Mike [Joh95]. mill [BKP64]. mind [CK98]. minded [VW06]. minds [NS05]. minidisk [Boz89]. Minimal [Woo69, Ano70k, Dim63]. minimize [Mar63]. mining [LI01, NN01, ORBdSB02, UMN^{+}04, WYB98]. mission [CDK08]. mission-critical [CDK08]. Missions [Qua70]. Mixed [CCJ78]. MMBase [BCvE^{+}05]. mobile [DEA99, GM03, KMSM^{+}07]. Mode [Cav95]. Model [ARNRSG06, BBCM06, BL76, Buc85, CC78, CBC^{+}06, CGP68, Dei82, GL74, HT06, Hei85, Kra78, Lar06, Lip68, Pad68, SNK^{+}08, ZLK^{+}07, AP90, Bat06, Bay98, Ber92b, BGL^{+}05, BDL^{+}07, BIJ06, Cha66, CBG^{+}07, CH06, ETD07, FHP02, Fer71, FS05, Fer91, Fer92, GB07, Hir90, HST06, KBB^{+}07, KHSW05, KBC^{+}07, MGM^{+}06, NA07, PV97, Rap04, Sar77, Sel06, Sel08, SAAF73c, SE08, SY71, SWS06, SBL^{+}94, WA02a, WKI03, Woo70, ZZC^{+}08, BR92, ETD07, EN07, HD88]. model-based [FHP02, Sel08, SWS06]. Model-driven [BBCM06, HT06, Lar06, SNK^{+}08, Bat06, BGL^{+}05, BIJ06, KBB^{+}07, KHSW05, KBC^{+}07, Sel06, ZZC^{+}08]. Modeling [AS82, App68, HKS71, Kan91, Rei76, Sea80, ABK^{+}06, ANPT02, AOFH92, BM68, BCK^{+}07, BS67, Bun00, Che69, Gor73, HBZ06, HAB^{+}92, Inm98, Kat90, LNH06, MM90, NNH^{+}05, Orw96, TRL^{+}08, Wat94a, WSG^{+}01, YC06]. modeller [BCG^{+}89]. Models [KRR80, VL05, WFS80, AW72, Bat06, Bov96, CdVF^{+}06, DF07, JEF04, Kin73, LMX07, Pic96, SR06]. modern [Lip96]. modifying [Ger62]. Modular [Gor80]. Moffat [Cra95]. molecular [SSH88, SM01]. molecules [Koi89]. Monitor [BCHM80, Bon69, Got92, Sud67, Tem92]. Monitoring [BGH76, KMSM^{+}07, BBC^{+}07, CT96, HR81]. Month [Cle95]. morphological [ESH89]. Morphologically [Wot93]. motion [Ver96]. motor [MFJ05]. mouse [LS05]. movement [RMC03]. Moving [JA00]. MPI [OIS^{+}06]. MPSX [SS78]. MPSX/370 [SS78]. multi [Pie68]. multi-item [Pie68]. multiagent [BSP^{+}02]. Multicast [BGG^{+}95a]. multidatabase [ABC^{+}95]. Multilevel [Bat06, BM68]. Multimedia [Bov96, DDMS97, KBB^{+}97, Moo90, SB00b, ABLM95, PRZ02, PWF05, RC91, WLE^{+}97, Nae94]. multiplayer [YC06]. Multiple [AS82, BHR89, DW00, GM79, GSN^{+}01, SGS89]. Multiple-System [GM79]. multiplicity [SM79]. Multiport [Syk88]. Multiport/2 [Syk88]. Multiprocessing [ACM74, HPSS79, Mac74, BGP62, BCE67, Dev67, Kee67, LR67, Sew67, Smi63, Sud67]. Multiprocessor [TB84, GMTW08, HMS92, KKS88]. multiprocessors [AGMS89, SPS^{+}91]. Multiprotocol [HLM95, PTE^{+}95]. multipurpose [HZ84]. Multistream [ZS99]. Multisystem [Bla64]. multitasking [Hav68]. Multithreaded [EFN^{+}02]. multivendor [Moo88, Ste92]. museum [SDP00]. music [JF91]. MVS [Ber78, CC78, Cla89, CM77, GTB90, LS91, MCE^{+}86, Rub89, Sch80a, Vos92]. MVS/ESA [Cla89]. MyMed [LRHH04]. Mythical [Cle95].

name [BMMR96]. namespace [ALE^{+}04]. Naming [ZABM92]. NAS [AAC^{+}95, Nai95]. National [Gra78]. native [BCH^{+}06]. Natural [Mil81, HKOS96]. Navigating [Sma96]. navigation [PWF05]. Navigator [DLP^{+}05]. NBBS [ACG^{+}95, BGG^{+}95b, MICG95, Owe95, PBCK95, TOV^{+}95]. Near [Zim96]. Near-field [Zim96]. Net [Orl73]. NetView [ACG88, CBH^{+}92, Got92, Tem92]. NetView/6000 [CBH^{+}92]. NetView/PC [ACG88]. NetVista [KRS^{+}98]. Network [Ahu79, AR76, Ano76j, Cul76, DSM83, EHJ77, Gar83, Got92, Gre79, Har80, Hob76, Irl92, McF76, MK73, Moo79, PJ86, SP78, Sul83, BGT96, BGG^{+}95a, CT96, EW66, Hob72, HAB^{+}92, Kan88, LGCB95, Owe95, PBCK95, RM88, SE08, SM92, Wei79, WB69, BS83, Got86, Hob76, Joh95]. network-centric [BGT96]. networked [Zim99]. Networking [Cra78, GM79, MLA83, CA95, HLM95, GCF^{+}87, IKS^{+}95, PTE^{+}95, SKG^{+}87]. Networks [BHWW83, Bux84, Fre73, GG87, GBK83, HHP79, Mar83, Ryd83, Sch87, Sul83, WI83, Alf91, AB92, BFH89, BR08a, BRM02, CHNS95, Dim63, HDS03, HS00, JMZ92, Lip00, Mon67, Moo88, OSK^{+}95, SVM^{+}03, Ste92, Vet95, Zim96]. neural [Alf91]. Neutron [Cre68]. News [ES00, BCE^{+}96, Eva96, GB00b, KB00]. Next [DDMS95, BTB95, BC04, DMB^{+}06, Sal88, Sin03]. NIL [PS83, PS99]. No [GM03]. node [BGG^{+}95b]. nonlinear [Mug62]. North [AOFH92]. Norwegian [AOFH92]. notation [Ive63, Ive99]. Note [FKS^{+}03, Ger62, Gom62, Hei63, AGMS89, ACO^{+}91, Bev00, Bla98, DCH^{+}03, HKOS96, LD99, LT89, LM93, dMR03, Mil99a, Pol98, Ric00, vdS00, Wat94a, Zha98]. Notes [Gin65]. Notification [NG05]. number [LGM69, Rec72]. Numbers [Ano65e, HLLW72]. Numeric [HG86]. Numerical [Alm72, HMS92, MMG^{+}00]. numerically [SSH88]. Nutt [Lor94a]. NWChem [SM01].

O [AT79, HS03, Maj81, McN93]. OASIS [NG05]. Object [Bec95, Hak90, HCGT90, Rad96, SFJ^{+}94, SC97, TK89, Alf94, BBN96, Bov96, CCHJ94, Coc93, Coc99, EDG^{+}96, Fli97, HP95, Hen98, LH97, McD96, Par96, PV97]. object-orientation [Coc93, Coc99]. Object-Oriented [Bec95, TK89, CCHJ94, EDG^{+}96, Fli97, HP95, McD96, PV97, SC97]. Objectives [Hum85, BCL^{+}04]. Objects [Ano96a, Rac96, HM93, PBv97, RLPG96]. ODC [BMK02]. Office [BBC82, Gar81, GB84, LCS82, Moo83, Som84, Ano85b, DeS81, EGLT79, GH79, Zlo82]. Office-by-Example [Zlo82]. OLAP [CMR02]. older [MDC^{+}05]. On-Demand [TLRH98]. On-Line [KW86, Dar70, Eva96, SBC97]. one [BHR89, Bou05]. Online [SR06, DMB^{+}06, LLS^{+}06, SSR^{+}06, YC06]. ontologies [SSM^{+}04]. ontology [AT06]. OPAS [LCS82]. Open [Asc86, Bou05, Col95, AAB^{+}05a, AAB^{+}05b, BCvE^{+}05, CFGS05, Cyp92, HBC^{+}05, NS05, OCK^{+}92, Sim05, YLG05, FFK92, MSZ^{+}01]. Open-source [Bou05, AAB^{+}05a, AAB^{+}05b, BCvE^{+}05, CFGS05, HBC^{+}05, NS05, Sim05, YLG05]. Opening [NS05, JMZ92]. operated [COC63]. Operating [AMP76, AJ73, Bir73, Kra78, Lor94a, McP74, RGH^{+}76, Sch73, SS69, Ano63b, AAB^{+}05b, AJ99, BHR89, Cla89, Dar70, Dor63, FSSS01, Hed63, Kie79, Lar63, LS87, LMPS02, Mue67, SRM^{+}98, SH68, Tal63, WT64, CRTW88, KR88, Miz88]. operating-system [AAB^{+}05b]. Operation [Har80, McG77a, Irl92]. Operational [KLS90, Sta69, BCK^{+}07, LR67, NC03]. Operations [MCE^{+}86, Ano65f, Ano65g, Bak65, Cal65a, Cal65b, CNL08, Dzi65, Gor65, Loe65, Sha65a]. Operator [Ant75]. Opportunities [ES05, Eva96, SCK02]. opportunity [KLS90]. optimal [EW66, Mug62]. Optimization [Gha75, WR92, ATJ92, BEH^{+}06, BG71, JK92b, FT92a, FT92b]. optimize [LDL^{+}08]. Optimized [AvKM^{+}07]. optimizer [MLR03]. Optimizing [MMS98, RG85, Ryd74, AAB^{+}07, HS00]. Optimum [Kos63a]. options [JAS08]. orbitals [SSH88]. order [Cal65b]. Orfali [Ano96a]. Organization [Buc63, Mat86, Bla64, CGP68, EN07, GTB90, Gla93b, HST06, Raz93, SB97, SAAF73b]. organizational [BBWL05, LS01, Mar01]. organizations [BVP93, CJ04, HV93, HV99]. organizing [BDL^{+}07]. orientation [CGH^{+}05, Coc93, Coc99]. Oriented [Bec95, Jon77, AGA^{+}08, ABW^{+}08, BBWL05, BCE67, BDE^{+}05, CCHJ94, CC05, CK05, CBC^{+}05, CDK^{+}05, Dev67, EDG^{+}96, FS05, Fli97, HOST05, HP95, KKLR05, KOH^{+}07, Kee67, KBC^{+}07, LR67, McD96, NLN^{+}07, NG05, PV97, SHLP05, Sew67, SC97, Sud67, TK89, VL05, Wal07]. originality [LHT^{+}07]. origins [WGB99]. OS/2 [CM88, Fer91, Fer92, Wat88]. OS/360 [Cla66, Mea66, Wit66]. OS/390 [DDMS97, DBC^{+}00, KBB^{+}97]. OS/VS1 [BHS74, Whe74]. OS/VS2 [Ber78, LP74, McP74, Sch73]. OS/VS2-2 [Sch73]. OSGi [GHM^{+}05]. OSI [JMZ92, SSY^{+}87]. OSI-SNA [SSY^{+}87]. OSL [Min92]. Other [Mac79, Bou05]. Outline [BB64]. outlines [Bak65]. Output [Bon69, Sar77]. Overview [Bro80, HJ84, Lor86, MICG95, MHMM75, PSLB97, SOT^{+}00, Sul83, Sum84, Tuc86, AG90, ACG^{+}95, Ano76j, ABD^{+}97, BS04, BAC^{+}81, BAC^{+}99, BDE^{+}04, CM88, GK07, Got98, Kah84, KBS94, LS87, McF76, ST89, ST99, Tod76, Wat88, YS91].

P [Cle95]. package [Rul68]. Page [She80, TBB^{+}87]. page-swapping [TBB^{+}87]. paging [SY71]. pair [Rob01]. pairing [ATJ92]. PAM [ALPC88]. PAM-CRASH [ALPC88]. panel [HRL87]. paradigm [Hei90, Zim99]. paradigms [BHW99]. paradox [LD08]. Parallel [BFG^{+}95, CFP^{+}95, JAS08, KDY97, Nai95, TPS^{+}88, Wil91, AAC^{+}95, BARM97, BEIW97, CLS88, SL67, SHFG95, WB95, Joh97, NMCB97]. Parallelism [MPTW94]. parallels [Ygl93]. Parameter [Gor65]. Parameters [KL74, Ano74j, Ano74l]. PARC [WGB99]. parsers [PMK^{+}06]. Part [Amd64, AJ99, Bak65, BB64, Bla64, BCE67, Blu64, BKP64, BD66, Cal65a, Cal65b, Cla66, CGP68, DeM68, Dev67, Dor63, Dye80, Dzi65, EG64, EW66, Gay66, Gor65, Hed63, Kee67, LR67, Lar63, Lin80, Lip68, Loe65, MW66, McG77a, McG77b, McG77c, McG77d, McG77e, Mea66, Men68, Mil80, Mil99b, O’N80, Pad64, Pad68, Qui80, Res00, SLW66, Sea66, SAAF73a, SAAF73b, SAAF73c, Sew67, Sha65a, Ste64, Sud67, Tal63, Vel64, WT64, Wit66]. part-period [DeM68, Men68]. Partial [PVW91, HH66]. partner [HFB^{+}08]. parts [Bau93]. Pass-Through [MLA83]. Past [Hen86, JMW83, Sen77]. path [Maj81, TOV^{+}95]. paths [Mon67]. Patient [MV80]. patients [KMSM^{+}07]. pattern [GKT63, LC01, MKGR07]. Patterns [Bec95, Mil08, PY08, Baa91, BFVY96, CN08, HMP02, Hes05, HBZ06, MGM^{+}06, Pau07]. Paving [SB95]. payment [AASW98]. PC [SH87, ACG88, Com90, Ghi85, HRL87]. PDAs [LC03]. PDM [MOCP85]. pedagogy [Pap00]. Peer [GCF^{+}87, SKG^{+}87]. Peer-to-Peer [GCF^{+}87, SKG^{+}87]. Penetrating [AMP76]. People [Dea82, ES05, KTBE05, LS05, MDC^{+}05]. Performance [Aga85, BS83, Bar71, Bar73, Bar75, Ber78, Bon69, Buc69, Bux84, Cal75, CLSW84, CC78, CBC^{+}06, CW86, DPR90, Dru69, Gro85, Kra78, Man74, Mor73, Sea80, Siw77, SA69, Ste79, TG84, TB84, AAC^{+}95, ABLU00, AJ05, BCF^{+}04, BDH75, CHL^{+}00, CC89, DMB^{+}06, DHPB03, DAK00, EOO^{+}06, HS00, JRD81, LGCB95, LS01, Mar01, MS75, MMG^{+}00, PBS97, Sam88, SG98, Sch80b, SY71, SKA86, SDP00, SSA^{+}95, Tem92, TB05, WA02b, WBC04, YC06]. performing [GBCW00]. period [DeM68, Men68]. Perlman [Joh95]. persistence [SC97]. Personal [Cla85, KMT85, Mur84, RG85, Rym90, TAR85, BBC^{+}03, CG77, Haa00b, HS03, Ive91, KMT86, KAN^{+}07, d’A89, BBC^{+}07, Zim96]. Personalization [PWF05, Ben96]. Personalizing [KBK^{+}03]. Perspective [CFS81, Dru69, Hop87, AR76, Aro67, Bra79, DK74, GCF^{+}87, GH79, KBS94, Rad96, Sch83]. Perspectives [RC91, Shu89, Shu99]. pervasive [AS99, DEA99, HBBWS01, LD99, Mar99]. petaflop [IBM01]. Peter [Wat94b]. Peterlee [Tod76]. pharmaceutical [BGL^{+}05, CF05]. phased [HLT94]. phases [Dav93, Dav99]. Philosophies [Sch73, Whe74]. philosophy [Cha89]. phonetic [Wot93]. Physical [Sme78]. Physically [PDI^{+}00]. Physics [Ger96a, MPH68]. piecewise [JK92b]. Pierson [McQ94]. PL [Sym68]. PL/I [Sym68]. place [SCO03]. places [Spo99]. Planning [Buc85, Dei82, Lan73, NORS80, Orl73, SB82a, Sme78, Bro80, Cal65a, Coo80, Dzi65, Dzi73, EKS92, HKS71, Sha65a, Zac82]. Planning-data [Lan73]. plans [SHP69]. plant [Omi81]. platform [BS05, GHM^{+}05, LGCB95, MBB^{+}07, MLDF08, RAK08, SMK^{+}02, Sel08, SSR^{+}06, STF^{+}01, ZZC^{+}08, dRW04, BDE^{+}05]. platforms [PGM^{+}92]. Playback [NS84, SR96]. playing [YC06]. plotter [Bre65]. point [FT92a, Pad68, Swe65]. points [All99, GS99, Gol99, Mar63, WB99]. Policies [WFS80]. Policy [ACGL04, Gha75, ABB^{+}03]. Policy-based [ACGL04, ABB^{+}03]. polymorphic [Wel65]. pool [BD66]. portable [GM91, BAI^{+}86]. portal [YLG05, SJL^{+}04, WRS04]. portals [MRB01]. portfolio [JK92b]. Porting [AG90, BPZ90, GL90]. Possible [JMW83, SS78]. Post [Lip96]. Post-modern [Lip96]. postgenomic [WPSL01]. Potential [Fra79]. Power [KMT85, Lai75, MLDF08, CBG^{+}07, KKL^{+}04, KMT86, Pap00, WB95]. powered [Sta96]. pp [Ano02a]. Practical [SS01, Wat94b]. practice [CN08, GR01, LS01, Mar01, MB96]. practices [Dye80, KAN^{+}07, Lin80, Qui80]. precedences [Gor71]. precision [KA71]. predicate [TVPL94]. predicate-based [TVPL94]. predictable [HYJK06]. Predicting [PBS97, Sea80, VPH^{+}08, WA02b]. prediction [SF01]. Predictive [VAH^{+}02, ANPT02]. Preface [BHS^{+}05, BR03, BR07, BR08b, DHS89, Ger85a, Ger85b, Ger86a, Ger86b, Ger86c, Ger87a, Ger87b, Ger87c, Ger87d, Ger88a, Ger88b, Ger88c, Ger88d, GBR06, HIP^{+}03, Hof89a, Hof89b, Hof89c, Hof90a, Hof90b, Hof90c, Hof90d, Hof91a, Hof91b, Hof91c, Hof91d, Hof92a, Hof92b, Hof92c, Hof92d, Hof93a, Hof93b, Hof93c, Hof93d, Hof94a, Hof94b, Hof94c, Hof94d, Hof95a, Hof95b, Hof95c, Hof95d, Hof96a, Hof96b, Hof96c, Hof97a, Hof97b, Hof97c, Hof97d, Hof98a, Hof98b, Hof98c, Hof98d, Hof99a, Hof99b, Hof99c, Hof00a, Hof00b, Hof00c, HRB02, JSR07, Lac81, Lac82a, Lac82b, Lac82c, Lac82d, Lac83a, Lac83c, Lac83d, Lac83b, Lac84a, Lac84b, Lac84c, Lac85, MSSR08, McQ75, McQ76a, McQ76b, McQ76c, McQ76d, McQ77a, McQ77b, McQ77c, MSR08, NSS^{+}06, PWSR07, RD01, RB01a, RB01b, RB02b, RD02, RB02a, RD03, RB04a, RS04a]. Preface [RS04b, RB04b, RS05, RB05a, RB05b, RB06, RS06, RB07, SR08, Thi77, Thi78c, Thi78d, Thi78a, Thi78b, Thi79a, Thi79b, Thi79c, Thi79d, Thi80a, Thi80b, Thi80c, Thi80d, Thi81a, Thi81b, Thi81c, VRS03]. premier [LMPS02]. Preparing [RR96, RW04]. presence [KTBE05]. Present [Gay95, Hen86, JMW83, Sen77]. Presentation [NJ80, Moo90, Paz89, dM93, FP90]. presented [Caf08]. President [Don03, Ken03, Klo08, Kov01, Mil02, Rom08, San07, Spe02, Unl08, Wei05]. Presidents [LK05]. Prevention [GTB90, Jon85, New85, MJHS90]. Price [Pal04]. Price-at-Risk [Pal04]. Pricing [Gha75, BKK^{+}02, Pal04]. Pricing-Priority [Gha75]. primary [LC01]. primer [San81]. Principals [GB84]. Principled [Mil99b]. Principles [Joh68, KKL06, Mil80, Hel69]. print [PWF05, dM93]. print-disabled [PWF05]. Printer [She80]. Printers [MDHM86]. Printing [HP93, dPBM88, dP88, dM93]. priorities [Hau68]. Priority [Gha75]. Privacy [Sch03, SS01]. Private [GG87, Joh95, KHC02, SMK^{+}02]. proactive [WPT03]. probabilistic [ANPT02, HAB^{+}92]. Probability [All74]. probing [BRM02]. Problem [CCJ78, For78, Fri70, Kol69, New85, WI83, ApS64, FS72, Gom62, Pie68]. problems [BBET71, CLS88, MW66]. procedural [van98]. Procedure [LCS82]. Procedures [Hir81]. Process [Fla85, HB85, Jon85, RROC85, RHMP85, SB93, BHC^{+}94, BCK^{+}94, BARM97, Bro80, CDM^{+}04, HBC^{+}94, Hum85, KPM01, KKL^{+}04, KHSW05, LRS02, LTL^{+}08, LMX07, Pau07, RHMP99, GTB90, Got86, MB96, Pau07]. Process-Integrated [Jon85]. Processes [Cha70, SG98, BGL^{+}05, FSSS01, HYJK06, JL00, JRST07, KKL06, LA94, Mug62, Sch93]. Processing [AS82, Aga85, AC68, Ahu68, AT79, Amd64, Ano68h, Ano68i, Ano68g, ADD68, App68, BM68, Bro68, CD68, Com68, Cre68, Dav78, DZ68, FG79, GTC68, GRW86, GHMWB84, Har80, Jac68, Joh68, Laf68, MPH68, McG77c, McG77e, Ohm86, Oka68, Rul68, Sch78, Sym68, AEE^{+}97, Bro80, CLS88, Gri63, Hir90, Kie79, Loe65, Lor79, MOB^{+}08, RS72, RC69, Sch88, Sch99, SE08, Tur63, WT64, ZLK^{+}07, Zlo82, CF90]. Processor [LC80, Maj81, Syk88, TB84, BHR89, McN93, OIS^{+}06, SH87]. producers [SBE^{+}00]. producing [Haz96]. Product [Cla81, Rof01, SL75, Bat06, Bla98, DP95, FS05, GMM^{+}06, Kan88, Dew90, HDH88]. product-based [DP95]. Production [Bak72, Cal65b, Bro80, Dzi65, HSY01, Sha65a]. Productivity [Fla85, Jon78, Jos81, Lam84, RMH08, Ste82, Str75, Tha81, Tha84, HYJK06, PV97]. Products [DH83, Sym86, DPR^{+}02, Kah84, Got86]. Professional [DW85, RG85]. Professionals [Hel81]. Profile [Lai75, SRM^{+}98]. Profile-directed [SRM^{+}98]. Profiler [VL00]. profiling [KJBH^{+}00]. profits [DHP02]. Program [ALP83, BL76, CM77, Cor89, Fag76, GHH^{+}83, HLM^{+}75, HG71, JD70, Lam84, Mor73, Pow83, Rog83, Ryd74, SZ88, Sta76, Van76, ABDD^{+}91, BESZ02, BHS^{+}65, Bro91, BDG^{+}94, Cle89, CMW87, DMMR92, Dev67, EFN^{+}02, Fag99, LR67, Lew66, McC78, McG65, Mon62, Mug62, O’N80, Rym90, Tal63, AGLR90, Hob76]. Program-To-Program [GHH^{+}83]. Programmable [RMSS96]. Programmer [Bak72, Tha84]. Programming [AT79, ACM74, Bak72, Ben94, BDK^{+}81, CCJ78, FS05, FG79, Fla85, Gor71, Ive63, Ive99, JDK^{+}91, Jon78, KRR80, Mil81, Mye76, PS83, PMM83, RROC85, RHMP85, Rog75, RS85, Sam88, Ste63, WF77b, Wil86, Alf91, All71, Bay98, CC05, Dor63, Dzi65, Fer91, FT92a, FMR02, GTB90, Hed63, HK65, JK92a, JN92, Lar63, Min92, MMG^{+}00, Neu71, OIS^{+}06, Orw00, PS99, Pom72, RHMP99, Sal88, Sha79, Shu89, Shu99, Tal63, TK89, WT64]. Programs [BEP80, KMM^{+}80, Qua70, Sew67, dL78, All71, BS67, FP90, Gin65, Kee68, MMS98, YFS94]. Project [DF63, Rac96, AAB^{+}05a, ABD^{+}97, ABC^{+}95, BDG^{+}94, DMMR92, Dim63, OCK^{+}92, Raz93, RCB98, AR92, Pra83]. projects [AAB^{+}07]. Prolog [BBGB92, Wil86]. promise [WB95]. prompted [Wat88]. proposal [Mil99a]. proposed [EPTD05]. proprietary [Bou05]. proproceptive [Ver96]. Prospects [CHPW07, GG87]. protection [JG07, MN06]. protein [DK01, IBM01, JRG^{+}01, Rob01, SF01, WMSW01]. proteins [LC01]. proteome [LI01]. Protocol [GJ87, MK73, DDK^{+}01, CGHK98, Che01, SVM^{+}03]. Protocols [Gre79, PTE^{+}95]. prototype [TBB^{+}87]. Prototyper [van00]. Prototyping [Ber84, VM03]. Providing [BS05]. provisioning [ACGL04]. Prycker [Cav95]. PS [HD88]. PS/2 [HD88]. pSeries [JBB03]. pseudo [ABDD^{+}91]. pseudo-random [ABDD^{+}91]. pseudorandom [LGM69]. Public [HHP79, Joh95, Ben01, LMJW93]. Publications [BDK^{+}81]. Pulsar [ABW^{+}08]. purpose [Blu64, BKP64, EG64, Gor62, Gou69, HS65, Vel64]. Putting [JB91]. puzzle [Fli97, TKE01].

quadratic [JK92a, JK92b]. Quake [DHB^{+}06]. Quality [Ano96a, BGH76, CH85, Jon78, Ben94, CCHJ94, HR81, HS08, Kan91, KDA^{+}94, KBS94, WM93, Wat94a]. quantifying [GKB^{+}03]. Quantitative [Nor93]. quantumcomputer [MBG00]. query [Cha02, MLR03, Wat88, Sor84, Zlo77]. Query-by-Example [Zlo77]. queue [Sha65b]. queued [BD66]. Queuing [All75, Cha70, Dei82, Rec72, Ano75j, Cha66].

R [Lor94b, Wat94b, BAC^{+}81, BAC^{+}99]. RACE [DMMR92, OCK^{+}92]. Radia [Joh95]. radiation [EKS92, HD88]. Radio [ES00]. Ralph [Ano96a, Bec95]. Random [Ano02a, BM83, ABDD^{+}91, BESZ02, Hei63, Rec72]. range [Cal65a]. Rapid [Hou93]. Raster [She80]. Rate [Bro85, Jen92]. Rational [BDE^{+}05, BIJ06, LNH06]. Rationale [Bot82, Bot89]. Re [Bri90, Dav93, Dav99, OT94]. RE-Analyzer [OT94]. Re-engineering [Bri90, Dav93, Dav99]. reactive [HMS92]. read [GMTW08]. read-copy-update [GMTW08]. readability [Hal77]. readable [GY67]. reading [Ano80h, Ano80i, Ano81i, Ano81j, Ano81k, Ano81l, Ano82i, Ano82j, Ano82k, Ano83b, Ano83c, Ano84d, Ano84e, Ano86b, Ano87d, Ano87e, Ano95b, Ano99d, Ano01c, Sti94]. Readings [Dav72]. Real [Aro67, BG71, Bur96, Mac79, NO70, RGH^{+}76, ABLM95, BM72, BGERS08, Gin65, GMTW08, HR81, HST08, MOB^{+}08, Mue67, PY08, SHLP05, SL95, Sel08, SH68, Wit63, HST08]. Real-Time [NO70, RGH^{+}76, Aro67, BG71, Bur96, ABLM95, BM72, BGERS08, Gin65, GMTW08, HR81, MOB^{+}08, Mue67, PY08, SL95, Sel08, SH68, Wit63, HST08]. real-world-aware [PY08]. Realizing [BDE^{+}05]. Realtime [Syk88]. REASON [PLGB90]. reasoning [MMS^{+}02]. recognition [KSWC93]. recognize [Pic00]. Recommendations [DH83, SSR87]. reconfigurable [WKI03]. reconfiguration [JBB03]. Record [GB07, Str97]. record-level [Str97]. records [Gla93a]. Recovery [Cru84, BCSW91, CBCS08, LSH^{+}08, BBMA91, CM77, MSS91]. Reduce [Fag76, Fag99]. Reduced [Hop87]. redundancy [SL67]. Referees [Ano81g, Ano85c, Ano89e, Ano93h, Ano97d, Ano01b, McQ77d]. Reference [BFH^{+}94, ETD07, HRL99, LG99, MMS98, SR93]. Reflections [MLA83]. Registers [Day70]. registration [ZABM92]. Related [DH83]. Relational [BE77, CMR02, Tod76, BCH^{+}06, CMCD94, FCM^{+}98, FKS^{+}02, FKS^{+}03, Hir90, Kah84, MPTW94, RLPG96, San81, THL^{+}03]. relationship [GOS05, RW04]. relationship-based [GOS05]. relationships [Jac68, WSG^{+}01]. Release [Got92, LP74, AGMS89]. Reliability [HPPS75, Mis83, BFSLT08, Wat94a, Wel65]. reliable [Bou05]. relocation [McG65]. Remote [BBC^{+}07, Irl92, Kra03]. reorganization [SBC97]. repair [CNS^{+}08]. Replacement [Ryd74, Bel66]. Replication [RMH08]. reply [DM76, WF77a]. reporting [RCPR07]. Repository [Ano90f, Sag90]. Representation [BEP80, HG86]. representations [SAAF73c]. Representing [Haa96, FG90]. Reprints [Ano76k, Ano77l, Ano78m, Ano82h, Ano90g, Ano75k, Ano78l]. reputation [Kes03]. request [BD66, LHJ^{+}07]. request-queued [BD66]. Requirements [Bjo75, Chu63, Dav82, Orl73, SSS^{+}87, WFS80, BFH^{+}05, CK98, DD64, DP95, MOCP85, Odo90]. Research [AAB^{+}05a, Ken03, Spe02, BGL^{+}05, BDCA00, DMMR92, GMM^{+}06, GH79, HCR^{+}07, LRHH04, JMP02, MKGR07, PVK08, SBL^{+}94, WGB99, CG77]. residue [Rob01]. resiliency [Caf08]. Resilient [SBC^{+}08, VSC^{+}03]. Resolution [Han80]. Resource [BBMA91, CW81, MSS91, ABW^{+}08, BSC^{+}01, DDMS97, LA94, PH01]. resource-control [ABW^{+}08]. Resources [GWM75, LP74, CHPW07, GC88, JL00]. respond [KBB^{+}05, KBB^{+}07, Pic00]. Response [Lam84, DK79a, Kos63a, Mor79]. Responsive [BMP08]. restart [BARM97]. restructuring [HG71, SRM^{+}98]. results [BBS02]. Retail [Ant75, BDH75, HPPS75, MHMM75, Ste93]. Retrieval [Jon77, Jac68, PRZ02, RM00]. Reusable [Bec95, Bau93]. reuse [Gri93, Lar06, PCH93, TG93, Was93, Ygl93]. reusing [HRL99]. Reverse [Aik98, BDG^{+}94]. review [Bon84]. Reviewed [Rac96, Ano96a]. Revisited [BBDT84]. Revolution [Bie97]. Revolutionary [SRCV06]. REXX [Cow84, Hoe89, PVW91]. RF [BL95]. rhythm [SLD89]. rich [PWF05]. Richard [Bec95]. Right [Ano96a, SCO03]. rights [Fah02]. Ring [DSM83]. RISC [ABDD^{+}91]. Risk [Ano96a, AvKM^{+}07, BEF^{+}07, JK92b, Pal04]. Robert [Ano96a]. Robotics [KIS87]. RODM [FC92]. Role [Ant75, EHJ77, Gro92, Hob76, NORS80, RCPR07, CGG90, Col95, FSSS01, GOS05, HAB^{+}92, RW04, Sea66, SSM^{+}04, YC06]. role-playing [YC06]. roles [CFL05, McN93]. ROMP [SH87]. route [Baa91]. Routing [Ahu79, JMW83, Mar83, Vet95]. RT [SH87].

Rubin [Rac96]. rule [JOZG02]. Running [DHB^{+}06, SZ88]. runtime [CFL05, CNL08]. runtime-adaptable [CNL08].

S [Rac96, KDY97, NMCB97]. S*P*A*R*K [GSBH^{+}89]. S/390 [KDY97, NMCB97]. SAA [Ber88, BDS88, Dei92, Dem88, DMRS88, Rei88, Sch88, Wol88]. Sales [Unl08, LPR^{+}07]. sales-force [LPR^{+}07]. Salient [MB96]. Sam [Ano02a]. Sampling [Tet79, SHP69]. San [Hir81, RCB98, GFS03, MPR^{+}03, ABD^{+}97, Boh98, CHL^{+}00, Inm98, Pol98]. sands [DM97]. SanFrancisco [JL00, JA00, van00]. Santa [McC78]. Satellite [FG83, GBK83]. satisfaction [BESZ02]. Saturn [JD70]. scalability [KDY97]. scalable [AAB^{+}05b, BGERS08, DAK00, MPR^{+}03, Nai95, VSC^{+}03]. scale [CC87, CHL^{+}00, CLS88, GCG^{+}04, PB90, WPSL01, Wim71]. scattered [HD88]. scheduler [Hel69]. Scheduling [CCJ78, Bre71, JL00]. Schema [PMK^{+}06]. Scheme [EMMT78, MLA91]. school [DB08]. schools [KRS^{+}98]. Science [CFS81, CF05, DB08, Glu08, IBM01, Lar08, LVW08, MMS^{+}04, Ano02a]. science-intensive [CF05]. Sciences [Kov01, HSK^{+}01, Lar08, Swo01]. Scientific [GRW86, KM89, MS88, AGMS89, CCT72, CLS88, Str72]. Scoring [HH08]. Screen [Mau83]. screening [WPSL01]. scrolling [ZS99]. SCSI [Fer92]. sculpture [LT89]. SDL [BBD^{+}92]. Sea [AOFH92]. search [BC04, Cas06, KPF^{+}04, WAV^{+}04]. Second [Cav95]. Secrets [Ano96a]. section [Cre68]. secure [Bou05, NNH^{+}05, SS01, Whi01]. Securing [HHSS01, HNN02]. Security [ADDS91, Ano91e, Att79, CPW03, GDD^{+}01, JD91, Joh95, KNN^{+}01, Len78, Mur84, Sum84, TB05, WFS80, BBN96, CJ04, CGHK98, KNNL98, NNH^{+}05, PCFY07]. Seeing [BEF^{+}07]. segmentation [EKS92]. Segmenting [FG90]. Seismic [KKS88]. selection [AAB^{+}07, DF63, TOV^{+}95, Yel03].

Self [BM83, CHPW07, Ver96]. Self-Checking [BM83]. self-managing [CHPW07]. Self-motion-sending [Ver96]. sell [SMK^{+}02]. sell-side [SMK^{+}02]. Selling [Fah02]. Semantic [HB05, Hir90, VL05]. sending [Ver96]. Senior [Don03, Mil02, Rom08, San07]. Sense [KBB^{+}07, KS03, KBB^{+}05]. Sense-and-respond [KBB^{+}07, KBB^{+}05]. Sense-making [KS03]. Sensing [Par96]. Sensor [PHS^{+}00, CCC^{+}08, Sha79]. Separation [BE01, Rob01]. sequence [IFM^{+}01, Rob01]. sequences [LC01, WMSW01]. sequencing [Cal65b, Gor65]. Sequential [Tur63, GW89]. Serial [Hai65]. series [Lew66, AJM^{+}83]. Series/1 [AJM^{+}83]. Series/1-Based [AJM^{+}83]. Server [BFGS98, Cha70, HHM04, ACO^{+}91, BDF^{+}00, DMB^{+}06, DHPB03, DBC^{+}00, DAK00, FK01, HS03, KJBH^{+}00, LMPS02, MPF97, SS01, STTA08, AJ05, Bay98, DDMS97, DCH^{+}03, Har03, KBB^{+}97, MST^{+}03, Sch03, WBC04]. servers [JBB03, TB05]. Service [Alt08, DB08, FS05, GK07, Gar83, GFG^{+}92, Gla78, JRST07, Lar08, Lew76, LMPR07, MSB^{+}07, dMR03, SHLP05, TVX^{+}04, AASW98, AM04, AGA^{+}08, ABW^{+}08, BR08a, BBWL05, BDL^{+}07, BCD^{+}07, BDE^{+}05, CNS^{+}08, CNL08, CGH^{+}05, CN08, CK05, CFL^{+}06, CBC^{+}05, CDK^{+}05, DHP02, Glu08, HDS03, HFB^{+}08, HS08, JHMO07, KKLR05, KOH^{+}07, KBC^{+}07, LD08, LHJ^{+}07, LLS^{+}06, LVW08, MBB^{+}07, Mil08, NLN^{+}07, NG05, RMP08, RCPR07, RAK08, VL05, Wal07, Zha98]. service-dominant [LVW08]. Service-oriented [FS05, AGA^{+}08, ABW^{+}08, BBWL05, BDE^{+}05, CBC^{+}05, CDK^{+}05, KKLR05, KOH^{+}07, KBC^{+}07, NLN^{+}07, NG05, SHLP05, VL05, Wal07]. service-oriented-architecture [CK05]. serviceability [HPPS75]. Services [AB92, BBE^{+}97, HS83, Ken03, LeB08, NG05, DLP^{+}05, Spe02, AHS90, AC89, CN08, DDK^{+}04, FK02, GGKS02, HNN02, HBM^{+}04, IKS^{+}95, LR02a, LRS02, MNQ^{+}02, NMB04, OCK^{+}92, Pal04, PVK08, Rap04, Rub89, SR06, Str72, VPH^{+}08, WLB^{+}07, WRS04, GR01, IKS^{+}95, KKL06, LK05, DLP^{+}05, Rom08]. serving [GNP^{+}04, KBB^{+}97]. Session [MOK78, FtCE87, Tem92]. Set [Hop87, KL74, Ano74j, Ano74l, MSZ^{+}01]. seventh [SD07]. shareable [van98]. Shared [Fri70, SNC91, CHNS95, GMTW08, LD99, PBS97]. shared-memory [GMTW08, PBS97]. Sharing [Bar75, MOK78, MS70, Bar71, Cha66, DAB^{+}00, Hel69, JMNT97, LD99, RS98, Str97]. shifting [DM97]. Should [Dea82]. sibling [KM06a]. Side [Bra81, SMK^{+}02]. Sidestream [LC80]. Signal [Ger96b, Yar65]. signatures [HN98]. significant [BAI^{+}86]. silver [May94, SBE^{+}00]. Simon [Lor94b]. Simple [Rog83, Cha66]. simplex [FT92b]. simplify [Mil99a]. simplifying [CHPW07]. SIMSCRIPT [DM64, DM99]. Simulating [SS69, MFJ05]. Simulation [BKP64, Cla81, Nah74, NORS80, SY71, Smi62, Vel64, Ano73e, BG64, Blu64, BL73, HMS92, KKLR05, Rec72, Sea66, SH68, SM01, Wah72]. simulations [Yag63]. Simulator [JD70, Blu64, BKP64, EG64, Gor62, Gou69, HS65, Vel64]. Single [Cha70, Ghi85, Fis91]. Single-Server [Cha70]. Single-System [Ghi85]. Sites [AS82, MDC^{+}05]. Six [Mye76]. sizes [Gor71]. sizing [DeM68, Men68, Pie68]. skill [ORBdSB02]. Sky [GBK83]. Skylab [Man74]. SLA [BCL^{+}04]. Small [ATC77, Hen86, LC80, BAI^{+}86, Dar70, Gor73, MBB^{+}07]. Small-System [Hen86, Dar70]. Smalltalk [Bur96]. smart [HHSS01, LLS^{+}06]. smart-call-center [LLS^{+}06]. SNA [AB92, Baa91, BFH89, BHWW83, CN79, GY82, GM79, GHH^{+}83, GBK83, Har80, HS83, JMW83, Mar83, Moo88, RM88, Ryd83, SB82b, Sul83, SSS^{+}87, SSY^{+}87, WI83]. SOA [LeB08, LTL^{+}08, SNK^{+}08, VCS^{+}08, ZZC^{+}08]. social [DEA99, Lar08, TKE01]. society [Pic96]. soft [Caf08]. Software [Ano96a, ALP83, Ber84, BDE^{+}05, CFS81, Cla81, Cle95, Gol86, Gri93, HS02, Har80, HB85, KBS94, Klo08, KW86, LNH06, Mil02, MV80, Mis83, NS84, Pea82, PJ86, Per02, SG98, Sme78, Spe02, Wei05, WCG^{+}80, Woo02, AT06, AHS^{+}03, BBS02, BCK^{+}94, Bou05, Bri90, BMK02, CCHJ94, CCO02, CK98, DMMR92, DGK96, Dye80, FHK^{+}97, GFS03, GM91, Gol99, HR81, HOST05, HLT94, HP95, HBC^{+}94, HGB05, Hum85, Kan91, KDA^{+}94, KPM01, Lin80, MM90, MOCP85, May94, Mil80, Mil99b, O’N80, PCFY07, PV97, PCH93, Qui80, Raz93, SB93, SB97, Sel08, STF^{+}01, Sim05, SHFG95, SLD89, TVPL94, TMF06, Was93, Wat94a, WKI03, WSKW^{+}03, YLG05, Ygl93, Bec95, CH85, LM96, Ran02]. solid [BCG^{+}89]. Solution [Cav95, Lei99, AP90, BESZ02, Con97, KRS^{+}98, LLS^{+}06, SSB^{+}00]. Solutions [LK05, Per02, Unl08, AGA^{+}08, BDE^{+}05, CBCS08, HCR^{+}07, KKLR05, LPR^{+}07, LSH^{+}08, PY08, PH99, Ric00, SNK^{+}08, Whi01, ZZC^{+}08, Kov01]. Solving [CCJ78]. SOM [BBN96]. SOMA [AGA^{+}08, ZZC^{+}08]. SOMA-ME [ZZC^{+}08]. Some [Bov96, Hel69]. sort [Rad66]. Sorting [Woo69, Ano70k, Lor71]. sound [GB00a]. Soup [Neg96]. source [AAB^{+}05a, AAB^{+}05b, BCvE^{+}05, Bou05, CFGS05, HBC^{+}05, LHT^{+}07, NS05, OT94, Sim05, YLG05, MSZ^{+}01]. sources [DCB^{+}01, GSN^{+}01, HSK^{+}01]. SP [CFP^{+}95]. SP1 [GL95]. SP2 [AAC^{+}95, AMM^{+}95, AMM^{+}99, Boz89, SHFG95, SSA^{+}95]. Space [Ryd83, SGS89]. Spaces [WMLF98, GH91, Mar99, SDP00]. specific [JRST07, LNH06, Sin03]. Specification [FtCE87, Bev00, NC03, Tho91]. specifications [BBD^{+}92, NG05, Odo90, Sag90]. specify [HKOS96]. Speciner [Joh95]. SPECS [DMMR92]. Speech [GB84, BSC^{+}01, FG90, THR^{+}05]. speech-based [THR^{+}05]. speed [Mon67]. SpeedTracer [WYB98]. spelling [LC03]. sphere [HD88]. Spheres [Dav78]. spline [Ahu68]. spline-like [Ahu68]. split [Fer92]. SQL [BCH^{+}06, FMR02, ÖCKM06]. SQL/XML [FMR02]. SRIRAM [VSC^{+}03]. staging [GW89]. stakeholders [BHT03]. Standard [Yeo05, FHP02, McD99, WCV69, YRPSK99, EMMT78]. standardization [FFK92]. Standardized [Cla85]. Standards [CN79, JHMO07, Sim05]. started [BC91]. State [Tet79, CCO02, MLDF08]. static [LMX07, PCFY07]. Statistical [Kos63b, Lew66]. Statistics [SH68, Sta69, Tho91]. Status [HHP79]. STCL [WSP^{+}02]. steel [BKP64]. Stenotype [NO70]. Stephen [Ano96a]. stills [MB96]. Storage [ARG75, Att79, AJ73, Bir73, BE77, BBDT84, CKB89, CM77, DD64, DCH^{+}03, GL74, Gra78, Gro85, Har03, MGST70, MPR^{+}03, MST^{+}03, Mis81, Mor73, Rog75, Sch73, SA69, Tri91, ABB^{+}03, AJ99, AFM03, Bel66, BN89, CCT72, Gel89, Gla93a, GFS03, GH03, Gro89, Gro92, HCGT90, HO07, KJP03, MPS71, McN93, MT03, PPTH72, SSH88, SVM^{+}03, SLW66, Sym68, Wit63]. Store [Ant75, BDH75, HPPS75, MHMM75, MS75]. Storing [RLPG96]. story [DM97, PDI^{+}00]. storyscapes [DAB^{+}00]. storyteller [DM97]. Strategic [HV93, HV99, Kon93, Nor93, PH01]. Strategies [Dav82, Mil81, New85, BVP93, Hel69, LLO93]. Strategy [Klo08, BW93, CFGS05, CA95, IJM93, KS03, MMNR90, SM92, TVPL94]. Stream [Mor78, Lev79]. streams [BM99, SKN06].

Streenstrup [Vet95]. Stringers [SBE^{+}00]. Structural [CGP68, Lip68, Pad68, HGW01]. Structure [AJ73, Bir73, McG77a, NO70, Sch73, Amd64, AJ99, BM68, BB64, Bla64, Cla66, GB00b, HS87, Hen69, Mea66, Oka68, Pad64, Paz89, PH99, SR96, SF01, Ste64, Sym68, WSG^{+}01, Wit66]. structure-function [WSG^{+}01]. Structured [BEP80, Rog75, SMC74, SMC99, HM93, MLH87, OT94]. Structures [Sch87, Sen77, Ano68h, BBWL05, Odo90, SAAF73a, SAAF73b, SAAF73c, Tag90]. Studies [SB82a, BMK02, MS75, TMF06, WAV^{+}04, PSLB97, SBL^{+}94]. Studio [LR02a, BDE^{+}04]. Study [AMP76, BBC82, FW75, Lam84, RHMP85, TB84, AOFH92, Bel66, Bri90, CNS^{+}08, CHL^{+}00, DAK00, HS08, KM06b, Kra03, PB90, RHMP99, Rob01, SB97, SM79, ZS99, Zie79, Zac82]. style [JRD81, Sam88]. Styles [Mil81]. Subject [Ano69k, Ano81h, Ano84c, Ano94c, Ano70m, Ano71k, Ano72k, Ano73l, Ano74n, Ano67m, Ano68j]. Subjects [Ano88e, HKOS96]. Subroutine [FT92a, FT92b, MS88, WR92, AGMS89, Rul68]. Subsystem [BHS74, Cul76, HH79, Ano76j, Gla93a, McN93, Tho88]. Subsystems [KW86, AFM03]. Succeeding [Rac96]. Success [Was93]. Suggested [Ano80h, Ano80i, Ano81i, Ano81j, Ano81k, Ano81l, Ano82i, Ano82j, Ano82k, Ano83b, Ano83c, Ano84d, Ano84e, Ano86b, Ano87d, Ano87e, Ano95b, Ano99d, Ano01c, Sti94]. suitability [JF91]. Suite [Rof01]. Summarizing [WAV^{+}04]. sun [TMF06]. supercomputer [IBM01]. superhighway [SB95]. Supermarket [Ant75, HPPS75, MHMM75, MS75]. SuperNews [DW00]. supertypes [HKOS96]. supply [GOS05, KBB^{+}07, Mar63]. Support [Boo85, CDF^{+}98, Har80, Qua70, AT06, Ano68i, BCSW91, BS05, BS94, CNL08, Cle89, Got92, HFB^{+}08, HP95, HBC^{+}94, IG04, JRG^{+}01, MSB^{+}07, MST^{+}03, SB00b, Tag90, Vos92, WAV^{+}04, Ahu88, KPF^{+}04, MV80]. Support/DL/I [MV80]. Support/DL/I-Patient [MV80]. Supporting [AR92, AFM03, HOST05, SB82a, BBE^{+}97, GMTW08]. surfaces [Alm72, PHS^{+}00]. Surf’N’Sign [HN98]. survey [CH06, GOS05, Hae03, Mea66, PCFY07, SD03]. Survival [Ano96a]. swapping [AHS^{+}03, TBB^{+}87]. switch [SSA^{+}95]. switched [CHNS95]. Switching [Kas86, Gay95]. Symbol [SL75]. symbolic [JOZG02]. Synchronous [DK74]. Synergistic [DAB^{+}00]. synthesis [SNK^{+}08]. Synthetic [Buc69, ES00]. sysplex [BDM97, JAS08, Joh97, KDY97, NMCB97]. System [Aga85, ATC77, All74, All75, AJM^{+}83, ARG75, AC89, AMP76, AJ73, Bar73, Boi74, Bon69, Buc69, Buc85, CW86, Col74, Dew90, DM75, Dru69, Enc81, Fla85, FW75, FS72, Gar81, Gel89, Ghi85, GB84, GM79, Gro85, Han80, HDH88, Haz96, HJM^{+}82, Hei85, Hel81, Hen86, HD77, JMV69, Jon77, KSWC93, Kas86, KRR80, Kra78, Lam84, LCS82, LP74, Man74, Mat86, McG77a, McG77b, McG77c, McG77d, McG77e, McP74, MS70, MV80, Par82, PMM83, RGH^{+}76, RS73, Sch80b, ST89, ST99, Siw77, SA69, Sta69, Ste64, SUW82, TAR85, VR86, YC06, Yel85, Yuv80, AM90, AMM^{+}95, AMM^{+}99, Alt08, Ano63b, Ano75j, AAB^{+}05b, AJ99, Bak65, BGP62, BM72, Bar71, BCE^{+}94, BCvE^{+}05, BBET71, BCFH65, BCE67, Boi99, BT84, BS67, Bre71, BD66, CNS^{+}08, CD68, Che69]. system [CC87, Cla89, COC63, CG77, CEH^{+}94, Dar70, DGK96, Dev67, DM88, Dor63, ESH89, EDG^{+}96, EGLT79, EW66, Fis91, Gay66, Gin65, GFS03, GSBH^{+}89, GKT63, Gri63, HSK^{+}01, HR81, HRL87, Hed63, Hei90, Hob72, HS00, IFM^{+}01, Irl92, JOZG02, Kee67, Kie79, KPF^{+}04, LR67, Lar63, LI01, LS91, Lev79, LRHH04, LS87, LMPS02, LDL^{+}08, LHT^{+}07, MW66, MQR05, MOCP85, MPR^{+}03, MO90, Mue67, NMB04, NN01, ÖLL07, RC69, SRM^{+}98, SLW66, Sea66, Sew67, SY71, SKA86, Smi63, SH68, SJL^{+}04, Sud67, Tal63, THL^{+}03, Tod76, UMN^{+}04, Vel64, WND07, WKI03, WT64, Wim71, Wit63, Zie79, d’A89, Amd64, BB64, Bla64, CGP68, FIS64, Fre66, LGM69, Lip68, Pad64, Pad68, Ste64, Tuc67, ADDS91, ABDD^{+}91, Ano91e, BDH75, BAC^{+}81, BAC^{+}99, Boo85, Bot82, Buc86, Cha89, CC89, Col74]. System [CRTW88, DO84, DPR90, Fla85, Gor80, Got86, GS04, HS87, HL76, HBM^{+}04, JD91, Jos81, KR88, KLS84, MHMM75, MS75, Mis81, Miz88, Pin89, PPDW01, SFJ^{+}94, ST89, ST99, SNC91, SKG^{+}87, Tuc86, Wat79]. System-independent [AC89]. System-managed [Gel89]. System-performance [YC06]. System/2 [CRTW88, KR88, Miz88]. System/36 [SKG^{+}87]. SYSTEM/360 [Amd64, BB64, Bla64, FIS64, Fre66, LGM69, Pad64, Ste64, Tuc67, CGP68, Lip68, Pad68]. System/370 [Buc86, Fla85, HL76, KLS84]. System/38 [Bot82]. System/400 [Cha89, CC89, Pin89, ST89, ST99]. System/6000 [ABDD^{+}91]. System/7 [Col74]. System/88 [HS87]. systematic [Min92]. Systems [Asc86, AA76, Bar75, BBMA91, BH80, BG64, Bro85, Buc85, Cha70, CKB89, DFHS90, DFS94, DK79b, Duk75, GJ87, Gro92, HBR86, LC80, Lor86, Mat86, Moo83, Mor73, PS83, Pec85, Pla89, Rei76, Rog75, Ryd74, SB82a, SGS89, Sch87, SS69, Sea80, Sen77, Sme78, Som84, Sym86, Zac82, Ake89, Ano68i, Ano85b, AHS^{+}03, Aro67, AJ99, BFSLT08, BBCM06, BFH^{+}94, BBE^{+}97, BGT96, BMP08, BBWL05, BSP^{+}02, BHR89, BARM97, BEIW97, BK63, BFH^{+}05, BAI^{+}86, CCT72, CNS^{+}08, CNL08, CCO02, CDK08, CA95, CFP^{+}95, DMB^{+}06, DM97, DEA99, DHK^{+}08, EAB^{+}04, Elo95, Fal91, GS99, Gor62, Gor73, GH03, GMTW08, HKOS96, Hav68, HZ92, Hem90, HM93, HKS08, Hum85, Inm98, Ive63, Ive99, Joh68, Joh97, Kee68, Kes03, KLS90, Lan73, LPPC03, LS00, McD96, MLH87, MPTW94, MBC08]. systems [MT03, Nai95, PHS^{+}00, PS99, PCFY07, PB90, Rei79, RC91, RMC03, Rym90, SL95, SSR87, SB00a, SAAF73a, SAAF73b, SAAF73c, SR93, Smi62, STTA08, SCK02, SZ92, Sow02, Ste93, Ste79, SSM^{+}04, Sum89, Tri91, THR^{+}05, VAH^{+}02, Wel65, Zac87, Zac99, ZABM92, Lor94a, Ahu79, Ahu88, AR76, Ano76j, Ant75, AJ73, Bir73, Cul76, EHJ77, HPPS75, Hob76, McF76, MLDF08, Sch73, WG88].

T [WMLF98]. table [Ger62, MBG00].

table-top [MBG00]. Tables

[Mar83, Mon62, FCM^{+}98, Spi66, Ano67n]. Tags [CBG^{+}82, BMMR96]. taking [COC63, LD99]. Taligent [DGK96]. Talk [Dea82, GB00a]. Tang [Ano96a]. tangible [UI00]. Tank [MPR^{+}03]. Tapping [CBG^{+}07]. targeting [LPR^{+}07]. Task [Ryd70, Wit66]. tasks [GW89]. TCP [BTB95]. TCP/IP [BTB95]. Team [Bak72]. Technical [ACMS06, AGMS89, ACO^{+}91, ABD^{+}97, Bau93, Ben94, Bev00, Bla98, Col95, Con97, DCH^{+}03, Fah02, FKS^{+}03, Got98, HKOS96, KPF^{+}04, LT89, LG99, LM93, dMR03, Mil99a, Pol98, Ric00, vdS00, TG93, Unl08, VW06, Wat94a, Ygl93, Zha98, CM88, HRL99, KBB^{+}05, PW88, WAV^{+}04]. Technique [ER70, Pea82, Ryd83, CCO02, DeM68, Gay66, Men68, Sha65b]. Techniques [AW72, BGML96, HD77, MGST70, Ais73, BFSLT08, BESZ02, Bou98, Cal65a, Dzi73, Hei63, KBB^{+}07, KHSW05, Kos63b, Nor93, Pom72, Rad66, RM67, WMSW01]. Technological [GH03]. technologies [BDG^{+}94, CH07, HBC^{+}94, KMSM^{+}07, LH97, PHG^{+}04, PG05, Smi01]. Technology [Ano96a, AJ73, Don03, JAS08, Klo08, MS85, Sel87, Wei05, Ano90f, ADD68, AJ99, BBGB92, BVP93, BC04, CCHJ94, CMCD94, CGRW07, DL89, EOO^{+}06, Fle96, Fra79, FKS^{+}02, FKS^{+}03, GSBH^{+}89, GHM^{+}05, Hel90, Hen98, HV93, HV99, Hes05, IJM93, KOH^{+}07, Kee93, LLO93, NMCB97, Rad96, Ric00, RWKN02, RHC^{+}06, Sag90, WB99, WLE^{+}97, San07]. technology-oriented [KOH^{+}07]. telecom [CNL08]. Telecommunications [AR76, DMMR92, Fra79, RAK08]. teleoncology [WLB^{+}07]. telepathology [WLB^{+}07]. Telephone [Kas86, Fah02, Vel64]. Teleprocessing [For78, BD66, EW66, Gay66, MW66, SLW66, Sea66, WCV69]. teleradiology [WLB^{+}07]. Teletex [Moo83]. television [BDCA00]. Temporal [KB00]. Teresa [McC78]. Terminal [Man74, MOK78]. Test [BM83, Han70, PJ86, Tod76, ABDD^{+}91, BBS02, BESZ02, EFN^{+}02, EPTD05, FHP02, Pin89, SWS06, WSP^{+}02, BR02, Woo02]. Testing [Duk75, LMPS02, Mau83, Ran02, BPZ90, BMK02, Gin65, HS02, KPM01, LTL^{+}08, TVPL94, DCH^{+}03]. Text [Cas06, DGK96, HJM^{+}82, MMS^{+}04, NN01, She80, GY67, GCG^{+}04, JOZG02, Sma96, UMN^{+}04, AYBDS06]. Text-Critiquing [HJM^{+}82]. text-mining [UMN^{+}04]. textile [PORG00]. textile-based [PORG00]. theater [SDP00]. Their [HHP79, BFSLT08, PBv97]. themes [SKN06]. theoretical [MKGR07]. Theory [All75, Ano75j]. thermodynamics [Ger96b]. Things [BMMR96, GB00a, Res96]. think [RMSS96, Res96]. thinking [ABK^{+}06]. ThinkPad [SCO03]. thirty [Bla98]. Three [Ano65d, Ano65e, BFSLT08, BMK02, JMV69, App68, Com68, Dav93, Dav99, Kah84, Ric00, CCHJ94]. three-dimensional [Com68, Ric00]. Three-Value [JMV69]. thresholding [HH08]. throughput [IFM^{+}01]. Tight [Che84]. Tightly [ACM74, Mac74]. Tightly-Coupled [ACM74, Mac74]. Time [Bar75, KB00, Lam84, MOK78, MS70, NO70, RGH^{+}76, SOT^{+}00, dL78, ABLM95, Aro67, ABW^{+}08, BM72, Bar71, BGERS08, BG71, Bur96, Cha66, Gin65, GMTW08, HR81, HST08, Hau68, Hel69, MOB^{+}08, Mue67, PY08, RM00, SCO03, SL95, Sel08, Sha65b, SH68, Wit63, HST08]. time-critical [ABW^{+}08]. Time-Sharing [Bar75, MOK78, MS70, Bar71, Hel69]. Timothy [Cra95]. to-dos [SDE^{+}06]. today [Bur80]. Together [KLRP86, Fli97, SCK02]. Token [DSM83]. Token-Ring [DSM83]. tolerant [HS87]. Tool [Cla81, DP86, Lai75, LC80, CCC^{+}08, KJBH^{+}00, LNH06, MSZ^{+}01, McI91, Sel06, Tag90, WYB98]. toolbox [RMS06]. toolkit [BSP^{+}02]. Tools [ALP83, Ben86, Cal75, Bou98, Bro91, Dzi73, EPTD05, GCS97, JRST07, KAN^{+}07, Pom72, SWS06, Ver96, WSP^{+}02, dRW04, van00, PPDW01]. Top [Van76, LC01, MBG00]. Top-Down [Van76, LC01]. topics [PVK08]. topologies [CDK08]. topology [CT96]. TOS [BFS67, Cen67]. TOS/360 [BFS67, Cen67]. total [KBS94]. TotalStorage [DCH^{+}03, Har03, MST^{+}03, RMH08]. tournaments [SF01]. toys [Ver96, RMSS96]. tpaML [DDK^{+}01]. TPC [HSY01]. Trace [Che69]. Trace-driven [Che69]. traceability [ARNRSG06]. tracking [LHT^{+}07]. trading [KHC02]. traditional [LH97]. traffic [ACG^{+}95, Baa91, BG71, Blu64, DDMS95, GFG^{+}92, HS03, Yar65]. traffic-signal [Yar65]. Trail [Bjo75]. train [ApS64]. training [ORBdSB02]. traits [PBv97]. Trajectory [Qua70]. Transaction [Bro85, DPR90, McG77e, Wat79, CGV^{+}02, ADDS91, Ano91e, JD91]. transactional [JR04]. Transcoding [BCC^{+}01, IMMB01]. Transcription [NO70, BBFK05, Wot93]. transduction [Fle96]. Transfer [Cav95]. Transformation [San07, AAB^{+}07, CH07, CH06, Dav93, Dav99, EBKT07, Hes05, VGW^{+}07, Wal07]. transformational [KBC^{+}07]. Transforming [LLO93, CdVF^{+}06, HV93, HV99]. Transition [Lor86]. transitional [LH97]. Translation [AJ73, AJ99]. Translator [ALP83]. Transmission [Cul76, Ano76j]. Transparent [GSN^{+}01]. Transport [PTE^{+}95, LGCB95, PBCK95]. transportation [Mar63]. treatment [EKS92]. tree [JOZG02]. trees [KM06a]. tremor [LS05]. Trends [Smi01, SKN06]. triage [HB05]. trim [Gom62]. Trust [HKLS03]. trusted [HFB^{+}08]. TSO [HZ84, Hoe89]. TSO/E [Hoe89]. Tuning [ARG75, Ber78, CLSW84, Sch80a, Yuv80]. Turing [BR02]. Turnaround [Hau68]. Turning [All99, GS99, Gol99, Mar99, WB99]. tutorial [dP88]. Twentieth [Ohm86]. Twenty [Ano02a]. Twenty-First [Ano02a]. Two [Ano65e, ABB^{+}03, BS67, HG86, Ber92a, vdS00]. typographic [GC88].

ubiquitous [WGB99, LD99]. ugly [HT06]. UIMA [GS04]. ULSI [dLMSC89]. ULSI/VLSI [dLMSC89]. UML [Sel06]. Uncontrolled [FW75]. Uncovering [SDE^{+}06]. Understanding [Miz88, Bro91, BDG^{+}94, CdVF^{+}06, Cle89, Cor89]. Unified [WRS04]. unifies [Zlo82]. unifying [ABLU00]. unique [HLLW72]. unit [Amd64, SH87]. Universal [ES05, ELM^{+}99, AM04, SL75]. universe [Ger00]. UNIX [Lev79]. Unstructured [FL04, MMS^{+}04]. unwanted [Fah02]. update [GMTW08]. upon [BCL^{+}04]. urban [Blu64]. USAA [PB90]. Usability [BDK^{+}81, GM03, Got86, Hel81, NS84, PJ86, THL^{+}03, DP95, KTBE05, Kra03, WBSK87]. Usable [Cla81, Ber88]. usage [DZ68, WYB98]. Use [BBC82, Pow83, Som84, BT84, GCS97, HBZ06, JL00, JMNT97, Rob01, RMS06, Spi66, Vre03, YLG05, Ano85b]. usefulness [SM79]. User [AHS90, Ben86, Ber88, BR92, Boi74, Boi99, Bot82, DK79b, HR81, HD77, Jon77, Mor73, Tha81, Bot89, DMRS88, KK03, KBK^{+}03, Lew76, Orw96, Pic00, Pin89, PVK08, PLGB90, RMDN03, Uhl88, UI00, Vre03, WRS04, WAV^{+}04, BHT03]. user-centered [KK03]. user-centric [PVK08]. User-definable [HR81]. User-Oriented [Jon77]. Users [Tet79, BHT03, BFH^{+}05, MFJ05, PWF05]. Uses [CCT72, GSBH^{+}89]. Using [AT06, Ano94d, BESZ02, Bon69, CdVF^{+}06, CCJ78, DHP02, EOO^{+}06, EAB^{+}04, FHP02, GB00a, Ghi85, KBC^{+}07, LC80, LR02b, Odo90, Pol98, SR96, TJ00, Van76, VMKB05, AJ05, Baa91, BBC^{+}07, GKT63, HHSS01, IBM01, JG07, KBB^{+}07, KMSM^{+}07, Kol69, MMS^{+}04, Moo90, Rec72, RLPG96, Sel08, Sha79, Tem92, VPH^{+}08, WCV69, YFS94]. utilities [ACO^{+}91]. Utility [BCL^{+}04, AM04, EAB^{+}04, GNP^{+}04, Pal04, Rap04, RW04]. utilization [KPF^{+}04, SSH88]. Utilizing [Moo88].

V [Bla64, Cal65b, Dor63, Gay66, McG77e, Qui80, Sud67]. validation [Pin89]. Value [JMV69, Alt08, BR08a, BHT03, CNS^{+}08, EN07, Gla93b, GKB^{+}03, Sim05]. values [Gor65]. Varieties [EBKT07]. Vector [Buc86, CW86, MS88, Sam88, AGMS89, HMS92, KKS88, SZ88]. vectorized [SZ88]. vectors [LM93, Mat91, MLA91]. Vehicle [Tod76]. vendor [BBS02]. vendor-developed [BBS02]. verbal [HLLW72]. Verification [ABDD^{+}91, JMV69, CCO02, GY67, HS02]. Version [Got92]. versus [Bou05]. vertical [HBZ06]. very [GCG^{+}04]. VF [ALPC88]. VGA [Tho88]. VI [Gor65, Sea66, Sew67, WT64]. via [BMK02]. Vice [Don03, Mil02, Rom08, San07, Ken03, Klo08, Kov01, LK05, Spe02, Unl08, Wei05]. Victor [Ano96a]. video [Lip96, Pic96, Tho88]. Videoconferencing [AJM^{+}83]. view [AM90, DEA99, Haa00b, Har03, Ive91, RMS06]. Viewer [Paz89]. Views [Due01, DW00]. VII [Sha65a]. Virtual [AAB^{+}05a, ARG75, Att79, AJ73, Bar75, Bir73, HH79, Jen79, Mac79, McG72, MS70, Mor73, PPTH72, Rog75, Ryd74, Sch73, TB05, VL00, AAB^{+}00, AJ99, Bel66, CCT72, CJ04, CS76, DBC^{+}00, GBCW00, HG71, PBS97, SSH88, SGS89, WPSL01, AR76, RMS06]. virtual-storage [Bel66]. virtualization [LDL^{+}08]. virtualized [SBC^{+}08]. vision [BFH^{+}05, IBM01, MFJ05]. Visual [Far87, Moo90, Shu89, Shu99, Wat94b, WSG^{+}01]. VisualAge [CLR98, Pol98]. visualization [BCG^{+}89, CEH^{+}94, Rei89, VW06, WSKW^{+}03, Wat94b]. visualizations [BEF^{+}07]. Visualizing [DLP^{+}05]. Vlissides [Bec95]. VLSI [dLMSC89]. VM [AMP76, Att79, BCSW91, BN89, Boz89, Cal75, CW81, DK79b, Fis91, GH91, HPSS79, MSS91, MLA83, OJG91, Rei79, SM79, SNC91, Tet79, TB84, TBB^{+}87, TB05]. VM/370 [AMP76, Att79, Cal75, CW81, HPSS79, SM79, Tet79, TB84]. VM/CMS [DK79b, Rei79]. VM/ESA [BCSW91, Fis91, MSS91, OJG91, SNC91]. VM/HPO [TBB^{+}87]. VM/Pass [MLA83]. VM/Pass-Through [MLA83]. VM/XA [BN89, Boz89]. Voice [CH07, Kas86, LC03]. Voice-enabled [CH07]. voice-spelling [LC03]. Voice/Data [Kas86]. Volume [Ano65d, Ano65e, Ano66i, Ano67j, Ano67l, Ano84b, Ano86a, Ano87c, Ano88d, Ano89d, Ano90e, Ano91d, Ano92d, Ano93f, Ano95a, Ano96b, Ano98b, Ano99c, Ano02b, Ano03b, Ano97c, Ano99b]. Volumes [Ano64d]. VS [AS82, Bir73, EHJ77, McG77a, McG77b, McG77c, McG77d, McG77e, Sea80, SUW82]. VS-Based [Kra78]. VS1 [BHS74, Whe74]. VS2 [Ber78, LP74, McP74]. VS2-2 [Sch73]. VSAM [Ano74j, Ano74l, KL74, Str97]. vulnerabilities [PCFY07]. waiting [Sha65b]. want [Orw96]. Warehouse [SS94]. warfare [Bou98]. was [Orw96]. WATINFO [ACO^{+}91]. way [Fah02, HKOS96]. wearable [BHW99, Sta96]. Web [HB05, NG05, CFL^{+}06, DDK^{+}04, DHPB03, FK02, GNP^{+}04, GGKS02, HBC^{+}05, HN98, HGB05, HNN02, HBM^{+}04, JR04, KKL06, KTBE05, LR02a, LRS02, MNQ^{+}02, dMR03, MDC^{+}05, DLP^{+}05, PCC^{+}98, PHG^{+}04, PG05, TLRH98, WYB98, Zha98]. Web-client [PHG^{+}04]. WebEntree [Zha98]. WebFountain [GCG^{+}04]. WebSphere [AJ05, Rof01, BCF^{+}04, Bay98, BS04, BDE^{+}04, DPR^{+}02, FK01, FNS04, HHM04, KKL^{+}04, LR02a, SRW04, WRS04, WBC04]. Westminster [Gra78]. Where [Pru01]. whom [Mac00]. whys [Bov96]. Wide [PCC^{+}98]. Williams [Ano02a]. Windows [GBK83]. WINSOM [BCG^{+}89]. Wireless [BL95, Zim99, HS00]. wires [GM03]. Within [Har80, SS94]. Witten [Cra95]. word [Zlo82]. Work [Alt08, CGG90, CN08, GH79, KM06b]. Workbench [ABK^{+}06]. Workers [Str75]. Workflow [LR97, BE01]. Workflow-based [LR97]. workflows [HYJK06]. Working [KLRP86]. workload [AEE^{+}97, WA02b]. Workloads [FW75, HSY01, HS03]. Workplace [Ber92b, BR92, MRB01]. works [Haa00b]. Workstation [AGLR90, BBC82]. Workstations [Che84, GHMWB84, KLRP86]. world [GM03, KS03, PY08, RR96, RMS06, RHC^{+}06, Joh95, PCC^{+}98]. Worldwide [Moo83, Pec85]. WORM [HO07]. Writing [CRTW88]. WSLA [DDK^{+}04]. WSLA-driven [DDK^{+}04].

X [Ano02a]. X.25 [DH83]. X.400 [SSR87]. XA [BN89, Boz89]. XC [GH91]. XML [ACMS06, BEH^{+}06, BCH^{+}06, Cas06, Cha02, FKS^{+}02, FMR02, FKS^{+}03, HBZ06, KM06a, LR02a, PMK^{+}06, RMS06, RHC^{+}06, SRCV06]. XML-centric [RHC^{+}06]. XPath [Cas06]. XQuery [AYBDS06, BCH^{+}06, Cas06, Cha02, FMR02, ÖCKM06]. XTABLES [FKS^{+}02, FKS^{+}03].

year [Ano70l, Ano70m, Ano71j, Ano71k, Ano72j, Ano72k, Ano73k, Ano73l, Ano74m, Ano74n]. years [Bla98, vdS00]. YES [MCE^{+}86]. YES/MVS [MCE^{+}86]. York [Ano02a].

z [GDD^{+}01, LMPS02, MST^{+}03, RMH08, TB05]. z/OS [GDD^{+}01, LMPS02, MST^{+}03, RMH08]. z/VM [TB05]. zSeries [BS05, LMPS02].

References

Ashany:1976:DBS

[AA76] R. Ashany and Michael Adamowicz. Data base systems. IBM Systems Journal, 15(3):253– 263, 1976. CODEN IBMSA7. ISSN 0018-8670.

Alpern:2000:JAV

[AAB^{+}00] B. Alpern, C. R. Attanasio, J. J. Barton, M. G. Burke, P. Cheng, J.-D. Choi, A. Cocchi, S. J. Fink, D. Grove, M. Hind, S. F. Hummel, D. Lieber, V. Litvinov, M. F. Mergen, T. Ngo, J. R. Russell, V. Sarkar, M. J. Serrano, J. C. Shepherd, S. E. Smith, V. C. Sreedhar, H. Srinivasan, and J. Whaley. The Jalapeño virtual machine. IBM Systems Journal, 39(1):211–238, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/391/alpern.html.

Alpern:2005:JRV

[AAB^{+}05a] B. Alpern, S. Augart, S. M. Blackburn, M. Butrico, A. Cocchi, P. Cheng, J. Dolby, S. Fink, D. Grove, M. Hind, K. S. McKinley, M. Mergen, J. E. B. Moss, T. Ngo, V. Sarkar, and M. Trapp. The Jikes Research Virtual Machine project: Building an open-source research community. IBM Systems Journal, 44 (2):399–417, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/alpern.html;   http:
//www.research.ibm.com/journal/
sj/442/alpern.pdf;   http:
//www.research.ibm.com/journal/
sj/442/alpern.txt.

Appavoo:2005:EKO

[AAB^{+}05b] J. Appavoo, M. Auslander, M. Butrico, D. da Silva, O. Krieger, M. Mergen, M. Ostrowski, B. Rosenburg, R. W. Wisniewski, and J. Xenidis. Experiences with K42, an opensource, Linux-compatible, scalable operating-system kernel. IBM Systems Journal, 44(2): 427–440, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/appavoo.html;   http:
//www.research.ibm.com/journal/
sj/442/appavoo.pdf;   http:

//www.research.ibm.com/journal/
sj/442/appavoo.txt.

Abe:2007:OSB

[AAB^{+}07] N. Abe, R. Akkiraju, S. Buckley, M. Ettl, P. Huang, D. Subramanian, and F. Tipu. On optimizing the selection of business transformation projects. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/464/abe.html.

Agarwal:1995:HPI

[AAC^{+}95] R. C. Agarwal, B. Alpern, L. Carter, F. G. Gustavson, D. J. Klepacki, R. Lawrence, and M. Zubair. Highperformance parallel implementations of the NAS kernel benchmarks on the IBM SP2. IBM Systems Journal, 34(2):263–272, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-2.html#eight.

AbadPeiro:1998:DGP

[AASW98] J. L. Abad Peiro, N. Asokan, M. Steiner, and M. Waidner. Designing a generic payment service. IBM Systems Journal, 37(1):72–88, 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/371/abadpeiro.html.

Allen:1992:SMS

[AB92] M. O. Allen and S. L. Benedict. SNA management services architecture for APPN networks. IBM Systems Journal, 31(2):336–352, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5477.

Ashton:2003:TDP

[ABB^{+}03] L. L. Ashton, E. A. Baker, A. J. Bariska, E. M. Dawson, R. L. Ferziger, S. M. Kissinger, T. A. Menendez, S. Shyam, J. P. Strickland, D. K. Thompson, G. R. Wilcock, and M. W. Wood. Two decades of policybased storage management for the IBM mainframe computer. IBM Systems Journal, 42(2): 302–321, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/422/ashton.pdf.

Attaluri:1995:CMP

[ABC^{+}95] G. K. Attaluri, D. P. Bradshaw, N. Coburn, P. A. Larson, P. Martin, A. Silberschatz, J. Slonim, and Q. Zhu. The CORDS multidatabase project. IBM Systems Journal, 34(1):39–62, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-1.html#four.

Arnold:1997:TFI

[ABD^{+}97] V. D. Arnold, R. J. Bosch, E. F. Dumstorff, P. J. Helfrich, T. C. Hung, V. M.

Johnson, R. F. Persik, and P. D. Whidden. Technical forum: IBM business frameworks: San Francisco project technical overview. IBM Systems Journal, 36(3):437–445, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/363/forum.html.

Aharon:1991:VIR

[ABDD^{+}91] A. Aharon, A. Bar-David, B. Dorfman, E. Gofman, M. Leibowitz, and V. Schwartzburd. Verification of the IBM RISC System/6000 by a dynamic biased pseudo-random test program generator. IBM Systems Journal, 30(4):527–538, 1991. CODEN IBMSA7. ISSN 0018-8670.

Abrams:2006:ATM

[ABK^{+}06] S. Abrams, B. Bloom, P. Keyser, D. Kimelman, E. Nelson, W. Neuberger, T. Roth, I. Simmonds, S. Tang, and J. Vlissides. Architectural thinking and modeling with the Architects’ Workbench. IBM Systems Journal, 45(3):481– 500, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/abrams.html.

Aldred:1995:AMC

[ABLM95] B. K. Aldred, G. W. Bonsall, H. S. Lambert, and H. D. Mitchell. An architecture for multimedia communication and real-time collaboration. IBM Systems Journal, 34(3):519–543, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-3.html#eleven.

Alexander:2000:UAP

[ABLU00] W. P. Alexander, R. F. Berry, F. E. Levine, and R. J. Urquhart. A unifying approach to performance analysis in the Java environment. IBM Systems Journal, 39(1): 118–134, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/391/alexander.html.

Abowd:1999:CEI

[Abo99] G. D. Abowd. Classroom 2000: An experiment with the instrumentation of a living educational environment. IBM Systems Journal, 38(4): 508–530, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/384/abowd.html.

Astley:2008:PRC

[ABW^{+}08] M. Astley, S. Bhola, M. J. Ward, K. Shagin, H. Paz, and G. Gershinsky. Pulsar: a resource-control architecture for time-critical serviceoriented applications. IBM Systems Journal, 47(2):265– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/astley.html.

Ahuja:1968:IGDa

[AC68] D. V. Ahuja and S. A. Coons.

Interactive graphics in data processing: Geometry for construction and display. IBM Systems Journal, 7(3/4):188–205, 1968. CODEN IBMSA7. ISSN 0018-8670.

Ashfield:1989:SFM

[AC89] J. C. Ashfield and D. B. Cybrynski. System-independent file management and distribution services. IBM Systems Journal, 28(2):241–259, 1989. CODEN IBMSA7. ISSN 0018-8670.

Ahmadi:1988:N

[ACG88] M. Ahmadi, J. H. Chou, and G. Gafka. NetView/PC. IBM Systems Journal, 27(1):32– 44, January 1988. CODEN IBMSA7. ISSN 0018-8670.

Ahmadi:1995:NTM

[ACG^{+}95] H. Ahmadi, P. F. Chimento, R. A. Guerin, L. Gun, B. Lin, R. O. Onvural, and T. E. Tedijanto. NBBS traffic management overview. IBM Systems Journal, 34(4):604–628, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#five.

Appleby:2004:PBA

[ACGL04] K. C. Appleby, S. B. Calo, J. R. Giles, and K.-W. Lee. Policy-based automated provisioning. IBM Systems Journal, 43(1):121–135, ???? 2004. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/431/appleby.pdf.

Arnold:1974:DTM

[ACM74] J. S. Arnold, D. P. Casey, and R. H. McKinstry. Design of tightly-coupled multiprocessing programming. IBM Systems Journal, 13(1):60–87, 1974. CODEN IBMSA7. ISSN 0018-8670.

Adler:2006:TCC

[ACMS06] S. Adler, R. Cochrane, J. F. Morar, and A. Spector. Technical context and cultural consequences of XML. IBM Systems Journal, 45(2):207–223, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/452/adler.html.

Appel:1991:TNW

[ACO^{+}91] A. Appel, G. A. Cuomo, E. A. Overly, J. A. Walicki, and R. E. Yozzo. Technical note: The WATINFO face server and associated utilities. IBM Systems Journal, 30(3):393–407, 1991. CODEN IBMSA7. ISSN 0018-8670.

Ambrosetti:1984:AA

[ACP84] R. Ambrosetti, T. A. Ciriani, and R. Pennacchi. An application analyzer. IBM Systems Journal, 23(4):336–350, 1984. CODEN IBMSA7. ISSN 0018-8670.

Appel:1968:IGDa

[ADD68] A. Appel, T. P. Dankowski, and R. L. Dougherty. Interactive graphics in data processing: Aspects of display technology. IBM Systems Journal, 7 (3/4):176–187, 1968. CODEN IBMSA7. ISSN 0018-8670.

Abraham:1991:TSS

[ADDS91] D. G. Abraham, G. M. Dolan, G. P. Double, and J. V. Stevens. Transaction Security System. IBM Systems Journal, 30(2):206–229, 1991. CODEN IBMSA7. ISSN 0018-8670. See erratum [Ano91e].

Aman:1997:AAM

[AEE^{+}97] J. Aman, C. K. Eilert, D. Emmes, P. Yocom, and D. Dillenberger. Adaptive algorithms for managing a distributed data processing workload. IBM Systems Journal, 36(2):242–283, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/362/aman.html.

Azagury:2003:AFS

[AFM03] A. C. Azagury, M. E. Factor, and W. F. Micka. Advanced functions for storage subsystems: Supporting continuous availability. IBM Systems Journal, 42(2):268–279, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/azagury.pdf.

Abraham:1990:PDI

[AG90] R. Abraham and B. F. Goodrich. Porting DPPX from the IBM 8100 to the IBM ES/9370: Feasibility and overview. IBM Systems Journal, 29(1):90–105, 1990. CO-DEN IBMSA7. ISSN 0018-8670.

Agassi:1985:PCD

[Aga85] S. Agassi. Performance considerations for a distributed data processing system designed for high availability. IBM Systems Journal, 24(3-4):200–212, 1985. CODEN IBMSA7. ISSN 0018-8670.

Arsanjani:2008:SMD

[AGA^{+}08] Ali Arsanjani, Shuvanker Ghosh, Abdul Allam, Tina Abdollah, Sella Ganapathy, and Kerrie Holley. SOMA: a method for developing service-oriented solutions. IBM Systems Journal, 47(3):377–396, ???? 2008. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/473/arsanjani.pdf.

Anderson:1990:IWP

[AGLR90] G. B. Anderson, B. P. Gross, S. M. Lewis, and J. A. Reimer.

ImagePlus Workstation Program. IBM Systems Journal, 29(3):398–407, 1990. CODEN IBMSA7. ISSN 0018-8670.

Agarwal:1989:TNE

[AGMS89] R. C. Agarwal, F. G. Gustavson, J. McComb, and S. Schmidt. Technical note: Engineering and scientific subroutine library release 3 for IBM ES/3090 vector multiprocessors. IBM Systems Journal, 28(2):345–350, 1989. CODEN IBMSA7. ISSN 0018-8670.

Artim:1990:UIS

[AHS90] J. M. Artim, J. M. Hary, and F. J. Spickhoff. User interface services in AD/Cycle. IBM Systems Journal, 29(2):236–249, 1990. CODEN IBMSA7. ISSN 0018-8670.

Appavoo:2003:EAB

[AHS^{+}03] J. Appavoo, K. Hui, C. A. N.

Soules, R. W. Wisniewski, D. M. Da Silva, O. Krieger, M. A. Auslander, D. J. Edelsohn, B. Gamsa, G. R. Ganger, P. McKenney, M. Ostrowski, B. Rosenburg, M. Stumm, and J. Xenidis. Enabling autonomic behavior in systems software with hot swapping. IBM Systems Journal, 42(1):60–76, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/appavoo.pdf.

Ahuja:1968:IGDb

[Ahu68] D. V. Ahuja. Interactive graphics in data processing: An algorithm for generating spline-like curves. IBM Systems Journal, 7(3/4):206–217, 1968. CODEN IBMSA7. ISSN 0018-8670.

Ahuja:1979:RFC

[Ahu79] Vijay Ahuja. Routing and flow control in Systems Network Architecture. IBM Systems Journal, 18(2):298–314, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Ahuja:1988:CCS

[Ahu88] V. Ahuja. Common Communications Support in Systems Application Architecture. IBM Systems Journal, 27(3):264– 280, 1988. CODEN IBMSA7. ISSN 0018-8670.

Aiken:1998:RED

[Aik98] P. H. Aiken. Reverse engineering of data. IBM Systems Journal, 37(2):246–269, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/372/aiken.html.

Aiso:1973:FT

[Ais73] M. Aiso. Forecasting techniques. IBM Systems Journal, 12(2):187–209, 1973. CODEN IBMSA7. ISSN 0018-8670.

Auslander:1973:FSI

[AJ73] M. A. Auslander and J. F. Jaffe. Function structure of IBM Virtual Storage Operating Systems — 1 influences of dynamic address translation on operating system technology. IBM Systems Journal, 12(4):368–381, 1973. CODEN IBMSA7. ISSN 0018-8670.

Auslander:1999:FSI

[AJ99] M. A. Auslander and J. F.

Jaffe. Functional structure of IBM virtual storage operating systems, part I: Influences of dynamic address translation on operating system technology. IBM Systems Journal, 38(2–3):340–353, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/382/auslander.pdf. Reprinted

from [AJ73].

Anand:2005:MPC

[AJ05] V. K. Anand and W. C. Jamison. A middleware performance characterization of Linux using IBM WebSphere Application Server. IBM Systems Journal, 44(2):353–367, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/anand.html;   http:
//www.research.ibm.com/journal/
sj/442/anand.pdf;   http:
//www.research.ibm.com/journal/
sj/442/anand.txt.

Anastassiou:1983:SVS

[AJM^{+}83] Dimitris Anastassiou, Hugh C.

Jones, J. L. Mitchell, William B. Pennebaker, Keith S. Pennington, and Marvin K. Brown. Series/1-based videoconferencing system. IBM Systems Journal, 22(1-2):97–110, 1983. CO-DEN IBMSA7. ISSN 0018-8670.

Aken:1989:LSE

[Ake89] B. R. Aken, Jr. Large systems and Enterprise Systems Architecture. IBM Systems Journal, 28(1):4–14, 1989. CODEN IBMSA7. ISSN 0018-8670.

Arnette:1995:CAD

[AKR^{+}95] W. S. Arnette, A. D. Kshemkalyani, W. B. Riley, J. P. Sanders, P. J. Schwaller, J. C. Terrien, and J. Q. Walker. CPI-C: An API for distributed applications. IBM Systems Journal, 34(3):501–518, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-3.html#ten.

Anderson:2004:GNF

[ALE^{+}04] O. T. Anderson, L. Luan, C. Everhart, M. Pereira, R. Sarkar, and J. Xu. Global namespace for files. IBM Systems Journal, 43(4):702– 722, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/434/anderson.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/434/anderson.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/434/anderson.txt.

Alfonseca:1991:AAA

[Alf91] M. Alfonseca. Advanced applications of APL: logic programming, neural networks, and hypertext. IBM Systems Journal, 30(4):543–553, 1991. CODEN IBMSA7. ISSN 0018-8670.

Alfred:1994:MLO

[Alf94] C. Alfred. Maximizing leverage from an object database. IBM Systems Journal, 33 (2):280–299, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-2.html#five.  G321-

5543.

Allen:1971:AFL

[All71] C. D. Allen. The application of formal logic to programs and programming. IBM Systems Journal, 10(1):2–38, 1971. CO-DEN IBMSA7. ISSN 0018-8670.

Allen:1974:EPS

[All74] A. O. Allen. Elements of probability for system design. IBM Systems Journal, 13(4):325– 348, 1974. CODEN IBMSA7. ISSN 0018-8670.

Allen:1975:EQT

[All75] Arnold O. Allen. Elements of queuing theory for system design. IBM Systems Journal, 14(2):161–187, 1975. CODEN IBMSA7. ISSN 0018-8670. See erratum [Ano75j].

Allen:1999:TPI

[All99] F. E. Allen. Turning points in interaction with computers. IBM Systems Journal, 38(2– 3):135–138, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/allen.pdf.

Almond:1972:NCM

[Alm72] D. B. Almond. Numerical control for machining complex surfaces. IBM Systems Journal, 11(2):150–168, 1972. CODEN IBMSA7. ISSN 0018-8670.

Archibald:1983:ADP

[ALP83] J. L. Archibald, B. M. Leavenworth, and L. R. Power. Abstract design and program translator: New tools for software design. IBM Systems Journal, 23(3):170–187, 1983. CODEN IBMSA7. ISSN 0018-8670.

Angeleri:1988:PII

[ALPC88] P. Angeleri, D. F. Lozupone, F. Piccolo, and J. Clinckemaillie. PAM-CRASH on the IBM 3090/VF: an integrated environment of crash analysis. IBM Systems Journal, 27(4): 541–560, November 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Alter:2008:SSF

[Alt08] S. Alter. Service system fundamentals: Work system, value chain, and life cycle. IBM Systems Journal, 47(1):71–??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/471/alter.html.

Addink:1990:AIS

[AM90] M. Addink and J. J. Mullen.

AS/400 ImagePlus system view. IBM Systems Journal, 29(3):451–466, 1990. CODEN IBMSA7. ISSN 0018-8670.

Albaugh:2004:UMS

[AM04] V. Albaugh and H. Madduri.

The utility metering service of the Universal Management Infrastructure. IBM Systems Journal, 43(1):179–189, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/431/albaugh.pdf.

Amdahl:1964:SSP

[Amd64] G. M. Amdahl. The structure of SYSTEM/360, part III: Processing unit design considerations. IBM Systems Journal, 3(2):144–164, 1964. CODEN IBMSA7. ISSN 0018-8670.

Agerwala:1995:SSA

[AMM^{+}95] T. Agerwala, J. L. Martin, J. H.

Mirza, D. C. Sadler, D. M. Dias, and M. Snir. SP2 system architecture. IBM Systems Journal, 34(2):152–184, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-2.html#three.

Agerwala:1999:SSA

[AMM^{+}99] T. Agerwala, J. L. Martin, J. H.

Mirza, D. C. Sadler, D. M. Dias, and M. Snir. SP2 system architecture. IBM Systems Journal, 38(2–3):414–448, ????

1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/382/agerwala.pdf. Reprinted

from [AMM^{+}95].

Attanasio:1976:POS

[AMP76] C. Richard Attanasio, Peter W. Markstein, and Ray J. Phillips. Penetrating an operating system: a study of VM/370 integrity. IBM Systems Journal, 15(1):102–116, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1963:A

[Ano63a] Anonymous. Authors. IBM Systems Journal, 2:328–330, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1963:DOS

[Ano63b] Anonymous. Design of an operating system. IBM Systems Journal, 2:297, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1964:Aa

[Ano64a] Anonymous. Authors. IBM Systems Journal, 3(1):112–114, 1964. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1964:Ab

[Ano64b] Anonymous. Authors. IBM Systems Journal, 3(2):196, 1964. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1964:Ac

[Ano64c] Anonymous. Authors. IBM Systems Journal, 3(3):263, 1964. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1964:CV

[Ano64d] Anonymous. Contents of volumes 1 & 2. IBM Systems Journal, 3(1):115–116, 1964. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:Aa

[Ano65a] Anonymous. Authors. IBM Systems Journal, 4(1):81–82, 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:Ab

[Ano65b] Anonymous. Authors. IBM Systems Journal, 4(2):171–172, 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:Ac

[Ano65c] Anonymous. Authors. IBM Systems Journal, 4(3):256–257, 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:CV

[Ano65d] Anonymous. Contents of volume three. IBM Systems Journal, 4(1):83–84, 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:EVT

[Ano65e] Anonymous. Errata in volume three, numbers two and three. IBM Systems Journal, 4(1):84, 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:FAOa

[Ano65f] Anonymous. Fabrication and assembly operations. IBM Systems Journal, 4(2):86, June 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1965:FAOb

[Ano65g] Anonymous. Fabrication and assembly operations. IBM Systems Journal, 4(3):224–244, 1965. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Ab

[Ano66a] Anonymous. Abstracts. IBM Systems Journal, 5(1):53–59, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Ad

[Ano66b] Anonymous. Abstracts. IBM Systems Journal, 5(2):127–131, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Af

[Ano66c] Anonymous. Abstracts. IBM Systems Journal, 5(3):192–200, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Ah

[Ano66d] Anonymous. Abstracts. IBM Systems Journal, 5(4):272–278, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Aa

[Ano66e] Anonymous. Authors. IBM Systems Journal, 5(1):52, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Ac

[Ano66f] Anonymous. Authors. IBM Systems Journal, 5(2):126, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Ae

[Ano66g] Anonymous. Authors. IBM Systems Journal, 5(3):190–191, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:Ag

[Ano66h] Anonymous. Authors. IBM Systems Journal, 5(4):271, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1966:CVF

[Ano66i] Anonymous. Contents of volume four. IBM Systems Journal, 5(1):60, 1966. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Aa

[Ano67a] Anonymous. Abstracts. IBM Systems Journal, 6(1):69–75, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Ab

[Ano67b] Anonymous. Abstracts. IBM Systems Journal, 6(2):135–140, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Ac

[Ano67c] Anonymous. Abstracts. IBM Systems Journal, 6(3):209–219, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Ad

[Ano67d] Anonymous. Abstracts. IBM Systems Journal, 6(4):305–312, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:AI

[Ano67e] Anonymous. Author Index: 1962–1967. IBM Systems Journal, 6(4):320–323, 1967. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Ae

[Ano67f] Anonymous. Authors. IBM Systems Journal, 6(1):68, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Af

[Ano67g] Anonymous. Authors. IBM Systems Journal, 6(2):133–134, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Ag

[Ano67h] Anonymous. Authors. IBM Systems Journal, 6(3):208, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:Ah

[Ano67i] Anonymous. Authors. IBM Systems Journal, 6(4):303–304, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:CVF

[Ano67j] Anonymous. Contents of volume five. IBM Systems Journal, 6(1):76, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:CI

[Ano67k] Anonymous. Cumulative Indices: 1962–1967. IBM Systems Journal, 6(4):313, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:EVF

[Ano67l] Anonymous. Errata in volume five. IBM Systems Journal, 6 (1):76, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:SI

[Ano67m] Anonymous. Subject Index:

1962–1967. IBM Systems Journal, 6(4):324–327, 1967. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1967:TC

[Ano67n] Anonymous. Tables of Contents: 1962–1967. IBM Systems Journal, 6(4):314–319, 1967. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:Aa

[Ano68a] Anonymous. Abstracts. IBM Systems Journal, 7(1):68–71, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:Ab

[Ano68b] Anonymous. Abstracts. IBM Systems Journal, 7(2):135–143, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:AI

[Ano68c] Anonymous. Author Index:

1962–1968. IBM Systems Journal, 7(3/4):386–390, 1968. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:Ac

[Ano68d] Anonymous. Authors. IBM Systems Journal, 7(1):67, 1968. CODEN IBMSA7. ISSN 0018-

Anonymous:1968:Ad

[Ano68e] Anonymous. Authors. IBM Systems Journal, 7(2):134, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:Ae

[Ano68f] Anonymous. Authors. IBM Systems Journal, 7(3/4):382– 385, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:IGDc

[Ano68g] Anonymous. Interactive graphics in data processing: Application areas. IBM Systems Journal, 7(3/4):309, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:IGDa

[Ano68h] Anonymous. Interactive graphics in data processing: Display concepts and structures. IBM Systems Journal, 7(3/4): 175, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:IGDb

[Ano68i] Anonymous. Interactive graphics in data processing: Language and systems support. IBM Systems Journal, 7(3/4): 247, 1968. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1968:SI

[Ano68j] Anonymous. Subject Index: 1962–1968. IBM Systems Journal, 7(3/4):391–395, 1968. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Aa

[Ano69a] Anonymous. Abstracts. IBM Systems Journal, 8(1):88–91, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Ab

[Ano69b] Anonymous. Abstracts. IBM Systems Journal, 8(2):172–175, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Ac

[Ano69c] Anonymous. Abstracts. IBM Systems Journal, 8(3):247–250, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Ad

[Ano69d] Anonymous. Abstracts. IBM Systems Journal, 8(4):336–337, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:AIa

[Ano69e] Anonymous. Author index:

1962–1969. IBM Systems Journal, 8(4):339–344, 1969. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Ae

[Ano69f] Anonymous. Authors. IBM Systems Journal, 8(1):87, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Af

[Ano69g] Anonymous. Authors. IBM Systems Journal, 8(2):170–171, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Ag

[Ano69h] Anonymous. Authors. IBM Systems Journal, 8(3):245–246, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:Ah

[Ano69i] Anonymous. Authors. IBM Systems Journal, 8(4):334–335, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:C

[Ano69j] Anonymous. Contents: 1969. IBM Systems Journal, 8(4): 338, 1969. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1969:AIb

[Ano69k] Anonymous. Subject index: 1962–1969. IBM Systems Journal, 8(4):345–349, 1969. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ac

[Ano70a] Anonymous. Abstracts. IBM Systems Journal, 9(1):73–75, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ae

[Ano70b] Anonymous. Abstracts. IBM Systems Journal, 9(2):160–162, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ag

[Ano70c] Anonymous. Abstracts. IBM Systems Journal, 9(3):235–238, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ai

[Ano70d] Anonymous. Abstracts. IBM Systems Journal, 9(4):319–322, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:AA

[Ano70e] Anonymous. Acknowledgment of authorship. IBM Systems Journal, 9(3):239, 1970. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ab

[Ano70f] Anonymous. Authors. IBM Systems Journal, 9(1):72, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ad

[Ano70g] Anonymous. Authors. IBM Systems Journal, 9(2):159, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Af

[Ano70h] Anonymous. Authors. IBM Systems Journal, 9(3):234, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:Ah

[Ano70i] Anonymous. Authors. IBM Systems Journal, 9(4):318, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:C

[Ano70j] Anonymous. Contents: 1970.

IBM Systems Journal, 9(4): 323, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:ESM

[Ano70k] Anonymous. Errata: “Internal sorting with minimal comparing”. IBM Systems Journal, 9(2):163, 1970. CODEN IBMSA7. ISSN 0018-8670. See [Woo69].

Anonymous:1970:FAI

[Ano70l] Anonymous. Five-year author index. IBM Systems Journal, 9(4):324–327, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1970:FSI

[Ano70m] Anonymous. Five-year subject index. IBM Systems Journal, 9(4):328–331, 1970. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Aa

[Ano71a] Anonymous. Abstracts. IBM Systems Journal, 10(1):81–83, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Ab

[Ano71b] Anonymous. Abstracts. IBM Systems Journal, 10(2):164– 165, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Ac

[Ano71c] Anonymous. Abstracts. IBM Systems Journal, 10(3):257– 258, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Ad

[Ano71d] Anonymous. Abstracts. IBM Systems Journal, 10(4):327– 328, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Ae

[Ano71e] Anonymous. Authors. IBM Systems Journal, 10(1):80, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Af

[Ano71f] Anonymous. Authors. IBM Systems Journal, 10(2):162– 163, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Ag

[Ano71g] Anonymous. Authors. IBM Systems Journal, 10(3):255– 256, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:Ah

[Ano71h] Anonymous. Authors. IBM Systems Journal, 10(4):325– 326, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:C

[Ano71i] Anonymous. Contents: 1971.

IBM Systems Journal, 10(4): 329, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:FAI

[Ano71j] Anonymous. Five-year author index. IBM Systems Journal, 10(4):330–333, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1971:FSI

[Ano71k] Anonymous. Five-year subject index. IBM Systems Journal, 10(4):334–337, 1971. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Ab

[Ano72a] Anonymous. Abstracts. IBM Systems Journal, 11(1):95–96, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Ad

[Ano72b] Anonymous. Abstracts. IBM Systems Journal, 11(2):183– 184, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Af

[Ano72c] Anonymous. Abstracts. IBM Systems Journal, 11(3):274– 275, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Ah

[Ano72d] Anonymous. Abstracts. IBM Systems Journal, 11(4):352– 353, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Aa

[Ano72e] Anonymous. Authors. IBM Systems Journal, 11(1):93–94, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Ac

[Ano72f] Anonymous. Authors. IBM Systems Journal, 11(2):181– 182, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Ae

[Ano72g] Anonymous. Authors. IBM Systems Journal, 11(3):272– 273, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:Ag

[Ano72h] Anonymous. Authors. IBM Systems Journal, 11(4):350– 351, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:C

[Ano72i] Anonymous. Contents: 1971.

IBM Systems Journal, 11(4): 354, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:FAI

[Ano72j] Anonymous. Five-year author index. IBM Systems Journal, 11(4):355–358, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1972:FSI

[Ano72k] Anonymous. Five-year subject index. IBM Systems Journal, 11(4):359–362, 1972. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Aa

[Ano73a] Anonymous. Abstracts. IBM Systems Journal, 12(1):108– 109, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Ab

[Ano73b] Anonymous. Abstracts. IBM Systems Journal, 12(2):212– 214, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Ad

[Ano73c] Anonymous. Abstracts. IBM Systems Journal, 12(3):330, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Ae

[Ano73d] Anonymous. Abstracts. IBM Systems Journal, 12(4):414– 415, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:AGM

[Ano73e] Anonymous. Addendum: “A general management business simulation in APL”. IBM Systems Journal, 12(1):109, 1973. CODEN IBMSA7. ISSN 0018-8670. See [Wah72].

Anonymous:1973:Af

[Ano73f] Anonymous. Authors. IBM Systems Journal, 12(1):106– 107, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Ag

[Ano73g] Anonymous. Authors. IBM Systems Journal, 12(2):210– 211, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Ac

[Ano73h] Anonymous. Authors. IBM Systems Journal, 12(3):328– 329, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:Ah

[Ano73i] Anonymous. Authors. IBM Systems Journal, 12(4):412– 413, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:C

[Ano73j] Anonymous. Contents: 1973. IBM Systems Journal, 12(4): 416, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:FAI

[Ano73k] Anonymous. Five-year author index. IBM Systems Journal, 12(4):417–419, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1973:FSI

[Ano73l] Anonymous. Five-year subject index. IBM Systems Journal, 12(4):420–423, 1973. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Aa

[Ano74a] Anonymous. Abstracts. IBM Systems Journal, 13(1):90–92, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Ab

[Ano74b] Anonymous. Abstracts. IBM Systems Journal, 13(2):181– 182, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Ac

[Ano74c] Anonymous. Abstracts. IBM Systems Journal, 13(3):271– 272, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Ad

[Ano74d] Anonymous. Abstracts. IBM Systems Journal, 13(4):350– 351, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Ae

[Ano74e] Anonymous. Authors. IBM Systems Journal, 13(1):88–89, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Af

[Ano74f] Anonymous. Authors. IBM Systems Journal, 13(2):179– 180, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Ag

[Ano74g] Anonymous. Authors. IBM Systems Journal, 13(3):270, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:Ah

[Ano74h] Anonymous. Authors. IBM Systems Journal, 13(4):349, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:C

[Ano74i] Anonymous. Contents: 1974. IBM Systems Journal, 13(4): 353, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:CDS

[Ano74j]  Anonymous.  Correction:

“VSAM data set design parameters”. IBM Systems Journal, 13(4):352, 1974. CODEN IBMSA7. ISSN 0018-8670. See [KL74].

Anonymous:1974:E

[Ano74k] Anonymous. Errata: “Indexing design considerations”. IBM Systems Journal, 13(2): 183, 1974. CODEN IBMSA7. ISSN 0018-8670. See [Wag73].

Anonymous:1974:EDS

[Ano74l] Anonymous. Erratum: “VSAM data set design parameters”. IBM Systems Journal, 13(4):

352, 1974. CODEN IBMSA7. ISSN 0018-8670. See [KL74].

Anonymous:1974:FAI

[Ano74m] Anonymous. Five-year author index. IBM Systems Journal, 13(4):354–356, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1974:FSI

[Ano74n] Anonymous. Five-year subject index. IBM Systems Journal, 13(4):357–360, 1974. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Aa

[Ano75a] Anonymous. Abstracts. IBM Systems Journal, 14(1):98–100, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Ab

[Ano75b] Anonymous. Abstracts. IBM Systems Journal, 14(2):205– 207, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Ac

[Ano75c] Anonymous. Abstracts. IBM Systems Journal, 14(3):306– 307, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Ad

[Ano75d] Anonymous. Abstracts. IBM Systems Journal, 14(3):308– 309, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Ae

[Ano75e] Anonymous. Abstracts. IBM Systems Journal, 14(4):409– 410, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Af

[Ano75f] Anonymous. Authors. IBM Systems Journal, 14(1):96–97, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Ag

[Ano75g] Anonymous. Authors. IBM Systems Journal, 14(2):203– 204, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:Ah

[Ano75h] Anonymous. Authors. IBM Systems Journal, 14(4):407– 408, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:C

[Ano75i] Anonymous. Contents: 1975. IBM Systems Journal, 14(4): 411, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1975:EQT

[Ano75j] Anonymous. Erratum: “Elements of queuing theory for system design”. IBM Systems Journal, 14(2):208, 1975. CO-DEN IBMSA7. ISSN 0018-8670. See [All75].

Anonymous:1975:IR

[Ano75k] Anonymous. Information on reprints. IBM Systems Journal, 14(3):310, 1975. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Aa

[Ano76a] Anonymous. Abstracts. IBM Systems Journal, 15(1):120, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Ab

[Ano76b] Anonymous. Abstracts. IBM Systems Journal, 15(2):173, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Ac

[Ano76c] Anonymous. Abstracts. IBM Systems Journal, 15(3):280, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Ad

[Ano76d] Anonymous. Abstracts. IBM Systems Journal, 15(4):380– 383, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Ae

[Ano76e] Anonymous. Authors. IBM Systems Journal, 15(1):117– 119, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Af

[Ano76f] Anonymous. Authors. IBM Systems Journal, 15(2):171– 172, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Ag

[Ano76g] Anonymous. Authors. IBM Systems Journal, 15(3):279, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:Ah

[Ano76h] Anonymous. Authors. IBM Systems Journal, 15(4):379, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:C

[Ano76i] Anonymous. Contents: 197. IBM Systems Journal, 15(4): 384, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1976:ENA

[Ano76j] Anonymous. Errata: “Systems Network Architecture: An overview” and “The transmission subsystem in Systems Network Architecture”. IBM Systems Journal, 15(2):174, 1976. CODEN IBMSA7. ISSN 0018-8670. See [Cul76, McF76].

Anonymous:1976:R

[Ano76k] Anonymous. Reprints. IBM Systems Journal, 15(2):175– 178, 1976. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Aa

[Ano77a] Anonymous. Abstracts. IBM Systems Journal, 16(1):76–77, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ab

[Ano77b] Anonymous. Abstracts. IBM Systems Journal, 16(2):198– 201, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ac

[Ano77c] Anonymous. Abstracts. IBM Systems Journal, 16(3):315, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ad

[Ano77d] Anonymous. Abstracts. IBM Systems Journal, 16(4):426– 427, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ae

[Ano77e] Anonymous. Authors. IBM Systems Journal, 16(1):74–75, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Af

[Ano77f] Anonymous. Authors. IBM Systems Journal, 16(2):197, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ag

[Ano77g] Anonymous. Authors. IBM Systems Journal, 16(3):313– 314, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ah

[Ano77h] Anonymous. Authors. IBM Systems Journal, 16(4):424– 425, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Ba

[Ano77i] Anonymous. Books. IBM Systems Journal, 16(2):202–203, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:Bb

[Ano77j] Anonymous. Books. IBM Systems Journal, 16(3):316–317, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:C

[Ano77k] Anonymous. Contents: 1977.

IBM Systems Journal, 16(4): 428, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1977:R

[Ano77l] Anonymous. Reprints. IBM Systems Journal, 16(1):78–80, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Aa

[Ano78a] Anonymous. Abstracts. IBM Systems Journal, 17(1):95–96, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ab

[Ano78b] Anonymous. Abstracts. IBM Systems Journal, 17(2):201– 202, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ac

[Ano78c] Anonymous. Abstracts. IBM Systems Journal, 17(3):316, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ad

[Ano78d] Anonymous. Abstracts. IBM Systems Journal, 17(4):465– 466, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ae

[Ano78e] Anonymous. Authors. IBM Systems Journal, 17(1):92–94, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Af

[Ano78f] Anonymous. Authors. IBM Systems Journal, 17(2):199– 200, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ag

[Ano78g] Anonymous. Authors. IBM Systems Journal, 17(3):314– 315, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ah

[Ano78h] Anonymous. Authors. IBM Systems Journal, 17(4):463– 464, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Ba

[Ano78i] Anonymous. Books. IBM Systems Journal, 17(3):317–320, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:Bb

[Ano78j] Anonymous. Books. IBM Systems Journal, 17(4):467–469, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:C

[Ano78k] Anonymous. Contents: 1978.

IBM Systems Journal, 17(4): 470, 1978. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1978:E

[Ano78l] Anonymous. Erratum: “Reprints”.

IBM Systems Journal, 17(2): 202, 1978. CODEN IBMSA7. ISSN 0018-8670. See [Ano78m].

Anonymous:1978:R

[Ano78m] Anonymous. Reprints. IBM Systems Journal, 17(1):97– 102, 1978. CODEN IBMSA7. ISSN 0018-8670. See erratum [Ano78l].

Anonymous:1979:Aa

[Ano79a] Anonymous. Abstracts. IBM Systems Journal, 18(2):351– 352, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:Ab

[Ano79b] Anonymous. Abstracts. IBM Systems Journal, 18(3):479, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:Ac

[Ano79c] Anonymous. Abstracts. IBM Systems Journal, 18(4):610– 611, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:Ad

[Ano79d] Anonymous. Authors. IBM Systems Journal, 18(1):181– 184, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:Ae

[Ano79e] Anonymous. Authors. IBM Systems Journal, 18(2):348– 350, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:Af

[Ano79f] Anonymous. Authors. IBM Systems Journal, 18(3):476– 478, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:Ag

[Ano79g] Anonymous. Authors. IBM Systems Journal, 18(4):608– 609, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:B

[Ano79h] Anonymous. Books. IBM Systems Journal, 18(4):612–614, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1979:C

[Ano79i] Anonymous. Contents: 1979.

IBM Systems Journal, 18(4): 615–616, 1979. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:Aa

[Ano80a] Anonymous. Abstracts. IBM Systems Journal, 19(1):163– 168, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:Ab

[Ano80b] Anonymous. Authors. IBM Systems Journal, 19(1):160– 162, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:Ac

[Ano80c] Anonymous. Authors. IBM Systems Journal, 19(2):284– 286, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:Ad

[Ano80d] Anonymous. Authors. IBM Systems Journal, 19(3):403– 407, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:Ae

[Ano80e] Anonymous. Authors. IBM Systems Journal, 19(4):554– 557, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:B

[Ano80f] Anonymous. Books. IBM Systems Journal, 19(2):287–288, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:C

[Ano80g] Anonymous. Contents: 1980. IBM Systems Journal, 19(4): 560–561, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:SRa

[Ano80h] Anonymous. Suggested reading. IBM Systems Journal, 19 (3):408–410, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1980:SRb

[Ano80i] Anonymous. Suggested reading. IBM Systems Journal, 19 (4):558–559, 1980. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:AI

[Ano81a] Anonymous. Author index: 1975–1981. IBM Systems Journal, 20(1):108–113, 1981. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:Aa

[Ano81b] Anonymous. Authors. IBM Systems Journal, 20(1):104– 105, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:Ab

[Ano81c] Anonymous. Authors. IBM Systems Journal, 20(2):264– 265, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:Ac

[Ano81d] Anonymous. Authors. IBM Systems Journal, 20(3):361– 362, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:Ad

[Ano81e] Anonymous. Authors. IBM Systems Journal, 20(4):441– 442, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:C

[Ano81f] Anonymous. Contents: 1981.

IBM Systems Journal, 20(4): 445–446, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:R

[Ano81g] Anonymous. Referees. IBM Systems Journal, 20(3):365– 367, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:SI

[Ano81h] Anonymous. Subject index:

1975–1981. IBM Systems Journal, 20(1):114–118, 1981. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:SRa

[Ano81i] Anonymous. Suggested reading. IBM Systems Journal, 20 (1):106–107, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:SRb

[Ano81j] Anonymous. Suggested reading. IBM Systems Journal, 20 (2):266–267, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:SRc

[Ano81k] Anonymous. Suggested reading. IBM Systems Journal, 20 (3):363–364, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1981:SRd

[Ano81l] Anonymous. Suggested reading. IBM Systems Journal, 20 (4):443–444, 1981. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:Aa

[Ano82a] Anonymous. Authors. IBM Systems Journal, 21(1):124– 125, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:Ab

[Ano82b] Anonymous. Authors. IBM Systems Journal, 21(2):245– 246, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:Ac

[Ano82c] Anonymous. Authors. IBM Systems Journal, 21(3):370– 374, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:Ad

[Ano82d] Anonymous. Authors. IBM Systems Journal, 21(4):511– 513, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:Ba

[Ano82e] Anonymous. Books. IBM Systems Journal, 21(3):378–379, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:Bb

[Ano82f] Anonymous. Books. IBM Systems Journal, 21(4):516–520, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:C

[Ano82g] Anonymous. Contents: 1982.

IBM Systems Journal, 21(4): 521, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:R

[Ano82h] Anonymous. Reprints. IBM Systems Journal, 21(4):522– 528, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:SRa

[Ano82i] Anonymous. Suggested reading. IBM Systems Journal, 21 (1):126–128, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:SRb

[Ano82j] Anonymous. Suggested reading. IBM Systems Journal, 21 (3):375–377, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1982:SRc

[Ano82k] Anonymous. Suggested reading. IBM Systems Journal, 21 (4):514–515, 1982. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1983:C

[Ano83a] Anonymous. Contents: 1983.

IBM Systems Journal, 22(4): 466, 1983. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1983:SRa

[Ano83b] Anonymous. Suggested reading. IBM Systems Journal, 22 (1/2):165–166, 1983. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1983:SRb

[Ano83c] Anonymous. Suggested reading. IBM Systems Journal, 22 (4):464–465, 1983. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1984:AI

[Ano84a] Anonymous. Author index: 1975–1984. IBM Systems Journal, 23(4):388–396, 1984. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1984:CV

[Ano84b] Anonymous. Contents of volume 23, 1984. IBM Systems Journal, 23(4):410, 1984. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1984:SI

[Ano84c] Anonymous. Subject index: 1975–1984. IBM Systems Journal, 23(4):397–409, 1984. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1984:SRa

[Ano84d] Anonymous. Suggested reading. IBM Systems Journal, 23(2):219, 1984. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1984:SRb

[Ano84e] Anonymous. Suggested reading. IBM Systems Journal, 23 (3):305–306, 1984. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1985:C

[Ano85a] Anonymous. Contents: 1985.

IBM Systems Journal, 24(3/4): 335, 1985. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1985:EIC

[Ano85b] Anonymous. Erratum: “Use of images in commercial and office systems”. IBM Systems Journal, 24(1):71, 1985. CODEN IBMSA7. ISSN 0018-8670. See [Som84].

Anonymous:1985:R

[Ano85c] Anonymous. Referees. IBM Systems Journal, 24(2):176– 177, 1985. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1986:CV

[Ano86a] Anonymous. Contents of volume 25, 1988. IBM Systems Journal, 25(3/4):417, 1986. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1986:SR

[Ano86b] Anonymous. Suggested reading. IBM Systems Journal, 25(1):129, 1986. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1987:Ba

[Ano87a] Anonymous. Books. IBM Systems Journal, 26(3):319–321, 1987. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1987:Bb

[Ano87b] Anonymous. Books. IBM Systems Journal, 26(4):453–454, 1987. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1987:CV

[Ano87c] Anonymous. Contents of volume 26, 1987. IBM Systems Journal, 26(4):455, 1987. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1987:SRa

[Ano87d] Anonymous. Suggested reading. IBM Systems Journal, 26(2):231, 1987. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1987:SRb

[Ano87e] Anonymous. Suggested reading. IBM Systems Journal, 26(3):322, 1987. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1988:Ba

[Ano88a] Anonymous. Books. IBM Systems Journal, 27(1):81–85, 1988. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1988:Bb

[Ano88b] Anonymous. Books. IBM Systems Journal, 27(2):246, 1988. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1988:Bc

[Ano88c] Anonymous. Books. IBM Systems Journal, 27(4):570–572, 1988. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1988:CV

[Ano88d] Anonymous. Contents of volume 27, 1988. IBM Systems Journal, 27(4):573, 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1988:CIA

[Ano88e] Anonymous. Cumulative index: Authors — subjects — abstracts. IBM Systems Journal, 27S(1):1–49, 1988. CO-DEN IBMSA7. ISSN 0018-8670. No volume or issue number, but bound with volume 27.

Anonymous:1989:Ba

[Ano89a] Anonymous. Books. IBM Systems Journal, 28(1):192, 1989. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1989:Bb

[Ano89b] Anonymous. Books. IBM Systems Journal, 28(2):351–356, 1989. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1989:Bc

[Ano89c] Anonymous. Books. IBM Systems Journal, 28(4):689–691, 1989. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1989:CV

[Ano89d] Anonymous. Contents of volume 28, 1989. IBM Systems Journal, 28(4):692, 1989. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1989:R

[Ano89e] Anonymous. Referees. IBM Systems Journal, 28(3):495– 496, 1989. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1990:Ba

[Ano90a] Anonymous. Books. IBM Systems Journal, 29(1):165–166, 1990. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1990:Bb

[Ano90b] Anonymous. Books. IBM Systems Journal, 29(2):297–298, 1990. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1990:Bc

[Ano90c] Anonymous. Books. IBM Systems Journal, 29(3):489–490, 1990. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1990:Bd

[Ano90d] Anonymous. Books. IBM Systems Journal, 29(4):575–576, 1990. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1990:CV

[Ano90e] Anonymous. Contents of volume 29, 1990. IBM Systems Journal, 29(4):577, 1990. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1990:EMT

[Ano90f] Anonymous. Erratum: “Repository Manager technology”. IBM Systems Journal, 29(4): 583, 1990. CODEN IBMSA7. ISSN 0018-8670. See [Sag90].

Anonymous:1990:R

[Ano90g] Anonymous. Reprints. IBM Systems Journal, 29(4):578– 582, 1990. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1991:Ba

[Ano91a] Anonymous. Books. IBM Systems Journal, 30(2):244–245, 1991. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1991:Bb

[Ano91b] Anonymous. Books. IBM Systems Journal, 30(3):408–412, 1991. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1991:Bc

[Ano91c] Anonymous. Books. IBM Systems Journal, 30(4):594–596, 1991. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1991:CV

[Ano91d] Anonymous. Contents of volume 30, 1991. IBM Systems Journal, 30(4):597, 1991. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1991:ESS

[Ano91e] Anonymous. Erratum: “Transaction Security System”. IBM Systems Journal, 30(4):598, 1991. CODEN IBMSA7. ISSN 0018-8670. See [ADDS91].

Anonymous:1992:Ba

[Ano92a] Anonymous. Books. IBM Systems Journal, 31(1):147–149, 1992. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1992:Bb

[Ano92b] Anonymous. Books. IBM Systems Journal, 31(2):409–410, 1992. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1992:Bc

[Ano92c] Anonymous. Books. IBM Systems Journal, 31(4):811–815, 1992. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1992:CV

[Ano92d] Anonymous. Contents of volume 31, 1992. IBM Systems Journal, 31(4):816–817, 1992. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:A

[Ano93a] Anonymous. Abstracts: 1989– 1992. IBM Systems Journal, 32(2):345–372, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:Ba

[Ano93b] Anonymous. Books. IBM Systems Journal, 32(1):222–227, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:Bb

[Ano93c] Anonymous. Books. IBM Systems Journal, 32(2):326–330, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:Bc

[Ano93d] Anonymous. Books. IBM Systems Journal, 32(3):541–542, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:Bd

[Ano93e] Anonymous. Books. IBM Systems Journal, 32(4):684–686, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:CV

[Ano93f] Anonymous. Contents of volume 32, 1993. IBM Systems Journal, 32(4):687, 1993. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:I

[Ano93g] Anonymous. Index: 1989– 1992. IBM Systems Journal, 32(2):331–344, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1993:R

[Ano93h] Anonymous. Referees. IBM Systems Journal, 32(3):543– 544, 1993. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1994:A

[Ano94a] Anonymous. Abstracts: 1962– 1994. IBM Systems Journal, 33(4):590–689, 1994. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1994:AI

[Ano94b] Anonymous. Author index:

1962–1994. IBM Systems Journal, 33(4):540–562, 1994. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1994:SI

[Ano94c] Anonymous. Subject index:

1962–1994. IBM Systems Journal, 33(4):564–587, 1994. CO-DEN IBMSA7. ISSN 0018-8670.

Anonymous:1994:UI

[Ano94d] Anonymous. Using the index.

IBM Systems Journal, 33(4): 537, 1994. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1995:CV

[Ano95a] Anonymous. Contents of volume 34, 1995. IBM Systems Journal, 34(4):758–759, 1995. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1995:SR

[Ano95b] Anonymous. Suggested reading. IBM Systems Journal, 34 (4):754–757, 1995. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1996:BIR

[Ano96a] Anonymous. Books: “Doing IT Right: Technology, Business, and Risk of Computing”, by Harold Lorin (ISBN 1-884777-09-0), reviewed by Robert M. Orfali. “The Essential Distributed Objects Survival Guide”, by Robert Orfali, Dan Harkey, and Jeri Edwards (ISBN 0-471-12993-3), reviewed by Harold Lorin. “Secrets of Software Quality: 40 Innovations from IBM”, by Craig Kaplan, Ralph Clark, and Victor Tang, (ISBN 0-07-911795-3). Reviewed by Stephen H. Kan. IBM Systems Journal, 35(1):113–116, 1996. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1996:CV

[Ano96b] Anonymous. Contents of volume 35, 1996. IBM Systems Journal, 35(3&4):651– 652, 1996. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1997:Ba

[Ano97a] Anonymous. Books. IBM Systems Journal, 36(3):464–466, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Anonymous:1997:Bb

[Ano97b] Anonymous. Books. IBM Systems Journal, 36(4):594–595, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/364/books.html.

Anonymous:1997:CV

[Ano97c] Anonymous. Contents of Volume 36, 1997. IBM Systems Journal, 36(4):596, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/364/v36cont.html.

Anonymous:1997:R

[Ano97d] Anonymous. Referees. IBM Systems Journal, 36(3):467– 471, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Anonymous:1998:B

[Ano98a] Anonymous. Books. IBM Systems Journal, 37(4):604– 605, ???? 1998. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/374/books.html.

Anonymous:1998:CV

[Ano98b] Anonymous. Contents of volume 37, 1998. IBM Systems Journal, 37(4):606–??, ???? 1998. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/.

Anonymous:1998:E

[Ano98c] Anonymous. Erratum. IBM Systems Journal, 37(1):152– 153, 1998. CODEN IBM-JAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL

                                        http://www.research.ibm.
                                        com/journal/sj/371/erratum.
                                        html.

Anonymous:1999:BP

[Ano99a] Anonymous. Books. IBM Systems Journal, 38(1):122– 124, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/381/books.html.

Anonymous:1999:CVP

[Ano99b] Anonymous. Contents of Volume 37, 1998. IBM Systems Journal, 38(4):606–628, ???? 1999. CODEN IBMSA7. ISSN 0018-8670.

Anonymous:1999:CV

[Ano99c] Anonymous. Contents of volume 38, 1999. IBM Systems Journal, 38(4):697–??, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/384/v38cont.html.

Anonymous:1999:SR

[Ano99d] Anonymous. Suggested reading. IBM Systems Journal, 38(1):125, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/381/books.html.

Anonymous:2000:BP

[Ano00] Anonymous. Books. IBM Systems Journal, 39(2):403– ??, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/books.html.

Anonymous:2001:B

[Ano01a] Anonymous. Books. IBM Systems Journal, 40(3):803– 805, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/books.html;   http:
//www.research.ibm.com/journal/
sj/403/books.pdf.

Anonymous:2001:R

[Ano01b] Anonymous. Referees. IBM Systems Journal, 40(3):806– 808, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/referees.html;   http:
//www.research.ibm.com/journal/
sj/403/referees.pdf.

Anonymous:2001:SR

[Ano01c] Anonymous. Suggested reading. IBM Systems Journal, 40 (4):1008–1011, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/404/suggest.html;   http:

                                        //www.research.ibm.com/journal/
                                        sj/404/suggest.pdf.

Anonymous:2002:BAB

[Ano02a] Anonymous. Books: Arguing A.I.: The Battle for Twenty-First Century Science, Sam Williams, Random House, New York, 2002. 128 pp. (ISBN 0-8129-91180-X). IBM Systems Journal, 41(3):540–541, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/413/books.html.

Anonymous:2002:CV

[Ano02b] Anonymous. Contents of volume 41, 2002. IBM Systems Journal, 41(4):767–768, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/414/chessell.pdf.

Anonymous:2003:B

[Ano03a] Anonymous. Books. IBM Systems Journal, 42(4):702– 705, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/424/books.pdf.

Anonymous:2003:CV

[Ano03b] Anonymous. Contents of volume 42, 2003. IBM Systems Journal, 42(4):706–708, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/424/books.pdf.

Anonymous:2004:E

[Ano04] Anonymous. Errata. IBM Systems Journal, 43(2):441, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/432/errata.html;   http:
//www.research.ibm.com/journal/
sj/432/errata.pdf;   http:
//www.research.ibm.com/journal/
sj/432/errata.txt.

Anonymous:2008:Eb

[Ano08a] Anonymous. Errata. IBM Systems Journal, 47(4):709, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/errata.pdf.  See

[SNK^{+}08].

Anonymous:2008:Ea

[Ano08b] Anonymous. Erratum. IBM Systems Journal, 47(2):351– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/erratum.html. See

[Mil08].

Apte:2002:PEF

[ANPT02] C. V. Apte, R. Natarajan, E. P. D. Pednault, and F. Tipu. A probabilistic estimation framework for predictive modeling analytics. IBM Systems Journal, 41(3):438–448, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/413/apte.html.

Antonelli:1975:ROS

[Ant75] Douglas C. Antonelli. The role of the operator in the Supermarket and Retail Store Systems. IBM Systems Journal, 14(1):35–45, 1975. CODEN IBMSA7. ISSN 0018-8670.

Aukrust:1992:MSN

[AOFH92] T. Aukrust, J. M. Oberhuber, E. J. Farrell, and P. M. Haugan. A modeling study of the North Atlantic with emphasis on the Greenland-Iceland-Norwegian Sea. IBM Systems Journal, 31(4):798–810, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5499.

Arjomandi:1995:ACI

[AOK^{+}95] E. Arjomandi, W. O’Farrell, I. Kalas, G. Koblents, F. C. Eigler, and G. R. Gao. ABC++: concurrency by inheritance in C++. IBM Systems Journal, 34(1):120–137, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-1.html#eight.

Avers:1990:IMA

[AP90] C. D. Avers and R. E. Probst. ImagePlus as a model for application solution development. IBM Systems Journal, 29(3):356–370, 1990. CODEN IBMSA7. ISSN 0018-8670.

Appel:1968:IGDb

[App68] A. Appel. Interactive graphics in data processing: Modeling in three dimensions. IBM Systems Journal, 7(3/4):310–321, 1968. CODEN IBMSA7. ISSN 0018-8670.

ApSimon:1964:AGP

[ApS64] H. G. ApSimon. Algorithm for a gear-train problem. IBM Systems Journal, 3(1):95–103, 1964. CODEN IBMSA7. ISSN 0018-8670.

Albrecht:1976:VTA

[AR76] Harold R. Albrecht and Keith D.

Ryder. The Virtual Telecommunications Access Method: a Systems Network Architecture perspective. IBM Systems Journal, 15(1):53–80, 1976. CODEN IBMSA7. ISSN 0018-8670.

Arfman:1992:PAS

[AR92] J. M. Arfman and P. Roden. Project Athena: Supporting distributed computing at MIT. IBM Systems Journal, 31(3):550–563, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5486.

Anderson:1975:TVS

[ARG75] Harold A. Anderson, Jr., Martin Reiser, and George L. Galati. Tuning a virtual storage system. IBM Systems Journal, 14(3):246–263, 1975. CODEN IBMSA7. ISSN 0018-8670.

Aizenbud-Reshef:2006:MT

[ARNRSG06] N. Aizenbud-Reshef, B. T.

Nolan, J. Rubin, and Y. Shaham-Gafni. Model traceability. IBM Systems Journal, 45(3):

515–526, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/aizenbud.html.

Aron:1967:RSP

[Aro67] J. D. Aron. Real-time systems in perspective. IBM Systems Journal, 6(1):49–67, 1967. CODEN IBMSA7. ISSN 0018-8670.

Arsanjani:2005:EBA

[Ars05] A. Arsanjani. Empowering the business analyst for on demand computing. IBM Systems Journal, 44(1):67–80, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/441/arsanjani.html;
http://www.research.ibm.
com/journal/sj/441/arsanjani.
pdf;   http://www.research.
ibm.com/journal/sj/441/arsanjani.
txt.

Acker:1982:MDP

[AS82] Robert D. Acker and Philip H. Seaman. Modeling distributed processing across multiple CICS/VS sites. IBM Systems Journal, 21(4):471–489, 1982. CODEN IBMSA7. ISSN 0018-8670.

Ark:1999:LHI

[AS99] W. S. Ark and T. Selker. A look at human interaction with pervasive computers. IBM Systems Journal, 38(4):504–507, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/384/ark.html.

Aschenbrenner:1986:OSI

[Asc86] J. R. Aschenbrenner. Open systems interconnection. IBM Systems Journal, 25(3-4):369–379, 1986. CODEN IBMSA7. ISSN 0018-8670.

Alfonseca:1991:AIG

[ASW91] M. Alfonseca, D. Selby, and R. Wilks. The APL IL Interpreter Generator. IBM Systems Journal, 30(4):490–497, 1991. CODEN IBMSA7. ISSN 0018-8670.

Albrecht:1979:IFD

[AT79] H. R. Albrecht and L. C.

Thomason. I/O facilities of the Distributed Processing Programming Executive (DPPX). IBM Systems Journal, 18(4):526–546, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Akerman:2006:UOS

[AT06] A. Akerman and J. Tyree. Using ontology to support development of software architectures. IBM Systems Journal, 45(4):813–825, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/454/akerman.html.

Alfonseca:1977:AIS

[ATC77] Manuel Alfonseca, Maria L.

Tavera, and Ramon Casajuana.

An APL interpreter and system for a small computer. IBM Systems Journal, 16(1):18–40, 1977. CODEN IBMSA7. ISSN 0018-8670.

Anbil:1992:GAC

[ATJ92] R. Anbil, R. Tanga, and E. L. Johnson. A global approach to crew-pairing optimization. IBM Systems Journal, 31(1): 71–78, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5462.

Attanasio:1979:VCS

[Att79] C. R. Attanasio. Virtual control storage — security measures in VM/370. IBM Systems Journal, 18(1):93–110, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Abrams:2007:OER

[AvKM^{+}07] C. Abrams, J. von Känel, S. Müller, B. Pfitzmann, and S. Ruschka-Taylor. Optimized enterprise risk management. IBM Systems Journal, 46(2): 219–??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/462/abrams.html.

Anthony:1972:TDA

[AW72] A. L. Anthony and H. K. Watson. Techniques for developing analytic models. IBM Systems Journal, 11(4):316–328, 1972. CODEN IBMSA7. ISSN 0018-8670.

Amer-Yahia:2006:XFT

[AYBDS06] S. Amer-Yahia, C. Botev, J. Dörre, and J. Shanmugasundaram. XQuery Full-Text extensions explained. IBM Systems Journal, 45(2):335– 351, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/amer.html.

Baade:1991:SRG

[Baa91] S. C. Baade. SNA route generation using traffic patterns. IBM Systems Journal, 30(3):250–258, 1991. CODEN IBMSA7. ISSN 0018-8670.

Blasgen:1981:SRA

[BAC^{+}81] M. W. Blasgen, M. M. Astrahan, D. D. Chamberlin, J. N. Gray, W. F. King, B. G. Lindsay, R. A. Lorie, J. W. Mehl, T. G. Price, G. R. Putzolu, M. Schkolnick, P. G. Selinger, D. R. Slutz, H. R. Strong, I. L. Traiger, B. W. Wade, and R. A. Yost. System R: an architectural overview. IBM Systems Journal, 20(1):41–62, 1981. CODEN IBMSA7. ISSN 0018-8670. URL https://doi.

org/10.1147/sj.201.0041.

Blasgen:1999:SRA

[BAC^{+}99] M. W. Blasgen, M. M. Astrahan, D. D. Chamberlin, J. N. Gray, W. F. King, B. G. Lindsay, R. A. Lorie, J. W. Mehl, T. G. Price, G. R. Putzolu, M. Schkolnick, P. G. Selinger, D. R. Slutz, H. R. Strong, I. L.

Traiger, B. W. Wade, and R. A. Yost. System R: An architectural overview. IBM Systems Journal, 38(2–3):375–397, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/382/blasgen.pdf. Reprinted

from [BAC^{+}81].

Burns:1986:PIE

[BAI^{+}86] N. A. Burns, T. J. Ashford, C. T. Iwaskiw, R. P. Starbird, and R. L. Flagg. The Portable Inference Engine: Fitting significant expertise into small systems. IBM Systems Journal, 25(2):236–243, 1986. CODEN IBMSA7. ISSN 0018-8670.

Baker:1965:FAO

[Bak65] C. T. Baker. Fabrication and assembly operations, part 1: The outlines of a control system. IBM Systems Journal, 4(2):87–93, 1965. CODEN IBMSA7. ISSN 0018-8670.

Baker:1972:CPT

[Bak72] F. T. Baker. Chief programmer team management of production programming. IBM Systems Journal, 11(1):56–73, 1972. CODEN IBMSA7. ISSN 0018-8670.

Baker:1980:LDA

[Bak80] Charles T. Baker. Logical distribution of applications and data. IBM Systems Journal, 19(2):171–191, 1980. CODEN IBMSA7. ISSN 0018-8670.

Bard:1971:PCM

[Bar71] Y. Bard. Performance criteria and measurement for a timesharing system. IBM Systems Journal, 10(3):193–216, 1971. CODEN IBMSA7. ISSN 0018-8670.

Bard:1973:EES

[Bar73] Y. Bard. Experimental evaluation of system performance. IBM Systems Journal, 12(3): 302–314, 1973. CODEN IBMSA7. ISSN 0018-8670.

Bard:1975:PAV

[Bar75] Yonathan Bard. Performance analysis of virtual memory time-sharing systems. IBM Systems Journal, 14(4):366– 384, 1975. CODEN IBMSA7. ISSN 0018-8670.

Bowen:1997:APS

[BARM97] N. S. Bowen, J. Antognini, R. D. Regan, and N. C. Matsakis. Availability in parallel systems: Automatic process restart. IBM Systems Journal, 36(2):284–300, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/362/antognini.html.

Batory:2006:MMM

[Bat06] D. Batory. Multilevel models in model-driven engineering, product lines, and metaprogramming. IBM Systems Journal, 45(3):527–539, ???? 2006.

CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/batory.html.

Bauer:1993:TFA

[Bau93] D. Bauer. Technical forum: a reusable parts center. IBM Systems Journal, 32(4):620–624, 1993. CODEN IBMSA7. ISSN 0018-8670.

Bayeh:1998:WAS

[Bay98] E. Bayeh. The WebSphere Application Server architecture and programming model. IBM Systems Journal, 37(3): 336–348, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/373/bayeh.html.

Blaauw:1964:SSPa

[BB64] G. A. Blaauw and F. P. Brooks, Jr. The structure of SYS-TEM/360, part I: Outline of the logical structure. IBM Systems Journal, 3(2):119–135, 1964. CODEN IBMSA7. ISSN 0018-8670.

Bullen:1982:CSO

[BBC82] Christine V. Bullen, John L. Bennett, and Eric D. Carlson. A case study of office workstation use. IBM Systems Journal, 21(3):351–369, 1982. CODEN IBMSA7. ISSN 0018-8670.

Bird:1995:AAA

[BBC^{+}95] B. R. Bird, C. Brotman, R. Case, G. Dudley, R. E.

Moore, and M. Peters. Advances in APPN architecture. IBM Systems Journal, 34(3):430–451, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-3.html#seven.

Bantz:2003:APC

[BBC^{+}03] D. F. Bantz, C. Bisdikian, D. Challener, J. P. Karidis, S. Mastrianni, A. Mohindra, D. G. Shea, and M. Vanover. Autonomic personal computing. IBM Systems Journal, 42 (1):165–176, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/421/bantz.pdf.

Blount:2007:RHC

[BBC^{+}07] M. Blount, V. M. Batra, A. N.

Capella, M. R. Ebling, W. F. Jerome, S. M. Martin, M. Nidd, M. R. Niemi, and S. P. Wright. Remote health-care monitoring using Personal Care Connect. IBM Systems Journal, 46 (1):95–113, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/461/blount.html.

Balmelli:2006:MDS

[BBCM06] L. Balmelli, D. Brown, M. Cantor, and M. Mott. Modeldriven systems development. IBM Systems Journal, 45(3): 569–585, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/453/balmelli.html.

Binding:1992:CCL

[BBD^{+}92] C. Binding, W. Bouma, M. Dauphin, G. Karjoth, and Y. Yang. A common compiler for LOTOS and SDL specifications. IBM Systems Journal, 31(4):668–690, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5492.

Bozman:1984:AFA

[BBDT84] Gerald Bozman, William Buco, Timothy P. Daly, and William H. Tetzlaff. Analysis of freestorage algorithms — revisited. IBM Systems Journal, 23(1): 44–64, 1984. CODEN IBMSA7. ISSN 0018-8670.

Bauer:1997:SSM

[BBE^{+}97] M. A. Bauer, R. B. Bunt, A. El Rayess, P. J. Finnigan, T. Kunz, H. L. Lutfiyya, A. D. Marshall, P. Martin, G. M. Oster, W. Powley, J. Rolia, D. Taylor, and M. Woodside. Services supporting management of distributed applications and systems. IBM Systems Journal, 36(4):508–526, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

                                        research.ibm.com/journal/
                                        sj/364/bauer.html.

Belady:1971:CGS

[BBET71] L. A. Belady, M. W. Blasgen, C. J. Evangelisti, and R. D.

Tennison. A computer graphics system for block diagram problems. IBM Systems Journal, 10(2):143–161, 1971. CODEN IBMSA7. ISSN 0018-8670.

Bain:2005:ATA

[BBFK05] K. Bain, S. Basson, A. Faisman, and D. Kanevsky. Accessibility, transcription, and access everywhere. IBM Systems Journal, 44(3):589–603, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/443/bain.html.

Bender:2000:ADH

[BBG^{+}00] W. Bender, W. Butera, D. Gruhl, R. Hwang, F. J. Paiz, and S. Pogreb. Applications for data hiding. IBM Systems Journal, 39(3/4):547–566, ???? 2000. CODEN IBMSA7. ISSN 0018-8670.

Benichou:1992:PAI

[BBGB92] M. Bénichou, H. Beringer, J.- M. Gauthier, and C. Beierle. Prolog at IBM: An advanced and evolving application development technology. IBM Systems Journal, 31(4):755–773, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5496.

Bennett:1991:SMC

[BBMA91] R. B. Bennett, W. J. Bitner, M. A. Musa, and M. K. Ainsworth. Systems management for Coordinated Resource Recovery. IBM Systems Journal, 30(1):90–106, 1991. CO-DEN IBMSA7. ISSN 0018-8670.

Benantar:1996:AOS

[BBN96] M. Benantar, B. Blakley, and A. J. Nadalin. Approach to object security in Distributed SOM. IBM Systems Journal, 35(2):192–203, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-2.html#six.

Bassin:2002:MEV

[BBS02] K. Bassin, S. Biyani, and P. Santhanam. Metrics to evaluate vendor-developed software based on test case execution results. IBM Systems Journal, 41(1):13–30, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/411/bassin.html;   http:
//www.research.ibm.com/journal/
sj/411/bassin.pdf.

Bieberstein:2005:ISO

[BBWL05] N. Bieberstein, S. Bose, L. Walker, and A. Lynch. Impact of service-oriented architecture on enterprise systems, organizational structures, and individuals. IBM Systems Journal, 44(4):691–708, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/bieberstein.html.

Brown:1991:AGS

[BC91] J. A. Brown and H. P. Crowder. APL2: Getting started. IBM Systems Journal, 30(4):433– 445, 1991. CODEN IBMSA7. ISSN 0018-8670.

Broder:2004:TNG

[BC04] A. Z. Broder and A. C. Ciccolo. Towards the next generation of enterprise search technology. IBM Systems Journal, 43(3):451–454, ???? 2004. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/433/broder.html;   http:
//www.research.ibm.com/journal/
sj/433/broder.pdf;   http:
//www.research.ibm.com/journal/
sj/433/broder.txt.

Britton:2001:TEB

[BCC^{+}01] K. H. Britton, R. Case, A. Citron, R. Floyd, Y. Li, C. Seekamp, B. Topol, and K. Tracey. Transcoding: Extending e-business to new environments. IBM Systems Journal, 40(1):153–178, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/401/britton.html.

Bainbridge:2001:CEJ

[BCCN01] A. Bainbridge, J. Colgrave, A. Colyer, and G. Normington. CICS and Enterprise JavaBeans. IBM Systems Journal, 40(1):46–67, ???? 2001. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/401/bainbridge.html.

Brittenham:2007:ISM

[BCD^{+}07] P. Brittenham, R. R. Cutlip, C. Draper, B. A. Miller, S. Choudhary, and M. Perazolo. IT service management architecture and autonomic computing. IBM Systems Journal, 46(3):565–??, July/September 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/463/brittenham.html.

Blakeney:1967:AMS

[BCE67] G. R. Blakeney, L. F. Cudney, and C. R. Eickhorn. An application-oriented multiprocessing system, part II: Design characteristics of the 9020 system. IBM Systems Journal, 6(2):80–94, 1967. CODEN IBMSA7. ISSN 0018-8670.

Bauer:1994:DSA

[BCE^{+}94] M. A. Bauer, N. Coburn, D. L. Erickson, P. J. Finnigan, J. W. Hong, P.-A. Larson, J. Pachl, J. Slonim, D. J. Taylor, and T. J. Teorey. A distributed system architecture for a distributed application environment. IBM Systems Journal, 33(3):399–425, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj33-3.html#three. G321-

5548.

Bender:1996:ECH

[BCE^{+}96] W. Bender, P. Chesnais, S. Elo, A. Shaw, and M. Shaw. Enriching communities: Harbingers of news in the future. IBM Systems Journal, 35(3&4):369– 380, 1996. CODEN IBMSA7. ISSN 0018-8670.

Bakalova:2004:WDC

[BCF^{+}04] R. Bakalova, A. Chow, C. Fricano, P. Jain, N. Kodali, D. Poirier, S. Sankaran, and D. Shupp. WebSphere Dynamic Cache: Improving J2EE application performance. IBM Systems Journal, 43(2):351–370, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/432/bakalova.html;   http:
//www.research.ibm.com/journal/
sj/432/bakalova.pdf;   http:
//www.research.ibm.com/journal/
sj/432/bakalova.txt.

Bitonti:1965:CLA

[BCFH65] F. Bitonti, D. W. Cooper, D. N.

Frayne, and H. H. Hansen. A computer-aided linkage analysis system. IBM Systems Journal, 4(3):200–223, 1965. CO-DEN IBMSA7. ISSN 0018-8670.

Burridge:1989:WSM

[BCG^{+}89] J. M. Burridge, B. M. Collins, B. N. Galton, A. R. Halbert, T. R. Heywood, W. H. Latham, R. W. Phippen, P. Quarendon, P. Reilly, M. W. Ricketts, J. Simmons, S. J. P. Todd, A. G. N. Walter, and J. R. Woodwark. The WIN-SOM solid modeller and its application to data visualization. IBM Systems Journal, 28(4):548–568, 1989. CODEN IBMSA7. ISSN 0018-8670.

Beyer:2006:DGH

[BCH^{+}06] K. Beyer, R. Cochrane, M. Hvizdos, V. Josifovski, J. Kleewein, G. Lapis, G. Lohman, R. Lyle, M. Nicola, F. Özcan, H. Pirahesh, N. Seemann, A. Singh, T. Truong, R. C. Van der Linden, B. Vickery, C. Zhang, and G. Zhang. DB2 goes hybrid: Integrating native XML and XQuery with relational data and SQL. IBM Systems Journal, 45(2):271–298, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/452/beyer.html.

Bleher:1980:GIA

[BCHM80] J. H. Bleher, P. G. Caspers, H. H. Henn, and K. Maerker. A graphic interactive application monitor. IBM Systems Journal, 19(3):382–402, 1980. CODEN IBMSA7. ISSN 0018-8670.

Billings:1994:JMS

[BCK^{+}94] C. Billings, J. Clifton, B. Kolkhorst, E. Lee, and W. B. Wingert. Journey to a mature software process. IBM Systems Journal, 33(1):46–61, 1994. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj33-1.html#four.  G321-

5532.

Bhattacharya:2007:ACO

[BCK^{+}07] K. Bhattacharya, N. S. Caswell, S. Kumaran, A. Nigam, and F. Y. Wu. Artifact-centered operational modeling: Lessons from customer engagements. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-8670.

Buco:2004:UCS

[BCL^{+}04] M. J. Buco, R. N. Chang, L. Z. Luan, C. Ward, J. L. Wolf, and P. S. Yu. Utility computing SLA management based upon business objectives. IBM Systems Journal, 43(1):159–178, ???? 2004. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/431/buco.pdf.

Barnes:1991:VES

[BCSW91] C. C. Barnes, A. Coleman, J. M. Showalter, and M. L. Walker. VM/ESA support for coordinated recovery of files. IBM Systems Journal, 30(1):107–125, 1991. CODEN IBMSA7. ISSN 0018-8670.

Becking:2005:MOS

[BCvE^{+}05] J. Becking, S. Course, G. van Enk, H. T. Hangyi, J. J. M. Lahaye, D. Ockeloen, R. Peters, H. Rosbergen, and R. van Wendel de Joode. MMBase: An open-source content management system. IBM Systems Journal, 44(2):381–397, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/becking.html;   http:
//www.research.ibm.com/journal/

                                        sj/442/becking.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/becking.txt.

Bricault:1966:TSD

[BD66] J. P. Bricault and I. Delgalvis. On teleprocessing system design, part III: An analysis of a request-queued buffer pool. IBM Systems Journal, 5(3):148–157, 1966. CODEN IBMSA7. ISSN 0018-8670.

Bove:2000:HTR

[BDCA00] V. M. Bove, Jr., J. Dakss, E. Chalom, and S. Agamanolis. Hyperlinked television research at the MIT Media Laboratory. IBM Systems Journal, 39(3/4):470–478, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part1/bove.html;
                                        http://www.research.ibm.
                                        com/journal/sj/393/part1/
                                        bove.pdf;   http://www.research.
                                        ibm.com/journal/sj/393/part1/
                                        bove.txt.

Budinsky:2004:WSO

[BDE^{+}04] F. Budinsky, G. DeCandio, R. Earle, T. Francis, J. Jones, J. Li, M. Nally, C. Nelin, V. Popescu, S. Rich, A. Ryman, and T. Wilson. WebSphere Studio overview. IBM Systems Journal, 43(2):384–419, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/432/budinsky.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/budinsky.pdf;   http:

//www.research.ibm.com/journal/
sj/432/budinsky.txt.

Brown:2005:RSO

[BDE^{+}05] A. W. Brown, M. Delbaere, P. Eeles, S. Johnston, and R. Weaver. Realizing serviceoriented solutions with the IBM Rational Software Development Platform. IBM Systems Journal, 44(4):727–752, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/brown.html.

Baylor:2000:JSB

[BDF^{+}00] S. J. Baylor, M. Devarakonda, S. Fink, E. Gluzberg, M. Kalantar, P. Muttineni, E. Barsness, R. Arora, R. Dimpsey, and S. J. Munroe. Java server benchmarks. IBM Systems Journal, 39(1):57–81, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/391/baylor.html.

Buss:1994:IRE

[BDG^{+}94] E. Buss, R. De Mori, W. M.

Gentleman, J. Henshaw, H. Johnson, K. Kontogiannis, E. Merlo, H. A. Muller, J. Mylopoulos, S. Paul, A. Prakash, M. Stanley, S. R. Tilley, J. Troster, and K. Wong. Investigating reverse engineering technologies for the CAS program understanding project. IBM Systems Journal, 33(3):477–500, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj33-3.html#seven. G321-

5552.

Berk:1975:DPC

[BDH75] Murray A. Berk, Curtis W. Dunbar, and G. Carder Hobson. Design and performance considerations for the Retail Store System. IBM Systems Journal, 14(1):64–80, 1975. CODEN IBMSA7. ISSN 0018-8670.

Bethke:1981:IUP

[BDK^{+}81] F. J. Bethke, W. M. Dean, P. H. Kaiser, E. Ort, and F. H. Pessin. Improving the usability of programming publications. IBM Systems Journal, 20(3):306–320, 1981. CODEN IBMSA7. ISSN 0018-8670.

Black:2007:IMO

[BDL^{+}07] J. Black, C. Draper, T. Lococo, F. Matar, and C. Ward. An integration model for organizing IT service management. IBM Systems Journal, 46(3):405–??, July/September 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/463/black.html.

Banks:1997:ECE

[BDM97] T. Banks, K. E. Davies, and C. Moxey. The evolution of CICS/ESA in the sysplex environment. IBM Systems Journal, 36(2):352–360, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.

research.ibm.com/journal/
sj/362/banks.html.

Buchwald:1988:IAS

[BDS88] L. A. Buchwald, R. W. Davison, and W. P. Stevens. Integrating applications with SAA. IBM Systems Journal, 27(3):315–324, 1988. CODEN IBMSA7. ISSN 0018-8670.

Blasgen:1977:SAR

[BE77] Michael W. Blasgen and Kapali P. Eswaran. Storage and access in relational data bases. IBM Systems Journal, 16(4):363–377, 1977. CODEN IBMSA7. ISSN 0018-8670.

Botha:2001:SDA

[BE01] R. A. Botha and J. H. P.

Eloff. Separation of duties for access control enforcement in workflow environments. IBM Systems Journal, 40(3):666– 682, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/botha.html;   http:
//www.research.ibm.com/journal/
sj/403/botha.pdf.

Beck:1995:BPE

[Bec95] Kent Beck. Books: “Design Patterns: Elements of Reusable Object-Oriented Software”, by Erich Gamma, Richard Helm, Ralph Johnson, and John Vlissides. IBM Systems Journal, 34(3):544–545, 1995. CODEN IBMSA7. ISSN 0018-8670.

Bellamy:2007:SBD

[BEF^{+}07] R. K. E. Bellamy, T. Erickson, B. Fuller, W. A. Kellogg, R. Rosenbaum, J. C. Thomas, and T. Vetting Wolf. Seeing is believing: Designing visualizations for managing risk and compliance. IBM Systems Journal, 46(2):205–??, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/462/bellamy.html.

Balmin:2006:CBO

[BEH^{+}06] A. Balmin, T. Eliaz, J. Hornibrook, L. Lim, G. M. Lohman, D. Simmen, M. Wang, and C. Zhang. Cost-based optimization in DB2 XML. IBM Systems Journal, 45(2):299– 319, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/balmin.html.

Bowen:1997:LFP

[BEIW97] N. S. Bowen, D. A. Elko, J. F. Isenberg, and G. W. Wang. A locking facility for parallel systems. IBM Systems Journal, 36(2):202–220, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/362/bowen.html.

Belady:1966:SRA

[Bel66] L. A. Belady. A study of replacement algorithms for a virtual-storage computer. IBM Systems Journal, 5(2):78–101, 1966. CODEN IBMSA7. ISSN 0018-8670.

Bennett:1986:TBA

[Ben86] J. L. Bennett. Tools for building advanced user interfaces. IBM Systems Journal, 25(3-4):354–368, 1986. CODEN IBMSA7. ISSN 0018-8670.

Bencher:1994:PQI

[Ben94] D. L. Bencher. Technical forum: Programming quality improvement in IBM. IBM Systems Journal, 33(1):215–219, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-1.html.

Bender:1996:CP

[Ben96] W. Bender. Community and personalization. IBM Systems Journal, 35(3&4):367– 368, 1996. CODEN IBMSA7. ISSN 0018-8670.

Bender:2000:LE

[Ben00] W. Bender. Learning and expressing. IBM Systems Journal, 39(3/4):683–684, ???? 2000. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part2/bender.html;
 http://www.research.ibm.
com/journal/sj/393/part2/
bender.pdf;   http://www.
research.ibm.com/journal/
sj/393/part2/bender.txt.

Benantar:2001:IPK

[Ben01]  M. Benantar.  The Inter-

net public key infrastructure. IBM Systems Journal, 40(3): 648–665, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/403/benantar.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/403/benantar.pdf.

Belady:1980:GED

[BEP80] Laszio A. Belady, Carlo J. Evangelisti, and Leigh R. Power. GREENPRINT — a graphic representation of structured programs. IBM Systems Journal, 19(4):542–553, 1980. CODEN IBMSA7. ISSN 0018-8670.

Beretvas:1978:PTO

[Ber78] Thomas Beretvas. Performance tuning in OS/VS2 MVS. IBM Systems Journal, 17(3):290– 313, 1978. CODEN IBMSA7. ISSN 0018-8670.

Beregi:1984:APS

[Ber84] William E. Beregi. Architecture prototyping in the software engineering environment. IBM Systems Journal, 23(1):4– 18, 1984. CODEN IBMSA7. ISSN 0018-8670.

Berry:1988:CUA

[Ber88] R. E. Berry. Common User Access — a consistent and usable human-computer interface for the SAA environments. IBM Systems Journal, 27(3):281– 300, 1988. CODEN IBMSA7. ISSN 0018-8670.

Berggren:1992:BMT

[Ber92a] C. Berggren. The BiProcessor: a merger of two architectures. IBM Systems Journal, 31(3):535–549, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5485.

Berry:1992:DMC

[Ber92b] R. E. Berry. The designer’s model of the CUA workplace. IBM Systems Journal, 31(3):429–458, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5481.

Bin:2002:UCS

[BESZ02] E. Bin, R. Emek, G. Shurek, and A. Ziv. Using a constraint satisfaction formulation and solution techniques for random test program generation. IBM Systems Journal, 41(3): 386–402, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/bin.html.

Brown:1972:CDA

[BET72] D. T. Brown, R. L. Eibsen, and C. A. Thorn. Channel and direct access device architecture. IBM Systems Journal, 11(3):186–199, 1972. CODEN IBMSA7. ISSN 0018-8670.

Bevington:2000:TNB

[Bev00] D. Bevington. Technical note — business function specification of commercial applications. IBM Systems Journal, 39(2):315–335, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/bevington.html.

Baru:1995:DPE

[BFG^{+}95] C. K. Baru, G. Fecteau, A. Goyal, H. Hsiao, A. Jhingran, S. Padmanabhan, G. P. Copeland, and W. G. Wilson. DB2 Parallel Edition. IBM Systems Journal, 34(2):292–322, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-2.html#ten.

Brackenbury:1998:IES

[BFGS98] I. F. Brackenbury, D. F. Ferguson, K. D. Gottschalk, and R. A. Storey. IBM’s Enterprise Server for Java. IBM Systems Journal, 37(3):323–335, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/373/brackenbury.html.

Ballard:1989:MCS

[BFH89] C. P. Ballard, L. Farfara, and B. J. Heldke. Managing changes in SNA networks. IBM Systems Journal, 28(2):260–273, 1989. CODEN IBMSA7. ISSN 0018-8670.

Bauer:1994:RAD

[BFH^{+}94] M. A. Bauer, P. J. Finnigan, J. W. Hong, J. A. Rolia, T. J. Teorey, and G. A.

Winters. Reference architecture for distributed systems management. IBM Systems Journal, 33(3):426–444, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-3.html#four.  G321-

5549.

Brunet:2005:ARS

[BFH^{+}05] P. Brunet, B. A. Feigenbaum, K. Harris, C. Laws, R. Schwerdtfeger, and L. Weiss. Accessibility requirements for systems design to accommodate users with vision impairments. IBM Systems Journal, 44(3): 445–466, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/443/brunet.html.

Bender:1967:FDD

[BFS67] G. Bender, D. N. Freeman, and J. D. Smith. Function and design of DOS/360 and TOS/360. IBM Systems Journal, 6(1):2– 21, 1967. CODEN IBMSA7. ISSN 0018-8670.

Bailey:2008:TRE

[BFSLT08] Dan Bailey, Erwin Frank-Schultz, Pieter Lindeque, and Joseph L. Temple III. Three reliability engineering techniques and their application to evaluating the availability of IT systems: An introduction. IBM Systems Journal, 47 (4):577–589, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.

                                        research.ibm.com/journal/
                                        sj/474/bailey.pdf.

Budinsky:1996:ACG

[BFVY96] F. J. Budinsky, M. A. Finnie, J. M. Vlissides, and P. S. Yu. Automatic code generation from design patterns. IBM Systems Journal, 35(2):151–171, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj35-2.html#four.

Blake:1964:SSD

[BG64] K. Blake and G. Gordon. Systems simulation with digital computers. IBM Systems Journal, 3(1):14–21, 1964. CODEN IBMSA7. ISSN 0018-8670.

Black:1971:RTF

[BG71] B. C. Black and D. C. Gazis. Real-time traffic flow optimization. IBM Systems Journal, 10(3):217–231, 1971. CODEN IBMSA7. ISSN 0018-8670.

Bauer:2008:TSR

[BGERS08] D. Bauer, L. Garcés-Erice, S. Rooney, and P. Scotton. Toward scalable real-time messaging. IBM Systems Journal, 47(2):237–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/472/bauer.html.

Budhiraja:1995:MNC

[BGG^{+}95a] N. Budhiraja, M. Gopal, M. Gupta, E. A. Hervatic, S. J.

Nadas, and P. A. Stirpe. Multicast network connection architecture. IBM Systems Journal, 34(4):590–603, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-4.html#four.

Budhiraja:1995:NAN

[BGG^{+}95b] N. Budhiraja, M. Gopal, M. Gupta, E. A. Hervatic, S. J. Nadas, P. A. Stirpe, L. A. Tomek, and D. C. Verma. The NBBS access node. IBM Systems Journal, 34(4):694–704, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#ten.

Bryant:1976:ELQ

[BGH76] Peter Bryant, Fred W. Giesin, Jr., and Roger M. Hayes. Experiments in line quality monitoring. IBM Systems Journal, 15(2):124–142, 1976. CODEN IBMSA7. ISSN 0018-8670.

Bourbonnais:2004:TII

[BGH^{+}04] S. Bourbonnais, V. M. Gogate, L. M. Haas, R. W. Horman, S. Malaika, I. Narang, and V. Raman. Towards an information infrastructure for the grid. IBM Systems Journal, 43 (4):665–688, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/434/bourbonnais.html;
 http://www.research.ibm.
com/journal/sj/434/bourbonnais.

                                        pdf;   http://www.research.
                                        ibm.com/journal/sj/434/bourbonnais.
                                        txt.

Bhattacharya:2005:MDA

[BGL^{+}05] K. Bhattacharya, R. Guttman, K. Lyman, F. F. Heath III, S. Kumaran, P. Nandi, F. Wu, P. Athma, C. Freiberg, L. Johannsen, and A. Staudt. A model-driven approach to industrializing discovery processes in pharmaceutical research. IBM Systems Journal, 44(1):145–162, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/441/bhattacharya.html;
                                         http://www.research.ibm.
                                        com/journal/sj/441/bhattacharya.
                                        pdf;   http://www.research.
                                        ibm.com/journal/sj/441/bhattacharya.
                                        txt.

Bender:1996:TDH

[BGML96] W. Bender, D. Gruhl, N. Morimoto, and A. Liu. Techniques for data hiding. IBM Systems Journal, 35(3&4):313– 336, 1996. CODEN IBMSA7. ISSN 0018-8670.

Baldwin:1962:MAL

[BGP62] F. R. Baldwin, W. B. Gibson, and C. B. Poland. A multiprocessing approach to a large computer system. IBM Systems Journal, 1:64–76, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Benantar:1996:ACS

[BGT96] M. Benantar, R. Guski, and K. M. Troidle. Access control systems: From hostcentric to network-centric computing. IBM Systems Journal, 35(1):94–112, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-1.html#six.

Bird:1980:SM

[BH80] R. A. Bird and C. A. Hofmann.

Systems management. IBM Systems Journal, 19(1):140– 159, 1980. CODEN IBMSA7. ISSN 0018-8670.

Bhandari:1994:IPI

[BHC^{+}94] I. Bhandari, M. J. Halliday, J. Chaar, R. Chillarege, K. Jones, J. S. Atkinson, C. Lepori-Costello, P. Y. Jasper, E. D. Tarver, C. C. Lewis, and M. Yonezawa. Inprocess improvement through defect data interpretation. IBM Systems Journal, 33 (1):182–214, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-1.html#ten. G321-5538.

Borden:1989:MOS

[BHR89] T. L. Borden, J. P. Hennessy, and J. W. Rymarczyk. Multiple operating systems on one processor complex. IBM Systems Journal, 28(1):104–123, 1989. CODEN IBMSA7. ISSN 0018-

Branin:1965:IPM

[BHS^{+}65] F. H. Branin, Jr., L. V. Hall, J. Suez, R. M. Carlitz, and T. C. Chen. An interpretive program for matrix arithmetic. IBM Systems Journal, 4(1):2– 24, 1965. CODEN IBMSA7. ISSN 0018-8670.

Baily:1974:JES

[BHS74] James H. Baily, John A. Howard, and Thomas J. Szczygielski. The job entry subsystem of OS/VS1. IBM Systems Journal, 13(3):253–269, 1974. CODEN IBMSA7. ISSN 0018-8670.

Basson:2005:P

[BHS^{+}05] Sara Basson, Vicki Hanson, Michael Strack, John J. Ritsko, and David I. Seidman. Preface. IBM Systems Journal, 44 (3):443–44, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/443/preface.html.

Berry:2003:DEV

[BHT03] D. Berry, C. Hungate, and T. Temple. Delivering expected value to users and stakeholders with User Engineering. IBM Systems Journal, 42 (4):542–567, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/424/berry.pdf.

Baber:1999:CPD

[BHW99] C. Baber, D. J. Haniff, and S. I. Woolley. Contrasting paradigms for the development of wearable computers. IBM Systems Journal, 38(4): 551–565, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/baber.html.

Benjamin:1983:ISN

[BHWW83] J. H. Benjamin, M. L. Hess, R. A. Weingarten, and W. R. Wheeler. Interconnecting SNA networks. IBM Systems Journal, 23(4):344–366, 1983. CO-DEN IBMSA7. ISSN 0018-8670.

Bieberstein:1997:ADE

[Bie97] N. Bieberstein. Application development as an engineering discipline: Revolution or evolution? IBM Systems Journal, 36(1):??, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/361/bieberstein.html.

Brown:2006:RAM

[BIJ06] A. W. Brown, S. Iyengar, and S. Johnston. A rational approach to model-driven development. IBM Systems Journal, 45(3):463–480, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/453/brown.html.

Birch:1973:FSI

[Bir73] J. P. Birch. Functional structure of IBM Virtual Storage Operating Systems — 3 architecture and design of DOS/ VS. IBM Systems Journal, 12 (4):401–411, 1973. CODEN IBMSA7. ISSN 0018-8670.

Bjork:1975:GAT

[Bjo75] Lawrence A. Bjork, Jr. Generalized audit trail requirements and concepts for data base applications. IBM Systems Journal, 14(3):229–245, 1975. CO-DEN IBMSA7. ISSN 0018-8670.

Boyd:1963:EEM

[BK63] D. F. Boyd and H. S. Krasnow. Economic evaluation of management information systems. IBM Systems Journal, 2:2–23, March 1963. CODEN IBMSA7. ISSN 0018-8670.

Bichler:2002:AFP

[BKK^{+}02] M. Bichler, J. Kalagnanam, K. Katircioglu, A. J. King, R. D. Lawrence, H. S. Lee, G. Y. Lin, and Y. Lu. Applications of flexible pricing in business-to-business electronic commerce. IBM Systems Journal, 41(2):287–302, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/412/bichler.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/412/bichler.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/412/bichler.txt.

Boyd:1964:GPD

[BKP64] D. F. Boyd, H. S. Krasnow, and A. C. R. Petit. A general purpose digital simulator and examples of its application, part IV: Simulation of an integrated steel mill. IBM Systems Journal, 3(1):51–56, 1964. CODEN IBMSA7. ISSN 0018-8670.

Brown:1973:ISB

[BL73] J. F. Brown and D. W. Low.

Interactive simulation for banking. IBM Systems Journal, 12 (2):172–186, 1973. CODEN IBMSA7. ISSN 0018-8670.

Belady:1976:MLP

[BL76] Laszio A. Belady and M. M.

Lehman. A model of large program development. IBM Systems Journal, 15(3):225–252, 1976. CODEN IBMSA7. ISSN 0018-8670.

Bauchot:1995:IWR

[BL95] F. J. Bauchot and F. Lanne.

IBM wireless RF LAN design and architecture. IBM Systems Journal, 34(3):390–408, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-3.html#five.

Blaauw:1964:SSPb

[Bla64] G. A. Blaauw. The structure of SYSTEM/360, part V: Multisystem organization. IBM Systems Journal, 3(2):181–195, 1964. CODEN IBMSA7. ISSN 0018-8670.

Blackman:1998:TNI

[Bla98]  K. R. Blackman.  Techni-

cal note — IMS celebrates thirty years as an IBM product. IBM Systems Journal, 37 (4):596–603, ???? 1998. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/374/blackman.html.

Blum:1964:GPD

[Blu64] A. M. Blum. A general purpose digital simulator and examples of its application, part III: Digital simulation of urban traffic. IBM Systems Journal, 3(1):41– 50, 1964. CODEN IBMSA7. ISSN 0018-8670.

Baskin:1968:IGD

[BM68] H. B. Baskin and S. P. Morse. Interactive graphics in data processing: a multilevel modeling structure for interactive graphic design. IBM Systems Journal, 7(3/4):218–228, 1968. CODEN IBMSA7. ISSN 0018-8670.

Banham:1972:DFR

[BM72] J. A. Banham and P. Mc-Clelland. Design features of a real-time check-clearing system. IBM Systems Journal, 11(4):329–349, 1972. CODEN IBMSA7. ISSN 0018-8670.

Bird:1983:AGR

[BM83] D. L. Bird and C. U. Munoz. Automatic generation of random self-checking test cases. IBM Systems Journal, 23(3):

228–245, 1983. CODEN IBMSA7. ISSN 0018-8670.

Bender:1996:DEI

[BM96] W. Bender and R. MacNeil.

Design of electronic information. IBM Systems Journal, 35 (3&4):497–498, 1996. CODEN IBMSA7. ISSN 0018-8670.

Barrett:1999:IAM

[BM99] R. Barrett and P. P. Maglio.

Intermediaries: An approach to manipulating information streams. IBM Systems Journal, 38(4):629–641, ???? 1999. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/barrett.html.

Butcher:2002:IST

[BMK02] M. Butcher, H. Munro, and T. Kratschmer. Improving software testing via ODC: Three case studies. IBM Systems Journal, 41(1):31–44, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/411/butcher.html;   http:
//www.research.ibm.com/journal/
sj/411/butcher.pdf.

Borovoy:1996:TBC

[BMMR96] R. Borovoy, M. McDonald, F. Martin, and M. Resnick. Things that blink: Computationally augmented name tags. IBM Systems Journal, 35 (3&4):488–496, 1996. CODEN IBMSA7. ISSN 0018-8670.

Berry:2008:RSI

[BMP08] R. F. Berry, P. E. McKenney, and F. N. Parr. Responsive systems: An introduction. IBM Systems Journal, 47(2):197–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/472/berry.html.

Blandy:1989:VSM

[BN89] G. O. Blandy and S. R. Newson. VM/XA storage management. IBM Systems Journal, 28(1):175–191, 1989. CODEN IBMSA7. ISSN 0018-8670.

Bohrer:1998:ASF

[Boh98] K. A. Bohrer. Architecture of the San Francisco frameworks. IBM Systems Journal, 37(2): 156–169, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/372/bohrer.html.

Boies:1974:UBI

[Boi74] S. J. Boies. User behavior on an interactive computer system. IBM Systems Journal, 13(1):2– 18, 1974. CODEN IBMSA7. ISSN 0018-8670.

Boies:1999:UBI

[Boi99] S. J. Boies. User behavior on an interactive computer system. IBM Systems Journal, 38(2–3):162–180, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/382/boies.pdf. Reprinted

from [Boi74].

Bonner:1969:USM

[Bon69] A. J. Bonner. Using system monitor output to improve performance. IBM Systems Journal, 8(4):290–298, 1969. CO-DEN IBMSA7. ISSN 0018-8670.

Bontempo:1984:BR

[Bon84] Charles J. Bontempo. Book review. IBM Systems Journal, 23(4):386–387, 1984. CODEN IBMSA7. ISSN 0018-8670.

Boos:1985:HIM

[Boo85] W. Boos. HONE: the IBM Marketing Support System. IBM Systems Journal, 24(3-4):189– 199, 1985. CODEN IBMSA7. ISSN 0018-8670.

Botterill:1982:DRS

[Bot82] J. Howard Botterill. The design rationale of the System/ 38 user interface. IBM Systems Journal, 21(4):384–423, 1982. CODEN IBMSA7. ISSN 0018-8670.

Botterill:1989:DRA

[Bot89] J. H. Botterill. Design rationale of the AS/400 user interface. IBM Systems Journal, 28(3):443–464, 1989. CODEN IBMSA7. ISSN 0018-8670.

Boulanger:1998:CGH

[Bou98] A. Boulanger. Catapults and grappling hooks: The tools and techniques of information warfare. IBM Systems Journal, 37(1):106–114, 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/371/boulanger.html.

Boulanger:2005:OSV

[Bou05] A. Boulanger. Open-source versus proprietary software: Is one more secure and reliable than the other? IBM Systems Journal, 44(2):239–248, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/boulanger.html;
http://www.research.ibm.
com/journal/sj/442/boulanger.
pdf;   http://www.research.
ibm.com/journal/sj/442/boulanger.
txt.

Bove:1996:MBO

[Bov96] V. M. Bove, Jr. Multimedia based on object models: Some whys and hows. IBM Systems Journal, 35(3&4):337– 348, 1996. CODEN IBMSA7. ISSN 0018-8670.

Bozman:1989:VSM

[Boz89] G. P. Bozman. VM/XA SP2 minidisk cache. IBM Systems Journal, 28(1):165–174, 1989. CODEN IBMSA7. ISSN 0018-8670.

Boehm:1990:PDI

[BPZ90] G. E. Boehm, A. M. Palmiotti, and D. P. Zingaretti. Porting DPPX from the IBM 8100 to the IBM ES/9370: Installation and testing. IBM Systems Journal, 29(1):124–140, 1990. CODEN IBMSA7. ISSN 0018-8670.

Berry:1992:ECU

[BR92] R. E. Berry and C. J. Reeves.

The evolution of the Common User Access Workplace Model. IBM Systems Journal, 31(3):414–428, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5480.

Brackenbury:2002:MIT

[BR02] I. Brackenbury and Y. Ravin.

Machine intelligence and the Turing Test. IBM Systems Journal, 41(3):524–529, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/413/forum.html.

Birman:2003:P

[BR03] Alex Birman and John J. Ritsko. Preface. IBM Systems Journal, 42(2):203–204, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/preface.pdf.

Birman:2007:P

[BR07] A. Birman and J. J. Ritsko. Preface. IBM Systems Journal, 46(4):??, October/December 2007. CODEN IBMSA7. ISSN 0018-8670.

Basole:2008:CSV

[BR08a] R. C. Basole and W. B. Rouse.

Complexity of service value networks: Conceptualization and empirical investigation. IBM Systems Journal, 47(1):53–??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/471/basole.html.

Birman:2008:P

[BR08b] A. Birman and J. J. Ritsko. Preface. IBM Systems Journal, 47(2):195–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/472/preface.html.

Branscomb:1979:CCP

[Bra79] L. M. Branscomb. Computing and communications—A perspective of the evolving environment. IBM Systems Journal, 18(2):189–201, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Branscomb:1981:FHS

[Bra81] L. M. Branscomb. Foreword: The human side of computers. IBM Systems Journal, 20 (2):120–121, 1981. CODEN IBMSA7. ISSN 0018-8670.

Bresenham:1965:ACC

[Bre65] J. E. Bresenham. Algorithm for computer control of a digital plotter. IBM Systems Journal, 4(1):25–30, 1965. CODEN IBMSA7. ISSN 0018-8670.

Brewer:1971:ISS

[Bre71] A. C. Brewer. Interactive scheduling system. IBM Systems Journal, 10(1):62–79, 1971. CODEN IBMSA7. ISSN 0018-8670.

Britcher:1990:RSC

[Bri90] R. N. Britcher. Re-engineering software: a case study. IBM Systems Journal, 29(4):551– 567, 1990. CODEN IBMSA7. ISSN 0018-8670.

Briggs:1998:LE

[Bri98]  B. Briggs.  Lotus eSuite.

IBM Systems Journal, 37(3): 372–385, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/373/briggs.html.

Brodie:2002:IPC

[BRM02] M. Brodie, I. Rish, and S. Ma.

Intelligent probing: a costeffective approach to fault diagnosis in computer networks. IBM Systems Journal, 41(3): 372–385, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/brodie.html.

Brown:1968:IGD

[Bro68] S. H. Brown. Interactive graphics in data processing: Conversational job control. IBM Systems Journal, 7(3/4):271–280, 1968. CODEN IBMSA7. ISSN 0018-8670.

Bronner:1980:OCP

[Bro80] L. Bronner. Overview of the capacity planning process for production data processing. IBM Systems Journal, 19(1):4–27, 1980. CODEN IBMSA7. ISSN 0018-8670.

Brooks:1985:AHA

[Bro85] R. C. Brooks. Approach to high availability in high-transaction-rate systems. IBM Systems Journal, 24(3-4):279–293, 1985. CODEN IBMSA7. ISSN 0018-8670.

Brown:1991:IHP

[Bro91] P. Brown. Integrated hypertext and program understanding tools. IBM Systems Journal, 30(3):363–392, 1991. CO-DEN IBMSA7. ISSN 0018-8670.

Brennan:1967:TCS

[BS67] R. D. Brennan and M. Y. Silberberg. Two continuous system modeling programs. IBM Systems Journal, 6(4):242–266, 1967. CODEN IBMSA7. ISSN 0018-8670.

Bailey:1983:PAM

[BS83] R. M. Bailey and R. C. Soucy. Performance and availability measurement of the IBM Information Network. IBM Systems Journal, 23(4):404–416, 1983. CODEN IBMSA7. ISSN 0018-8670.

Bustamente:1994:DSA

[BS94] G. G. Bustamente and K. Sorenson. Decision support at Land’s End — an evolution. IBM Systems Journal, 33(2):228–238, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-2.html#two. G321-5540.

Bhaskaran:2004:WBI

[BS04] K. Bhaskaran and M.-T.

Schmidt. WebSphere business integration: An architectural overview. IBM Systems Journal, 43(2):238–253, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/432/bhaskaran.html;
http://www.research.ibm.
com/journal/sj/432/bhaskaran.
pdf;   http://www.research.
ibm.com/journal/sj/432/bhaskaran.
txt.

Borntrager:2005:PLS

[BS05] C. Bornträger and M. Schwidefsky. Providing Linux 2.6 support for the zSeries platform. IBM Systems Journal, 44(2): 331–340, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/borntraeger.html;
 http://www.research.ibm.
com/journal/sj/442/borntraeger.
pdf;   http://www.research.
ibm.com/journal/sj/442/borntraeger.
txt.

Brown:2001:TSK

[BSC^{+}01] E. W. Brown, S. Srinivasan, A. Coden, D. Ponceleon, J. W. Cooper, and A. Amir. Toward speech as a knowledge resource. IBM Systems Journal, 40(4): 985–1001, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/404/brown.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/404/brown.pdf.

Bigus:2002:ATB

[BSP^{+}02] J. P. Bigus, D. A. Schlosnagle, J. R. Pilgrim, W. N. Mills III, and Y. Diao. ABLE: a toolkit for building multiagent autonomic systems. IBM Systems Journal, 41(3):350–371, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/413/bigus.html.

Branscomb:1984:EUS

[BT84] Lewis M. Branscomb and John C. Thomas. Ease of use: a system design challenge. IBM Systems Journal, 23(3):224– 235, 1984. CODEN IBMSA7. ISSN 0018-8670.

Britton:1995:TNG

[BTB95] E. G. Britton, J. Tavs, and R. Bournas. TCP/IP: The next generation. IBM Systems Journal, 34(3):452–471, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj34-3.html#eight.

Buchholz:1963:FOA

[Buc63] Werner Buchholz. File organization and addressing. IBM Systems Journal, 2(1):86–111, June 1963. CODEN IBMSA7. ISSN 0018-8670. Comprehensive survey of hashing, with a good discussion of hash functions.

Buchholz:1969:SJM

[Buc69] W. Buchholz. A synthetic job for measuring system performance. IBM Systems Journal, 8(4):309–318, 1969. CODEN IBMSA7. ISSN 0018-8670.

Buckelew:1985:SPG

[Buc85] B. R. Buckelew. The system planning grid: a model for building integrated information systems. IBM Systems Journal, 24(3-4):294–306, 1985. CODEN IBMSA7. ISSN 0018-8670.

Buchholz:1986:ISV

[Buc86] W. Buchholz. The IBM System/370 vector architecture. IBM Systems Journal, 25(1): 51–62, 1986. CODEN IBMSA7. ISSN 0018-8670.

Bunting:2000:BFM

[Bun00] R. Bunting. Bridging the framework modeling and implementation gap. IBM Systems Journal, 39(2):267–284, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/392/bunting.html.

Burchi:1980:IGT

[Bur80] R. S. Burchi. Interactive graphics today. IBM Systems Journal, 19(3):292–313, 1980. CO-DEN IBMSA7. ISSN 0018-

Burbeck:1996:RCM

[Bur96] S. L. Burbeck. Real-time complexity metrics for Smalltalk methods. IBM Systems Journal, 35(2):204–226, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-2.html#seven.

Bux:1984:PIL

[Bux84] W. Bux. Performance issues in local-area networks. IBM Systems Journal, 23(4):351–374, 1984. CODEN IBMSA7. ISSN 0018-8670.

Boynton:1993:NCS

[BVP93] A. C. Boynton, B. Victor, and B. J. Pine, II. New competitive strategies: challenges to organizations and information technology. IBM Systems Journal, 32(1):40–64, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5502.

Broadbent:1993:IBI

[BW93] M. Broadbent and P. Weill. Improving business and information strategy alignment: Learning from the banking industry. IBM Systems Journal, 32 (1):162–179, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5508.

Cohen:1995:ASI

[CA95] M. E. Cohen and D. S. Abensour. An ATM strategy for IBM networking systems. IBM Systems Journal, 34(4):554–563, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#two.

Caffrey:2008:RCP

[Caf08] James M. Caffrey. The resiliency challenge presented by soft failure incidents. IBM Systems Journal, 47(4):641– 652, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/caffrey.pdf.

Calica:1965:FAOa

[Cal65a] A. B. Calica. Fabrication and assembly operations, part II: Long-range planning techniques. IBM Systems Journal, 4(2):94–104, 1965. CODEN IBMSA7. ISSN 0018-8670.

Calica:1965:FAOb

[Cal65b] A. B. Calica. Fabrication and assembly operations, part V: Production order sequencing. IBM Systems Journal, 4 (3):225–240, 1965. CODEN IBMSA7. ISSN 0018-8670.

Callaway:1975:PMT

[Cal75] Peter H. Callaway. Performance measurement tools for VM/370. IBM Systems Journal, 14(2):134–160, 1975. CO-DEN IBMSA7. ISSN 0018-8670.

Case:2006:EXS

[Cas06] P. Case. Enhancing XML search with XQuery 1.0 and XPath 2.0 full-text. IBM Systems Journal, 45(2):353– 360, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/case.html.

Cavin:1995:BMG

[Cav95] J. M. Cavin. Books: “Asynchronous Transfer Mode: Solution for Broadband ISDN”, Second Edition, by Martin de Prycker. IBM Systems Journal, 34(2):324–325, 1995. CODEN IBMSA7. ISSN 0018-8670.

Cavallo:2000:EDL

[Cav00] D. Cavallo. Emergent design and learning environments: Building on indigenous knowledge. IBM Systems Journal, 39(3/4):768–781, ???? 2000. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part2/cavallo.html;
 http://www.research.ibm.
com/journal/sj/393/part2/
cavallo.pdf;   http://www.
research.ibm.com/journal/
sj/393/part2/cavallo.txt.

Crawford:2005:TDS

[CBC^{+}05] C. H. Crawford, G. P. Bate, L. Cherbakov, K. Holley, and C. Tsocanos. Toward an on demand service-oriented architecture. IBM Systems Journal, 44(1):81–107, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/441/crawford.html;   http:

//www.research.ibm.com/journal/
sj/441/crawford.pdf;   http:
//www.research.ibm.com/journal/
sj/441/crawford.txt.

Chowdhary:2006:MDD

[CBC^{+}06] P. Chowdhary, K. Bhaskaran, N. S. Caswell, H. Chang, T. Chao, S.-K. Chen, M. Dikun, H. Lei, J.-J. Jeng, S. Kapoor, C. A. Lang, G. Mihaila, I. Stanoi, and L. Zeng. Model driven development for business performance management. IBM Systems Journal, 45(3): 587–605, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/chowdhary.html.

Clitherow:2008:CHA

[CBCS08] David Clitherow, Michael Brookbanks, Nicolas Clayton, and Gail Spear. Combining high availability and disaster recovery solutions for critical IT environments. IBM Systems Journal, 47(4):563–575, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/clitherow.pdf.

Chamberlin:1982:JID

[CBG^{+}82] D. D. Chamberlin, O. P.

Bertrand, M. J. Goodfellow, J. C. King, D. R. Slutz, S. J. P. Todd, and B. W. Wade. JANUS: an interactive document formatter based on declarative tags. IBM Systems Journal, 21(3):250–271, 1982.

CODEN IBMSA7. ISSN 0018-8670.

Cherbakov:2007:CCI

[CBG^{+}07] L. Cherbakov, A. Bravery, B. D. Goodman, A. Pandya, and J. Baggett. Changing the corporate IT development model: Tapping the power of grassroots computing. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/464/cherbakov.html.

Chou:1992:AN

[CBH^{+}92] J. H. Chou, C. R. Buckman, T. Hemp, A. Himwich, and F. Niemi. AIX NetView/ 6000. IBM Systems Journal, 31(2):270–285, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5473.

Chiu:1978:PMM

[CC78] Willy W. Chiu and We-Min Chow. A performance model of MVS. IBM Systems Journal, 17(4):444–462, 1978. CODEN IBMSA7. ISSN 0018-8670.

Chess:1987:LCC

[CC87] David M. Chess and Mike Cowlishaw. A large-scale computer conferencing system. IBM Systems Journal, 26(1): 138–153, January 1987. CO-DEN IBMSA7. ISSN 0018-8670.

Clark:1989:ASP

[CC89] B. E. Clark and M. J. Corrigan. Application System/ 400 performance characteristics. IBM Systems Journal, 28(3):407–423, 1989. CODEN IBMSA7. ISSN 0018-8670.

Colyer:2005:AOP

[CC05] A. Colyer and A. Clement.

Aspect-oriented programming with AspectJ. IBM Systems Journal, 44(2):301–308, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/colyer.html;   http:
//www.research.ibm.com/journal/
sj/442/colyer.pdf;   http:
//www.research.ibm.com/journal/
sj/442/colyer.txt.

Chen:2008:DTD

[CCC^{+}08] H. Chen, P. B. Chou, N. H.

Cohen, S. S. Duri, and C. W. Jung. DRIVE: a tool for developing, deploying, and managing distributed sensor and actuator applications. IBM Systems Journal, 47(2):289– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/chen.html.

Capper:1994:IOO

[CCHJ94] Nigel P. Capper, Roger J.

Colgate, Jim C. Hunter, and Martin F. James. The impact of object-oriented technology on software quality:

Three case histories. IBM Systems Journal, 33(1):131–157, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-1.html#eight. G321-

5536.

Chen:1978:SIS

[CCJ78] Richard L. Chen, Harlan Crowder, and Ellis L. Johnson. Solving the installation scheduling problem using mixed integer linear programming. IBM Systems Journal, 17(1):82–91, 1978. CODEN IBMSA7. ISSN 0018-8670.

Cobleigh:2002:FFS

[CCO02] J. M. Cobleigh, L. A. Clarke, and L. J. Osterweil. FLAVERS: a finite state verification technique for software systems. IBM Systems Journal, 41(1): 140–165, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/411/cobleigh.html;   http:
//www.research.ibm.com/journal/
sj/411/cobleigh.pdf.

Callaway:1972:UVS

[CCT72] P. H. Callaway, J. P. Considine, and C. H. Thompson. Uses of virtual storage systems in a scientific environment. IBM Systems Journal, 11(3):200–218, 1972. CODEN IBMSA7. ISSN 0018-8670.

Chen:1968:IGD

[CD68] F. C. Chen and R. L. Dougherty. Interactive graphics in data processing: a system for implementing interactive applications. IBM Systems Journal, 7(3/4):257–270, 1968. CODEN IBMSA7. ISSN 0018-8670.

Codella:1998:SEJ

[CDF^{+}98] C. F. Codella, D. N. Dillenberger, D. F. Ferguson, R. D. Jackson, T. A. Mikalsen, and I. Silva-Lepe. Support for Enterprise JavaBeans in Component Broker. IBM Systems Journal, 37(4):502–538, ???? 1998. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/374/codella.html.

Curbera:2005:CLM

[CDK^{+}05] F. Curbera, M. J. Duftler, R. Khalaf, W. A. Nagy, N. Mukhi, and S. Weerawarana. Colombo: Lightweight middleware for service-oriented computing. IBM Systems Journal, 44(4):799–820, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/444/curbera.html.

Cocchiara:2008:DCT

[CDK08] Richard Cocchiara, Hugh Davis, and Doug Kinnaird. Data center topologies for mission-critical business systems. IBM Systems Journal, 47(4):695–706, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/cocchiara.pdf.

Chang:2004:DEP

[CDM^{+}04] K. Chang, A. Dasari, H. Madduri, A. Mendoza, and J. Mims. Design of an enablement process for on demand applications. IBM Systems Journal, 43(1):190–203, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/431/chang.pdf.

Chandra:2006:ULD

[CdVF^{+}06] S. Chandra, J. de Vries, J. Field, H. Hess, M. Kalidasan, K. V. Raghavan, F. Nieuwerth, G. Ramalingam, and J. Xue. Using logical data models for understanding and transforming legacy business applications. IBM Systems Journal, 45(3):647–655, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/453/chandra.html.

Consens:1994:AAH

[CEH^{+}94] M. P. Consens, F. Ch. Eigler, M. Z. Hasan, A. O. Mendelzon, E. G. Noik, A. G. Ryman, and D. Vista. Architecture and applications of the hy+ visualization system. IBM Systems Journal, 33(3):458–476, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj33-3.html#six. G321-5551.

Cenfetelli:1967:DMC

[Cen67] A. R. Cenfetelli. Data management concepts for DOS/360 and TOS/360. IBM Systems Journal, 6(1):22–37, 1967. CO-DEN IBMSA7. ISSN 0018-8670.

Casey:1990:IFP

[CF90] R. G. Casey and D. R. Ferguson. Intelligent Forms Processing. IBM Systems Journal, 29(3):435–450, 1990. CODEN IBMSA7. ISSN 0018-8670.

Cortada:2005:MFS

[CF05] J. W. Cortada and H. E.

Fraser. Mapping the future in science-intensive industries: Lessons from the pharmaceutical industry. IBM Systems Journal, 44(1):163–183, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/441/cortada.html;   http:
//www.research.ibm.com/journal/
sj/441/cortada.pdf;   http:
//www.research.ibm.com/journal/
sj/441/cortada.txt.

Capek:2005:HIO

[CFGS05] P. G. Capek, S. P. Frank, S. Gerdt, and D. Shields. A history of IBM’s opensource involvement and strategy. IBM Systems Journal, 44 (2):249–257, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/capek.html;   http:
//www.research.ibm.com/journal/
sj/442/capek.pdf;   http:
//www.research.ibm.com/journal/
sj/442/capek.txt.

Cabri:2005:IRJ

[CFL05] G. Cabri, L. Ferrari, and L. Leonardi. Injecting roles in Java agents through runtime bytecode manipulation. IBM Systems Journal, 44(1): 185–208, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/441/cabri.html;   http:
//www.research.ibm.com/journal/
sj/441/cabri.pdf;   http:
//www.research.ibm.com/journal/
sj/441/cabri.txt.

Cozzi:2006:AMW

[CFL^{+}06] A. Cozzi, S. Farrell, T. Lau, B. A. Smith, C. Drews, J. Lin, B. Stachel, and T. P. Moran. Activity management as a Web service. IBM Systems Journal, 45(4):695–712, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/454/cozzi.html.

Corbett:1995:PFS

[CFP^{+}95] P. F. Corbett, D. G. Feitelson, J.-P. Prost, G. S. Almasi, S. J. Baylor, A. S. Bolmarcich, Y. Hsu, J. Satran, M. Snir, R. Colao, B. D. Herr, J. Kavaky, T. R. Morgan, and A. Zlotek. Parallel file systems for the IBM SP computers. IBM Systems Journal, 34(2):222–248, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-2.html#six.

Christensen:1981:PSS

[CFS81] Ken Christensen, George P.

Fitsos, and Charles P. Smith. A perspective on software science. IBM Systems Journal, 20(4):372–387, 1981. CODEN IBMSA7. ISSN 0018-8670.

Cole:1977:IRD

[CG77] Henderson Cole and Anthony A. Guido. The IBM 5100 and the Research Device Coupler—A personal laboratory automation system. IBM Systems Journal, 16(1):41–53, 1977. CODEN IBMSA7. ISSN 0018-8670.

Chroust:1990:RWM

[CGG90] G. Chroust, H. Goldmann, and O. Gschwandtner. The role of work management in application development. IBM Systems Journal, 29(2):189–208, 1990. CODEN IBMSA7. ISSN 0018-8670.

Cherbakov:2005:ISO

[CGH^{+}05] L. Cherbakov, G. Galambos, R. Harishankar, S. Kalyana, and G. Rackham. Impact of service orientation at the business level. IBM Systems Journal, 44(4):653–668, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/cherbakov.html.

Cheng:1998:SAI

[CGHK98] P.-C. Cheng, J. A. Garay, A. Herzberg, and H. Krawczyk. A security architecture for the Internet Protocol. IBM Systems Journal, 37(1):42–60, 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/371/cheng.html.

Conti:1968:SAS

[CGP68] C. J. Conti, D. H. Gibson, and S. H. Pitkowsky. Structural aspects of the SYSTEM/360 model 85, part I: General organization. IBM Systems Journal, 7(1):2–14, 1968. CODEN IBMSA7. ISSN 0018-8670.

Chow:2007:ECC

[CGRW07] A. W. Chow, B. D. Goodman, J. W. Rooney, and C. D. Wyble. Engaging a corporate community to manage technology and embrace innovation. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-8670.

Chessell:2002:ECT

[CGV^{+}02] M. Chessell, C. Griffin, D. Vines, M. Butler, C. Ferreira, and P. Henderson. Extending the concept of transaction compensation. IBM Systems Journal, 41(4):743–758, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/414/chessell.pdf.

Carpenter:1985:QEI

[CH85] M. B. Carpenter and H. K. Hallman. Quality emphasis at IBM’s Software Engineering Institute. IBM Systems Journal, 24(2):121–133, 1985. CODEN IBMSA7. ISSN 0018-8670.

Czarnecki:2006:FBS

[CH06] K. Czarnecki and S. Helsen.

Feature-based survey of model transformation approaches. IBM Systems Journal, 45(3): 621–645, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/czarnecki.html.

Christensen:2007:VEI

[CH07] J. Christensen and B. Hughes.

Voice-enabled IT transformation: The new voice technologies. IBM Systems Journal, 46(4):??, October/December 2007. CODEN IBMSA7. ISSN 0018-8670.

Chang:1966:QMS

[Cha66] W. Chang. A queuing model for a simple case of time sharing. IBM Systems Journal, 5 (2):115–125, 1966. CODEN IBMSA7. ISSN 0018-8670.

Chang:1970:SQP

[Cha70] W. Chang. Single-server queuing processes in computing systems. IBM Systems Journal, 9(1):36–71, 1970. CODEN IBMSA7. ISSN 0018-8670.

Charland:1989:ASH

[Cha89] D. A. Charland. The Application System/400 help facility— design philosophy and considerations. IBM Systems Journal, 28(3):424–442, 1989. CODEN IBMSA7. ISSN 0018-8670.

Chamberlin:2002:XXQ

[Cha02] D. Chamberlin. XQuery: An XML query language. IBM Systems Journal, 41(4):597– 615, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/414/chamberlin.pdf.

Cheng:1969:TDS

[Che69] P. S. Cheng. Trace-driven system modeling. IBM Systems Journal, 8(4):280–289, 1969. CODEN IBMSA7. ISSN 0018-8670.

Chess:1984:TCW

[Che84] David M. Chess. A tight coupling of workstations. IBM Systems Journal, 23(3):255–263, 1984. CODEN IBMSA7. ISSN 0018-8670.

Cheng:2001:AIK

[Che01] P.-C. Cheng. An architecture for the Internet Key Exchange Protocol. IBM Systems Journal, 40(3):721–746, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/403/cheng.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/403/cheng.pdf.

Christ:2000:SFP

[CHL^{+}00] R. Christ, S. L. Halter, K. Lynne, S. Meizer, S. J. Munroe, and M. Pasch. San Francisco performance: a case study in performance for large-scale Java applications. IBM Systems Journal, 39(1): 4–20, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/391/christ.html.

Christensen:1995:LAN

[CHNS95] K. J. Christensen, L. C. Haas, F. E. Noel, and N. C. Strole. Local area networks — evolving from shared to switched access. IBM Systems Journal, 34(3):347–374, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-3.html#three.

Chess:2007:PSI

[CHPW07] D. M. Chess, J. E. Hanson, J. A. Pershing, Jr., and S. R. White. Prospects for simplifying ITSM-based management through self-managing resources. IBM Systems Journal, 46(3):599–??, July/September 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/463/chess.html.

Church:1963:RGE

[Chu63] F. L. Church. Requirements generation, explosions, and bills of material. IBM Systems Journal, 2:268–287, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Carpenter:2004:AIS

[CJ04] B. E. Carpenter and P. A. Janson. Abstract interdomain security assertions: a basis for extra-grid virtual organizations. IBM Systems Journal, 43(4):689–701, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/434/carpenter.html;
                                        http://www.research.ibm.
                                        com/journal/sj/434/carpenter.
                                        pdf;   http://www.research.
                                        ibm.com/journal/sj/434/carpenter.
                                        txt.

Crowston:1998:CCM

[CK98] K. Crowston and E. E. Kammerer. Coordination and collective mind in software requirements development. IBM Systems Journal, 37(2):227– 245, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/372/crowston.html.

Cox:2005:MSO

[CK05] D. E. Cox and H. Kreger. Management of the serviceoriented-architecture life cycle. IBM Systems Journal, 44(4): 709–726, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/444/cox.html.

Cohen:1989:SHE

[CKB89] E. I. Cohen, G. M. King, and J. T. Brady. Storage hierarchies (Enterprise Systems Architecture/370). IBM Systems Journal, 28(1):62–76, 1989. CO-DEN IBMSA7. ISSN 0018-8670.

Cody:2002:IBI

[CKKS02] W. F. Cody, J. T. Kreulen, V. Krishna, and W. S. Spangler. The integration of business intelligence and knowledge management. IBM Systems Journal, 41(4):697–713, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/414/cody.pdf.

Clark:1966:FSO

[Cla66] W. A. Clark. The functional structure of OS/360, part III: Data management. IBM Systems Journal, 5(1):30–51, 1966. CODEN IBMSA7. ISSN 0018-8670.

Clark:1981:SST

[Cla81] I. A. Clark. Software simulation as a tool for usable product design. IBM Systems Journal, 20(3):272–293, 1981. CODEN IBMSA7. ISSN 0018-8670.

Clarkson:1985:SGI

[Cla85] T. B. Clarkson, III. Standardized graphics on the IBM Personal Computer. IBM Systems Journal, 24(1):3–13, 1985. CO-DEN IBMSA7. ISSN 0018-

Clark:1989:FEM

[Cla89] C. E. Clark. The facilities and evolution of MVS/ESA (operating system). IBM Systems Journal, 28(1):124–150, 1989. CODEN IBMSA7. ISSN 0018-8670.

Cleveland:1989:PUS

[Cle89] L. Cleveland. A program understanding support environment. IBM Systems Journal, 28(2):324–344, 1989. CODEN IBMSA7. ISSN 0018-8670.

Clements:1995:BDP

[Cle95] Paul C. Clements. Books: “The Mythical Man-Month — Essays on Software Engineering”, 20th anniversary edition, by Frederick P. Brooks, Jr. IBM Systems Journal, 34(3):545–547, 1995. CODEN IBMSA7. ISSN 0018-8670.

Chamberland:1998:IVJ

[CLR98] L. A. Chamberland, S. F. Lymer, and A. G. Ryman. IBM VisualAge for Java. IBM Systems Journal, 37(3):386– 408, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/373/chamberland.html.

Clementi:1988:IPP

[CLS88] E. Clementi, D. Logan, and J. Saarinen. ICAP/3090: parallel processing for large-scale scientific and engineering problems. IBM Systems Journal, 27(4):475–509, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Cheng:1984:IDP

[CLSW84] Josephine M. Cheng, Christopher R. Loosley, Akira Shibamiya, and Patricia S. Worthington. IBM database 2 performance: Design, implementation, and tuning. IBM Systems Journal, 23(2):189–210, 1984. CODEN IBMSA7. ISSN 0018-8670.

Considine:1977:MMA

[CM77] James P. Considine and James J.

Myers. MARC: MVS Archival Storage and Recovery program. IBM Systems Journal, 16(4):378–397, 1977. CODEN IBMSA7. ISSN 0018-8670.

Chang:1988:OED

[CM88] P. Y. Chang and W. W. Myre.

OS/2 EE database manager overview and technical highlights. IBM Systems Journal, 27(2):105–118, June 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Cheng:1994:ERD

[CMCD94] J. M. Cheng, N. M. Mattos, D. D. Chamberlin, and L. G. DeMichiel. Extending relational database technology for new applications. IBM Systems Journal, 33(2):264–279, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-2.html#four.  G321-

Colossi:2002:REO

[CMR02] N. Colossi, W. Malloy, and B. Reinwald. Relational extensions for OLAP. IBM Systems Journal, 41(4):714–731, ???? 2002. CODEN IBMSA7. ISSN 0018-8670.

Cordell:1987:AIE

[CMW87] R. Q. Cordell, II, M. Misra, and R. F. Wolfe. Advanced Interactive Executive program development environment. IBM Systems Journal, 26(4):361–382, 1987. CODEN IBMSA7. ISSN 0018-8670.

Corr:1979:SEI

[CN79] Frank P. Corr and Don H. Neal. SNA and emerging international standards. IBM Systems Journal, 18(2):244–262, 1979. CODEN IBMSA7. ISSN 0018-8670.

Clarke:2008:BSC

[CN08] R. J. Clarke and A. G. Nilsson. Business services as communication patterns: a work practice approach for analyzing service encounters. IBM Systems Journal, 47(1):129– ??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/471/clarke.html.

Chen:2008:RAS

[CNL08] Ing-Yi (Eric) Chen, Guo-Kai (Ken) Ni, and Chau-Young (Ivan) Lin. A runtimeadaptable service bus design for telecom operations support systems. IBM Systems Journal, 47(3):445–456, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/473/chen.pdf.

Caswell:2008:EVS

[CNS^{+}08] N. S. Caswell, C. Nikolaou, J. Sairamesh, M. Bitsaki, G. D. Koutras, and G. Iacovidis. Estimating value in service systems: a case study of a repair service system. IBM Systems Journal, 47(1):87–??, January/March 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/471/caswell.html.

Cole:1963:CLD

[COC63] H. Cole, Y. Okaya, and F. W.

Chambers. A computeroperated laboratory datataking system. IBM Systems Journal, 2:240–247, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Cockburn:1993:IOA

[Coc93] A. A. R. Cockburn. The impact of object-orientation on application development. IBM Systems Journal, 32(3):420–444, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5519.

Cockburn:1999:IOO

[Coc99] A. A. R. Cockburn. The impact of object-orientation on application development. IBM Systems Journal, 38(2– 3):308–334, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/382/cockburn.pdf. Reprinted

from [Coc93].

Cole:1974:SHL

[Col74] H. Cole. System/7 in a hierarchical laboratory automation system. IBM Systems Journal, 13(4):307–324, 1974. CODEN IBMSA7. ISSN 0018-8670.

Colosimo:1995:TFT

[Col95] J. A. Colosimo. Technical forum: The role of IBM’s Open Blueprint approach in distributed computing. IBM Systems Journal, 34(1):138–141, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj34-1.html.

Comba:1968:IGD

[Com68] P. G. Comba. Interactive graphics in data processing: a language for three-dimensional geometry. IBM Systems Journal, 7(3/4):292–307, 1968. CO-DEN IBMSA7. ISSN 0018-8670.

Comba:1990:ECI

[Com90] P. G. Comba. Exponentiation cryptosystems on the IBM PC. IBM Systems Journal, 29(4):526–538, 1990. CODEN IBMSA7. ISSN 0018-8670.

Connolly:1997:TFI

[Con97] D. Connolly. Technical forum: IBM’s electronic commerce solution: Commerce-POINT. IBM Systems Journal, 36(1):??, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/361/connolly.html.

Cooper:1980:CPM

[Coo80] J. C. Cooper. A capacity planning methodology. IBM Systems Journal, 19(1):28–45, 1980. CODEN IBMSA7. ISSN 0018-8670.

Corbi:1989:PUC

[Cor89] T. A. Corbi. Program understanding: Challenge for the 1990s. IBM Systems Journal, 28(2):294–306, 1989. CODEN IBMSA7. ISSN 0018-8670.

Cowlishaw:1984:DRL

[Cow84] M. F. Cowlishaw. The design of the REXX language. IBM Systems Journal, 23(4):326–335, 1984. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/234/ibmsj2304D.pdf. First

published as IBM Hursley Technical Report TR12.223, October 1983.

Chess:2003:SAC

[CPW03] D. M. Chess, C. C. Palmer, and S. R. White. Security in an autonomic computing environment. IBM Systems Journal, 42(1):107–118, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/421/chess.pdf.

Crabtree:1978:JN

[Cra78] Robert P. Crabtree. Job networking. IBM Systems Journal, 17(3):206–220, 1978. CODEN IBMSA7. ISSN 0018-8670.

Craft:1995:BGC

[Cra95] D. J. Craft. Books: “Managing Gigabytes: Compressing and Indexing Documents and Images”, by Ian H. Witten, Alistair Moffat, and Timothy C. Bell. IBM Systems Journal, 34(2):323–324, 1995. CODEN IBMSA7. ISSN 0018-8670.

Creasy:1968:IGD

[Cre68] R. J. Creasy. Interactive graphics in data processing: Neutron cross-section evaluation. IBM Systems Journal, 7(3/4):355– 364, 1968. CODEN IBMSA7. ISSN 0018-8670.

Cook:1988:WOS

[CRTW88] R. L. Cook, F. L. Rawson, III, J. A. Tunkel, and R. L. Williams. Writing an Operating System/2 application. IBM Systems Journal, 27(2): 134–157, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Crus:1984:DRI

[Cru84] Richard A. Crus. Data recovery in IBM Database 2. IBM Systems Journal, 23(2):178–188, 1984. CODEN IBMSA7. ISSN 0018-8670.

Chandersekaran:1976:FVM

[CS76] C. S. Chandersekaran and K. S.

Shankar. Forum: On virtual machine integrity. IBM Systems Journal, 15(3):264–269, 1976. CODEN IBMSA7. ISSN 0018-8670. See [DM75, DM76].

Chao:1996:ENT

[CT96] W. Chao and W. Tsun. Eventdriven network topology monitoring function. IBM Systems Journal, 35(1):25–36, 1996. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-1.html#three.

Cullum:1976:TSS

[Cul76] Phillip G. Cullum. The transmission subsystem in Systems Network Architecture. IBM Systems Journal, 15(1):24– 38, 1976. CODEN IBMSA7. ISSN 0018-8670. See erratum [Ano76j].

Chang:1965:CCI

[CW65] W. Chang and D. J. Wong.

Computer channel interference analysis. IBM Systems Journal, 4(2):162–170, 1965. CODEN IBMSA7. ISSN 0018-8670.

Chess:1981:VRL

[CW81] David M. Chess and Gerald Waldbaum. The VM/370 Resource Limiter. IBM Systems Journal, 20(4):424–437, 1981. CODEN IBMSA7. ISSN 0018-

Clark:1986:VSP

[CW86] R. S. Clark and T. L. Wilson. Vector system performance of the IBM 3090. IBM Systems Journal, 25(1):63–82, 1986. CODEN IBMSA7. ISSN 0018-8670.

Cypser:1992:EOC

[Cyp92] R. J. Cypser. Evolution of an open communications architecture. IBM Systems Journal, 31(2):161–188, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5468.

dArielli:1989:MMS

[d’A89] L. d’Arielli. A message management system for personal computers. IBM Systems Journal, 28(3):479–493, 1989. CODEN IBMSA7. ISSN 0018-8670.

Davenport:2000:SSC

[DAB^{+}00] G. Davenport, S. Agamanolis, B. Barry, B. Bradley, and K. Brooks. Synergistic storyscapes and constructionist cinematic sharing. IBM Systems Journal, 39(3/4):456– 469, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/davenport.html;
 http://www.research.ibm.
com/journal/sj/393/part1/
davenport.pdf;   http://
www.research.ibm.com/journal/
sj/393/part1/davenport.txt.

Dimpsey:2000:JSP

[DAK00] R. Dimpsey, R. Arora, and K. Kuiper. Java server performance: a case study of building efficient, scalable JVMs. IBM Systems Journal, 39(1): 151–174, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/391/dimpsey.html.

Darga:1970:OIU

[Dar70] K. Darga. On-line inquiry under a small-system operating system. IBM Systems Journal, 9(1):2–11, 1970. CODEN IBMSA7. ISSN 0018-8670.

Davies:1972:RM

[Dav72] P. M. Davies. Readings in microprogramming. IBM Systems Journal, 11(1):16–40, 1972. CODEN IBMSA7. ISSN 0018-8670.

Davies:1978:DPS

[Dav78] Charles T. Davies, Jr. Data processing spheres of control. IBM Systems Journal, 17(2):179–198, 1978. CODEN IBMSA7. ISSN 0018-8670.

Davis:1982:SIR

[Dav82] G. B. Davis. Strategies for information requirements determination. IBM Systems Journal, 21(1):4–30, 1982. CODEN IBMSA7. ISSN 0018-8670.

Davidson:1993:BRT

[Dav93] W. H. Davidson. Beyond reengineering: The three phases of business transformation. IBM Systems Journal, 32(1): 65–79, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5503.

Davidson:1999:BRE

[Dav99] W. H. Davidson. Beyond reengineering: The three phases of business transformation. IBM Systems Journal, 38(2– 3):485–501, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/davidson.pdf. Reprinted

from [Dav93].

Day:1970:CAD

[Day70] W. H. E. Day. Compiler assignment of data items to registers. IBM Systems Journal, 9(4):281–317, 1970. CODEN IBMSA7. ISSN 0018-8670.

Davis:2008:SSC

[DB08] M. M. Davis and I. Berdrow. Service science: Catalyst for change in business school curricula. IBM Systems Journal, 47(1):29–??, January/March 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/471/davis.html.

Dillenberger:2000:BJV

[DBC^{+}00] D. Dillenberger, R. Bordawekar, C. W. Clark, D. Durand, D. Emmes, O. Gohda, S. Howard, M. F. Oliver, F. Samuel, and R. W. St. John. Building a Java virtual machine for server applications: The JVM on OS/390. IBM Systems Journal, 39(1):194–210, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/391/dillenberger.html.

Davidson:2001:KKG

[DCB^{+}01] S. B. Davidson, J. Crabtree, B. P. Brunk, J. Schug, V. Tannen, G. C. Overton, and C. J. Stoeckert, Jr. K2/ Kleisli and GUS: Experiments in integrated access to genomic data sources. IBM Systems Journal, 40(2):512–531, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/davidson.html;   http:
//www.research.ibm.com/journal/
sj/402/davidson.pdf;   http:
//www.research.ibm.com/journal/
sj/402/davidson.txt.

Dillon:2003:TNI

[DCH^{+}03] M. R. Dillon, P. M. Collins, P. R. Hurley, J. E. Handlin, and S. A. Reed. Technical note — the IBM TotalStorage Enterprise Storage Server: Testing for general availability and beyond. IBM Systems Journal, 42(2):373–382, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/dillon.pdf.

Delgalvis:1964:SRD

[DD64] I. Delgalvis and G. Davison.

Storage requirements for a data exchange. IBM Systems Journal, 3(1):2–13, 1964. CODEN IBMSA7. ISSN 0018-8670.

Dan:2001:BBI

[DDK^{+}01] A. Dan, D. M. Dias, R. Kearney, T. C. Lau, T. N. Nguyen, F. N. Parr, M. W. Sachs, and H. H. Shaikh. Businessto-business integration with tpaML and a business-to-business protocol framework. IBM Systems Journal, 40(1): 68–90, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/401/dan.html.

Dan:2004:WSD

[DDK^{+}04] A. Dan, D. Davis, R. Kearney, A. Keller, R. King, D. Kuebler, H. Ludwig, M. Polan, M. Spreitzer, and A. Youssef. Web services on demand: WSLAdriven automated management. IBM Systems Journal, 43(1):136–158, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/431/dan.pdf.

Debelack:1995:NGA

[DDMS95] A. S. Debelack, J. D. Dehn, L. L. Muchinsky, and D. M. Smith. Next generation air traffic control automation. IBM Systems Journal, 34(1):63–77, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj34-1.html#five.

Dan:1997:MRM

[DDMS97] A. Dan, S. Dulin, S. Marcotte, and D. Sitaram. Multimedia resource management in OS/390 LAN Server. IBM Systems Journal, 36(3):393–410, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/363/dan.html.

Dean:1982:HCS

[Dea82] Morris Dean. How a computer should talk to people. IBM Systems Journal, 21(4):424–453, 1982. CODEN IBMSA7. ISSN 0018-8670.

Dryer:1999:WCP

[DEA99] D. C. Dryer, C. Eisbach, and W. S. Ark. At what cost pervasive? A social computing view of mobile computing systems. IBM Systems Journal, 38(4):652–676, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/384/dryer.html.

Deitch:1982:AQM

[Dei82] Martin Deitch. Analytic queuing model for CICS capacity planning. IBM Systems Journal, 21(4):454–470, 1982. CO-DEN IBMSA7. ISSN 0018-8670.

Deinhart:1992:SDF

[Dei92] K. Deinhart. SAA distributed file access to the CICS environment. IBM Systems Journal, 31(3):516–534, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5484.

DeMatteis:1968:ELT

[DeM68] J. J. DeMatteis. An economic lot-sizing technique, part I: The part-period algorithm. IBM Systems Journal, 7(1):30– 38, 1968. CODEN IBMSA7. ISSN 0018-8670.

Demers:1988:DFS

[Dem88] R. A. Demers. Distributed files for SAA. IBM Systems Journal, 27(3):348–361, 1988. CODEN IBMSA7. ISSN 0018-8670.

Denneau:2003:B

[Den03] Monty M. Denneau. Books. IBM Systems Journal, 42(1): 196–197, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/421/books.pdf.

DeSousa:1981:EII

[DeS81] M. R. DeSousa. Electronic information interchange in an office environment. IBM Systems Journal, 20(1):4–22, 1981. CODEN IBMSA7. ISSN 0018-8670.

Devereaux:1967:AMS

[Dev67] J. A. Devereaux. An applicationoriented multiprocessing system, part III: Control program features. IBM Systems Journal, 6(2):95–102, 1967. CODEN IBMSA7. ISSN 0018-8670.

Dewell:1990:CSP

[Dew90] M. E. Dewell. Cross System Product application generator: Application design. IBM Systems Journal, 29(2):265–273, 1990. CODEN IBMSA7. ISSN 0018-8670.

Dimsdale:1963:PES

[DF63] B. Dimsdale and H. P. Flatt.

Project evaluation and selection. IBM Systems Journal, 2: 200–217, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Delbaere:2007:ADA

[DF07] M. Delbaere and R. Ferreira.

Addressing the data aspects of compliance with industry models. IBM Systems Journal, 46 (2):319–??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/462/delbaere.html.

Demers:1992:IID

[DFGS92] R. A. Demers, J. D. Fisher, S. S. Gaitonde, and R. R. Sanders. Inside IBM’s Distributed Data Management architecture. IBM Systems Journal, 31(3):459–487, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5482.

Davis:1990:GAI

[DFHS90] A. G. Davis, J. R. Friedman, M. J. Haims, and G. C. Stierhoff. Guidelines for authors of the IBM Systems Journal. IBM Systems Journal, 29 (4):568–574, 1990. CODEN IBMSA7. ISSN 0018-8670.

Davis:1994:AGI

[DFS94] A. G. Davis, J. R. Friedman, and C. R. Seddon. Author guidelines for the IBM Systems Journal. IBM Systems Journal, 33(4):692–699, 1994. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sjaulong.

html. G321-5554.

Davis:1996:CGS

[DGK96] M. E. Davis, J. D. Grimes, and D. J. Knoles. Creating global software: Text handling and localization in Taligent’s CommonPoint application system. IBM Systems Journal, 35(2):227–243, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj35-2.html#eight.

Deaton:1983:XRR

[DH83] George A. Deaton, Jr. and Richard A. Hippert, Jr. X.25 and related recommendations in IBM products. IBM Systems Journal, 23(1-2):11–29, 1983. CODEN IBMSA7. ISSN 0018-8670.

Deen:2006:RQI

[DHB^{+}06] G. Deen, M. Hammer, J. Bethencourt, I. Eiron, J. Thomas, and J. H. Kaufman. Running Quake II on a grid. IBM Systems Journal, 45(1):21–44, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/451/deen.html.

Dube:2008:HHM

[DHK^{+}08] P. Dube, N. Halim, K. Karenos, M. Kim, Z. Liu, S. Parthasarathy, D. Pendarakis, and H. Yang. Harmony: Holistic messaging middleware for event-driven systems. IBM Systems Journal, 47(2):281–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/472/dube.html.

Diao:2002:UFC

[DHP02] Y. Diao, J. L. Hellerstein, and S. Parekh. Using fuzzy control to maximize profits in service level management. IBM Systems Journal, 41(3): 403–420, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/diao.html.

Diao:2003:MWS

[DHPB03] Y. Diao, J. L. Hellerstein, S. Parekh, and J. P. Bigus. Managing Web server performance with AutoTune agents. IBM Systems Journal, 42(1): 136–149, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/421/diao.pdf.

Davis:1989:P

[DHS89] Al Davis, Murray Heims, and George Stierhoff. Preface. IBM Systems Journal, 28(1): 2–3, 1989. CODEN IBMSA7. ISSN 0018-8670.

Dimsdale:1963:CCM

[Dim63] B. Dimsdale. Computer construction of minimal project networks. IBM Systems Journal, 2:24–36, March 1963. CO-DEN IBMSA7. ISSN 0018-8670.

Donnan:1974:SDL

[DK74] R. A. Donnan and J. R. Kersey. Synchronous data link control: a perspective. IBM Systems Journal, 13(2):140–162, 1974. CODEN IBMSA7. ISSN 0018-8670.

Doherty:1979:FAR

[DK79a] W. J. Doherty and R. P. Kelisky. Forum: Authors’ response. IBM Systems Journal, 18(4):605–607, 1979. CODEN IBMSA7. ISSN 0018-8670. See [DK79b, Rei79].

Doherty:1979:MVS

[DK79b] W. J. Doherty and R. P. Kelisky. Managing VM/CMS systems for user effectiveness. IBM Systems Journal, 18(1):143– 163, 1979. CODEN IBMSA7. ISSN 0018-8670. See letters [Rei79, DK79a].

Duan:2001:CPF

[DK01] Y. Duan and P. A. Kollman. Computational protein folding: From lattice to all-atom. IBM Systems Journal, 40(2): 297–308, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/402/duan.html;   http:
//www.research.ibm.com/journal/
sj/402/duan.pdf;   http:/
/www.research.ibm.com/journal/
sj/402/duan.txt.

deFreitas:1978:MTA

[dL78] Simplicio Lopes de Freitas and Pierre J. Lavell. A method for the time analysis of programs. IBM Systems Journal, 17(1): 26–38, 1978. CODEN IBMSA7. ISSN 0018-8670.

Dadam:1989:AIM

[DL89] P. Dadam and V. Linnemann. Advanced Information Management (AIM): advanced database technology for integrated applications. IBM Systems Journal, 28(4):661–681, 1989. CODEN IBMSA7. ISSN 0018-8670.

deLima:1989:GIE

[dLMSC89] A. H. V. de Lima, R. C. B. Martins, R. Stern, and L. M. F. Carneiro. GARDEN—an integrated and evolving environment for ULSI/VLSI CAD applications. IBM Systems Journal, 28(4):580–599, 1989. CO-DEN IBMSA7. ISSN 0018-8670.

Pauw:2005:WSN

[DLP^{+}05] W. De Pauw, M. Lei, E. Pring, L. Villard, M. Arnold, and J. F.

Morar. Web Services Navigator: Visualizing the execution of Web Services. IBM Systems Journal, 44(4):821–845, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/444/depauw.html.

Dimsdale:1964:DSL

[DM64] B. Dimsdale and H. M. Markowitz. A description of the SIMSCRIPT language. IBM Systems Journal, 3(1):57–67, 1964. CODEN IBMSA7. ISSN 0018-8670.

Donovan:1975:HAC

[DM75] John J. Donovan and S. E. Madnick. Hierarchical approach to computer system integrity. IBM Systems Journal, 14(2):188–202, 1975. CODEN IBMSA7. ISSN 0018-8670. See letters [CS76, DM76].

Donovan:1976:FAR

[DM76] J. J. Donovan and S. E. Madnick. Forum: Authors’ reply. IBM Systems Journal, 15 (3):270–278, 1976. CODEN IBMSA7. ISSN 0018-8670. See [DM75, CS76].

Devlin:1988:ABI

[DM88] B. A. Devlin and P. T. Murphy. An architecture for a business and information system. IBM Systems Journal, 27(1): 60–80, January 1988. CODEN IBMSA7. ISSN 0018-8670.

deBry:1993:AFP

[dM93] R. K. deBry and M. W.

Munger. Advanced Function Printing—from print to presentation. IBM Systems Journal, 32(4):647–664, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5528.

Davenport:1997:ASS

[DM97] G. Davenport and M. Murtaugh. Automatist storyteller systems and the shifting sands of story. IBM Systems Journal, 36(3):446–456, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/363/davenport.html.

Dimsdale:1999:DSL

[DM99] B. Dimsdale and H. M.

Markowitz. A description of the SIMSCRIPT language. IBM Systems Journal, 38(2– 3):151–161, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/dimsdale.pdf. Reprinted

from [DM64].

DAmora:2006:HPS

[DMB^{+}06] B. D’Amora, K. Magerlein, A. Binstock, A. Nanda, and B. Yee. High-performance server systems and the next generation of online games. IBM Systems Journal, 45(1): 103–118, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.

                                        research.ibm.com/journal/
                                        sj/451/damora.html.

Dauphin:1992:ETR

[DMMR92] M. Dauphin, M. M. Marques, A. P. Mullery, and P. Rodier. The European telecommunications research and development program RACE and its software project SPECS. IBM Systems Journal, 31(4):649–667, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5491.

Mes:2003:TNW

[dMR03] A. de Mes and E. Rongen. Technical note—Web Service Credentials. IBM Systems Journal, 42(3):532–537, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/423/demes.pdf.

Dunfee:1988:DSA

[DMRS88] W. P. Dunfee, J. D. McGehe, R. C. Rauf, and K. O. Shipp. Designing SAA applications and user interfaces. IBM Systems Journal, 27(3):325–347, 1988. CODEN IBMSA7. ISSN 0018-8670.

Dash:1984:IDI

[DO84] Jnan R. Dash and Robert N. Ojala. IBM Database 2 in an Information Management System environment. IBM Systems Journal, 23(2):165–177, 1984. CODEN IBMSA7. ISSN 0018-8670.

Donofrio:2003:MSV

[Don03] Nicholas M. Donofrio. Message from the Senior Vice President, Technology and Manufacturing. IBM Systems Journal, 42(4):??, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/424/message.pdf.

Dorrance:1963:DIP

[Dor63] R. T. Dorrance. Design of an integrated programming and operating system, part V: The system’s COBOL compiler. IBM Systems Journal, 2: 322–327, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Doherty:1986:CTH

[DP86] W. J. Doherty and W. G. Pope.

Computing as a tool for human augmentation. IBM Systems Journal, 25(3-4):306–320, 1986. CODEN IBMSA7. ISSN 0018-8670.

deBry:1988:AFP

[dP88] R. K. deBry and B. G. Platte.

Advanced Function Printing: a tutorial. IBM Systems Journal, 27(2):219–233, June 1988. CODEN IBMSA7. ISSN 0018-8670.

DiAngelo:1995:CPB

[DP95] M. F. DiAngelo and C. J.

Petrun. Collecting productbased usability requirements. IBM Systems Journal, 34 (1):4–19, 1995. CODEN IBMSA7. ISSN 0018-8670.

URL http://www.research.
ibm.com/journal/sj34-1.html#
two.

deBry:1988:AAF

[dPBM88] R. K. deBry, B. G. Platte, C. L. Berinato, and J. W. Marlin. Architectures of Advanced Function Printing. IBM Systems Journal, 27(2):234–245, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Dinan:1990:IHP

[DPR90] R. F. Dinan, L. D. Painter, and R. R. Rodite. ImagePlus High Performance Transaction System. IBM Systems Journal, 29(3):421–434, 1990. CODEN IBMSA7. ISSN 0018-8670.

Dias:2002:CII

[DPR^{+}02] D. M. Dias, S. L. Palmer, J. T. Rayfield, H. H. Shaikh, and T. K. Sreeram. E-commerce interoperability with IBM’s Web-Sphere Commerce products. IBM Systems Journal, 41(2): 272–286, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/412/dias.html;   http:
//www.research.ibm.com/journal/
sj/412/dias.pdf;   http:/
/www.research.ibm.com/journal/
sj/412/dias.txt.

Drummond:1969:PSP

[Dru69] M. E. Drummond, Jr. A perspective on system performance evaluation. IBM Systems Journal, 8(4):252–263, 1969. CO-DEN IBMSA7. ISSN 0018-8670.

desRivieres:2004:EPI

[dRW04] J. des Riviêres and J. Wiegand. Eclipse: a platform for integrating development tools. IBM Systems Journal, 43(2): 371–383, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/432/desrivieres.html;
 http://www.research.ibm.
com/journal/sj/432/desrivieres.
pdf;   http://www.research.
ibm.com/journal/sj/432/desrivieres.
txt.

Dixon:1983:TNL

[DSM83] Roy C. Dixon, Norman C. Strole, and James D. Markov. Token-ring network for local data communications. IBM Systems Journal, 23(1-2):47– 62, 1983. CODEN IBMSA7. ISSN 0018-8670.

Dueck:2001:VKH

[Due01] G. Dueck. Views of knowledge are human views. IBM Systems Journal, 40(4):885–888, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/404/dueck.html;   http:
//www.research.ibm.com/journal/
sj/404/dueck.pdf.

Duke:1975:TCS

[Duk75] Melvin O. Duke. Testing in a complex systems environment. IBM Systems Journal, 14(4):353–365, 1975. CODEN IBMSA7. ISSN 0018-8670.

Duke:1985:PGC

[DW85] K. A. Duke and W. A. Wall. A professional graphics controller. IBM Systems Journal, 24(1): 14–25, 1985. CODEN IBMSA7. ISSN 0018-8670.

Dean:2000:SMF

[DW00] S. Elo Dean and L. Weitzman. SuperNews: Multiple feeds for multiple views. IBM Systems Journal, 39(3/4):633–645, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part1/elodean.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part1/
                                        elodean.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part1/elodean.txt.

Demers:1992:DDC

[DY92] R. A. Demers and K. Yamaguchi. Data description and conversion architecture. IBM Systems Journal, 31(3):488– 515, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5483.

Dyer:1980:MSE

[Dye80] M. Dyer. The management of software engineering, part IV: Software development practices. IBM Systems Journal, 19(4):451–465, 1980. CODEN IBMSA7. ISSN 0018-8670.

Day:1968:IGD

[DZ68] C. W. Day and L. L. Zimmerman. Interactive graphics in data processing: Implementation and usage. IBM Systems Journal, 7(3/4):373–381, 1968. CODEN IBMSA7. ISSN 0018-8670.

Dzielinski:1965:FAO

[Dzi65] B. P. Dzielinski. Fabrication and assembly operations, part IV: Linear programming in production planning. IBM Systems Journal, 4(2):122–135, 1965. CODEN IBMSA7. ISSN 0018-8670.

Dzielinski:1973:GFP

[Dzi73] B. P. Dzielinski. A guide to financial planning tools and techniques. IBM Systems Journal, 12(2):126–144, 1973. CODEN IBMSA7. ISSN 0018-8670.

Eilam:2004:UUC

[EAB^{+}04] T. Eilam, K. Appleby, J. Breh, G. Breiter, H. Daur, S. A. Fakhouri, G. D. H. Hunt, T. Lu, S. D. Miller, L. B. Mummert, J. A. Pershing, and H. Wagner. Using a utility computing framework to develop utility systems. IBM Systems Journal, 43(1):97–120, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/431/eilam.html.

Eckman:2007:VIT

[EBKT07] B. A. Eckman, C. A. Bennett, J. H. Kaufman, and J. W. Tenner. Varieties of interoperability in the transformation of the health-care information infrastructure. IBM Systems Journal, 46(1):19–41, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/461/eckman.html.

Elliott:1996:OSM

[EDG^{+}96] P. J. Elliott, J. Diedrichsen, K. J. Goodson, R. Riste-Smith, and G. J. Sivewright. An object-oriented system for 3D medical image analysis. IBM Systems Journal, 35(1):4–24, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj35-1.html#two.

Edelstein:2002:MJP

[EFN^{+}02] O. Edelstein, E. Farchi, Y. Nir, G. Ratsaby, and S. Ur. Multithreaded Java program test generation. IBM Systems Journal, 41(1):111–125, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/411/edelstein.html;
                                        http://www.research.ibm.
                                        com/journal/sj/411/edelstein.
                                        pdf.

Efron:1964:GPD

[EG64] R. Efron and G. Gordon. A general purpose digital simulator and examples of its application, part I: Description of the simulator. IBM Systems Journal, 3(1):22–34, 1964. CODEN IBMSA7. ISSN 0018-8670.

Edelsohn:2005:CGC

[EGH^{+}05] D. Edelsohn, W. Gellerich, M. Hagog, D. Naishlos, M. Namolaru, E. Pasch, H. Penner, U. Weigand, and A. Zaks. Contributions to the GNU Compiler Collection. IBM Systems Journal, 44(2):259–278, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/edelsohn.html;   http:
//www.research.ibm.com/journal/
sj/442/edelsohn.pdf;   http:
//www.research.ibm.com/journal/
sj/442/edelsohn.txt.

Engel:1979:OCS

[EGLT79] G. H. Engel, J. Groppuso, R. A.

Lowenstein, and W. G. Traub. An office communications system. IBM Systems Journal, 18(3):402–431, 1979. CODEN IBMSA7. ISSN 0018-8670.

Eade:1977:CRS

[EHJ77] D. J. Eade, P. Homan, and J. H.

Jones. CICS/VS and its role in Systems Network Architecture. IBM Systems Journal, 16(3):258–286, 1977. CODEN IBMSA7. ISSN 0018-8670.

Eisenbies:1967:CDD

[Eis67] J. L. Eisenbies. Conventions for digital data communication link design. IBM Systems Journal, 6(4):267–302, 1967. CO-DEN IBMSA7. ISSN 0018-

Elliott:1992:IIS

[EKS92] P. J. Elliott, J. M. Knapman, and W. Schlegel. Interactive image segmentation for radiation treatment planning. IBM Systems Journal, 31(4):620–634, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5489.

Eustice:1999:UIA

[ELM^{+}99] K. F. Eustice, T. J. Lehman, A. Morales, M. C. Munson, S. Edlund, and M. Guillen. A universal information appliance. IBM Systems Journal, 38 (4):575–601, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/eustice.html.

Elofson:1995:IAE

[Elo95] G. Elofson. Intelligent agents extend knowledge-based systems feasibility. IBM Systems Journal, 34(1):78–95, 1995. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-1.html#six.

Ehrsam:1978:CKM

[EMMT78] William F. Ehrsam, Stephen M. Matyas, Carl H. Meyer, and Walter L. Tuchman. A cryptographic key management scheme for implementing the Data Encryption Standard. IBM Systems Journal, 17(2):106–125, 1978. CODEN IBMSA7. ISSN 0018-8670.

Ernest:2007:AVI

[EN07] M. Ernest and J. M. Nisavic.

Adding value to the IT organization with the Component Business Model. IBM Systems Journal, 46(3):387–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/ernest.html.

Encarnacao:1981:FGK

[Enc81] J. Encarnacao. Forum: Graphical kernel system. IBM Systems Journal, 20(4):438–440, 1981. CODEN IBMSA7. ISSN 0018-8670.

Eichenberger:2006:UAC

[EOO^{+}06] A. E. Eichenberger, J. K.

O’Brien, K. M. O’Brien, P. Wu, T. Chen, P. H. Oden, D. A. Prener, J. C. Shepherd, B. So, Z. Sura, A. Wang, T. Zhang, P. Zhao, M. K. Gschwind, R. Archambault, Y. Gao, and R. Koo. Using advanced compiler technology to exploit the performance of the Cell Broadband Engine^{T M} architecture. IBM Systems Journal, 45(1): 59–84, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/451/eichenberger.html.

Englefield:2005:PAI

[EPTD05] P. Englefield, C. Paddison, M. Tibbits, and I. Damani. A proposed architecture for integrating accessibility test tools.

IBM Systems Journal, 44(3): 537–555, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/443/englefield.html.

Elson:1970:CTL

[ER70] M. Elson and S. T. Rake. Code-generation technique for large-language compilers. IBM Systems Journal, 9(3):166–188, 1970. CODEN IBMSA7. ISSN 0018-8670.

Emnett:2000:SNR

[ES00] K. Emnett and C. Schmandt. Synthetic news radio. IBM Systems Journal, 39(3/4):646– 659, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/393/part1/emnett.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part1/
                                        emnett.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part1/emnett.txt.

Emiliani:2005:UAA

[ES05] P. L. Emiliani and C. Stephanidis. Universal access to ambient intelligence environments: Opportunities and challenges for people with disabilities. IBM Systems Journal, 44(3): 605–619, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/443/emiliani.html.

ElSadany:1989:AMS

[ESH89] T. A. El-Sadany and M. A.

Hashish. An Arabic morphological system. IBM Systems Journal, 28(4):600–612, 1989. CODEN IBMSA7. ISSN 0018-8670.

Eggebraaten:2007:HCD

[ETD07] T. J. Eggebraaten, J. W. Tenner, and J. C. Dubbels. A health-care data model based on the HL7 Reference Information Model. IBM Systems Journal, 46(1):5–18, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/461/eggebraaten.html.

Evard:1996:CIN

[Eva96] M. Evard. Children’s interests in news: On-line opportunities. IBM Systems Journal, 35 (3&4):417–430, 1996. CODEN IBMSA7. ISSN 0018-8670.

Evernden:1996:IF

[Eve96]  R. Evernden.  The Infor-

mation FrameWork. IBM Systems Journal, 35(1):37–68, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj35-1.html#four.

Esau:1966:TSD

[EW66] L. R. Esau and K. C. Williams.

On teleprocessing system design, part II: a method for approximating the optimal network. IBM Systems Journal, 5(3):142–147, 1966. CODEN IBMSA7. ISSN 0018-8670.

Fagan:1976:DCI

[Fag76] Michael E. Fagan. Design and code inspections to reduce errors in program development. IBM Systems Journal, 15(3):182–211, 1976. CODEN IBMSA7. ISSN 0018-8670.

Fagan:1999:DCI

[Fag99] M. E. Fagan. Design and code inspections to reduce errors in program development. IBM Systems Journal, 38(2– 3):258–288, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/fagan.pdf. Reprinted

from [Fag76].

Fahlman:2002:TFS

[Fah02] S. E. Fahlman. Technical forum: Selling interrupt rights: a way to control unwanted e-mail and telephone calls. IBM Systems Journal, 41(4): 759–766, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/414/chessell.pdf.

Falkoff:1991:IFA

[Fal91] A. D. Falkoff. The IBM family of APL systems. IBM Systems Journal, 30(4):416–432, 1991. CODEN IBMSA7. ISSN 0018-8670.

Farrell:1987:VIC

[Far87] E. J. Farrell. Visual interpretation of complex data. IBM Systems Journal, 26(2):174–200, 1987. CODEN IBMSA7. ISSN 0018-8670.

Finkel:1992:RCI

[FC92] A. J. Finkel and S. B. Calo.

RODM: a control information base. IBM Systems Journal, 31(2):252–269, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5472.

Fuh:1998:LDT

[FCM^{+}98] G. Y. Fuh, J.-H. Chow, N. M.

Mattos, B. T. Tran, and T. C. Truong. On the linkage of dynamic tables in relational DBMSs. IBM Systems Journal, 37(4):539–551, ???? 1998. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/374/fuh.html.

Ferdinand:1971:AMI

[Fer71] A. E. Ferdinand. An analysis of the machine interference model. IBM Systems Journal, 10(2):129–142, 1971. CODEN IBMSA7. ISSN 0018-8670.

Feriozi:1991:CPM

[Fer91] D. T. Feriozi. A C programming model for OS/2 device drivers. IBM Systems Journal, 30(3):322–335, 1991. CO-DEN IBMSA7. ISSN 0018-8670. G321-5438.

Feriozi:1992:SMO

[Fer92] D. T. Feriozi. A split model for OS/2 SCSI device drivers. IBM Systems Journal, 31(1):114– 122, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5465.

Fanderl:1992:ODA

[FFK92] H. Fanderl, K. Fischer, and J. Kamper. The Open Document Architecture: From standardization to the market. IBM Systems Journal, 31(4):728– 754, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5495.

Fitzgerald:1979:DMD

[FG79] A. K. Fitzgerald and B. F. Goodrich. Data management for the Distributed Processing Programming Executive (DPPX). IBM Systems Journal, 18(4):547–564, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Fennel:1983:SCC

[FG83] John W. Fennel, Jr. and Bruce D. Gobioff. Satellite communications controller. IBM Systems Journal, 23(1-2): 81–95, 1983. CODEN IBMSA7. ISSN 0018-8670.

Faulk:1990:SDD

[FG90] R. D. Faulk and F. Goertzel Gustavson. Segmenting discrete data representing continuous speech input. IBM Systems Journal, 29(2):287–296, 1990. CODEN IBMSA7. ISSN 0018-8670.

Finnigan:1997:SB

[FHK^{+}97] P. J. Finnigan, R. C. Holt, I. Kalas, S. Kerr, K. Kontogiannis, H. A. Müller, J. Mylopoulos, S. G. Perelgut, M. Stanley, and K. Wong. The software bookshelf. IBM Systems Journal, 36(4):564–593, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/364/finnigan.html.

Farchi:2002:UMB

[FHP02] E. Farchi, A. Hartman, and S. S. Pinter. Using a modelbased test generator to test for standard conformance. IBM Systems Journal, 41(1):89– 110, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/411/farchi.html;   http:
//www.research.ibm.com/journal/
sj/411/farchi.pdf.

Falkoff:1964:FDS

[FIS64] A. D. Falkoff, K. E. Iverson, and E. H. Sussenguth. A formal description of SYS-TEM/360. IBM Systems Journal, 3(3):198–261, 1964. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/032/falkoff.pdf.

Fischofer:1991:VSS

[Fis91] W. T. Fischofer. VM/ESA:

a single system for centralized and distributed computing. IBM Systems Journal, 30(1):4–13, 1991. CODEN IBMSA7. ISSN 0018-8670.

Ferguson:2001:WBS

[FK01] D. F. Ferguson and R. Kerth. WebSphere as an e-business server. IBM Systems Journal, 40(1):25–45, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/401/ferguson.html.

Farrell:2002:WSM

[FK02] J. A. Farrell and H. Kreger. Web services management approaches. IBM Systems Journal, 41(2):212–227, ???? 2002. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/412/farrell.html;   http:
//www.research.ibm.com/journal/
sj/412/farrell.pdf;   http:
//www.research.ibm.com/journal/
sj/412/farrell.txt.

Funderburk:2002:XBR

[FKS^{+}02] J. E. Funderburk, G. Kiernan, J. Shanmugasundaram, E. Shekita, and C. Wei. XTA-BLES: Bridging relational technology and XML. IBM Systems Journal, 41(4):616–641, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/414/funderburk.pdf.

Funderburk:2003:TNX

[FKS^{+}03] J. E. Funderburk, G. Kiernan, J. Shanmugasundaram, E. Shekita, and C. Wei. Technical note—XTABLES: Bridging relational technology and XML. IBM Systems Journal, 42(3):538–541, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/funderburk.pdf.

Ferrucci:2004:BEA

[FL04] D. Ferrucci and A. Lally. Building an example application with the Unstructured Information Management Architecture. IBM Systems Journal, 43 (3):455–475, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/433/ferrucci.html;   http:
//www.research.ibm.com/journal/
sj/433/ferrucci.pdf;   http:
//www.research.ibm.com/journal/
sj/433/ferrucci.txt.

Flaherty:1985:PPP

[Fla85] M. J. Flaherty. Programming process productivity measurement system for System/

370. IBM Systems Journal, 24(2):168–175, 1985. CODEN IBMSA7. ISSN 0018-8670.

Fletcher:1996:FTM

[Fle96]  R. Fletcher.  Force trans-

duction materials for humantechnology interfaces. IBM Systems Journal, 35(3&4):630– 638, 1996. CODEN IBMSA7. ISSN 0018-8670.

Flint:1997:CJP

[Fli97] E. S. Flint. The COBOL jigsaw puzzle: Fitting objectoriented and legacy applications together. IBM Systems Journal, 36(1):??, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/361/flint.html.

Funderburk:2002:XPS

[FMR02] J. E. Funderburk, S. Malaika, and B. Reinwald. XML programming with SQL/XML and XQuery. IBM Systems Journal, 41(4):642–665, ???? 2002. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/414/reinwald.pdf.

Fontes:2004:WCA

[FNS04] S. M. Fontes, C. J. Nordstrom, and K. W. Sutter. WebSphere connector architecture evolution. IBM Systems Journal, 43(2):316–326, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/432/fontes.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/fontes.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/fontes.txt.

Ford:1978:EPD

[For78] Julian B. Ford. Enhanced problem determination capability for teleprocessing. IBM Systems Journal, 17(3):276–289, 1978. CODEN IBMSA7. ISSN 0018-8670.

Franklin:1990:EAD

[FP90] S. M. Franklin and A. M.

Peters. Effective application development for Presentation Manager programs. IBM Systems Journal, 29(1):44–58, 1990. CODEN IBMSA7. ISSN 0018-8670.

Frazer:1979:PTI

[Fra79] W. D. Frazer. Potential technology implications for computers and telecommunications in the 1980s. IBM Systems Journal, 18(2):333–347, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Freeman:1966:MLD

[Fre66] D. N. Freeman. Macro language design for SYSTEM/360. IBM Systems Journal, 5(2):62– 77, 1966. CODEN IBMSA7. ISSN 0018-8670.

Fredericksen:1973:DDC

[Fre73] D. H. Fredericksen. Describing data in computer networks. IBM Systems Journal, 12(3):257–282, 1973. CODEN IBMSA7. ISSN 0018-8670.

Friedman:1970:APS

[Fri70] T. D. Friedman. The authorization problem in shared files. IBM Systems Journal, 9(4):258–280, 1970. CODEN IBMSA7. ISSN 0018-8670.

Fromm:1972:SAL

[FS72] J. E. Fromm and D. E.

Schreiber. System aspects of large-problem computation and display. IBM Systems Journal, 11(1):41–55, 1972. CODEN IBMSA7. ISSN 0018-8670.

Ferguson:2005:SOA

[FS05] D. F. Ferguson and M. L. Stockton. Service-oriented architecture: Programming model and product architecture. IBM Systems Journal, 44 (4):753–780, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/444/ferguson.html.

Fahey:2001:LBO

[FSSS01] L. Fahey, R. Srivastava, J. S. Sharon, and D. E. Smith. Linking e-business and operating processes: The role of knowledge management. IBM Systems Journal, 40(4):889– 907, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/404/fahey.html;   http:
//www.research.ibm.com/journal/
sj/404/fahey.pdf.

Forrest:1992:IIP

[FT92a] J. J. H. Forrest and J. A. Tomlin. Implementing interior point linear programming methods in the Optimization Subroutine Library. IBM Systems Journal, 31(1):26–38, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5458.

Forrest:1992:ISM

[FT92b] J. J. H. Forrest and J. A. Tomlin. Implementing the simplex method for the Optimization Subroutine Library. IBM Systems Journal, 31(1):11–25, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5457.

Fleischmann:1987:SII

[FtCE87] Albert Fleischmann, Seung tae Chin, and Wolfgang Effelsberg. Specification and implementation of an ISO session layer. IBM Systems Journal, 26(3):255–275, 1987. CODEN IBMSA7. ISSN 0018-8670.

Flurry:2001:IAF

[FV01] G. Flurry and W. Vicknair.

The IBM Application Framework for e-business. IBM Systems Journal, 40(1):8–24, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/401/flurry.html.

Friedman:1975:ESC

[FW75] Herman P. Friedman and Gerald Waldbaum. Evaluating system changes under uncontrolled workloads: a case study. IBM Systems Journal, 14(4):340– 352, 1975. CODEN IBMSA7. ISSN 0018-8670.

Gardner:1981:SAO

[Gar81] P. C. Gardner, Jr. A system for the automated office environment. IBM Systems Journal, 20(3):321–345, 1981. CODEN IBMSA7. ISSN 0018-8670.

Garrigues:1983:ANM

[Gar83] Richard D. Garrigues. Application of network management at a large computing service. IBM Systems Journal, 23(1-2):143– 164, 1983. CODEN IBMSA7. ISSN 0018-8670.

Gay:1966:TSD

[Gay66] T. W. Gay, Jr. On teleprocessing system design, part V: a technique for estimating channel interference. IBM Systems Journal, 5(3):171–174, 1966. CODEN IBMSA7. ISSN 0018-8670.

Gayek:1995:DLS

[Gay95]  P. W. Gayek.  Data link

switching: Present and future. IBM Systems Journal, 34(3):409–429, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-3.html#six.

Gould:1984:SFO

[GB84] John D. Gould and Stephen J. Boies. Speech filing — an office system for principals. IBM Systems Journal, 23(1):65–81, 1984. CODEN IBMSA7. ISSN 0018-8670.

Gerasimov:2000:TTU

[GB00a] V. Gerasimov and W. Bender. Things that talk: Using sound for device-to-device and device-to-human communication. IBM Systems Journal, 39(3/4):530–546, ???? 2000.

CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/gerasimov.html;
 http://www.research.ibm.
com/journal/sj/393/part1/
gerasimov.pdf;   http://
www.research.ibm.com/journal/
sj/393/part1/gerasimov.txt.

Gruhl:2000:NSN

[GB00b] D. Gruhl and W. Bender. A new structure for news editing. IBM Systems Journal, 39 (3/4):569–588, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/gruhl.html;
 http://www.research.ibm.
com/journal/sj/393/part1/
gruhl.pdf;   http://www.
research.ibm.com/journal/
sj/393/part1/gruhl.txt.

Gold:2007:HRB

[GB07] J. D. Gold and M. J. Ball. The Health Record Banking imperative: a conceptual model. IBM Systems Journal, 46(1): 43–55, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/461/gold.html.

Gu:2000:EHP

[GBCW00] W. Gu, N. A. Burns, M. T.

Collins, and W. Y. P. Wong. The evolution of a highperforming Java virtual machine. IBM Systems Journal, 39(1):135–150, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/391/gu.html.

Grover:1983:WSF

[GBK83] G. A. Grover and K. Bharath-Kumar. Windows in the sky — flow control in SNA networks with satellite links. IBM Systems Journal, 23(4):451–463, 1983. CODEN IBMSA7. ISSN 0018-8670.

Greif:2006:P

[GBR06] Irene Greif, Alex Birman, and John J. Ritsko. Preface. IBM Systems Journal, 45(4): 659–661, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/454/preface.html.

Griffee:1988:ITF

[GC88] A. W. Griffee and C. A. Casey. An introduction to typographic fonts and digital font resources. IBM Systems Journal, 27(2): 206–218, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Ganek:2003:DAC

[GC03] A. G. Ganek and T. A. Corbi. The dawning of the autonomic computing era. IBM Systems Journal, 42(1):5–18, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/421/verma.pdf.

Green:1987:PAP

[GCF^{+}87] P. E. Green, Jr., R. J. Chappuis, J. D. Fisher, P. S. Frosch, and C. E. Wood. A perspective on Advanced Peer-to-Peer Networking. IBM Systems Journal, 26(4):414–428, December 1987. CODEN IBMSA7. ISSN 0018-8670.

Gruhl:2004:HBW

[GCG^{+}04] D. Gruhl, L. Chavet, D. Gibson, J. Meyer, P. Pattanayak, A. Tomkins, and J. Zien. How to build a WebFountain: An architecture for very largescale text analytics. IBM Systems Journal, 43(1):64–77, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/431/gruhl.pdf.

Guinan:1997:EUA

[GCS97] P. J. Guinan, J. G. Cooprider, and S. Sawyer. The effective use of automated application development tools. IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/361/guinan.html.

Guski:2001:SZC

[GDD^{+}01] R. Guski, J. C. Dayka, L. N.

Distel, W. B. Farrell, K. A. Gdaniec, M. J. Kelly, M. A. Nelson, L. H. Overby, and L. G. Robinson. Security on z/OS: Comprehensive, current, and flexible. IBM Systems Journal, 40(3):696–720, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/403/guski.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/403/guski.pdf.

Gelb:1989:SS

[Gel89] J. P. Gelb. System-managed storage. IBM Systems Journal, 28(1):77–103, 1989. CODEN IBMSA7. ISSN 0018-8670.

Gerson:1962:NMA

[Ger62] G. Gerson. Note—on modifying the 1620 ADD table. IBM Systems Journal, 1:82–83, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1985:Pa

[Ger85a] Gary Gershon. Preface. IBM Systems Journal, 24(2):74–75, 1985. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1985:Pb

[Ger85b] Gary Gershon. Preface. IBM Systems Journal, 24(3/4):180– 181, 1985. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1986:Pa

[Ger86a] Gary Gershon. Preface. IBM Systems Journal, 25(1):2–3, 1986. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1986:Pb

[Ger86b] Gary Gershon. Preface. IBM Systems Journal, 25(2):132– 133, 1986. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1986:Pc

[Ger86c] Gary Gershon. Preface. IBM Systems Journal, 25(3):254– 255, 1986. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1987:Pa

[Ger87a] Gary Gershon. Preface. IBM Systems Journal, 26(1):2–3, 1987. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1987:Pb

[Ger87b] Gary Gershon. Preface. IBM Systems Journal, 26(2):156, 1987. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1987:Pc

[Ger87c] Gary Gershon. Preface. IBM Systems Journal, 26(3):234, 1987. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1987:Pd

[Ger87d] Gary Gershon. Preface. IBM Systems Journal, 26(4):324– 325, 1987. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1988:Pa

[Ger88a] Gary Gershon. Preface. IBM Systems Journal, 27(1):2, 1988. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1988:Pb

[Ger88b] Gary Gershon. Preface. IBM Systems Journal, 27(2):88–89, 1988. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1988:Pc

[Ger88c] Gary Gershon. Preface. IBM Systems Journal, 27(3):248– 249, 1988. CODEN IBMSA7. ISSN 0018-8670.

Gershon:1988:Pd

[Ger88d] Gary Gershon. Preface. IBM Systems Journal, 27(4):392, 1988. CODEN IBMSA7. ISSN 0018-8670.

Gershenfeld:1996:PM

[Ger96a] N. Gershenfeld. Physics and media. IBM Systems Journal, 35(3&4):575–576, 1996. CO-DEN IBMSA7. ISSN 0018-8670.

Gershenfeld:1996:SET

[Ger96b] N. Gershenfeld. Signal entropy and the thermodynamics of computation. IBM Systems Journal, 35(3&4):577– 586, 1996. CODEN IBMSA7. ISSN 0018-8670.

Gershenfeld:2000:EUL

[Ger00] N. Gershenfeld. Everything, the universe, and life. IBM Systems Journal, 39(3/4):932– 934, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part3/gershenfeld.
html;   http://www.research.
ibm.com/journal/sj/393/part3/
gershenfeld.pdf;   http:/
/www.research.ibm.com/journal/
sj/393/part3/gershenfeld.
txt.

Geihs:1992:STM

[GFG^{+}92] K. Geihs, P. François, D. Griffin, C. Kaas-Petersen, and A. Mann. Service and traffic management for IBCN. IBM Systems Journal, 31(4):711– 727, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5494.

Glider:2003:SAS

[GFS03] J. S. Glider, C. F. Fuente, and W. J. Scales. The software architecture of a SAN storage control system. IBM Systems Journal, 42(2):232–249, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/glider.pdf.

Green:1987:PDC

[GG87] Paul E. Green, Jr. and Dominique N. Godard. Prospects and design choices for integrated private networks. IBM Systems Journal, 26(1):37– 54, January 1987. CODEN IBMSA7. ISSN 0018-8670.

Gottschalk:2002:IWS

[GGKS02] K. Gottschalk, S. Graham, H. Kreger, and J. Snell. Introduction to Web services architecture. IBM Systems Journal, 41(2):168–177, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/412/gottschalk.html;
http://www.research.ibm.
com/journal/sj/412/gottschalk.
pdf;   http://www.research.

                                        ibm.com/journal/sj/412/gottschalk.
                                        txt.

Goodwin:2002:ILA

[GGW02] R. Goodwin, S. F. Goh, and F. Y. Wu. Instance-level access control for business-to-business electronic commerce. IBM Systems Journal, 41(2): 303–317, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/412/goodwin.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/412/goodwin.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/412/goodwin.txt.

Gruhn:1979:RPC

[GH79] A. M. Gruhn and A. C. Hohl. A research perspective on computer-assisted office work. IBM Systems Journal, 18(3):432–456, 1979. CODEN IBMSA7. ISSN 0018-8670.

Gdaniec:1991:VDS

[GH91] J. M. Gdaniec and J. P. Hennessy. VM data spaces and ESA/XC facilities. IBM Systems Journal, 30(1):14–33, 1991. CODEN IBMSA7. ISSN 0018-8670.

Grochowski:2003:TIM

[GH03] E. Grochowski and R. D. Halem. Technological impact of magnetic hard disk drives on storage systems. IBM Systems Journal, 42(2):338–346, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/grochowski.pdf.

Ghanem:1975:CCO

[Gha75] Sawsan B. Ghanem. Computing center optimization by a pricing-priority policy. IBM Systems Journal, 14(3):272– 291, 1975. CODEN IBMSA7. ISSN 0018-8670.

Gray:1983:APC

[GHH^{+}83] J. P. Gray, P. J. Hansen, P. Homan, M. A. Lerner, and M. Pozefsky. Advanced program-to-program communication in SNA. IBM Systems Journal, 23(4):298–318, 1983. CODEN IBMSA7. ISSN 0018-8670.

Ghiotti:1985:SIU

[Ghi85] M. M. Ghiotti. Single-system interface using the IBM 3270-PC. IBM Systems Journal, 24 (3-4):236–247, 1985. CODEN IBMSA7. ISSN 0018-8670.

Gruber:2005:EPA

[GHM^{+}05] O. Gruber, B. J. Hargrave, J. McAffer, P. Rapicault, and T. Watson. The Eclipse 3.0 platform: Adopting OSGi technology. IBM Systems Journal, 44(2):289–299, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/gruber.html;   http:
//www.research.ibm.com/journal/
sj/442/gruber.pdf;   http:
//www.research.ibm.com/journal/
sj/442/gruber.txt.

Goldstein:1984:DCP

[GHMWB84] Barry C. Goldstein, Andrew R. Heller, Franklin H. Moss, and Irving Wladawsky-Berger. Directions in cooperative processing between workstations and hosts. IBM Systems Journal, 23(3):236–244, 1984. CODEN IBMSA7. ISSN 0018-8670.

Ginzberg:1965:NTR

[Gin65] M. G. Ginzberg. Notes on testing real-time system programs. IBM Systems Journal, 4(1):58– 72, 1965. CODEN IBMSA7. ISSN 0018-8670.

Goldstein:1987:DCI

[GJ87] Barry C. Goldstein and Jeffrey M. Jaffe. Data communications: the implications of communication systems for protocol design. IBM Systems Journal, 26(1):122–137, January 1987. CODEN IBMSA7. ISSN 0018-8670.

Greene:2003:IDF

[GJMT03] S. L. Greene, L. Jones, P. Matchen, and J. C. Thomas. Iterative development in the field. IBM Systems Journal, 42 (4):594–612, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/424/greene.pdf.

Gladney:1975:CIA

[GJS75] Henry M. Gladney, Dion L. Johnson, and Richard L. Stone. Computer installation accounting. IBM Systems Journal, 14 (4):314–339, 1975. CODEN IBMSA7. ISSN 0018-8670.

Ganek:2007:OIS

[GK07] A. Ganek and K. Kloeckner. An overview of IBM Service Management. IBM Systems Journal, 46(3):375–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/ganek.html.

Grey:2003:AAQ

[GKB^{+}03] W. Grey, K. Katircloglu, S. Bagchi, D. Shi, G. Gallego, D. Seybold, and S. Stefanis. An analytical approach for quantifying the value of e-business initiatives. IBM Systems Journal, 42(3):484–497, ???? 2003. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/grey.pdf.

Griffin:1963:PIS

[GKT63] J. S. Griffin, Jr., J. H. King, Jr., and C. J. Tunis. A pattern identification system using linear decision functions. IBM Systems Journal, 2: 248–267, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Gecsei:1974:MES

[GL74] J. Gecsei and J. A. Lukes. A model for the evaluation of storage hierarchies. IBM Systems Journal, 13(2):163–178, 1974. CODEN IBMSA7. ISSN 0018-

Goodrich:1990:PDI

[GL90] C. Goodrich and M. B. Loughlin. Porting DPPX from the IBM 8100 to the IBM ES/9370: Migration. IBM Systems Journal, 29(1):106–123, 1990. CO-DEN IBMSA7. ISSN 0018-8670.

Gropp:1995:EIS

[GL95] W. D. Gropp and E. Lusk. Experiences with the IBM SP1. IBM Systems Journal, 34(2):249–262, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-2.html#seven.

Gladney:1978:ACC

[Gla78] Henry M. Gladney. Administrative control of computing service. IBM Systems Journal, 17(2):151–178, 1978. CODEN IBMSA7. ISSN 0018-8670.

Gladney:1993:SSI

[Gla93a] H. M. Gladney. A storage subsystem for image and records management. IBM Systems Journal, 32(3):512–540, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5523.

Glazer:1993:MVI

[Gla93b] R. Glazer. Measuring the value of information: The information-intensive organization. IBM Systems Journal, 32(1):99–110, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5505.

Glushko:2008:DSS

[Glu08] R. J. Glushko. Designing a service science discipline with discipline. IBM Systems Journal, 47(1):15–??, January/March 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/471/glushko.html.

Gray:1979:SMN

[GM79] James P. Gray and Tony B.

McNeill. SNA multiple-system networking. IBM Systems Journal, 18(2):263–297, 1979. CODEN IBMSA7. ISSN 0018-8670.

Goldberg:1991:BPC

[GM91] S. H. Goldberg and J. A. Mouton, Jr. A base for portable communications software. IBM Systems Journal, 30(3):259– 279, 1991. CODEN IBMSA7. ISSN 0018-8670.

Gorlenko:2003:NWA

[GM03] L. Gorlenko and R. Merrick.

No wires attached: Usability challenges in the connected mobile world. IBM Systems Journal, 42(4):639–651, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/424/gorlenko.pdf.

Geyer:2006:AEA

[GMM^{+}06] W. Geyer, M. J. Muller, M. T.

Moore, E. Wilcox, L.-T. Cheng, B. Brownholtz, C. Hill, and D. R. Millen. Activity Explorer: Activity-centric collaboration from research to product. IBM Systems Journal, 45 (4):713–738, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/454/geyer.html.

Guniguntala:2008:RCU

[GMTW08] D. Guniguntala, P. E. McKenney, J. Triplett, and J. Walpole. The read-copy-update mechanism for supporting real-time applications on shared-memory multiprocessor systems with Linux. IBM Systems Journal, 47(2):221–??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/472/guniguntala.html.

Gayek:2004:WCS

[GNP^{+}04] P. Gayek, R. Nesbitt, H. Pearthree, A. Shaikh, and B. Snitzer. A Web content serving utility. IBM Systems Journal, 43 (1):43–63, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/431/gayek.pdf.

Goldberg:1986:SEE

[Gol86] R. Goldberg. Software engineering: an emerging discipline. IBM Systems Journal, 25 (3-4):334–353, 1986. CODEN IBMSA7. ISSN 0018-8670.

Goldberg:1999:TPS

[Gol99] R. Goldberg. Turning points on software development. IBM Systems Journal, 38(2–3):225– 230, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/goldberg.pdf.

Gomory:1962:NTP

[Gom62] R. E. Gomory. Note—the trim problem. IBM Systems Journal, 1:77–81, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Gordon:1962:GPS

[Gor62] G. Gordon. A general purpose systems simulator. IBM Systems Journal, 1:18–32, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Gorenstein:1965:FAO

[Gor65] S. Gorenstein. Fabrication and assembly operations, part VI: Parameter values for sequencing control. IBM Systems Journal, 4(3):241–249, 1965. CO-DEN IBMSA7. ISSN 0018-8670.

Gorenstein:1971:PEL

[Gor71] S. Gorenstein. Programming for economic lot-sizes with precedences between items. IBM Systems Journal, 10(3): 232–243, 1971. CODEN IBMSA7. ISSN 0018-8670.

Gordon:1973:FMS

[Gor73] R. J. Gordon. Financial modeling on small systems. IBM Systems Journal, 12(2):161–171, 1973. CODEN IBMSA7. ISSN 0018-8670.

Gordon:1980:MAC

[Gor80] R. Douglas Gordon. The Modular Application Customizing System. IBM Systems Journal, 19(4):521–541, 1980. CODEN IBMSA7. ISSN 0018-8670.

Grey:2005:RMR

[GOS05] W. Grey, T. Olavson, and D. Shi. The role of emarketplaces in relationshipbased supply chains: a survey. IBM Systems Journal, 44 (1):109–123, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/441/grey.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/441/grey.pdf;   http:/
                                        /www.research.ibm.com/journal/
                                        sj/441/grey.txt.

Gottschalk:1986:SUP

[Got86] K. D. Gottschalk. The System Usability Process for Network Management Products. IBM Systems Journal, 25(1):83–91, 1986. CODEN IBMSA7. ISSN 0018-8670.

Gottschalk:1992:NVR

[Got92] K. D. Gottschalk. NetView Version 2 Release 3 Graphic Monitor Facility: Network management graphics support for the 1990s. IBM Systems Journal, 31(2):223–251, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5471.

Gottschalk:1998:TOI

[Got98] K. D. Gottschalk. Technical overview of IBM’s Java initiatives. IBM Systems Journal, 37 (3):308–322, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/373/gottschalk.html.

Gould:1969:GIG

[Gou69] R. L. Gould. GPSS/360—an improved general purpose simulator. IBM Systems Journal, 8(1):16–27, 1969. CODEN IBMSA7. ISSN 0018-8670.

Gongla:2001:ECP

[GR01] P. Gongla and C. R. Rizzuto. Evolving communities of practice: IBM Global Services experience. IBM Systems Journal, 40(4):842–862, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/404/gongla.html;   http:
//www.research.ibm.com/journal/
sj/404/gongla.pdf.

Greif:2006:I

[GR06] Irene Greif and Michael Rhodin.

Introduction. IBM Systems Journal, 45(4):??, ???? 2006. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/454/intro.html.

Gravina:1978:NWB

[Gra78] Chris M. Gravina. National Westminster Bank mass storage archiving. IBM Systems Journal, 17(4):344–358, 1978. CODEN IBMSA7. ISSN 0018-8670.

Green:1979:INA

[Gre79] Paul E. Green. An introduction to network architectures and protocols. IBM Systems Journal, 18(2):202–222, 1979. CODEN IBMSA7. ISSN 0018-8670.

Griffith:1963:IAP

[Gri63] J. E. Griffith. An intrinsically addressed processing system. IBM Systems Journal, 2: 182–199, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Griss:1993:SRL

[Gri93] M. L. Griss. Software reuse: From library to factory. IBM Systems Journal, 32(4):548– 566, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5524.

Grossman:1985:CSD

[Gro85] C. P. Grossman. Cache-DASD storage design for improving system performance. IBM Systems Journal, 24(3-4):316–334, 1985. CODEN IBMSA7. ISSN 0018-8670.

Grossman:1989:EDS

[Gro89] C. P. Grossman. Evolution of the DASD storage control. IBM Systems Journal, 28(2):196–226, 1989. CODEN IBMSA7. ISSN 0018-8670.

Grossman:1992:RDS

[Gro92] C. P. Grossman. Role of the DASD storage control in an Enterprise Systems Connection environment. IBM Systems Journal, 31(1):123–146, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5466.

Gibson:1986:ESP

[GRW86] D. H. Gibson, D. W. Rain, and H. F. Walsh. Engineering and scientific processing on the IBM 3090. IBM Systems Journal, 25(1):36–50, 1986. CODEN IBMSA7. ISSN 0018-8670.

Ganek:1999:TPS

[GS99] A. G. Ganek and E. H. Sussenguth. Turning points in systems architecture. IBM Systems Journal, 38(2–3):335–339, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/382/ganek.pdf.

Gotz:2004:DIU

[GS04] T. Götz and O. Suhre. Design and implementation of the UIMA Common Analysis System. IBM Systems Journal, 43(3):476–489, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/gotz.html;   http:
//www.research.ibm.com/journal/
sj/433/gotz.pdf;   http:/
/www.research.ibm.com/journal/
sj/433/gotz.txt.

Gongla:1989:SKS

[GSBH^{+}89] P. Gongla, G. Sakamoto, A. Back-Hock, P. Goldweic, L. Ramos, R. C. Sprowls, and C.-K. Kim. S*P*A*R*K: a knowledge-based system for identifying competitive uses of information technology. IBM Systems Journal, 28(4):628– 645, 1989. CODEN IBMSA7. ISSN 0018-8670.

Goble:2001:TAM

[GSN^{+}01] C. A. Goble, R. Stevens, G. Ng, S. Bechhofer, N. W. Paton, P. G. Baker, M. Peim, and A. Brass. Transparent access to multiple bioinformatics information sources. IBM Systems Journal, 40(2):532–551, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/goble.html;   http:
//www.research.ibm.com/journal/
sj/402/goble.pdf;   http:
//www.research.ibm.com/journal/
sj/402/goble.txt.

Gale:1990:IDP

[GTB90] J. L. Gale, J. R. Tirso, and C. A. Burchfield. Implementing the Defect Prevention Process in the MVS interactive programming organization. IBM Systems Journal, 29(1):33–43, 1990. CODEN IBMSA7. ISSN 0018-8670.

Gagliano:1968:IGD

[GTC68] F. W. Gagliano, H. W. Thombs, and R. E. Cornish. Interactive graphics in data processing: a conversational display capability. IBM Systems Journal, 7(3/4):281–291, 1968. CODEN IBMSA7. ISSN 0018-8670.

Gazdag:1989:CCS

[GW89] J. Gazdag and H.-H. Wang.

Concurrent computing by sequential staging of tasks. IBM Systems Journal, 28(4):646– 660, 1989. CODEN IBMSA7. ISSN 0018-8670.

Gladney:1975:ACM

[GWM75] Henry M. Gladney, Eldon L.

Worley, and James J. Myers. An access control mechanism for computing resources. IBM Systems Journal, 14(3):212– 228, 1975. CODEN IBMSA7. ISSN 0018-8670.

Galli:1967:ADV

[GY67] E. J. Galli and H. Yamada.

An automatic dictionary and the verification of machinereadable text. IBM Systems Journal, 6(3):192–207, 1967. CODEN IBMSA7. ISSN 0018-8670.

George:1982:SFC

[GY82] Frederick D. George and Gerald E. Young. SNA flow control: Architecture and implementation. IBM Systems Journal, 21(2):179–210, 1982. CODEN IBMSA7. ISSN 0018-8670.

Haase:1996:FRK

[Haa96] K. Haase. FramerD: Representing knowledge in the large. IBM Systems Journal, 35(3&4):381– 397, 1996. CODEN IBMSA7. ISSN 0018-8670.

Haase:2000:IB

[Haa00a] K. Haase. Interlingual BRICO. IBM Systems Journal, 39(3/4): 589–596, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/haase.html;
 http://www.research.ibm.
com/journal/sj/393/part1/
haase.pdf;   http://www.
research.ibm.com/journal/
sj/393/part1/haase.txt.

Haase:2000:WML

[Haa00b] K. Haase. Why the Media Lab works — A personal view. IBM Systems Journal, 39(3/4): 419–431, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/haase.html;   http:
//www.research.ibm.com/journal/
sj/393/haase.pdf;   http:
//www.research.ibm.com/journal/
sj/393/haase.txt.

Hovorka:1992:CPN

[HAB^{+}92] R. Hovorka, S. Andreassen, J. J. Benn, K. G. Olesen, and E. R. Carson. Causal probabilistic network modeling —an illustration of its role in the management of chronic diseases. IBM Systems Journal, 31(4):635–648, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5490.

Haeckel:2003:SMF

[Hae03] S. H. Haeckel. A survey of manufacturing flexibility: Implications for e-business flexibility. IBM Systems Journal, 42(3):405–413, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/haeckel.pdf.

Haines:1965:SCF

[Hai65] L. H. Haines. Serial compilation and the 1401 FORTRAN compiler. IBM Systems Journal, 4(1):73–80, 1965. CODEN IBMSA7. ISSN 0018-8670.

Haims:1989:B

[Hai89] Murray J. Haims. Books.

IBM Systems Journal, 28(3): 494, 1989. CODEN IBMSA7. ISSN 0018-8670.

Hakeda:1990:IOC

[Hak90] Y. Hakeda. The Image Object Content Architecture. IBM Systems Journal, 29(3):333– 342, 1990. CODEN IBMSA7. ISSN 0018-8670.

Halstead:1977:FLC

[Hal77] M. H. Halstead. Forum: On lines of code and readability. IBM Systems Journal, 16 (4):421–422, 1977. CODEN IBMSA7. ISSN 0018-8670. See [WF77b, WF77a].

Hammond:1982:MCI

[Ham82] Lynn W. Hammond. Management considerations for an Information Center. IBM Systems Journal, 21(2):131–161, 1982. CODEN IBMSA7. ISSN 0018-8670.

Hanford:1970:AGT

[Han70] K. V. Hanford. Automatic generation of test cases. IBM Systems Journal, 9(4):242–257, 1970. CODEN IBMSA7. ISSN 0018-8670.

Handelman:1980:HCG

[Han80] S. W. Handelman. A highresolution computer graphics system. IBM Systems Journal, 19(3):356–366, 1980. CODEN IBMSA7. ISSN 0018-8670.

Harrison:1980:DPC

[Har80] Edward S. Harrison. Distributed processing communications software support for operation within an SNA network. IBM Systems Journal, 19(2):192–207, 1980. CODEN IBMSA7. ISSN 0018-8670.

Hartung:2003:ITE

[Har03] M. Hartung. IBM TotalStorage Enterprise Storage Server: a designer’s view. IBM Systems Journal, 42(2):383–396, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/hartung.pdf.

Hauth:1968:TTM

[Hau68] C. Hauth. Turnaround time for messages of differing priorities. IBM Systems Journal, 7(2):103–122, 1968. CODEN IBMSA7. ISSN 0018-8670.

Havender:1968:ADM

[Hav68] J. W. Havender. Avoiding deadlock in multitasking systems. IBM Systems Journal, 7(2):74–84, 1968. CODEN IBMSA7. ISSN 0018-8670.

Hazony:1996:SGP

[Haz96]  Y. Hazony.  System gen-

erator for producing manufacturing applications. IBM Systems Journal, 35(1):69–93, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj35-1.html#five.

Hoffnagle:1985:ASD

[HB85] G. F. Hoffnagle and W. E.

Beregi. Automating the software development process. IBM Systems Journal, 24(2):102– 120, 1985. CODEN IBMSA7. ISSN 0018-8670.

Harper:2005:STI

[HB05] S. Harper and S. Bechhofer.

Semantic triage for increased web accessibility. IBM Systems Journal, 44(3):637–648, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/443/harper.html.

Hild:2001:AHP

[HBBWS01] S. G. Hild, C. Binding, D. Bourges-Waldegg, and C. Steenkeste. Application hosting for pervasive computing. IBM Systems Journal, 40 (1):193–219, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/401/hild.html.

Heineman:1994:ETS

[HBC^{+}94] G. T. Heineman, J. E. Botsford, G. Caldiera, G. E. Kaiser, M. I. Kellner, and N. H. Madhavji. Emerging technologies that support a software process life cycle. IBM Systems Journal, 33(3):501–529, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-3.html#eight. G321-

5553.

Hanson:2005:IWA

[HBC^{+}05] V. L. Hanson, J. P. Brezin, S. Crayne, S. Keates, R. Kjeldsen, J. T. Richards, C. Swart, and S. Trewin. Improving Web accessibility through an enhanced open-source browser. IBM Systems Journal, 44(3): 573–588, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/443/hanson.html.

Horn:2004:LSB

[HBM^{+}04] B. Horn, H. Balakrishnan, B. T. Maniampadavathu, J. Warnes, and D. A. Elko. A Logger System based on Web services. IBM Systems Journal, 43(4): 723–733, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/434/horn.html;   http:

//www.research.ibm.com/journal/
sj/434/horn.pdf;   http:/
/www.research.ibm.com/journal/
sj/434/horn.txt.

Hodil:1986:KSC

[HBR86] E. D. Hodil, C. W. Butler, and G. L. Richardson. Knowledgebased systems in the commercial environment. IBM Systems Journal, 25(2):147–158, 1986. CODEN IBMSA7. ISSN 0018-8670.

Hinkelman:2006:EPU

[HBZ06] S. Hinkelman, D. Buddenbaum, and L.-J. Zhang. Emerging patterns in the use of XML for information modeling in vertical industries. IBM Systems Journal, 45(2):373–378, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/452/hinkelman.html.

Harding:1990:OSH

[HCGT90] W. B. Harding, C. M. Clark, C. L. Gallo, and H. Tang. Object storage hierarchy management. IBM Systems Journal, 29(3):384–397, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hehenberger:2007:ISI

[HCR^{+}07] M. Hehenberger, A. Chatterjee, U. Reddy, J. Hernandez, and J. Sprengel. IT solutions for imaging biomarkers in biopharmaceutical research and development. IBM Systems Journal, 46(1):183–198, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/461/hehenberger.html.

Heyne:1977:DTU

[HD77] George F. Heyne and Chester J. Daniel. Design techniques for a user controlled DB/DC system. IBM Systems Journal, 16(4):344–362, 1977. CODEN IBMSA7. ISSN 0018-8670.

Halpern:1988:ICD

[HD88] P. Halpern and J. V. Dave. Interactive computations and display of characteristics of the radiation scattered by a sphere: A demonstration for PS/2 Model 80. IBM Systems Journal, 27(4):561–569, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Haynes:1988:CSP

[HDH88] W. K. Haynes, M. E. Dewell, and P. J. Herman. The Cross System Product application generator: An evolution. IBM Systems Journal, 27(3):384–390, 1988. CODEN IBMSA7. ISSN 0018-8670.

Haas:2003:ASD

[HDS03] R. Haas, P. Droz, and B. Stiller. Autonomic service deployment in networks. IBM Systems Journal, 42(1):150–164, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/421/haas.pdf.

Hedberg:1963:DIP

[Hed63] R. Hedberg. Design of an integrated programming and operating system, part III: The expanded function of the loader. IBM Systems Journal, 2:298–310, September/ December 1963. CODEN IBMSA7. ISSN 0018-8670.

Hedger:2003:B

[Hed03] Richard J. Hedger. Books.

IBM Systems Journal, 42(2): 397–399, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/422/books.pdf.

Heising:1963:NRA

[Hei63] W. P. Heising. Note on random addressing techniques. IBM Systems Journal, 2:112–116, June 1963. CODEN IBMSA7. ISSN 0018-8670.

Hein:1985:ISM

[Hei85] K. P. Hein. Information system model and architecture generator. IBM Systems Journal, 24 (3-4):213–235, 1985. CODEN IBMSA7. ISSN 0018-8670.

Hein:1990:DNP

[Hei90] K. P. Hein. DevelopMate: a new paradigm for information system enabling. IBM Systems Journal, 29(2):250–264, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hellerman:1969:SPT

[Hel69] H. Hellerman. Some principles of time-sharing scheduler strategies. IBM Systems Journal, 8(2):94–117, 1969. CO-DEN IBMSA7. ISSN 0018-8670.

Helander:1981:ISU

[Hel81] G. A. Helander. Improving system usability for business professionals. IBM Systems Journal, 20(3):294–305, 1981. CO-DEN IBMSA7. ISSN 0018-8670.

Helms:1990:IIT

[Hel90] R. M. Helms. Introduction to image technology. IBM Systems Journal, 29(3):313–332, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hembry:1990:KSA

[Hem90] D. M. Hembry. Knowledgebased systems in the AD/Cycle environment. IBM Systems Journal, 29(2):274–286, 1990. CODEN IBMSA7. ISSN 0018-8670.

Henry:1969:HSD

[Hen69] W. R. Henry. Hierarchical structure for data management. IBM Systems Journal, 8(1):2– 15, 1969. CODEN IBMSA7. ISSN 0018-8670.

Henry:1986:ISA

[Hen86] G. G. Henry. IBM small-system architecture and design — past, present, and future. IBM Systems Journal, 25(3-4):321–333, 1986. CODEN IBMSA7. ISSN 0018-8670.

Henders:1998:EAA

[Hen98] R. A. Henders. An evolutionary approach to application development with object technology. IBM Systems Journal, 37 (2):181–188, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/372/henders.html.

Hess:2005:ATB

[Hes05] H. M. Hess. Aligning technology and business: Applying patterns for legacy transformation. IBM Systems Journal, 44(1):25–45, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/441/hess.html;   http:
//www.research.ibm.com/journal/
sj/441/hess.pdf;   http:/
/www.research.ibm.com/journal/
sj/441/hess.txt.

Haataia:2008:EAS

[HFB^{+}08] Bob Haataia, Scott Fredericksen, Eliot Burris, Lauren Bragg, Kent W. Eastley, and John Bird. Enhancing availability with service automation and a trusted support partner. IBM Systems Journal, 47 (4):667–680, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/haataia.pdf.

Hatfield:1971:PRV

[HG71] D. J. Hatfield and J. Gerald. Program restructuring for virtual memory. IBM Systems Journal, 10(3):168–192, March 1971. CODEN IBMSA7. ISSN 0018-8670.

Hagamen:1986:NRK

[HG86] W. D. Hagamen and M. Gardy. The numeric representation of knowledge and logic — two artificial intelligence applications in medical education. IBM Systems Journal, 25(2):207–235, 1986. CODEN IBMSA7. ISSN 0018-8670.

Hoffman:2005:DSA

[HGB05] D. Hoffman, E. Grivel, and L. Battle. Designing software architectures to facilitate accessible Web applications. IBM Systems Journal, 44(3):467–483, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/443/hoffman.html.

Head-Gordon:2001:CCS

[HGW01] T. Head-Gordon and J. C. Wooley. Computational challenges in structural and functional genomics. IBM Systems Journal, 40(2):265–296, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/402/headgordon.html;
                                        http://www.research.ibm.
                                        com/journal/sj/402/headgordon.
                                        pdf;   http://www.research.

ibm.com/journal/sj/402/headgordon.
txt.

Hahn:1966:KAE

[HH66] S. G. Hahn and E. V. Hankam.

Kernel analysis of elliptic partial differential equations. IBM Systems Journal, 5(4):248–270, 1966. CODEN IBMSA7. ISSN 0018-8670.

Hendricks:1979:EVM

[HH79] E. C. Hendricks and T. C. Hartmann. Evolution of a virtual machine subsystem. IBM Systems Journal, 18(1):111–142, 1979. CODEN IBMSA7. ISSN 0018-8670.

Heisig:2008:STA

[HH08] Stephen Heisig and Jonathan R. M. Hosking. Scoring and thresholding for availability. IBM Systems Journal, 47(4): 653–666, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/heisig.pdf.

Herness:2004:WAS

[HHM04] E. N. Herness, R. J. High, and J. R. McGee. WebSphere application server: a foundation for on demand computing. IBM Systems Journal, 43 (2):213–237, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/432/herness.html;   http:
//www.research.ibm.com/journal/
sj/432/herness.pdf;   http:

                                        //www.research.ibm.com/journal/
                                        sj/432/herness.txt.

Halsey:1979:PDN

[HHP79] John R. Halsey, Leslie E. Hardy, and Leo F. Powning. Public data networks: Their evolution, interfaces, and status. IBM Systems Journal, 18 (2):223–243, 1979. CODEN IBMSA7. ISSN 0018-8670.

Hamann:2001:SBA

[HHSS01] E.-M. Hamann, H. Henn, T. Schäck, and F. Seliger. Securing e-business applications using smart cards. IBM Systems Journal, 40(3):635– 647, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/403/hamann.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/403/hamann.pdf.

Herger:2003:P

[HIP^{+}03] Lorraine Herger, Kazuo Iwano, Pratap Pattnaik, John J. Ritsko, and Alfred G. Davis. Preface. IBM Systems Journal, 42(1):3–4, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/421/preface.pdf.

Hirsch:1981:PHF

[Hir81] R. S. Hirsch. Procedures of the Human Factors Center at San Jose. IBM Systems Journal, 20(2):123–171, 1981. CODEN IBMSA7. ISSN 0018-8670.

Hirao:1990:ERD

[Hir90] T. Hirao. Extension of the relational database semantic processing model. IBM Systems Journal, 29(4):539–550, 1990. CODEN IBMSA7. ISSN 0018-8670.

Haderle:1984:IDO

[HJ84] Donald J. Haderle and Robert D.

Jackson. IBM database 2 overview. IBM Systems Journal, 23(2):112–125, 1984. CO-DEN IBMSA7. ISSN 0018-8670.

Heidorn:1982:ETS

[HJM^{+}82] G. E. Heidorn, K. Jensen, L. A.

Miller, R. J. Byrd, and M. S. Chodorow. The EPISTLE textcritiquing system. IBM Systems Journal, 21(3):305–326, 1982. CODEN IBMSA7. ISSN 0018-8670.

Held:1965:CDD

[HK65] M. Held and R. M. Karp.

The construction of discrete dynamic programming algorithms. IBM Systems Journal, 4(2):136–147, 1965. CODEN IBMSA7. ISSN 0018-8670.

Huang:2003:TID

[HKLS03] H. Huang, C. Keser, J. Leland, and J. Shachat. Trust, the Internet, and the digital divide. IBM Systems Journal, 42(3):507–518, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/423/huang.pdf.

Harrison:1996:TND

[HKOS96] W. H. Harrison, H. Kilov, H. L. Ossher, and I. Simmonds. Technical note — from dynamic supertypes to subjects: a natural way to specify and develop systems. IBM Systems Journal, 35(2):244–256, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj35-2.html#nine.

Hanssmann:1971:MCC

[HKS71] F. Hanssmann, W. Kistler, and H. Schulz. Modeling for computing center planning. IBM Systems Journal, 10(4):305– 324, 1971. CODEN IBMSA7. ISSN 0018-8670.

High:2008:CMC

[HKS08] Rob High, Jr., Gopal Krishnan, and Matt Sanchez. Creating and maintaining coherency in loosely coupled systems. IBM Systems Journal, 47(3): 357–376, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/473/high.pdf.

Hassitt:1976:AES

[HL76] A. Hassitt and L. E. Lyon. An APL emulator on System/ 370. IBM Systems Journal, 15(4):358–378, 1976. CODEN IBMSA7. ISSN 0018-8670.

Hagamen:1972:EVI

[HLLW72] W. D. Hagamen, D. J. Linden, H. S. Long, and J. C. Weber.

Encoding verbal information as unique numbers. IBM Systems Journal, 11(4):278–315, 1972. CODEN IBMSA7. ISSN 0018-8670.

Hagamen:1975:PG

[HLM^{+}75] Wilbur D. Hagamen, D. J. Linden, K. F. Mai, S. M. Newell, and J. C. Weber. A program generator. IBM Systems Journal, 14(2):102–133, 1975. CO-DEN IBMSA7. ISSN 0018-8670.

Hess:1995:MNB

[HLM95] M. L. Hess, J. A. Lorrain, and G. R. McGee. Multiprotocol networking — a blueprint. IBM Systems Journal, 34(3):330–346, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-3.html#two.

Haas:2002:DIT

[HLR02] L. M. Haas, E. T. Lin, and M. A. Roth. Data integration through database federation. IBM Systems Journal, 41 (4):578–596, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/414/haas.pdf.

Hausler:1994:ACS

[HLT94] P. A. Hausler, R. C. Linger, and C. J. Trammell. Adopting Cleanroom software engineering with a phased approach. IBM Systems Journal, 33(1):89–109, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-1.html#six. G321-5534.

Hevner:1993:BMS

[HM93] A. R. Hevner and H. D. Mills. Box-structured methods for systems development with objects. IBM Systems Journal, 32(2):232–251, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5511.

Horn:2006:I

[HM06] Paul Horn and Steven Mills. Introduction. IBM Systems Journal, 45(2):??, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/intro.html.

Hellerstein:2002:DAP

[HMP02] J. L. Hellerstein, S. Ma, and C.-S. Perng. Discovering actionable patterns in event data. IBM Systems Journal, 41(3): 475–493, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/hellerstein.html.

Hebeker:1992:NSR

[HMS92] F. K. Hebeker, R. R. Maly, and S. U. Schoeffel. Numerical simulation of reactive flow on the IBM ES/3090 vector multiprocessor. IBM Systems Journal, 31(4):788–797, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5498.

Herzberg:1998:SCS

[HN98] A. Herzberg and D. Naor.

Surf’N’Sign: Client signatures on Web documents. IBM Systems Journal, 37(1):61–71, 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/371/herzberg.html.

Hondo:2002:SWS

[HNN02] M. Hondo, N. Nagaratnam, and A. Nadalin. Securing Web services. IBM Systems Journal, 41(2):228–241, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/412/hondo.html;   http:
//www.research.ibm.com/journal/
sj/412/hondo.pdf;   http:
//www.research.ibm.com/journal/
sj/412/hondo.txt.

Hsu:2007:WSE

[HO07] W. W. Hsu and S. Ong.

WORM storage is not enough. IBM Systems Journal, 46(2): 363–??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/462/hsu.html.

Hobgood:1972:EIS

[Hob72] W. S. Hobgood. Evaluation of an interactive-batch system network. IBM Systems Journal, 11(1):2–15, 1972. CODEN IBMSA7. ISSN 0018-8670.

Hobgood:1976:RNC

[Hob76] W. Sands Hobgood. The role of the Network Control Program in Systems Network Architecture. IBM Systems Journal, 15(1):39–52, 1976. CO-DEN IBMSA7. ISSN 0018-8670.

Hoernes:1989:RT

[Hoe89] G. E. Hoernes. REXX on TSO/E. IBM Systems Journal, 28(2):274–293, 1989. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1989:Pa

[Hof89a] Gene Hoffnagle. Preface. IBM Systems Journal, 28(2):194– 195, 1989. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1989:Pb

[Hof89b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 28(3):358–359, 1989. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1989:Pc

[Hof89c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 28(4):500–501, 1989. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1990:Pa

[Hof90a] Gene F. Hoffnagle. Preface. IBM Systems Journal, 29(1): 2–3, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1990:Pb

[Hof90b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 29(2):168–169, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1990:Pc

[Hof90c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 29(3):302–303, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1990:Pd

[Hof90d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 29(4):492–493, 1990. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1991:Pa

[Hof91a] Gene F. Hoffnagle. Preface.

IBM Systems Journal, 30(1): 2–3, 1991. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1991:Pb

[Hof91b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 30(2):128–129, 1991. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1991:Pc

[Hof91c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 30(3):248–249, 1991. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1991:Pd

[Hof91d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 30(4):414–415, 1991. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1992:Pa

[Hof92a] Gene F. Hoffnagle. Preface.

IBM Systems Journal, 31(1): 2–3, 1992. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1992:Pb

[Hof92b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 31(2):152–153, 1992. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1992:Pc

[Hof92c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 31(3):412–413, 1992. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1992:Pd

[Hof92d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 31(4):618–619, 1992. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1993:Pa

[Hof93a] Gene F. Hoffnagle. Preface. IBM Systems Journal, 32(1): 2–3, 1993. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1993:Pb

[Hof93b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 32(2):230–231, 1993. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1993:Pc

[Hof93c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 32(3):374–375, 1993. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1993:Pd

[Hof93d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 32(4):546–547, 1993. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1994:Pa

[Hof94a]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 33(1):2–3, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-1.html#one.

Hoffnagle:1994:Pb

[Hof94b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 33(2):226–227, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-2.html#one.

Hoffnagle:1994:Pc

[Hof94c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 33(3):380–381, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-3.html#one.

Hoffnagle:1994:Pd

[Hof94d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 33(4):536, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-4.html#one.

Hoffnagle:1995:Pa

[Hof95a]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 34(1):2–3, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.

research.ibm.com/journal/
sj34-1.html#one.

Hoffnagle:1995:Pb

[Hof95b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 34(2):144–145, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-1.html#one. Issue Order

No. G321-0120.

Hoffnagle:1995:Pc

[Hof95c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 34(3):328–329, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-3.html#one.

Hoffnagle:1995:Pd

[Hof95d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 34(4):552–553, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-4.html#one. Issue Order

No. G321-0122.

Hoffnagle:1996:Pa

[Hof96a]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 35(1):2–3, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-1.html#one. Issue Order

No. G321-0123.

Hoffnagle:1996:Pb

[Hof96b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 35(2):122–123, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-2.html#one.

Hoffnagle:1996:Pc

[Hof96c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 35(3&4):265, 1996. CODEN IBMSA7. ISSN 0018-8670.

Hoffnagle:1997:Pa

[Hof97a] Gene F. Hoffnagle. Preface.

IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/361/preface.html.

Hoffnagle:1997:Pb

[Hof97b]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 36(2):170–171, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/362/preface.html.

Hoffnagle:1997:Pc

[Hof97c]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 36(3):372–373, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/363/preface.html.

Hoffnagle:1997:Pd

[Hof97d]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 36(4):472–473, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/364/preface.html.

Hoffnagle:1998:Pa

[Hof98a]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 37(1):2–3, 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/371/preface.html.

Hoffnagle:1998:Pb

[Hof98b] Gene F. Hoffnagle. Preface. IBM Systems Journal, 37(2): 154–155, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/372/preface.html.

Hoffnagle:1998:Pc

[Hof98c] Gene F. Hoffnagle. Preface. IBM Systems Journal, 37(3): 306–307, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/373/preface.html.

Hoffnagle:1998:Pd

[Hof98d] Gene F. Hoffnagle. Preface. IBM Systems Journal, 37(4):

482–483, ???? 1998. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/374/preface.html.

Hoffnagle:1999:Pa

[Hof99a]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 38(1):2–??, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/381/preface.html.

Hoffnagle:1999:Pb

[Hof99b] Gene F. Hoffnagle. Preface.

IBM Systems Journal, 38(2– 3):2–134, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/preface.pdf.

Hoffnagle:1999:Pc

[Hof99c] Gene F. Hoffnagle. Preface.

IBM Systems Journal, 38(4): 502–503, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/preface.html.

Hoffnagle:2000:Pa

[Hof00a]  Gene F. Hoffnagle.  Pref-

ace. IBM Systems Journal, 39(1):2–3, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/391/preface.html.

Hoffnagle:2000:Pb

[Hof00b] Gene F. Hoffnagle. Preface. IBM Systems Journal, 39(2): 242–243, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/392/preface.html.

Hoffnagle:2000:Pc

[Hof00c] Gene F. Hoffnagle. Preface. IBM Systems Journal, 39(3/4): 415–??, ???? 2000. CODEN IBMSA7. ISSN 0018-8670.

Hopkins:1987:PIS

[Hop87] Martin E. Hopkins. A perspective on the 801/Reduced Instruction Set Computer. IBM Systems Journal, 26(1):107– 121, January 1987. CODEN IBMSA7. ISSN 0018-8670.

Harrison:2005:SAO

[HOST05] W. Harrison, H. Ossher, S. Sutton, and P. Tarr. Supporting aspect-oriented software development with the Concern Manipulation Environment. IBM Systems Journal, 44(2):309–318, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/442/harrison.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/harrison.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/harrison.txt.

Hough:1993:RDE

[Hou93] D. Hough. Rapid Delivery: An evolutionary approach for application development. IBM Systems Journal, 32(3):397– 419, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5518.

Howarth:1993:CEA

[HP93] R. J. Howarth and B. G.

Platte. The continuing evolution of Advanced Function Printing. IBM Systems Journal, 32(4):665–683, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5529.

Hayne:1995:EOG

[HP95] S. C. Hayne and M. Pendergast. Experiences with objectoriented group support software development. IBM Systems Journal, 34(1):96–119, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-1.html#seven.

Hippert:1975:RAS

[HPPS75] Richard O. Hippert, Jr., Lubos R. Palounek, John Provetero, and Ralph O. Skatrud. Reliability, availability, and serviceability design considerations for the Supermarket and Retail Store Systems. IBM Systems Journal, 14(1):81–95, 1975. CODEN IBMSA7. ISSN 0018-8670.

Holley:1979:VAM

[HPSS79] L. H. Holley, R. P. Parmelee, C. A. Salisbury, and D. N. Saul. VM/370 asymmetric multiprocessing. IBM Systems Journal, 18(1):47–70, 1979. CODEN IBMSA7. ISSN 0018-8670.

Halpern:1981:USA

[HR81] P. Halpern and J. W. Rettberg. User-definable software applied to a real-time ambient air quality monitoring system. IBM Systems Journal, 20(1):86–103, 1981. CODEN IBMSA7. ISSN 0018-8670.

Hong:2002:P

[HRB02] Se June Hong, John J. Ritsko, and Alex Birman. Preface. IBM Systems Journal, 41 (3):328–330, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/413/preface.html.

Halpern:1987:IPS

[HRL87] P. Halpern, S. M. Roberts, and L. Lopez. An incidence-matrix-driven panel system for the IBM PC. IBM Systems Journal, 26(2):201–214, 1987. CO-DEN IBMSA7. ISSN 0018-8670.

Harris:1999:ERT

[HRL99] T. Harris, J. W. Rothwell, and P. T. L. Lloyd. Experiences in reusing technical reference architectures. IBM Systems Journal, 38(1):98–117, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/381/harris.html.

Herscovitch:1965:GEG

[HS65] H. Herscovitch and T. H. Schneider. GPSS III—an expanded general purpose simulator. IBM Systems Journal, 4(3):174–183, 1965. CODEN IBMSA7. ISSN 0018-8670.

Housel:1983:SDS

[HS83] B. C. Housel and C. J.

Scopinich. SNA distribution services. IBM Systems Journal, 23(4):319–343, 1983. CODEN IBMSA7. ISSN 0018-8670.

Harrison:1987:SSF

[HS87] E. S. Harrison and E. J.

Schmitt. The structure of System/88, a fault-tolerant computer. IBM Systems Journal, 26(3):293–318, 1987. CODEN IBMSA7. ISSN 0018-8670.

Housel:2000:EES

[HS00] B. C. Housel and I. Shields.

Emulator Express: a system for optimizing emulator performance for wireless networks. IBM Systems Journal, 39(2): 384–402, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/housel.html.

Hailpern:2002:SDT

[HS02] B. Hailpern and P. Santhanam.

Software debugging, testing, and verification. IBM Systems Journal, 41(1):4–12, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/411/hailpern.html;   http:
//www.research.ibm.com/journal/
sj/411/hailpern.pdf.

Hsu:2003:CTP

[HS03] W. W. Hsu and A. J. Smith. Characteristics of I/O traffic in personal computer and server workloads. IBM Systems Journal, 42(2):347–372, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/hsu.pdf.

Hickey:2008:ISD

[HS08] J. Hickey and J. Siegel. Improving service delivery through integrated quality initiatives: a case study. IBM Systems Journal, 47(1):167–??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/471/hickey.html.

Haas:2001:DSI

[HSK^{+}01] L. M. Haas, P. M. Schwarz, P. Kodali, E. Kotlar, J. E. Rice, and W. C. Swope. DiscoveryLink: a system for integrated access to life sciences data sources. IBM Systems Journal, 40(2):489–511, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/haas.html;   http:
//www.research.ibm.com/journal/
sj/402/haas.pdf;   http:/
/www.research.ibm.com/journal/
sj/402/haas.txt.

Hirschheim:2006:MMM

[HST06] R. Hirschheim, A. Schwarz, and P. Todd. A marketing maturity model for IT: Building a customer-centric IT organization. IBM Systems Journal, 45(1):181–199, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/451/hirschheim.html.

Hart:2008:RTL

[HST08] D. Hart, J. Stultz, and T. Ts’o.

Real-time Linux in real time. IBM Systems Journal, 47(2): 207–??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/hart.html.

Hsu:2001:CPD

[HSY01] W. W. Hsu, A. J. Smith, and H. C. Young. Characteristics of production database workloads and the TPC benchmarks. IBM Systems Journal, 40(3):781–802, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/hsu.html;   http:/
/www.research.ibm.com/journal/
sj/403/hsu.pdf.

Hailpern:2006:MDD

[HT06] B. Hailpern and P. Tarr.

Model-driven development: The good, the bad, and the ugly. IBM Systems Journal, 45 (3):451–461, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/hailpern.html.

Huang:1998:CIA

[Hua98] K.-T. Huang. Capitalizing on intellectual assets. IBM Systems Journal, 37(4):570– 583, ???? 1998. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/374/huang.html.

Humphrey:1985:ILS

[Hum85] W. S. Humphrey. The IBM large-systems software development process: Objectives and direction. IBM Systems Journal, 24(2):76–78, 1985. CO-DEN IBMSA7. ISSN 0018-8670.

Henderson:1993:SAL

[HV93] J. C. Henderson and N. Venkatraman. Strategic alignment: Leveraging information technology for transforming organizations. IBM Systems Journal, 32(1):4–16, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5500.

Henderson:1999:SAL

[HV99] J. C. Henderson and H. Venkatraman. Strategic alignment: Leveraging information technology for transforming organizations. IBM Systems Journal, 38(2–3):472–484, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/382/henderson.pdf. Reprinted

from [HV93].

Horn:2005:I

[HW05] Paul Horn and Frances West.

Introduction. IBM Systems Journal, 44(3):??, ???? 2005. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/443/intro.html.

Hill:2006:BPW

[HYJK06] C. Hill, R. Yates, C. Jones, and S. L. Kogan. Beyond predictable workflows: Enhancing productivity in artful business processes. IBM Systems Journal, 45(4):663–682, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/454/hill.html.

Hammond:1984:TAM

[HZ84] Kenneth R. Hammond and Melvin R. Zimowski. TSO Attach: a multipurpose communication channel to IBM Database 2. IBM Systems Journal, 23(2):151–164, 1984. CODEN IBMSA7. ISSN 0018-8670.

Hazony:1992:CSE

[HZ92] Y. Hazony and L. Zeidner. Customized systems for engineering applications. IBM Systems Journal, 31(1):94–113, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5464.

IBGT:2001:BGV

[IBM01] IBM Blue Gene team. Blue Gene: a vision for protein science using a petaflop supercomputer. IBM Systems Journal, 40(2):310–327, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/allen.html;   http:
//www.research.ibm.com/journal/
sj/402/allen.pdf;   http:
//www.research.ibm.com/journal/
sj/402/allen.txt.

Inman:2001:HTD

[IFM^{+}01] J. T. Inman, H. R. Flores, G. D. May, J. W. Weller, and C. J. Bell. A highthroughput distributed DNA sequence analysis and database system. IBM Systems Journal, 40(2):464–486, ???? 2001. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/402/inman.html;   http:
//www.research.ibm.com/journal/
sj/402/inman.pdf;   http:
//www.research.ibm.com/journal/
sj/402/inman.txt.

Iyer:2004:FDA

[IG04] B. Iyer and R. Gottlieb. The Four-Domain Architecture: An approach to support enterprise architecture design. IBM Systems Journal, 43(3):587– 597, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/433/iyer.html;   http:
//www.research.ibm.com/journal/
sj/433/iyer.pdf;   http:/

/www.research.ibm.com/journal/
sj/433/iyer.txt.

Ives:1993:GBD

[IJM93] B. Ives, S. L. Jarvenpaa, and R. O. Mason. Global business drivers: Aligning information technology to global business strategy. IBM Systems Journal, 32(1):143–161, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5507.

Immanuel:1995:ASN

[IKS^{+}95] C. P. Immanuel, G. M. Kump, H. J. Sandick, D. A. Sinicrope, and K. V. Vu. Access services for the Networking BroadBand Services architecture. IBM Systems Journal, 34(4):659–671, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#eight.

Ihde:2001:IBT

[IMMB01] S. C. Ihde, P. P. Maglio, J. Meyer, and R. Barrett. Intermediary-based transcoding framework. IBM Systems Journal, 40(1):179–192, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/401/ihde.html.

Inman:1998:EMA

[Inm98] E. E. Inman. Enterprise modeling advantages of San Francisco for general ledger systems. IBM Systems Journal, 37(2):170–180, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/372/inman.html.

Irlbeck:1992:NSA

[Irl92] B. W. Irlbeck. Network and system automation and remote system operation. IBM Systems Journal, 31(2):206–222, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5470.

Inokuchi:2007:MCI

[ITIW07] A. Inokuchi, K. Takeda, N. Inaoka, and F. Wakao. MedTAKMI-CDI: Interactive knowledge discovery for clinical decision intelligence. IBM Systems Journal, 46(1):115– 133, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/461/inokuchi.html.

Iverson:1963:PNS

[Ive63]  K. E. Iverson.  Program-

ming notation in systems design. IBM Systems Journal, 2: 117–128, June 1963. CODEN IBMSA7. ISSN 0018-8670.

Iverson:1991:PVA

[Ive91] K. E. Iverson. A personal view of APL. IBM Systems Journal, 30(4):582–593, 1991. CODEN IBMSA7. ISSN 0018-8670.

Iverson:1999:PNS

[Ive99] K. E. Iverson. Programming notation in systems design. IBM Systems Journal, 38(2– 3):139–150, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/iverson.pdf. Reprinted

from [Ive63].

Johnson:2000:ISM

[JA00] V. M. Johnson and D. K.

Austreng. IBM SanFrancisco: Moving into the marketplace. IBM Systems Journal, 39(2): 244–247, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/johnson.html.

Jacobsen:1968:IGD

[Jac68] J. D. Jacobsen. Interactive graphics in data processing: Geometric relationships for retrieval of geographic information. IBM Systems Journal, 7 (3/4):331–341, 1968. CODEN IBMSA7. ISSN 0018-8670.

Jacobsen:1977:FCC

[Jac77] Tom Jacobsen. Forum: On compilability and composite design. IBM Systems Journal, 16(2):196, 1977. CODEN IBMSA7. ISSN 0018-8670.

Jews:2008:IPS

[JAS08] Curts Jews, Riaz Ahmad, and David H. Surman. IBM Parallel Sysplex clustering: Technology options for continuous availability. IBM Systems Journal, 47(4):505–518, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/jews.pdf.

Jensen:1991:PNF

[JB91] J. R. Jensen and K. A. Beaty. Putting a new face on APL2. IBM Systems Journal, 30(4): 469–489, 1991. CODEN IBMSA7. ISSN 0018-8670.

Jacobson:1996:CDC

[JB96] N. Jacobson and W. Bender. Color as a determined communication. IBM Systems Journal, 35(3&4):526–538, 1996. CO-DEN IBMSA7. ISSN 0018-8670.

Jann:2003:DRB

[JBB03] J. Jann, L. M. Browning, and R. S. Burugula. Dynamic reconfiguration: Basic building blocks for autonomic computing on IBM pSeries servers. IBM Systems Journal, 42(1): 29–37, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/421/jann.pdf.

Jacobson:1997:LB

[JCT^{+}97] J. Jacobson, B. Comiskey, C. Turner, J. Albert, and P. Tsao. The last book. IBM Systems Journal, 36 (3):457–463, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/363/jacobson.html.

Jacobs:1970:ISF

[JD70] J. H. Jacobs and T. J. Dillon. Interactive Saturn flight program simulator. IBM Systems Journal, 9(2):145–158, 1970. CODEN IBMSA7. ISSN 0018-8670.

Johnson:1991:TSS

[JD91] D. B. Johnson and G. M.

Dolan. Transaction Security System extensions to the Common Cryptographic Architecture. IBM Systems Journal, 30(2):230–243, 1991. CODEN IBMSA7. ISSN 0018-8670.

Johnson:1991:CCA

[JDK^{+}91] D. B. Johnson, G. M. Dolan, M. J. Kelly, A. V. Le, and S. M. Matyas. Common Cryptographic Architecture Cryptographic Application Programming Interface. IBM Systems Journal, 30(2):130–150, 1991. CODEN IBMSA7. ISSN 0018-8670.

Joseph:2004:EGC

[JEF04] J. Joseph, M. Ernest, and C. Fellenstein. Evolution of grid computing architecture and grid adoption models. IBM Systems Journal, 43(4): 624–645, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/434/joseph.html;   http:
//www.research.ibm.com/journal/
sj/434/joseph.pdf;   http:
//www.research.ibm.com/journal/
sj/434/joseph.txt.

Jensen:1979:FAC

[Jen79] R. M. Jensen. A formal approach for communication between logically isolated virtual machines. IBM Systems Journal, 18(1):71–92, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Jennings:1992:EFR

[Jen92] T. Jennings. Estimating the fault rate function. IBM Systems Journal, 31(2):300–312, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5475.

Jordan:1991:FSA

[JF91] Stanley Jordan and Erik S. Friis. The foundations of suitability of APL2 for music. IBM Systems Journal, 30 (4):513–526, 1991. CODEN IBMSA7. ISSN 0018-8670.

Johnson:2007:CDP

[JG07] C. M. Johnson and T. W. A. Grandison. Compliance with data protection laws using Hippocratic Database active enforcement and auditing. IBM Systems Journal, 46(2):255– ??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/462/johnson.html.

Johnson:2007:ESI

[JHMO07] M. W. Johnson, A. Hately, B. A. Miller, and R. Orr. Evolving standards for IT service management. IBM Systems Journal, 46(3):583–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.

research.ibm.com/journal/
sj/463/johnson.html.

Jensen:1992:DMQ

[JK92a] D. L. Jensen and A. J. King.

A decomposition method for quadratic programming. IBM Systems Journal, 31(1):39–48, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5459.

Jensen:1992:FGI

[JK92b] D. L. Jensen and A. J. King.

Frontier: a graphical interface for portfolio optimization in a piecewise linear-quadratic risk framework. IBM Systems Journal, 31(1):62–70, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5461.

Jaufmann:2000:UIS

[JL00] E. J. Jaufmann, Jr. and D. C.

Logan. The use of IBM San-Francisco core business processes in human resources scheduling. IBM Systems Journal, 39(2):285–292, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/392/jaufmann.html.

Josten:1997:DUC

[JMNT97] J. W. Josten, C. Mohan, I. Narang, and J. Z. Teng. DB2’s use of the coupling facility for data sharing. IBM Systems Journal, 36(2):327–351, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/362/josten.html.

Mattos:2002:IIR

[JMP02] A. D. Jhingran, N. Mattos, and H. Pirahesh. Information integration: a research agenda. IBM Systems Journal, 41(4): 555–562, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/414/preface.pdf.

Jephson:1969:TCD

[JMV69] J. S. Jephson, R. P. Mcquarrie, and R. E. Vogelsberg. A threevalue computer design verification system. IBM Systems Journal, 8(3):178–188, 1969. CODEN IBMSA7. ISSN 0018-8670.

Jaffe:1983:SRP

[JMW83] J. M. Jaffe, F. H. Moss, and R. A. Weingarten. SNA routing: Past, present, and possible future. IBM Systems Journal, 23(4):417–434, 1983. CODEN IBMSA7. ISSN 0018-8670.

Janson:1992:ADO

[JMZ92] P. Janson, R. Molva, and S. Zatti. Architectural directions for opening IBM networks: The case of OSI. IBM Systems Journal, 31(2):313– 335, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5476.

Johnson:1992:RDF

[JN92] E. L. Johnson and G. L. Nemhauser. Recent developments and future directions in mathematical programming. IBM Systems Journal, 31(1):

79–93, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5463.

Johnson:1968:IGD

[Joh68] C. I. Johnson. Interactive graphics in data processing: Principles of interactive systems. IBM Systems Journal, 7 (3/4):147–173, 1968. CODEN IBMSA7. ISSN 0018-8670.

Johnson:1995:BSP

[Joh95] D. B. Johnson. Books: “Network Security: Private Communication in a Public World”, by Charlie Kaufman, Radia Perlman, and Mike Speciner. IBM Systems Journal, 34(4): 752–753, 1995. CODEN IBMSA7. ISSN 0018-8670.

Johnson:1997:ISM

[Joh97] P. Johnson. The importance of systems management for a Parallel Sysplex. IBM Systems Journal, 36(2):301–326, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/362/johnson.html.

Jones:1977:UDR

[Jon77]  Alice U. Jones.  A user-

oriented data-base retrieval system. IBM Systems Journal, 16(1):4–17, 1977. CODEN IBMSA7. ISSN 0018-8670.

Jones:1978:MPQ

[Jon78] T. Capers Jones. Measuring programming quality and productivity. IBM Systems Journal, 17(1):39–63, 1978. CO-DEN IBMSA7. ISSN 0018-8670.

Jones:1985:PAD

[Jon85]  C. L. Jones.  A process-

integrated approach to defect prevention. IBM Systems Journal, 24(2):150–167, 1985. CO-DEN IBMSA7. ISSN 0018-8670.

Joslin:1981:SPF

[Jos81] Philip H. Joslin. System Productivity Facility. IBM Systems Journal, 20(4):388–406, 1981. CODEN IBMSA7. ISSN 0018-8670.

Johnson:2002:DTB

[JOZG02] D. E. Johnson, F. J. Oles, T. Zhang, and T. Goetz. A decision-tree-based symbolic rule induction system for text categorization. IBM Systems Journal, 41(3):428–437, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/413/johnson.html.

Johnson:2004:IDT

[JR04] R. D. Johnson and D. Reimer. Issues in the development of transactional Web applications. IBM Systems Journal, 43(2): 430–440, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/432/johnson.html;   http:
//www.research.ibm.com/journal/
sj/432/johnson.pdf;   http:
//www.research.ibm.com/journal/
sj/432/johnson.txt.

Judisch:1981:EMS

[JRD81] J. M. Judisch, B. A. Rupp, and R. A. Dassinger. Effects of manual style on performance in education and machine maintenance. IBM Systems Journal, 20(2):172–183, 1981. CODEN IBMSA7. ISSN 0018-8670.

Jurisica:2001:IDS

[JRG^{+}01] I. Jurisica, P. Rogers, J. I.

Glasgow, S. Fortier, J. R. Luft, J. R. Wolfley, M. A. Bianca, D. R. Weeks, and G. T. DeTitta. Intelligent decision support for protein crystal growth. IBM Systems Journal, 40(2):394–409, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/jurisica.html;   http:
//www.research.ibm.com/journal/
sj/402/jurisica.pdf;   http:
//www.research.ibm.com/journal/
sj/402/jurisica.txt.

Joshi:2007:IDS

[JRST07] N. Joshi, W. Riley, J. Schneider, and Y.-S. Tan. Integration of domain-specific IT processes and tools in IBM Service Management. IBM Systems Journal, 46(3):497–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/joshi.html.

Jasinski:2007:P

[JSR07] Joseph M. Jasinski, David I.

Seidman, and John J. Ritsko. Preface. IBM Systems Journal, 46(1):3–??, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/461/preface.html.

Kuki:1971:FEL

[KA71] H. Kuki and J. Ascoly. FOR-TRAN extended-precision library. IBM Systems Journal, 10(1):39–61, 1971. CODEN IBMSA7. ISSN 0018-8670.

Kahn:1984:OTR

[Kah84] S. Kahn. An overview of three relational data base products. IBM Systems Journal, 23(2):100–111, 1984. CODEN IBMSA7. ISSN 0018-8670.

Kanyuh:1988:INM

[Kan88] D. Kanyuh. An integrated network management product. IBM Systems Journal, 27(1): 45–59, January 1988. CODEN IBMSA7. ISSN 0018-8670.

Kan:1991:MSD

[Kan91] S. H. Kan. Modeling and software development quality. IBM Systems Journal, 30(3):351– 362, 1991. CODEN IBMSA7. ISSN 0018-8670.

Kudo:2007:BPT

[KAN^{+}07] M. Kudo, Y. Araki, H. Nomiyama, S. Saito, and Y. Sohda. Best practices and tools for personal information compliance management. IBM Systems Journal, 46(2):235–??, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/462/kudo.html.

Kasson:1986:AVT

[Kas86] J. M. Kasson. Advanced voice/ data telephone switching system. IBM Systems Journal, 25 (3-4):380–398, 1986. CODEN IBMSA7. ISSN 0018-8670.

Katz:1990:BEM

[Kat90] R. L. Katz. Business/enterprise modeling. IBM Systems Journal, 29(4):509–525, 1990. CO-DEN IBMSA7. ISSN 0018-8670.

Koen:2000:TFT

[KB00] D. B. Koen and W. Bender. Time frames: Temporal augmentation of the news. IBM Systems Journal, 39(3/4): 597–616, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/koen.html;
http://www.research.ibm.
com/journal/sj/393/part1/
koen.pdf;   http://www.research.
ibm.com/journal/sj/393/part1/
koen.txt.

Kienzle:1997:MFS

[KBB^{+}97] M. G. Kienzle, R. R. Berbec, G. P. Bozman, C. K. Eilert, M. Eshel, and R. Mansell. Multimedia file serving with the OS/390 LAN Server. IBM Systems Journal, 36(3):374–392, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/363/kienzle.html.

Kapoor:2005:TFS

[KBB^{+}05] S. Kapoor, K. Bhattacharya, S. Buckley, P. Chowdhary, M. Ettl, K. Katircioglu, E. Mauch, and L. Phillips. A technical framework for senseand-respond business management. IBM Systems Journal, 44(1):5–24, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/441/kapoor.html;   http:
//www.research.ibm.com/journal/
sj/441/kapoor.pdf;   http:
//www.research.ibm.com/journal/
sj/441/kapoor.txt.

Kapoor:2007:SRS

[KBB^{+}07] S. Kapoor, B. Binney, S. Buckley, H. Chang, T. Chao, M. Ettl, E. N. Luddy, R. K. Ravi, and J. Yang. Senseand-respond supply chain using model-driven techniques. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/464/kapoor.html.

Kumaran:2007:UMD

[KBC^{+}07] S. Kumaran, P. Bishop, T. Chao, P. Dhoolia, P. Jain, R. Jaluka, H. Ludwig, A. Moyer, and A. Nigam. Using a model-driven transformational approach and service-oriented architecture for service delivery management. IBM Systems Journal, 46(3):513–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/kumaran.html.

Karat:2003:PUE

[KBK^{+}03] C. M. Karat, C. Brodie, J. Karat, J. Vergo, and S. R. Alpert. Personalizing the user experience on ibm.com. IBM Systems Journal, 42(4): 686–701, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/424/brodie.pdf.

Kan:1994:SQO

[KBS94] S. H. Kan, V. R. Basili, and L. N. Shapiro. Software quality: An overview from the perspective of total quality management. IBM Systems Journal, 33(1):4–19, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-1.html#two. G321-5530.

Kan:1994:ASQ

[KDA^{+}94] S. H. Kan, S. D. Dull, D. N.

Amundson, R. J. Lindner, and R. J. Hedger. AS/400 software quality management. IBM Systems Journal, 33(1):62–88, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/

                                        sj33-1.html#five.  G321-

5533.

King:1997:CAP

[KDY97] G. M. King, D. M. Dias, and P. S. Yu. Cluster architectures and S/390 Parallel Sysplex scalability. IBM Systems Journal, 36(2):221–241, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/362/king.html.

Keeley:1967:AMS

[Kee67] J. F. Keeley. An applicationoriented multiprocessing system, part I: Introduction. IBM Systems Journal, 6(2):78–79, 1967. CODEN IBMSA7. ISSN 0018-8670.

Keefe:1968:HCP

[Kee68] D. D. Keefe. Hierarchical control programs for systems evaluation. IBM Systems Journal, 7(2):123–133, 1968. CODEN IBMSA7. ISSN 0018-8670.

Keen:1993:ITM

[Kee93] P. G. W. Keen. Information technology and the management difference: A fusion map. IBM Systems Journal, 32(1): 17–39, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5501.

Kennelly:2003:MVP

[Ken03] Peggy Kennelly. Message from the Vice President, On Demand Innovation Services, Research Division. IBM Systems Journal, 42(3):??, ????

2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/423/message.pdf.

Keser:2003:EGD

[Kes03] C. Keser. Experimental games for the design of reputation management systems. IBM Systems Journal, 42(3):498– 506, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/keser.pdf.

Kumaran:2002:FBA

[KHC02] S. Kumaran, Y. Huang, and J.-Y. Chung. A frameworkbased approach to building private trading exchanges. IBM Systems Journal, 41(2):253– 271, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/412/kumaran.html;   http:
//www.research.ibm.com/journal/
sj/412/kumaran.pdf;   http:
//www.research.ibm.com/journal/
sj/412/kumaran.txt.

Koehler:2005:DTM

[KHSW05] J. Koehler, R. Hauser, S. Sendall, and M. Wahler. Declarative techniques for model-driven business process integration. IBM Systems Journal, 44(1): 47–65, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/441/koehler.html;   http:

                                        //www.research.ibm.com/journal/
                                        sj/441/koehler.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/441/koehler.txt.

Kiely:1979:OSD

[Kie79] S. C. Kiely. An operating system for distributed processing —DPPX. IBM Systems Journal, 18(4):507–525, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Kingston:1973:CFM

[Kin73] P. L. Kingston. Concepts of financial models. IBM Systems Journal, 12(2):113–125, 1973. CODEN IBMSA7. ISSN 0018-8670.

Korein:1987:R

[KIS87] James U. Korein and Jahuda Ish-Shalom. Robotics. IBM Systems Journal, 26(1):55– 95, January 1987. CODEN IBMSA7. ISSN 0018-8670.

Kazi:2000:JCS

[KJBH^{+}00] I. H. Kazi, D. P. Jose, B. Ben-Hamida, C. J. Hescott, C. Kwok, J. A. Konstan, D. J. Lilja, and P.-C Yew. JaViz: a client/server Java profiling tool. IBM Systems Journal, 39(1):96–117, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/391/kazi.html.

Kaczmarski:2003:BBT

[KJP03] M. Kaczmarski, T. Jiang, and D. A. Pease. Beyond backup toward storage management.

IBM Systems Journal, 42(2): 322–337, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/422/kaczmarski.pdf.

Karat:2003:EUC

[KK03] J. Karat and C. M. Karat. The evolution of user-centered focus in the human-computer interaction field. IBM Systems Journal, 42(4):532–541, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/424/karat.pdf.

Kloppmann:2004:BPC

[KKL^{+}04] M. Kloppmann, D. König, F. Leymann, G. Pfau, and D. Roller. Business process choreography in WebSphere: Combining the power of BPEL and J2EE. IBM Systems Journal, 43(2):270–296, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/432/kloppmann.html;
http://www.research.ibm.
com/journal/sj/432/kloppmann.
pdf;   http://www.research.
ibm.com/journal/sj/432/kloppmann.
txt.

Khalaf:2006:BPW

[KKL06] R. Khalaf, A. Keller, and F. Leymann. Business processes for Web Services: Principles and applications. IBM Systems Journal, 45(2):425–

446, ????  2006.  CO-

DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/452/khalaf.html.

Kano:2005:ASB

[KKLR05] M. Kano, A. Koide, T.-K. Liu, and B. Ramachandran. Analysis and simulation of business solutions in a service-oriented architecture. IBM Systems Journal, 44(4):669–690, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/444/kano.html.

Kamel:1988:SCI

[KKS88] A. Kamel, M. Kindelan, and P. Sguazzero. Seismic computations on the IBM 3090 Vector Multiprocessor. IBM Systems Journal, 27(4):510– 527, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Keehn:1974:VDS

[KL74] Daniel G. Keehn and John O. Lacy. VSAM data set design parameters. IBM Systems Journal, 13(3):186–212, 1974. CODEN IBMSA7. ISSN 0018-8670. See correction [Ano74j] and errata [Ano74l].

Kloeckner:2008:MVP

[Klo08] Kristof Kloeckner. Message from the Vice President of Strategy and Technology, IBM Software Group. IBM Systems Journal, 47(2):??, April/June 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/472/message.html.

Kravitz:1986:WMC

[KLRP86] J. K. Kravitz, D. Lieber, F. H. Robbins, and J. M. Palermo. Workstations and mainframe computers working together. IBM Systems Journal, 25(1):116–128, 1986. CO-DEN IBMSA7. ISSN 0018-8670.

Kozuh:1984:SCD

[KLS84] Frank T. Kozuh, David L. Livingston, and Thomas C. Spillman. System/370 capability in a desktop computer. IBM Systems Journal, 23(3):245–254, 1984. CODEN IBMSA7. ISSN 0018-8670.

Kingman:1990:OIS

[KLS90] L. C. Kingman, III, R. E. Lambert, and R. P. Steen. Operational image systems: a new opportunity. IBM Systems Journal, 29(3):304–312, 1990. CO-DEN IBMSA7. ISSN 0018-8670.

Kolsky:1989:HCI

[KM89] H. G. Kolsky and R. A. MacKinnon. History and contributions of the IBM Scientific Centers. IBM Systems Journal, 28(4):502–524, 1989. CODEN IBMSA7. ISSN 0018-8670.

Kanne:2006:ISC

[KM06a] C. C. Kanne and G. Moerkotte. The importance of sibling clustering for efficient bulkload of XML document trees.

IBM Systems Journal, 45(2): 321–334, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/kanne.html.

Kogan:2006:ESC

[KM06b] S. L. Kogan and M. J. Muller. Ethnographic study of collaborative knowledge work. IBM Systems Journal, 45(4): 759–771, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/454/kogan.html.

Konheim:1980:ICP

[KMM^{+}80] Alan G. Konheim, Marian H. Mack, Robert K. McNeill, Bryant Tuckerman, and Gerald Waldbaum. The IPS cryptographic programs. IBM Systems Journal, 19(2):253–283, 1980. CODEN IBMSA7. ISSN 0018-8670.

Kirsch:2007:MCI

[KMSM^{+}07] C. Kirsch, M. Mattingley-Scott, C. Muszynski, F. Schaefer, and C. Weiss. Monitoring chronically ill patients using mobile technologies. IBM Systems Journal, 46(1):85–93, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/461/kirsch.html.

Korn:1985:EPC

[KMT85] P. A. Korn, J. P. McAdaragh, and C. L. Tondo. Expanded personal computing power and capability. IBM Systems Journal, 24(1):26–36, 1985. CO-DEN IBMSA7. ISSN 0018-8670. See erratum [KMT86].

Korn:1986:EEP

[KMT86] P. A. Korn, J. P. McAdaragh, and C. L. Tondo. Erratum: “Expanded personal computing power and capability”. IBM Systems Journal, 25(1): 130, 1986. CODEN IBMSA7. ISSN 0018-8670. See [KMT85].

Koved:2001:SCE

[KNN^{+}01] L. Koved, A. Nadalin, N. Nagaratnam, M. Pistoia, and T. Shrader. Security challenges for Enterprise Java in an e-business environment. IBM Systems Journal, 40(1): 130–152, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/401/koved.html.

Koved:1998:EJS

[KNNL98] L. Koved, A. J. Nadalin, D. Neal, and T. Lawson. The evolution of Java security. IBM Systems Journal, 37 (3):349–364, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/373/koved.html.

Keel:2007:TOS

[KOH^{+}07] A. J. Keel, M. A. Orr, R. R.

Hernandez, E. A. Patrocinio, and J. Bouchard. From a technology-oriented to a service-oriented approach to IT management. IBM Systems Journal, 46(3):549–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/463/keel.html.

Koide:1989:DMC

[Koi89] A. Koide. Designing molecules and crystals by computer. IBM Systems Journal, 28(4):613– 627, 1989. CODEN IBMSA7. ISSN 0018-8670.

Kolsky:1969:PFU

[Kol69] H. G. Kolsky. Problem formulation using APL. IBM Systems Journal, 8(3):204–219, 1969. CODEN IBMSA7. ISSN 0018-8670.

Konzynski:1993:SCE

[Kon93] B. R. Konsynski. Strategic control in the extended enterprise. IBM Systems Journal, 32(1):111–142, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5506.

Kossack:1963:ORA

[Kos63a] C. F. Kossack. Optimum response analysis. IBM Systems Journal, 2:49–56, March 1963. CODEN IBMSA7. ISSN 0018-8670.

Kossack:1963:SCT

[Kos63b] C. F. Kossack. Statistical classification techniques. IBM Systems Journal, 2:136–161, June 1963. CODEN IBMSA7. ISSN 0018-8670.

Kovac:2001:MVP

[Kov01] Caroline A. Kovac. Message from the Vice President, Life Sciences Solutions. IBM Systems Journal, 40(2):??, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/message.html;   http:
//www.research.ibm.com/journal/
sj/402/message.pdf;   http:
//www.research.ibm.com/journal/
sj/402/message.txt.

Kozakov:2004:GEU

[KPF^{+}04] L. Kozakov, Y. Park, T. Fin, Y. Drissi, Y. Doganata, and T. Cofino. Glossary extraction and utilization in the information search and delivery system for IBM Technical Support. IBM Systems Journal, 43(3):546–563, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/kozakov.html;   http:
//www.research.ibm.com/journal/
sj/433/kozakov.pdf;   http:
//www.research.ibm.com/journal/
sj/433/kozakov.txt.

Kan:2001:PMS

[KPM01] S. H. Kan, J. Parrish, and D. Manlove. In-process metrics for software testing. IBM Systems Journal, 40(1):220– 241, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.

                                        research.ibm.com/journal/
                                        sj/401/kan.html.

Kogan:1988:DOS

[KR88] Michael S. Kogan and Freeman L. Rawson, III. The design of Operating System/2. IBM Systems Journal, 27(2):90–104, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Kraemer:1978:PID

[Kra78] Wolfgang Kraemer. Performance investigations with a DOS/VS-based operating system model. IBM Systems Journal, 17(4):409–443, 1978. CO-DEN IBMSA7. ISSN 0018-8670.

Krauss:2003:MRU

[Kra03] F. S. H. Krauss. Methodology for remote usability activities: a case study. IBM Systems Journal, 42(4):582–593, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/424/krauss.pdf.

Kreger:2001:JME

[Kre01] H. Kreger. Java Management Extensions for application management. IBM Systems Journal, 40(1):104–129, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/401/kreger.html.

Katz:1980:SCL

[KRR80] Shumel Katz, Lawrence J. Risman, and Michael Rodeh. A system for constructing linear programming models. IBM Systems Journal, 19(4):505– 520, 1980. CODEN IBMSA7. ISSN 0018-8670.

Kellogg:1998:NGI

[KRS^{+}98] W. A. Kellogg, J. T. Richards, C. Swart, P. Malkin, M. Laff, V. Hanson, and B. Hailpern. NetVista: Growing an Internet solution for schools. IBM Systems Journal, 37(1):19–41, 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/371/kellogg.html.

Kurtz:2003:NDS

[KS03] C. F. Kurtz and D. J. Snowden.

The new dynamics of strategy: Sense-making in a complex and complicated world. IBM Systems Journal, 42(3): 462–483, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/kurtz.pdf.

Kamel:1993:SRH

[KSWC93] M. S. Kamel, H. C. Shen, A. K. C. Wong, and R. I. Campeanu. System for the recognition of human faces. IBM Systems Journal, 32(2): 307–320, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5515.

King:2005:MUP

[KTBE05] M. King, J. W. Thatcher, P. M. Bronstad, and R. Easton.

Managing usability for people with disabilities in a large Web presence. IBM Systems Journal, 44(3):519–535, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/443/king.html.

Koved:1986:IAS

[KW86] L. Koved and G. Waldbaum. Improving availability of software subsystems through online error detection. IBM Systems Journal, 25(1):105–115, 1986. CODEN IBMSA7. ISSN 0018-8670.

Leymann:1994:MBP

[LA94] F. Leymann and W. Altenhuber. Managing business processes as an information resource. IBM Systems Journal, 33(2):326–348, 1994. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj33-2.html#seven. G321-

5545.

Lacy:1981:P

[Lac81] John Lacy. Preface. IBM Systems Journal, 20(4):370–371, 1981. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1982:Pa

[Lac82a] John Lacy. Preface. IBM Systems Journal, 21(1):2–3, 1982. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1982:Pb

[Lac82b] John Lacy. Preface. IBM Systems Journal, 21(2):130, 1982. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1982:Pc

[Lac82c] John Lacy. Preface. IBM Systems Journal, 21(3):248–249, 1982. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1982:Pd

[Lac82d] John Lacy. Preface. IBM Systems Journal, 21(4):382–383, 1982. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1983:Pa

[Lac83a] John Lacy. Preface. IBM Systems Journal, 23(1/2):2–3, 1983. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1983:Pd

[Lac83b] John Lacy. Preface. IBM Systems Journal, 23(1):2–3, 1983. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1983:Pb

[Lac83c] John Lacy. Preface. IBM Systems Journal, 23(3):168–169, 1983. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1983:Pc

[Lac83d] John Lacy. Preface. IBM Systems Journal, 23(4):296–297, 1983. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1984:Pa

[Lac84a] John Lacy. Preface. IBM Systems Journal, 23(2):98–99, 1984. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1984:Pb

[Lac84b] John Lacy. Preface. IBM Systems Journal, 23(3):222–223, 1984. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1984:Pc

[Lac84c] John Lacy. Preface. IBM Systems Journal, 23(4):308, 1984. CODEN IBMSA7. ISSN 0018-8670.

Lacy:1985:P

[Lac85] John Lacy. Preface. IBM Systems Journal, 24(1):2, 1985. CODEN IBMSA7. ISSN 0018-8670.

Lafuente:1968:IGD

[Laf68] J. M. Lafuente. Interactive graphics in data processing: Cam design on a graphics console. IBM Systems Journal, 7 (3/4):365–372, 1968. CODEN IBMSA7. ISSN 0018-8670.

Laird:1975:PPE

[Lai75] Jack A. Laird. Power profile — an installation management tool. IBM Systems Journal, 14(3):264–271, 1975. CODEN IBMSA7. ISSN 0018-8670.

Lambert:1984:CSS

[Lam84] Geoffrey N. Lambert. A comparative study of system response time on program developer productivity. IBM Systems Journal, 23(1):36–43, 1984. CODEN IBMSA7. ISSN 0018-8670.

Lande:1973:PS

[Lan73] H. F. Lande. Planning-data systems. IBM Systems Journal, 12(2):145–160, 1973. CODEN IBMSA7. ISSN 0018-8670.

Larner:1963:DIP

[Lar63]  R. Larner.  Design of an

integrated programming and operating system, part IV: The system’s FORTRAN compiler. IBM Systems Journal, 2: 311–321, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Larsen:2006:MDD

[Lar06] G. Larsen. Model-driven development: Assets and reuse. IBM Systems Journal, 45(3): 541–553, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/larsen.html.

Larson:2008:SSI

[Lar08] R. C. Larson. Service science:

At the intersection of management, social, and engineering sciences. IBM Systems Journal, 47(1):41–??, January/March 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/471/larson.html.

Leach:1980:SAU

[LC80] Jack R. Leach and Robert D.

Campenni. A sidestream approach using a small processor as a tool for managing communication systems. IBM Systems Journal, 19(1):120–139, 1980. CODEN IBMSA7. ISSN 0018-8670.

Liu:2001:FCP

[LC01] A. H. Liu and A. Califano. Functional classification of proteins by pattern discovery and top-down clustering of primary sequences. IBM Systems Journal, 40(2):379–393, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/liu.html;   http:/
/www.research.ibm.com/journal/
sj/402/liu.pdf;   http://
www.research.ibm.com/journal/
sj/402/liu.txt.

Lewis:2003:DVS

[LC03] J. R. Lewis and P. M. Commarford. Developing a voicespelling alphabet for PDAs. IBM Systems Journal, 42(4): 624–638, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/424/lewis.pdf.

Lum:1982:OOP

[LCS82] V. Y. Lum, D. M. Choy, and N. C. Shu. OPAS: an office procedure automation system. IBM Systems Journal, 21(3):327–350, 1982. CODEN IBMSA7. ISSN 0018-8670.

Landay:1999:MSP

[LD99] J. A. Landay and R. C.

Davis. Making sharing pervasive: Ubiquitous computing for shared note taking. IBM Systems Journal, 38(4): 531–550, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/landay.html.

Leon:2008:MSP

[LD08] N. Leon and A. Davies. Managed service paradox. IBM Systems Journal, 47(1):153– ??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/471/leon.html.

Loveland:2008:LVO

[LDL^{+}08] Scott Loveland, Eli M. Dow, Frank LeFevre, Duane Beyer, and Philip F. Chan. Leveraging virtualization to optimize high-availability system configurations. IBM Systems Journal, 47(4):591–604, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/loveland.pdf.

LeBlanc:2008:MGM

[LeB08]  Robert LeBlanc.  Message

from the General Manager, Global Consulting Services and SOA. IBM Systems Journal, 47(3):??, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/473/message.pdf.

Leishman:1999:SC

[Lei99] D. A. Leishman. Solution customization. IBM Systems Journal, 38(1):76–97, ???? 1999. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/381/leishman.html.

Lennon:1978:CAI

[Len78] Richard E. Lennon. Cryptography architecture for information security. IBM Systems Journal, 17(2):138–150, 1978. CODEN IBMSA7. ISSN 0018-8670.

Levine:1979:FDS

[Lev79] John R. Levine. Forum: Data stream linkage and the UNIX system. IBM Systems Journal, 18(3):470–472, 1979. CODEN IBMSA7. ISSN 0018-8670. See [Mor78, Mor79].

Lewis:1966:CPS

[Lew66] P. A. W. Lewis. A computer program for the statistical analysis of series of events. IBM Systems Journal, 5(4):202–225, 1966. CODEN IBMSA7. ISSN 0018-8670.

Lewis:1976:SLC

[Lew76] Leslie J. Lewis. Service levels: a concept for the user and the computer center. IBM Systems Journal, 15(4):328–357, 1976. CODEN IBMSA7. ISSN 0018-8670.

Lloyd:1999:TRA

[LG99] P. T. L. Lloyd and G. M.

Galambos. Technical reference architectures. IBM Systems Journal, 38(1):51–75, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/381/lloyd.html.

Lebizay:1995:HTN

[LGCB95] G. Lebizay, C. Galand, D. Chevalier, and F. Barre. A highperformance transport network platform. IBM Systems Journal, 34(4):705–724, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-4.html#eleven.

Lewis:1969:PNG

[LGM69] P. A. W. Lewis, A. S. Goodman, and J. M. Miller. A pseudorandom number generator for the SYSTEM/360. IBM Systems Journal, 8(2):136–146, 1969. CODEN IBMSA7. ISSN 0018-8670.

Lindsey:1997:BTO

[LH97] A. H. Lindsey and P. R. Hoffman. Bridging traditional and object technologies: Creating transitional applications. IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/361/lindsey.html.

Ludwig:2007:CBS

[LHJ^{+}07] H. Ludwig, J. Hogan, R. Jaluka, D. Loewenstern, S. Kumaran, A. Gilbert, A. Roy, T. R. Nellutla, and M. Surendra. Catalog-based service request management. IBM Systems Journal, 46(3):531–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/ludwig.html.

Luo:2007:AEB

[LHT^{+}07] L. Luo, D. M. Hao, Z. Tian, Y. B. Dang, B. Hou, P. Malkin, and S. X. Yang. Ariadne: An Eclipse-based system for tracking originality of source code. IBM Systems Journal, 46(2):289–??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/462/luo.html.

Lee:2001:GSG

[LI01] C. Lee and K. Irizarry. The GeneMine system for genome/ proteome annotation and collaborative data mining. IBM Systems Journal, 40(2):592– 603, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/402/lee.html;   http:/
/www.research.ibm.com/journal/
sj/402/lee.pdf;   http://
www.research.ibm.com/journal/
sj/402/lee.txt.

Lickel:2007:I

[Lic07] Charles W. Lickel. Introduction. IBM Systems Journal, 46(2):??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/462/intro.html.

Linger:1980:MSE

[Lin80] R. C. Linger. The management of software engineering, part III: Software design practices. IBM Systems Journal, 19(4):432–450, 1980. CODEN IBMSA7. ISSN 0018-8670.

Liptay:1968:SAS

[Lip68] J. S. Liptay. Structural aspects of the SYSTEM/360 model 85, part II: The cache. IBM Systems Journal, 7(1):15–21, 1968. CODEN IBMSA7. ISSN 0018-

8670. URL https://doi.org/
10.1147/sj.71.0015.

Lippman:1996:PV

[Lip96] A. Lippman. Post-modern video. IBM Systems Journal, 35(3&4):269–271, 1996. CO-DEN IBMSA7. ISSN 0018-8670.

Lippman:2000:MNC

[Lip00] A. Lippman. Media, networks, and content. IBM Systems Journal, 39(3/4):435–437, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/393/part1/lippman.html;
http://www.research.ibm.
com/journal/sj/393/part1/

lippman.pdf;   http://www.
research.ibm.com/journal/
sj/393/part1/lippman.txt.

Lippman:1996:MBE

[LK96] A. Lippman and R. Kermode. Media banks: Entertainment and the Internet. IBM Systems Journal, 35(3&4):272– 291, 1996. CODEN IBMSA7. ISSN 0018-8670.

Lasser:2005:MVP

[LK05] Catherine Lasser and Margaret Kennelly. Message from the Vice Presidents, Industry Solutions and Emerging Businesses and On Demand Innovation Services. IBM Systems Journal, 44(1):??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/441/message.html;   http:
//www.research.ibm.com/journal/
sj/441/message.pdf;   http:
//www.research.ibm.com/journal/
sj/441/message.txt.

Luftman:1993:TEA

[LLO93] J. N. Luftman, P. R. Lewis, and S. H. Oldach. Transforming the enterprise: The alignment of business and information technology strategies. IBM Systems Journal, 32(1):198–221, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5510.

Luo:2006:CAS

[LLS^{+}06] L. Luo, J. Liu, L. Shao, W. Lu, and M. Ye. A context-aware smart-call-center solution: Improving customer service for online games. IBM Systems Journal, 45(1):145–160, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/451/luo.html.

Longley:1993:TNC

[LM93] D. Longley and S. M. Matyas.

Technical note: Complementarity attacks and control vectors. IBM Systems Journal, 32(2):321–325, 1993. CODEN IBMSA7. ISSN 0018-8670.

Lieberman:1996:IAS

[LM96] H. Lieberman and D. Maulsby.

Instructible agents: Software that just keeps getting better. IBM Systems Journal, 35 (3&4):539–556, 1996. CODEN IBMSA7. ISSN 0018-8670.

Le:1993:PKE

[LMJW93] A. V. Le, S. M. Matyas, D. B.

Johnson, and J. D. Wilkins. A public key extension to the Common Cryptographic Architecture. IBM Systems Journal, 32(3):461–485, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5521.

Lindquist:2007:ISM

[LMPR07] D. Lindquist, H. Madduri, C. J.

Paul, and B. Rajaraman. IBM Service Management architecture. IBM Systems Journal, 46(3):423–??, July/September 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/463/lindquist.html.

Loveland:2002:TZP

[LMPS02] S. Loveland, G. Miller, R. Prewitt, and M. Shannon. Testing z/OS: The premier operating system for IBM’s zSeries server. IBM Systems Journal, 41(1):55–73, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/411/loveland.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/411/loveland.pdf.

Liu:2007:SCC

[LMX07] Y. Liu, S. Müller, and K. Xu. A static compliance-checking framework for business process models. IBM Systems Journal, 46(2):335–??, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/462/liu.html.

Leroux:2006:RSA

[LNH06] D. Leroux, M. Nally, and K. Hussey. Rational Software Architect: a tool for domainspecific modeling. IBM Systems Journal, 45(3):555–568, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/453/leroux.html.

Loewner:1965:FAO

[Loe65] P. G. Loewner. Fabrication and assembly operations, part III: Matrix methods for processing configuration data. IBM Systems Journal, 4(2):105–121, 1965. CODEN IBMSA7. ISSN 0018-8670.

Lorin:1971:GBS

[Lor71] H. Lorin. A guided bibliography to sorting. IBM Systems Journal, 10(3):244–254, March 1971. CODEN IBMSA7. ISSN 0018-8670.

Lorin:1979:DPA

[Lor79] H. Lorin. Distributed processing: An assessment. IBM Systems Journal, 18(4):582–603, 1979. CODEN IBMSA7. ISSN 0018-8670.

Lorin:1986:SAT

[Lor86] H. Lorin. Systems architecture in transition — an overview. IBM Systems Journal, 25(3-4):256–273, 1986. CODEN IBMSA7. ISSN 0018-8670.

Lorin:1994:BCD

[Lor94a] Harold Lorin. Books: Centralized and Distributed Operating Systems by Gary J. Nutt. IBM Systems Journal, 33(2):375–377, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-2.html.

Lorin:1994:BEC

[Lor94b] Harold Lorin. Books: Enterprise Computing and Implementing the Enterprise by Alan R. Simon. IBM Systems Journal, 33(2):373–375, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-2.html.

Lynch:1974:ORS

[LP74] H. W. Lynch and J. B. Page. The OS/VS2 Release 2 System Resources Manager. IBM Systems Journal, 13(4):274–291, 1974. CODEN IBMSA7. ISSN 0018-8670.

Lanfranchi:2003:TNL

[LPPC03] G. Lanfranchi, P. Della Peruta, A. Perrone, and D. Calvanese. Toward a new landscape of systems management in an autonomic computing environment. IBM Systems Journal, 42(1):119–128, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/lanfranchi.pdf.

Lawrence:2007:ADS

[LPR^{+}07] R. Lawrence, C. Perlich, S. Rosset, J. Arroyo, M. Callahan, J. M. Collins, A. Ershov, S. Feinzig, I. Khabibrakhmanov, S. Mahatma, M. Niemaszyk, and S. M. Weiss. Analyticsdriven solutions for customer targeting and sales-force allocation. IBM Systems Journal, 46(4):??, October/December 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/464/lawrence.html.

Lancto:1967:AMS

[LR67] D. C. Lancto and R. L. Rockefeller. An application-oriented multiprocessing system, part IV: The operational error analysis program. IBM Systems Journal, 6(2):103–115, 1967. CODEN IBMSA7. ISSN 0018-8670.

Leymann:1997:WBA

[LR97] F. Leymann and D. Roller.

Workflow-based applications. IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/361/leymann.html.

Lau:2002:DXW

[LR02a] C. Lau and A. Ryman. Developing XML Web services with WebSphere Studio Application Developer. IBM Systems Journal, 41(2):178–197, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/412/lau.html;   http:/
/www.research.ibm.com/journal/
sj/412/lau.pdf;   http://
www.research.ibm.com/journal/
sj/412/lau.txt.

Leymann:2002:UFI

[LR02b] F. Leymann and D. Roller. Using flows in information integration. IBM Systems Journal, 41(4):732–742, ???? 2002. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/414/leymann.pdf.

Lewis:2004:MDS

[LRHH04] K. N. Lewis, M. D. Robinson, T. R. Hughes, and C. W. V. Hogue. MyMed: a database system for biomedical research on MEDLINE data. IBM Systems Journal, 43(4):756– 767, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/434/lewis.html;   http:
//www.research.ibm.com/journal/
sj/434/lewis.pdf;   http:
//www.research.ibm.com/journal/
sj/434/lewis.txt.

Leymann:2002:WSB

[LRS02] F. Leymann, D. Roller, and M.-T. Schmidt. Web services and business process management. IBM Systems Journal, 41(2):198–211, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/412/leymann.html;   http:
//www.research.ibm.com/journal/
sj/412/leymann.pdf;   http:
//www.research.ibm.com/journal/
sj/412/leymann.txt.

Loucks:1987:AIE

[LS87] L. K. Loucks and C. H. Sauer. Advanced Interactive Executive (AIX) operating system overview. IBM Systems Journal, 26(4):326–345, December 1987. CODEN IBMSA7. ISSN 0018-8670. URL http:/

/technologists.com/sauer/
Advanced%20Interactive%20Executive%

20(AIX)%20Operating%20System%
20Overview.pdf.

Lenz:1991:KSM

[LS91] N. G. Lenz and S. F. L. Saelens.

A knowledge-based system for MVS dump analysis. IBM Systems Journal, 30(3):336–350, 1991. CODEN IBMSA7. ISSN 0018-8670.

Lieberman:2000:CCS

[LS00] H. Lieberman and T. Selker.

Out of context: Computer systems that adapt to, and learn from, context. IBM Systems Journal, 39(3/4):617–632, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/393/part1/lieberman.html;
http://www.research.ibm.
com/journal/sj/393/part1/
lieberman.pdf;   http://
www.research.ibm.com/journal/
sj/393/part1/lieberman.txt.

Lesser:2001:CPO

[LS01] E. L. Lesser and J. Storck.

Communities of practice and organizational performance. IBM Systems Journal, 40(4): 831–841, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/404/lesser.html;   http:
//www.research.ibm.com/journal/
sj/404/lesser.pdf.

Levine:2005:MAP

[LS05] J. L. Levine and M. A. Schappert. A mouse adapter for people with hand tremor. IBM Systems Journal, 44(3):621– 628, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/443/levine.html.

Lasser:2007:I

[LS07] Catherine Lasser and Michael Svinte. Introduction. IBM Systems Journal, 46(1):??, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/461/intro.html.

Lumpp:2008:HAD

[LSH^{+}08] Thomas Lumpp, Juergen Schneider, Juergen Holtz, Markus Mueller, Norbert Lenz, Ana Biazetti, and David Petersen. From high availability and disaster recovery to business continuity solutions. IBM Systems Journal, 47(4):605–619, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/lumpp.pdf.

Latham:1989:TNC

[LT89] W. H. Latham and S. J. P. Todd. Technical note: Computer sculpture. IBM Systems Journal, 28(4):682–688, 1989. CODEN IBMSA7. ISSN 0018-8670.

Li:2008:BPD

[LTL^{+}08] Zhong Jie Li, Hua Fang Tan, He Heui Liu, Jun Zhu, and Naomi M. Mitsumori. Business-process-driven graybox SOA testing. IBM Systems Journal, 47(3):457–472, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/473/li.pdf.

Lucente:1996:CHB

[Luc96] M. Lucente. Computational holographic bandwidth compression. IBM Systems Journal, 35(3&4):349–366, 1996. CO-DEN IBMSA7. ISSN 0018-8670.

Luetje:1970:IAC

[Lue70] J. H. Luetje. Interactive aeronautical charting. IBM Systems Journal, 9(3):219–233, 1970. CODEN IBMSA7. ISSN 0018-8670.

Lusch:2008:TCF

[LVW08] R. F. Lusch, S. L. Vargo, and G. Wessels. Toward a conceptual foundation for service science: Contributions from service-dominant logic. IBM Systems Journal, 47(1):5–??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/471/lusch.html.

MacKinnon:1974:AFE

[Mac74] R. A. MacKinnon. Advanced function extended with tightlycoupled multiprocessing. IBM Systems Journal, 13(1):32–59, 1974. CODEN IBMSA7. ISSN 0018-8670.

MacKinnon:1979:CVM

[Mac79] R. A. MacKinnon. The changing virtual machine environment: Interfaces to real hardware, virtual hardware, and other virtual machines. IBM Systems Journal, 18(1):18–46, 1979. CODEN IBMSA7. ISSN 0018-8670.

Learning:2000:FAW

[Mac00] T. Machover. Future arts: By and for whom, and for what? IBM Systems Journal, 39(3/4): 678–679, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/393/part1/machover.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part1/
                                        machover.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part1/machover.txt.

Major:1981:PIP

[Maj81] J. B. Major. Processor, I/O path, and DASD configuration capacity. IBM Systems Journal, 20(1):63–85, 1981. CO-DEN IBMSA7. ISSN 0018-8670.

Mancini:1974:PAS

[Man74] R. J. Mancini. Performance analysis for the Skylab terminal system. IBM Systems Journal, 13(2):94–114, 1974. CODEN IBMSA7. ISSN 0018-8670.

Maranzana:1963:LSP

[Mar63] F. E. Maranzana. On the location of supply points to minimize transportation costs. IBM Systems Journal, 2:129– 135, June 1963. CODEN IBMSA7. ISSN 0018-8670.

Marks:1973:DCC

[Mar73] B. L. Marks. Design of a checkout compiler. IBM Systems Journal, 12(3):315–326, 1973. CODEN IBMSA7. ISSN 0018-8670.

Maruyama:1983:DRT

[Mar83] K. Maruyama. Defining routing tables for SNA networks. IBM Systems Journal, 23(4):435– 450, 1983. CODEN IBMSA7. ISSN 0018-8670.

Mark:1999:TPC

[Mar99] W. Mark. Turning pervasive computing into mediated spaces. IBM Systems Journal, 38(4):677, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/mark.html.

Marwick:2001:CPO

[Mar01] A. D. Marwick. Communities of practice and organizational performance. IBM Systems Journal, 40(4):814–830, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/404/marwick.html;   http:
//www.research.ibm.com/journal/
sj/404/marwick.pdf.

Matick:1986:IMS

[Mat86] R. E. Matick. Impact of memory systems on computer architecture and system organization. IBM Systems Journal, 25 (3-4):274–305, 1986. CODEN IBMSA7. ISSN 0018-8670.

Matyas:1991:KHC

[Mat91] S. M. Matyas. Key handling with control vectors. IBM Systems Journal, 30(2):151–174, 1991. CODEN IBMSA7. ISSN 0018-8670.

Maurer:1983:FTI

[Mau83] M. E. Maurer. Full-screen testing of interactive applications. IBM Systems Journal, 23(3):246–261, 1983. CODEN IBMSA7. ISSN 0018-8670.

Mays:1994:FSB

[May94] R. G. Mays. Forging a silver bullet from the essence of software. IBM Systems Journal, 33(1):20–45, 1994. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj33-1.html#three. G321-

5531.

Massey:1996:SSP

[MB96] M. Massey and W. Bender. Salient stills: Process and practice. IBM Systems Journal, 35 (3&4):557–574, 1996. CODEN IBMSA7. ISSN 0018-8670.

Mastrianni:2007:IAF

[MBB^{+}07] S. Mastrianni, D. F. Bantz, K. A. Beaty, T. Chefalas, S. Jalan, G. Kar, A. Kochut, D. J. Lan, L. O’Connell, A. Sailer, G. Wang, Q. B. Wang, and D. G. Shea. IT Autopilot: a flexible IT service management and delivery platform for small and medium business. IBM Systems Journal, 46(3):609–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/mastrianni.html.

Mendonca:1998:AIE

[MBBD98] M. G. Mendonca, V. R. Basili, I. S. Bhandari, and J. Dawson. An approach to improving existing measurement frameworks. IBM Systems Journal, 37(4):484–501, ???? 1998. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/374/mendonca.html.

Morrill:2008:ACA

[MBC08] Harriet Morrill, Mark Beard, and David Clitherow. Achieving continuous availability of IBM systems infrastructures. IBM Systems Journal, 47(4): 493–503, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/morrill.pdf.

Maguire:2000:TTT

[MBG00] Y. Maguire, E. Boyden, and N. Gershenfeld. Toward a table-top quantumcomputer. IBM Systems Journal, 39(3/4): 823–839, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part3/maguire.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part3/
                                        maguire.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part3/maguire.txt.

Musgrave:1996:EDG

[MC96] J. F. Musgrave and M. R. Cooper. Experiments in digital graphic design. IBM Systems Journal, 35(3&4):499– 513, 1996. CODEN IBMSA7. ISSN 0018-8670.

McCue:1978:IST

[McC78] Gerald M. McCue. IBM’s Santa Teresa Laboratory— architectural design for program development. IBM Systems Journal, 17(1):4–25, 1978. CODEN IBMSA7. ISSN 0018-8670.

McDavid:1996:BLA

[McD96] D. W. McDavid. Business language analysis for objectoriented information systems. IBM Systems Journal, 35 (2):128–150, 1996. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj35-2.html#three.

McDavid:1999:SBA

[McD99] D. W. McDavid. A standard for business architecture description. IBM Systems Journal, 38(1):12–31, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/381/mcdavid.html.

Milliken:1986:YAO

[MCE^{+}86] K. R. Milliken, A. V. Cruise, R. L. Ennis, A. J. Finkel, J. L. Hellerstein, D. J. Loeb, D. A. Klein, M. J. Masullo, H. M. Van Woerkom, and N. B. Waite. YES/MVS and the automation of operations for large computer complexes. IBM Systems Journal, 25(2):159–180, 1986. CODEN IBMSA7. ISSN 0018-8670.

McFadyen:1976:SNA

[McF76]  John H. McFadyen.  Sys-

tems Network Architecture: An overview. IBM Systems Journal, 15(1):4–23, 1976. CODEN IBMSA7. ISSN 0018-8670. See erratum [Ano76j].

McGee:1965:DPR

[McG65] W. C. McGee. On dynamic program relocation. IBM Systems Journal, 4(3):184–199, 1965. CODEN IBMSA7. ISSN 0018-8670.

McGrath:1972:VMC

[McG72] M. McGrath. Virtual machine computing in an engineering environment. IBM Systems Journal, 11(2):131–149, 1972. CODEN IBMSA7. ISSN 0018-8670.

McGee:1977:IMSa

[McG77a] William C. McGee. The information management system IMS/VS, part I: General structure and operation. IBM Systems Journal, 16(2):84–95, 1977. CODEN IBMSA7. ISSN 0018-8670.

McGee:1977:IMSb

[McG77b] William C. McGee. The information management system IMS/VS, part II: Data base facilities. IBM Systems Journal, 16(2):96–122, 1977. CODEN IBMSA7. ISSN 0018-8670.

McGee:1977:IMSc

[McG77c] William C. McGee. The information management system IMS/VS, part III: Batch processing facilities. IBM Systems Journal, 16(2):123–135, 1977. CODEN IBMSA7. ISSN 0018-8670.

McGee:1977:IMSd

[McG77d] William C. McGee. The information management system IMS/VS, part Iv: Data communication facilities. IBM Systems Journal, 16(2):136–147, 1977. CODEN IBMSA7. ISSN 0018-8670.

McGee:1977:IMSe

[McG77e] William C. McGee. The information management system IMS/VS, part V: Transaction processing facilities. IBM Systems Journal, 16(2):148–168, 1977. CODEN IBMSA7. ISSN 0018-8670.

McIntyre:1991:LIT

[McI91] D. B. McIntyre. Language as an intellectual tool: From hieroglyphics to APL. IBM Systems Journal, 30(4):554–581, 1991.

CODEN IBMSA7. ISSN 0018-8670.

McNutt:1993:ISC

[McN93] B. McNutt. I/O subsystem configurations for ESA. New roles for processor storage. IBM Systems Journal, 32(2):252–264, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5512.

McPhee:1974:OSI

[McP74] William S. McPhee. Operating system integrity in OS/ VS2. IBM Systems Journal, 13(3):230–252, 1974. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1975:P

[McQ75] George McQuilken. Preface.

IBM Systems Journal, 14(4): 313, 1975. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1976:Pa

[McQ76a] George McQuilken. Preface.

IBM Systems Journal, 15(1): 2–3, 1976. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1976:Pb

[McQ76b] George McQuilken. Preface.

IBM Systems Journal, 15(2): 123, 1976. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1976:Pc

[McQ76c]  George McQuilken.  Pref-

ace. IBM Systems Journal, 15(3):180–181, 1976. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1976:Pd

[McQ76d]  George McQuilken.  Pref-

ace. IBM Systems Journal, 15(4):282–284, 1976. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1977:Pa

[McQ77a] George McQuilken. Preface. IBM Systems Journal, 16(1): 2–3, 1977. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1977:Pb

[McQ77b] George McQuilken. Preface. IBM Systems Journal, 16(2): 82–83, 1977. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1977:Pc

[McQ77c]  George McQuilken.  Pref-

ace. IBM Systems Journal, 16(3):206–207, 1977. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1977:R

[McQ77d]  George McQuilken.  Refer-

ees. IBM Systems Journal, 16 (3):318–319, 1977. CODEN IBMSA7. ISSN 0018-8670.

McQuilken:1994:BDC

[McQ94] George McQuilken. Books: Distributed Computing Environments by Daniel Cerutti and Donna Pierson. IBM Systems Journal, 33(2):372–373, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-2.html.

Milne:2005:GEI

[MDC^{+}05] S. Milne, A. Dickinson, A. Carmichael, D. Sloan, R. Eisma, and P. Gregor. Are guidelines enough? An introduction to designing Web sites accessible to older people. IBM Systems Journal, 44(3):557–571, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/443/milne.html.

Mayadas:1986:EPD

[MDHM86] A. F. Mayadas, R. C. Durbeck, W. D. Hinsberg, and J. M. Mc-Crossin. The evolution of printers and displays. IBM Systems Journal, 25(3-4):399–416, 1986. CODEN IBMSA7. ISSN 0018-8670.

Mealy:1966:FSO

[Mea66] G. H. Mealy. The functional structure of OS/360, part I: Introductory survey. IBM Systems Journal, 5(1):3–11, 1966. CODEN IBMSA7. ISSN 0018-8670.

Meggitt:1964:CCH

[Meg64] J. E. Meggitt. A character computer for high-level language interpretation. IBM Systems Journal, 3(1):68–78, 1964. CO-DEN IBMSA7. ISSN 0018-8670.

Mendoza:1968:ELT

[Men68] A. G. Mendoza. An economic lot-sizing technique, part II: Mathematical analysis of the part-period algorithm. IBM Systems Journal, 7(1):39–46, 1968. CODEN IBMSA7. ISSN 0018-8670.

Mankoff:2005:EAS

[MFJ05] J. Mankoff, H. Fait, and R. Juang. Evaluating accessibility by simulating the experiences of users with vision or motor impairments. IBM Systems Journal, 44(3): 505–517, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/443/mankoff.html.

Moody:2006:BAP

[MGM^{+}06] P. Moody, D. Gruen, M. J. Muller, J. Tang, and T. P. Moran. Business activity patterns: a new model for collaborative business applications. IBM Systems Journal, 45(4): 683–694, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/454/moody.html.

Mattson:1970:ETS

[MGST70] R. L. Mattson, J. Gecsei, D. R. Slutz, and I. L. Traiger. Evaluation techniques for storage hierarchies. IBM Systems Journal, 9(2):78–117, 1970. CO-DEN IBMSA7. ISSN 0018-8670.

McEnroe:1975:OSS

[MHMM75] Paul V. McEnroe, Henry T. Huth, Jr., Eugene A. Moore, and Walter W. Morris, III. Overview of the Supermarket System and the Retail Store System. IBM Systems Journal, 14(1):3–15, 1975. CODEN IBMSA7. ISSN 0018-8670.

Marin:1995:ONA

[MICG95] G. A. Marin, C. P. Immanuel, P. F. Chimento, and I. S. Gopal. Overview of the NBBS architecture. IBM Systems Journal, 34(4):564–589, 1995. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-4.html#three.

Mills:1980:MSE

[Mil80] H. D. Mills. The management of software engineering, part I: Principles of software engineering. IBM Systems Journal, 19(4):414–420, 1980. CODEN IBMSA7. ISSN 0018-8670.

Miller:1981:NLP

[Mil81] L. A. Miller. Natural language programming: Styles, strategies, and contrasts. IBM Systems Journal, 20(2):184–215, 1981. CODEN IBMSA7. ISSN 0018-8670.

Millet:1999:TNP

[Mil99a] I. Millet. Technical note — A proposal to simplify data flow diagrams. IBM Systems Journal, 38(1):118–121, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/381/millet.html.

Mills:1999:MSE

[Mil99b] H. D. Mills. The management of software engineering, part I: Principled of software engineering. IBM Systems Journal, 38(2–3):289–296, ????

1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/382/mills.pdf. Reprinted

from [Mil80].

Mills:2002:MSV

[Mil02] Steve Mills. Message from the Senior Vice President and Group Executive, IBM Software Group. IBM Systems Journal, 41(2):??, ???? 2002. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/412/execletter.html.

Miles:2008:PIS

[Mil08] I. Miles. Patterns of innovation in service industries. IBM Systems Journal, 47(1):115– ??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/471/miles.html. See er-

ratum [Ano08b] with corrected Figure 1.

Minkoff:1992:SAO

[Min92] A. S. Minkoff. A systematic approach to OSL application programming. IBM Systems Journal, 31(1):49–61, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5460.

Misra:1981:CAM

[Mis81] P. N. Misra. Capacity analysis of the Mass Storage System. IBM Systems Journal, 20(3):346–360, 1981. CODEN IBMSA7. ISSN 0018-8670.

Misra:1983:SRA

[Mis83] P. N. Misra. Software reliability analysis. IBM Systems Journal, 23(3):262–270, 1983. CODEN IBMSA7. ISSN 0018-8670.

Mizell:1988:UDD

[Miz88] A. M. Mizell. Understanding device drivers in Operating System/2. IBM Systems Journal, 27(2):170–184, June 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Mays:1990:EDP

[MJHS90] R. G. Mays, C. L. Jones, G. J.

Holloway, and D. P. Studinski. Experiences with Defect Prevention. IBM Systems Journal, 29(1):4–32, 1990. CODEN IBMSA7. ISSN 0018-8670.

McKay:1973:PCN

[MK73] D. B. McKay and D. P. Karp.

Protocol for a computer network. IBM Systems Journal, 12(1):94–105, 1973. CODEN IBMSA7. ISSN 0018-8670.

Mushlin:2007:GTA

[MKGR07] R. A. Mushlin, A. Kershenbaum, S. T. Gallagher, and T. R. Rebbeck. A graphtheoretical approach for pattern discovery in epidemiological research. IBM Systems Journal, 46(1):135–149, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/461/mushlin.html.

Matick:1984:AID

[ML84] Richard E. Matick and Daniel T. Ling. Architecture implications in the design of microprocessors. IBM Systems Journal, 23(3):264–280, 1984. CODEN IBMSA7. ISSN 0018-8670.

Mendelsohn:1983:RVF

[MLA83] Noah Mendelsohn, Mark H. Linehan, and William J. Anzick. Reflections on VM/Pass-Through: a facility for interactive networking. IBM Systems Journal, 23(1-2):63–79, 1983. CODEN IBMSA7. ISSN 0018-8670.

Matyas:1991:KSB

[MLA91] S. M. Matyas, A. V. Le, and D. G. Abraham. A keymanagement scheme based on control vectors. IBM Systems Journal, 30(2):175–191, 1991. CODEN IBMSA7. ISSN 0018-8670.

McLaughlin:2008:IPS

[MLDF08] Grant T. McLaughlin, Leo Y. Liu, Daniel J. DeGroff, and Kenneth W. Fleck. IBM Power Systems platform: Advancements in the state of the art in IT availability. IBM Systems Journal, 47(4):519–533, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/mclaughlin.pdf.

Mills:1987:BSI

[MLH87] H. D. Mills, R. C. Linger, and A. R. Hevner. Box structured information systems. IBM Systems Journal, 26(4):395– 413, December 1987. CODEN IBMSA7. ISSN 0018-8670.

Markl:2003:LAQ

[MLR03] V. Markl, G. M. Lohman, and V. Raman. LEO: An autonomic query optimizer for DB2. IBM Systems Journal, 42(1): 98–106, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/421/markl.pdf.

Matyas:1978:GDI

[MM78] Stephen M. Matyas and Carl H.

Meyer. Generation, distribution, and installation of cryptographic keys. IBM Systems Journal, 17(2):126–137, 1978. CODEN IBMSA7. ISSN 0018-8670.

Matthews:1990:DMS

[MM90] R. W. Matthews and W. C.

McGee. Data modeling for software development. IBM Systems Journal, 29(2):228–235, 1990. CODEN IBMSA7. ISSN 0018-8670.

Moreira:2000:JPH

[MMG^{+}00] J. E. Moreira, S. P. Midkiff, M. Gupta, P. V. Artigas, M. Snir, and R. D. Lawrence. Java programming for high-performance numerical computing. IBM Systems Journal, 39(1):21–56, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/391/moreira.html.

Mercurio:1990:ACS

[MMNR90] V. J. Mercurio, B. F. Meyers, A. M. Nisbet, and G. Radin. AD/Cycle strategy and architecture. IBM Systems Journal, 29(2):170–188, 1990. CODEN IBMSA7. ISSN 0018-8670.

Midkiff:1998:OAR

[MMS98] S. P. Midkiff, J. E. Moreira, and M. Snir. Optimizing array reference checking in Java programs. IBM Systems Journal, 37(3):409–453, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

                                        tronic). URL http://www.
                                        research.ibm.com/journal/
                                        sj/373/midkiff.html.

Martin:2000:MDK

[MMS00] F. Martin, B. Mikhak, and B. Silverman. MetaCricket: a designer’s kit for making computational devices. IBM Systems Journal, 39(3/4):795– 815, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/martin.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part2/
                                        martin.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/martin.txt.

McCarthy:2002:ADC

[MMS^{+}02] J. McCarthy, M. Minsky, A. Sloman, L. Gong, T. Lau, L. Morgenstern, E. T. Mueller, D. Riecken, M. Singh, and P. Singh. An architecture of diversity for commonsense reasoning. IBM Systems Journal, 41(3):530–539, ???? 2002. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/forum.html#part2.

Mack:2004:TAL

[MMS^{+}04] R. Mack, S. Mukherjea, A. Soffer, N. Uramoto, E. Brown, A. Coden, J. Cooper, A. Inokuchi, B. Iyer, Y. Mass, H. Matsuzawa, and L. V. Subramaniam. Text analytics for life science using the Unstructured Information Management Architecture. IBM Systems Journal, 43(3):490–515, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/mack.html;   http:
//www.research.ibm.com/journal/
sj/433/mack.pdf;   http:/
/www.research.ibm.com/journal/
sj/433/mack.txt.

Myles:2006:CPG

[MN06] G. Myles and S. Nusser. Content protection for games. IBM Systems Journal, 45(1): 119–143, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/451/myles.html.

Malaika:2002:DWS

[MNQ^{+}02] S. Malaika, C. J. Nelin, R. Qu, B. Reinwald, and D. C. Wolfson. DB2 and Web services. IBM Systems Journal, 41(4): 666–685, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/414/malaika.pdf.

Morris:1990:ISC

[MO90] H. M. Morris and R. H. Orth. Image system communications. IBM Systems Journal, 29(3):371–383, 1990. CODEN IBMSA7. ISSN 0018-8670.

Magid:2008:GRT

[MOB^{+}08] Y. Magid, D. Oren, D. Botzer, A. Adi, B. Shulman, E. Rabinovich, and M. Barnea. Generating real-time complex eventprocessing applications. IBM Systems Journal, 47(2):251– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/472/magid.html.

Mays:1985:PRM

[MOCP85] R. G. Mays, L. S. Orzech, W. A. Ciarfella, and R. W. Phillips. PDM: a requirements methodology for software system enhancements. IBM Systems Journal, 24(2):134–149, 1985. CODEN IBMSA7. ISSN 0018-8670.

McCrossin:1978:TDT

[MOK78] James M. McCrossin, Robert P. O’Hara, and L. R. Koster. Time-sharing display terminal session manager. IBM Systems Journal, 17(3):260–275, 1978. CODEN IBMSA7. ISSN 0018-8670.

Montalbano:1962:TFC

[Mon62] M. Montalbano. Tables, flow charts and program logic. IBM Systems Journal, 1:51– 63, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Montalbano:1967:HCC

[Mon67] M. Montalbano. High-speed calculation of the critical paths of large networks. IBM Systems Journal, 6(3):163–191, 1967. CODEN IBMSA7. ISSN 0018-8670.

Moore:1979:ELC

[Moo79] Roy S. Moore. Evolution of a laboratory communication network. IBM Systems Journal, 18(2):315–332, 1979. CODEN IBMSA7. ISSN 0018-8670.

Moore:1983:TWL

[Moo83] Daniel J. Moore. Teletex — a worldwide link among office systems for electronic document exchange. IBM Systems Journal, 23(1-2):30–45, 1983. CODEN IBMSA7. ISSN 0018-8670.

Moore:1988:USA

[Moo88] R. E. Moore. Utilizing the SNA alert in the management of multivendor networks. IBM Systems Journal, 27(1): 15–31, January 1988. CODEN IBMSA7. ISSN 0018-8670.

Moore:1990:MPD

[Moo90]  D. J. Moore.  Multimedia

presentation development using the Audio Visual Connection. IBM Systems Journal, 29(4):494–508, 1990. CODEN IBMSA7. ISSN 0018-8670.

Morrison:1973:UPP

[Mor73] J. E. Morrison. User program performance in virtual storage systems. IBM Systems Journal, 12(3):216–237, 1973. CODEN IBMSA7. ISSN 0018-8670.

Morrison:1978:DSL

[Mor78] J. Paul Morrison. Data stream linkage mechanism. IBM Systems Journal, 17(4):383–408, 1978. CODEN IBMSA7. ISSN 0018-8670. See letters [Lev79, Mor79].

Morrison:1979:FAR

[Mor79] J. Paul Morrison. Forum: Author’s response. IBM Systems Journal, 18(3):473–475, 1979. CODEN IBMSA7. ISSN 0018-8670. See [Mor78, Lev79].

Meliksetian:2004:DIE

[MPB^{+}04] D. S. Meliksetian, J.-P. Prost, A. S. Bahl, I. Boutboul, D. P. Currier, S. Fibra, J.-Y. Girard, K. M. Kassab, J.-L. Lepesant, C. Malone, and P. Manesco. Design and implementation of an enterprise grid. IBM Systems Journal, 43 (4):646–664, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/

sj/434/meliksetian.html;
http://www.research.ibm.
com/journal/sj/434/meliksetian.
pdf;   http://www.research.
ibm.com/journal/sj/434/meliksetian.
txt.

Meier:1997:DDC

[MPF97] M. Meier, H. Pan, and G. Y.

Fuh. Debugging DB2/CS client/server applications. IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/361/meier.html.

McGee:1968:IGD

[MPH68] W. C. McGee, H. R. Penafiel, and S. K. Howry. Interactive graphics in data processing: Analysis and display of physics data. IBM Systems Journal, 7(3/4):342–354, 1968. CODEN IBMSA7. ISSN 0018-8670.

Menon:2003:IST

[MPR^{+}03] J. Menon, D. A. Pease, R. Rees, L. Duyanovich, and B. Hillsberg. IBM Storage Tank — A heterogeneous scalable SAN file system. IBM Systems Journal, 42(2):250–267, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/menon.pdf.

Margolin:1971:AFA

[MPS71] B. H. Margolin, R. P. Parmelee, and M. Schatzoff. Analysis of free-storage algorithms. IBM Systems Journal, 10(4):283– 304, 1971. CODEN IBMSA7. ISSN 0018-8670.

Mohan:1994:PRD

[MPTW94] C. Mohan, H. Pirahesh, W. G. Tang, and Y. Wang. Parallelism in relational database management systems. IBM Systems Journal, 33(2):349– 371, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://

www.research.ibm.com/journal/
sj33-2.html#eight. G321-

5546.

Martin:2005:ILM

[MQR05] K. Martin, M. A. Quigley, and S. Rogers. Implementing a learning management system globally: An innovative change management approach. IBM Systems Journal, 44(1): 125–143, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/441/martin.html;   http:
//www.research.ibm.com/journal/
sj/441/martin.pdf;   http:
//www.research.ibm.com/journal/
sj/441/martin.txt.

Mack:2001:KPE

[MRB01] R. Mack, Y. Ravin, and R. J. Byrd. Knowledge portals and the emerging digital knowledge workplace. IBM Systems Journal, 40(4):925–955, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/404/mack.html;   http:

//www.research.ibm.com/journal/
sj/404/mack.pdf.

Meyer:1970:VMT

[MS70] R. A. Meyer and L. H. Seawright. A virtual machine timesharing system. IBM Systems Journal, 9(3):199–218, 1970. CODEN IBMSA7. ISSN 0018-8670.

Metz:1975:SPS

[MS75] Walter C. Metz, Jr. and David Savir. Store performance studies for the Supermarket System. IBM Systems Journal, 14(1): 46–63, 1975. CODEN IBMSA7. ISSN 0018-8670.

McManigal:1980:AIG

[MS80] D. F. McManigal and D. A.

Stevenson. Architecture of the IBM 3277 graphics attachment. IBM Systems Journal, 19(3):331–344, 1980. CODEN IBMSA7. ISSN 0018-8670.

Mudie:1985:ITA

[MS85] M. W. Mudie and D. J. Schafer.

Information technology architecture for change. IBM Systems Journal, 24(3-4):307–315, 1985. CODEN IBMSA7. ISSN 0018-8670.

McComb:1988:ESS

[MS88] J. McComb and S. Schmidt.

Engineering and Scientific Subroutine Library for the IBM 3090 Vector Facility. IBM Systems Journal, 27(4):404– 415, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Madduri:2007:CMD

[MSB^{+}07] H. Madduri, S. S. B. Shi, R. Baker, N. Ayachitula, L. Shwartz, M. Surendra, C. Corley, M. Benantar, and S. Patel. A configuration management database architecture in support of IBM Service Management. IBM Systems Journal, 46(3):441–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/madduri.html.

Morrill:2008:P

[MSR08] Harriet Morrill, David Seidman, and John J. Ritsko. Preface. IBM Systems Journal, 47(4):491, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/preface.pdf.

Maslak:1991:CRR

[MSS91] B. A. Maslak, J. M. Showalter, and T. J. Szczygielski. Coordinated Resource Recovery in VM/ESA. IBM Systems Journal, 30(1):72–89, 1991. CO-DEN IBMSA7. ISSN 0018-8670.

Maglio:2008:P

[MSSR08] P. P. Maglio, J. Spohrer, D. I. Seidman, and J. J. Ritsko. Preface. IBM Systems Journal, 47 (1):3–??, January/March 2008. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.

research.ibm.com/journal/
sj/471/preface.html.

Meritt:2003:ZSI

[MST^{+}03] A. S. Meritt, J. A. Staubi, K. M. Trowell, G. Whistance, and H. M. Yudenfriend. z/OS support of the IBM TotalStorage Enterprise Storage Server. IBM Systems Journal, 42(2): 280–301, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/422/meritt.pdf.

Mangalam:2001:GOS

[MSZ^{+}01] H. Mangalam, J. Stewart, J. Zhou, K. Schlauch, M. Waugh, G. Chen, A. D. Farmer, G. Colello, and J. W. Weller. GeneX: An Open Source gene expression database and integrated tool set. IBM Systems Journal, 40(2):552–569, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/mangalam.html;   http:
//www.research.ibm.com/journal/
sj/402/mangalam.pdf;   http:
//www.research.ibm.com/journal/
sj/402/mangalam.txt.

Morris:2003:ESS

[MT03] R. J. T. Morris and B. J.

Truskowski. The evolution of storage systems. IBM Systems Journal, 42(2):205–217, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/422/preface.pdf.

Mueller:1967:AGR

[Mue67] J. H. Mueller. Aspects of the Gemini real-time operating system. IBM Systems Journal, 6(3):150–162, 1967. CODEN IBMSA7. ISSN 0018-8670.

Mugele:1962:POC

[Mug62] R. A. Mugele. A program for optimal control of nonlinear processes. IBM Systems Journal, 1:2–17, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Murray:1984:SCP

[Mur84] William H. Murray. Security considerations for personal computers. IBM Systems Journal, 23(3):297–304, 1984. CO-DEN IBMSA7. ISSN 0018-8670.

Musgrave:1978:ECG

[Mus78] Joan Forrester Musgrave. Experiments in computer-aided graphic expression. IBM Systems Journal, 17(3):241–259, 1978. CODEN IBMSA7. ISSN 0018-8670.

Mishelevich:1980:ADS

[MV80] David J. Mishelevich and Don Van Slyke. Application development system — the software architecture of the IBM Health Care Support/DL/Ipatient care system. IBM Systems Journal, 19(4):478–504, 1980. CODEN IBMSA7. ISSN 0018-8670.

Margopoulos:1966:TSD

[MW66] W. P. Margopoulos and R. J.

Williams. On teleprocessing system design, part I: Characteristic problems. IBM Systems Journal, 5(3):134–141, 1966. CODEN IBMSA7. ISSN 0018-8670.

Myers:1976:CDF

[Mye76] Glenford J. Myers. Composite design facilities of six programming languages. IBM Systems Journal, 15(3):212–224, 1976. CODEN IBMSA7. ISSN 0018-8670.

Newbold:2007:MCL

[NA07] D. L. Newbold and M. C. Azua.

A model for CIO-led innovation. IBM Systems Journal, 46(4):??, October/December 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/464/newbold.html.

Naeger:1994:BDI

[Nae94] R. D. Naeger. Books: Designing Interactive Multimedia by Arch C. Luther. IBM Systems Journal, 33(1):223–224, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-2.html.

Nahouraii:1974:DDS

[Nah74] E. Nahouraii. Direct-access device simulation. IBM Systems Journal, 13(1):19–31, 1974. CODEN IBMSA7. ISSN 0018-

Naik:1995:SIN

[Nai95] V. K. Naik. A scalable implementation of the NAS Parallel Benchmark BT on distributed memory systems. IBM Systems Journal, 34(2):273–291, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-2.html#nine.

Nigam:2003:BAA

[NC03] A. Nigam and N. S. Caswell. Business artifacts: An approach to operational specification. IBM Systems Journal, 42 (3):428–445, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/nigam.pdf.

Negroponte:1996:AS

[Neg96] N. Negroponte. Asparagus soup. IBM Systems Journal, 35 (3&4):266–268, 1996. CODEN IBMSA7. ISSN 0018-8670.

Negroponte:2000:BDD

[Neg00] N. Negroponte. From being digital to digital beings. IBM Systems Journal, 39(3/4): 417–418, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/negroponte.html;
http://www.research.ibm.
com/journal/sj/393/negroponte.
pdf;   http://www.research.
ibm.com/journal/sj/393/negroponte.
txt.

Neuhold:1971:FDP

[Neu71] E. J. Neuhold. The formal description of programming languages. IBM Systems Journal, 10(2):86–112, 1971. CODEN IBMSA7. ISSN 0018-8670.

Newman:1982:TID

[New82] P. S. Newman. Towards an integrated development environment. IBM Systems Journal, 21(1):81–107, 1982. CODEN IBMSA7. ISSN 0018-8670.

Newton:1985:SPP

[New85] J. Newton. Strategies for problem prevention. IBM Systems Journal, 24(3-4):248–263, 1985. CODEN IBMSA7. ISSN 0018-8670.

Niblett:2005:ESO

[NG05] P. Niblett and S. Graham.

Events and service-oriented architecture: The OASIS Web Services Notification specifications. IBM Systems Journal, 44(4):869–886, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/444/niblett.html.

Niehoff:1980:AAP

[NJ80] W. H. Niehoff and A. L. Jones.

An APL approach to presentation graphics. IBM Systems Journal, 19(3):367–381, 1980. CODEN IBMSA7. ISSN 0018-8670.

Nayak:2007:CBA

[NLN^{+}07] N. Nayak, M. Linehan, A. Nigam, D. Marston, J.-J. Jeng, F. Y.

Wu, D. Boullery, L. F. White, P. Nandi, and J. L. C. Sanz. Core business architecture for a service-oriented enterprise. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/464/nayak.html.

Naik:2004:ACS

[NMB04] V. K. Naik, A. Mohindra, and D. F. Bantz. An architecture for the coordination of system management services. IBM Systems Journal, 43(1): 78–96, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/431/naik.pdf.

Nick:1997:CTP

[NMCB97] J. M. Nick, B. B. Moore, J.- Y. Chung, and N. S. Bowen. S/390 cluster technology: Parallel Sysplex. IBM Systems Journal, 36(2):172–201, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/362/nick.html.

Nasukawa:2001:TAK

[NN01] T. Nasukawa and T. Nagano. Text analysis and knowledge mining system. IBM Systems Journal, 40(4):967–984, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/

sj/404/nasukawa.html;   http:
//www.research.ibm.com/journal/
sj/404/nasukawa.pdf.

Nagaratnam:2005:BDA

[NNH^{+}05] N. Nagaratnam, A. Nadalin, M. Hondo, M. McIntosh, and P. Austel. Business-driven application security: From modeling to managing secure applications. IBM Systems Journal, 44(4):847–867, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/nagaratnam.html.

Newitt:1970:SRS

[NO70] J. W. Newitt and A. Odarchenko.

A structure for real-time Stenotype transcription. IBM Systems Journal, 9(1):24–35, 1970. CODEN IBMSA7. ISSN 0018-8670.

Norden:1993:QTS

[Nor93] P. V. Norden. Quantitative techniques in strategic alignment. IBM Systems Journal, 32(1):180–197, 1993. CODEN IBMSA7. ISSN 0018-8670.

Norman:2003:AMD

[NOR03] D. A. Norman, A. Ortony, and D. M. Russell. Affect and machine design: Lessons for the development of autonomous machines. IBM Systems Journal, 42(1):38–44, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/norman.pdf.

Nguyen:1980:RDS

[NORS80] Hung C. Nguyen, Arnold Ockene, Richard Revell, and Walter J. Skwish. The role of detailed simulation in capacity planning. IBM Systems Journal, 19(1):81–101, 1980. CO-DEN IBMSA7. ISSN 0018-8670.

Neal:1984:PME

[NS84] Alan S. Neal and Roger M. Simons. Playback: a method for evaluating the usability of software and its documentation. IBM Systems Journal, 23(1): 82–96, 1984. CODEN IBMSA7. ISSN 0018-8670.

Neus:2005:OMC

[NS05] A. Neus and P. Scherf. Opening minds: Cultural change with the introduction of opensource collaboration methods. IBM Systems Journal, 44(2): 215–225, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/neus.html;   http:
//www.research.ibm.com/journal/
sj/442/neus.pdf;   http:/
/www.research.ibm.com/journal/
sj/442/neus.txt.

Neti:2006:P

[NSS^{+}06] Chalapathy Neti, Anees Shaikh, Chris Sharp, Randy Moulic, Patty Fry, Dick Anderson, John J. Ritsko, and David I. Seidman. Preface. IBM Systems Journal, 45(1):3–6, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/451/preface.html.

Oshisanwo:1992:ROS

[OCK^{+}92] A. O. Oshisanwo, M. D. Chapman, M. Key, A. P. Mullery, and J. Saint-Blancat. The RACE open services architecture project. IBM Systems Journal, 31(4):691–710, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5493.

Ozcan:2006:ISX

[ÖCKM06] F. Özcan, D. Chamberlin, K. Kulkarni, and J.-E. Michels. Integration of SQL and XQuery in IBM DB2. IBM Systems Journal, 45(2):245–270, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/452/ozcan.html.

Odom:1990:UBS

[Odo90] J. E. Odom. Using box structures for definition of requirements specifications. IBM Systems Journal, 29(1):59–78, 1990. CODEN IBMSA7. ISSN 0018-8670.

Ohms:1986:CPD

[Ohm86] B. G. Ohms. Computer processing of dates outside the Twentieth Century. IBM Systems Journal, 25(2):244–251, 1986. CODEN IBMSA7. ISSN 0018-8670.

Ohara:2006:MMP

[OIS^{+}06] M. Ohara, H. Inoue, Y. Sohda, H. Komatsu, and T. Nakatani.

MPI microtask for programming the Cell Broadband Engine^{T M} processor. IBM Systems Journal, 45(1):85– 102, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/451/ohara.html.

Osisek:1991:EIA

[OJG91] D. L. Osisek, K. M. Jackson, and P. H. Gum. ESA/390 interpretive-execution architecture, foundation for VM/ESA. IBM Systems Journal, 30(1): 34–51, 1991. CODEN IBMSA7. ISSN 0018-8670.

Okaya:1968:IGD

[Oka68] Y. Okaya. Interactive graphics in data processing: Interactive aspects of crystal structure analysis. IBM Systems Journal, 7(3/4):322–330, 1968. CODEN IBMSA7. ISSN 0018-8670.

Olund:2007:BIM

[ÖLL07] G. Ölund, P. Lindqvist, and J.-E. Litton. BIMS: An information management system for biobanking in the 21st Century. IBM Systems Journal, 46(1):171–182, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/461/olund.html.

Ominsky:1981:HFD

[Omi81] M. Ominsky. Human factors in the development of a family of plant data communications. IBM Systems Journal, 20(2):216–236, 1981. CODEN IBMSA7. ISSN 0018-8670.

ONeill:1980:MSE

[O’N80] D. O’Neill. The management of software engineering, part II: Software engineering program. IBM Systems Journal, 19(4):421–431, 1980. CODEN IBMSA7. ISSN 0018-8670.

Omojola:2000:IIF

[OPH^{+}00] O. Omojola, E. R. Post, M. D. Hancher, Y. Maguire, R. Pappu, B. Schoner, P. R. Russo, R. Fletcher, and N. Gershenfeld. An installation of interactive furniture. IBM Systems Journal, 39(3/4):861– 879, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part3/omojola.html;
 http://www.research.ibm.
com/journal/sj/393/part3/
omojola.pdf;   http://www.
research.ibm.com/journal/
sj/393/part3/omojola.txt.

Oblinger:2002:CTA

[ORBdSB02] D. A. Oblinger, M. Reid, M. Brodie, and R. de Salvo Braz. Cross training and its application to skill mining. IBM Systems Journal, 41(3):449– 460, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/oblinger.html.

Orlicky:1973:NCM

[Orl73] Joseph A. Orlicky. Net change material requirements planning. IBM Systems Journal, 12(1):2–29, 1973. CODEN IBMSA7. ISSN 0018-8670.

Orwant:1996:WBU

[Orw96] J. Orwant. For want of a bit the user was lost: Cheap user modeling. IBM Systems Journal, 35 (3&4):398–416, 1996. CODEN IBMSA7. ISSN 0018-8670.

Orwant:2000:EAP

[Orw00] J. Orwant. EGGG: Automated programming for game generation. IBM Systems Journal, 39(3/4):782–794, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part2/orwant.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part2/
                                        orwant.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/orwant.txt.

Onvural:1995:EAN

[OSK^{+}95] R. O. Onvural, H. J. Sandick, G. M. Kump, E. A. Hervatic, and P. F. Chimento. Enabling ATM networks. IBM Systems Journal, 34(4):672–693, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj34-4.html#nine.

OHare:1994:RAS

[OT94] A. B. O’Hare and E. W. Troan. RE-Analyzer: From source code to structured analysis. IBM Systems Journal, 33(1):110–130, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-1.html#seven.

Owen:1995:NNM

[Owe95]  S. A. Owen.  NBBS net-

work management. IBM Systems Journal, 34(4):725–750, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#twelve.

Padegs:1964:SSP

[Pad64] A. Padegs. The structure of SYSTEM/360, part IV: Channel design considerations. IBM Systems Journal, 3(2):165–180, 1964. CODEN IBMSA7. ISSN 0018-8670.

Padegs:1968:SAS

[Pad68] A. Padegs. Structural aspects of the SYSTEM/360 model 85, part III: Extensions to floatingpoint architecture. IBM Systems Journal, 7(1):22–29, 1968. CODEN IBMSA7. ISSN 0018-8670.

Palmer:2001:EH

[Pal01] C. C. Palmer. Ethical hacking. IBM Systems Journal, 40 (3):769–780, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/palmer.html;   http:
//www.research.ibm.com/journal/
sj/403/palmer.pdf.

Paleologo:2004:PRM

[Pal04] G. A. Paleologo. Price-at-Risk: a methodology for pricing utility computing services. IBM Systems Journal, 43(1): 20–31, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/431/paleologo.pdf.

Papert:2000:WBI

[Pap00] S. Papert. What’s the big idea? toward a pedagogy of idea power. IBM Systems Journal, 39(3/4):720–729, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part2/papert.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part2/
                                        papert.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/papert.txt.

Parker:1982:EIA

[Par82] M. M. Parker. Enterprise information analysis: Cost-benefit analysis and the data-managed system. IBM Systems Journal, 21(1):108–123, 1982. CODEN IBMSA7. ISSN 0018-8670.

Paradiso:1996:IBS

[Par96] J. A. Paradiso. The interactive balloon: Sensing, actuation and behavior in a common object. IBM Systems Journal, 35 (3&4):473–487, 1996. CODEN IBMSA7. ISSN 0018-8670.

Paul:2007:PBP

[Pau07] C. J. Paul. The process of building a Process Manager: Architecture and design patterns. IBM Systems Journal, 46(3):479–??, July/September 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/463/paul.html.

Pazel:1989:DIG

[Paz89] D. P. Pazel. DS-Viewer—an interactive graphical data structure presentation facility. IBM Systems Journal, 28(2):307– 323, 1989. CODEN IBMSA7. ISSN 0018-8670.

Plesums:1990:LIS

[PB90] C. A. Plesums and R. W. Bartels. Large-scale image systems: USAA case study. IBM Systems Journal, 29(3):343–355, 1990. CODEN IBMSA7. ISSN 0018-8670.

Peyravian:1995:ETD

[PBCK95] M. Peyravian, R. Bodner, C.- S. Chow, and M. Kaplan. Efficient transport and distribution of network control information in NBBS. IBM Systems Journal, 34(4):640–658, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#seven.

Parsons:1997:PPD

[PBS97] E. W. Parsons, M. Brorsson, and K. C. Sevcik. Predicting the performance of distributed virtual shared-memory applications. IBM Systems Journal, 36(4):527–549, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/364/parsons.html.

Prins:1997:FTB

[PBv97] R. Prins, A. Blokdijk, and N. E. van Oosterom. Family traits in business objects and their applications. IBM Systems Journal, 36(1):??, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/361/prins.html.

Peng:1998:AEB

[PCC^{+}98] C.-S. Peng, S.-K. Chen, J.- Y. Chung, A. Roy-Chowdhury, and V. Srinivasan. Accessing existing business data from the World Wide Web. IBM Systems Journal, 37(1):115–132, 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/371/peng.html.

Pistoia:2007:SSA

[PCFY07] M. Pistoia, S. Chandra, S. J. Fink, and E. Yahav. A survey of static analysis methods for identifying security vulnerabilities in software systems. IBM Systems Journal, 46(2): 265–??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/462/pistoia.html.

Poulin:1993:BCS

[PCH93] J. S. Poulin, J. M. Caruso, and D. R. Hancock. The business case for software reuse. IBM Systems Journal, 32(4):567– 594, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5525.

Pinhanez:2000:PIS

[PDI^{+}00] C. S. Pinhanez, J. W. Davis, S. Intille, M. P. Johnson, A. D. Wilson, A. F. Bobick, and B. Blumberg. Physically interactive story environments. IBM Systems Journal, 39(3/4): 438–455, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/pinhanez.html;
 http://www.research.ibm.
com/journal/sj/393/part1/
pinhanez.pdf;   http://www.
research.ibm.com/journal/
sj/393/part1/pinhanez.txt.

Pearsall:1982:TAE

[Pea82] Ronald J. Pearsall. Technique for assessing external design of software. IBM Systems Journal, 21(2):211–219, 1982. CO-DEN IBMSA7. ISSN 0018-8670.

Peck:1985:WSE

[Pec85] T. G. Peck. Worldwide systems engineering. IBM Systems Journal, 24(3-4):182–188, 1985. CODEN IBMSA7. ISSN 0018-

Pentland:2000:A

[Pen00] A. Pentland. It’s alive! IBM Systems Journal, 39(3/4):821– 822, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/393/part3/pentland.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part3/
                                        pentland.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part3/pentland.txt.

Perna:2002:MGM

[Per02] Janet Perna. Message from the general manager, data management solutions, IBM software group. IBM Systems Journal, 41(4):??, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/414/message.pdf.

Ponzo:2005:EIW

[PG05] J. Ponzo and O. Gruber. Eclipse: Integrating Web technologies. IBM Systems Journal, 44(2):279–288, ???? 2005. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/442/ponzo.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/ponzo.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/ponzo.txt.

Papetti:1992:ICE

[PGM^{+}92] F. Papetti, S. Golini, M. Maggiore, S. Succi, P. Gaillard, and J.-M. Perez. Internal combustion engine design on IBM platforms. IBM Systems Journal, 31(4):774–787, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5497.

Plachy:1999:ESS

[PH99] E. C. Plachy and P. A. Hausler.

Enterprise solutions structure. IBM Systems Journal, 38(1): 4–11, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/381/plachy.html.

Parise:2001:KRE

[PH01] S. Parise and J. C. Henderson. Knowledge resource exchange in strategic alliances. IBM Systems Journal, 40(4): 908–924, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/404/parise.html;   http:
//www.research.ibm.com/journal/
sj/404/parise.pdf.

Paradiso:2000:DIE

[PHBT00] J. A. Paradiso, K. Hsiao, A. Y. Benbasat, and Z. Teegarden. Design and implementation of expressive footwear. IBM Systems Journal, 39(3/4): 511–529, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/paradiso.html;
 http://www.research.ibm.
com/journal/sj/393/part1/

                                        paradiso.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part1/paradiso.txt.

Ponzo:2004:DWC

[PHG^{+}04] J. Ponzo, L. D. Hasson, J. George, G. Thomas, O. Gruber, R. Konuru, A. Purakayastha, R. D. Johnson, J. Colson, and R. A. Pollak. On demand Web-client technologies. IBM Systems Journal, 43(2):297–315, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/432/ponzo.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/ponzo.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/ponzo.txt.

Paradiso:2000:SSI

[PHS^{+}00] J. A. Paradiso, K. Hsiao, J. Strickon, J. Lifton, and A. Adler. Sensor systems for interactive surfaces. IBM Systems Journal, 39(3/4):892– 914, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/393/part3/paradiso.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part3/
                                        paradiso.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part3/paradiso.txt.

Picard:1996:SMV

[Pic96] R. W. Picard. A society of models for video and image libraries. IBM Systems Journal, 35(3&4):292–312, 1996. CO-DEN IBMSA7. ISSN 0018-8670.

Picard:2000:TCR

[Pic00] R. W. Picard. Toward computers that recognize and respond to user emotion. IBM Systems Journal, 39(3/4):705–719, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/393/part2/picard.html;
 http://www.research.ibm.
com/journal/sj/393/part2/
picard.pdf;   http://www.
research.ibm.com/journal/
sj/393/part2/picard.txt.

Pierce:1968:MEL

[Pie68] J. F. Pierce. A multi-item economic lot-sizing problem. IBM Systems Journal, 7(1):47–66, 1968. CODEN IBMSA7. ISSN 0018-8670.

Pine:1989:DTV

[Pin89] B. J. Pine, II. Design, test, and validation of the Application System/400 through early user involvement. IBM Systems Journal, 28(3):376–385, 1989. CODEN IBMSA7. ISSN 0018-8670.

Percival:1986:NMS

[PJ86] L. C. Percival and S. K. Johnson. Network management software usability test design and implementation. IBM Systems Journal, 25(1):92–104, 1986. CODEN IBMSA7. ISSN 0018-

Plambeck:1989:CES

[Pla89] K. E. Plambeck. Concepts of Enterprise Systems Architecture/370. IBM Systems Journal, 28(1):39–61, 1989. CO-DEN IBMSA7. ISSN 0018-8670.

Prager:1990:RIU

[PLGB90] J. M. Prager, D. M. Lamberti, D. L. Gardner, and S. R. Balzac. REASON: An intelligent user assistant for interactive environments. IBM Systems Journal, 29(1):141–164, 1990. CODEN IBMSA7. ISSN 0018-8670.

Perkins:2006:GEP

[PMK^{+}06] E. Perkins, M. Matsa, M. G. Kostoulas, A. Heifets, and N. Mendelsohn. Generation of efficient parsers through direct compilation of XML Schema grammars. IBM Systems Journal, 45(2):225–244, ???? 2006. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/452/perkins.html.

Pazel:1983:SAE

[PMM83] D. P. Pazel, A. Malhotra, and H. M. Markowitz. The system architecture of EAS-E: An integrated programming and data base language. IBM Systems Journal, 23(3):188–198, March 1983. CODEN IBMSA7. ISSN 0018-8670.

Polan:1998:TNU

[Pol98] M. G. Polan. Technical note:

Using the San Francisco frameworks with VisualAge for Java. IBM Systems Journal, 37(2): 215–226, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/372/polan.html.

Pomeroy:1972:GPT

[Pom72] J. W. Pomeroy. A guide to programming tools and techniques. IBM Systems Journal, 11(3):234–254, 1972. CODEN IBMSA7. ISSN 0018-8670.

Post:2000:BDF

[PORG00] E. R. Post, M. Orth, P. R.

Russo, and N. Gershenfeld. Ebroidery: Design and fabrication of textile-based computing. IBM Systems Journal, 39 (3/4):840–860, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part3/post.html;
http://www.research.ibm.
com/journal/sj/393/part3/
post.pdf;   http://www.research.
ibm.com/journal/sj/393/part3/
post.txt.

Power:1983:DUP

[Pow83] L. R. Power. Design and use of a program execution analyzer. IBM Systems Journal, 23(3):271–294, 1983. CODEN IBMSA7. ISSN 0018-8670.

Pohs:2001:LKD

[PPDW01] W. Pohs, G. Pinder, C. Dougherty, and M. White. The Lotus Knowledge Discovery System: Tools and experiences. IBM Systems Journal, 40(4): 956–966, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/404/pohs.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/404/pohs.pdf.

Parmelee:1972:VSV

[PPTH72] R. P. Parmelee, T. I. Peterson, C. C. Tillman, and D. J. Hatfield. Virtual storage and virtual machine concepts. IBM Systems Journal, 11(2):99–130, 1972. CODEN IBMSA7. ISSN 0018-8670.

Prager:1983:PAL

[Pra83] J. M. Prager. The Project Automated Librarian. IBM Systems Journal, 23(3):214–228, 1983. CODEN IBMSA7. ISSN 0018-8670.

Prusak:2001:WDK

[Pru01] L. Prusak. Where did knowledge management come from? IBM Systems Journal, 40(4): 1002–1007, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/404/prusak.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/404/prusak.pdf.

Perrone:2002:MLM

[PRZ02] M. P. Perrone, G. F. Russell, and A. Ziq. Machine learning in a multimedia document retrieval framework. IBM Systems Journal, 41(3): 494–503, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/perrone.html.

Parr:1983:NHL

[PS83] Francis N. Parr and Robert E.

Strom. NIL: a high-level language for distributed systems programming. IBM Systems Journal, 23(1-2):111–127, 1983. CODEN IBMSA7. ISSN 0018-8670.

Parr:1999:NHL

[PS99] F. N. Parr and R. E. Strom.

NIL: a high-level language for distributed systems programming. IBM Systems Journal, 38(2–3):181–198, ???? 1999. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/parr.pdf. Reprinted

from [PS83].

Perna:2004:I

[PS04] Janet Perna and Alfred Spector. Introduction. IBM Systems Journal, 43(3):??, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/intro.html;   http:
//www.research.ibm.com/journal/
sj/433/intro.pdf;   http:

//www.research.ibm.com/journal/
sj/433/intro.txt.

Perelgut:1997:OCA

[PSLB97] S. G. Perelgut, G. M. Silberman, K. A. Lyons, and K. L. Bennet. Overview: The Centre for Advanced Studies. IBM Systems Journal, 36(4):474–488, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/364/perelgut.html.

Pozefsky:1995:MTN

[PTE^{+}95] D. Pozefsky, R. Turner, A. K. Edwards, S. Sarkar, J. Mathew, G. Bollella, K. Tracey, D. Poirier, J. Fetvedt, W. S. Hobgood, W. A. Doeringer, and D. Dykeman. Multiprotocol Transport Networking: Eliminating application dependencies on communications protocols. IBM Systems Journal, 34(3):472–500, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-3.html#nine.

Potok:1997:EBM

[PV97] T. E. Potok and M. A. Vouk. The effects of the business model on object-oriented software development productivity. IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/361/potok.html.

Pitkanen:2008:LRT

[PVK08] O. Pitkänen, P. Virtanen, and J. Kemppinen. Legal research topics in user-centric services. IBM Systems Journal, 47(1):143–??, January/March 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/471/pitkanen.html.

Pinter:1991:PCR

[PVW91] R. Y. Pinter, P. Vortman, and Z. Weiss. Partial compilation of REXX. IBM Systems Journal, 30(3):312–321, 1991. CODEN IBMSA7. ISSN 0018-8670.

Prairie:1988:IDT

[PW88] P. L. Prairie and A. H. Weis.

IBM’s directions in technical computing. IBM Systems Journal, 27(4):393–403, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Perry:1990:EGI

[PWBK90] B. T. Perry, B. A. Wester, W. W. Baker, and J. F. Kemmis. Experience gained in implementing ImagePlus. IBM Systems Journal, 29(3):467– 488, 1990. CODEN IBMSA7. ISSN 0018-8670.

Petrie:2005:PIN

[PWF05] H. L. Petrie, G. Weber, and W. Fisher. Personalization, interaction, and navigation in rich multimedia documents for print-disabled users. IBM Systems Journal, 44(3):629–

635, ????  2005.  CO-

DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/443/petrie.html.

Paul:2007:P

[PWSR07] C. J. Paul, C. Ward, D. I. Seidman, and J. J. Ritsko. Preface. IBM Systems Journal, 46(3):373–??, July/September 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/463/preface.html.

Parr:2008:PRW

[PY08] F. N. Parr and L. Yusuf. Patterns for real-world-aware and real-time solutions. IBM Systems Journal, 47(2):335– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/parr.html.

Quarles:1970:TCP

[Qua70] D. R. Quarles. Trajectory control programs in support of Apollo missions. IBM Systems Journal, 9(1):12–23, 1970. CO-DEN IBMSA7. ISSN 0018-8670.

Quinnan:1980:MSE

[Qui80] R. E. Quinnan. The management of software engineering, part V: Software engineering management practices. IBM Systems Journal, 19(4):466– 477, 1980. CODEN IBMSA7. ISSN 0018-8670.

Rackmales:1996:BSO

[Rac96] Bernard A. Rackmales. Books:

Succeeding with Objects: Decision Frameworks for Project Management, by Adele Goldberg and Kenneth S. Rubin (ISBN 0-201-62878-3) reviewed. IBM Systems Journal, 35(2):257–258, 1996. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj35-2.html.

Radke:1966:MAM

[Rad66] C. E. Radke. Merge-sort analysis by matrix techniques. IBM Systems Journal, 5(4):226–247, 1966. CODEN IBMSA7. ISSN 0018-8670.

Radin:1996:OTP

[Rad96] G. Radin. Object technology in perspective. IBM Systems Journal, 35(2):124–127, 1996. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj35-2.html#two.

Ramanathan:2008:ITS

[RAK08] Sriram Ramanathan, Michael Alexander, and Gordon Kerr. The IBM telecommunications service delivery platform. IBM Systems Journal, 47(3):433– 443, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/473/ramanathan.pdf.

Rankin:2002:STA

[Ran02] C. Rankin. The Software Testing Automation Framework. IBM Systems Journal, 41(1): 126–139, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/411/rankin.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/411/rankin.pdf.

Rappa:2004:UBM

[Rap04] M. A. Rappa. The utility business model and the future of computing services. IBM Systems Journal, 43(1):32–42, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/431/rappa.pdf.

Raz:1993:IPM

[Raz93] T. Raz. Introduction of the project management discipline in a software development organization. IBM Systems Journal, 32(2):265–277, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5513.

Ritsko:2001:Pb

[RB01a] John J. Ritsko and Marilyn L. Bates. Preface. IBM Systems Journal, 40(3):612–613, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/403/preface.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/403/preface.pdf.

Ritsko:2001:Pc

[RB01b] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 40(4):812–813, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/404/preface.html;   http:
//www.research.ibm.com/journal/
sj/404/preface.pdf.

Ritsko:2002:Pc

[RB02a] John J. Ritsko and Marilyn Bates. Preface. IBM Systems Journal, 41(4):553–554, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/414/preface.pdf.

Ritsko:2002:Pa

[RB02b] John J. Ritsko and Marilyn L.

Bates. Preface. IBM Systems Journal, 41(1):2–3, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/411/preface.html;   http:
//www.research.ibm.com/journal/
sj/411/preface.pdf.

Ritsko:2004:Pa

[RB04a] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 43(1):3–4, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/431/preface.pdf.

Ritsko:2004:Pd

[RB04b] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 43(4):622–623, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/434/preface.html;   http:
//www.research.ibm.com/journal/
sj/434/preface.pdf;   http:
//www.research.ibm.com/journal/
sj/434/preface.txt.

Ritsko:2005:Pb

[RB05a] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 44(2):213–214, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/preface.html;   http:
//www.research.ibm.com/journal/
sj/442/preface.pdf;   http:
//www.research.ibm.com/journal/
sj/442/preface.txt.

Ritsko:2005:Pc

[RB05b] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 44(4):651–652, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/preface.html.

Ritsko:2006:Pa

[RB06] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 45(2):205, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/452/preface.html.

Ritsko:2007:P

[RB07] John J. Ritsko and Alex Birman. Preface. IBM Systems Journal, 46(2):203–??, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/462/preface.html.

Ruggiero:1969:APS

[RC69] J. F. Ruggiero and D. A.

Coryell. An auxiliary processing system for array calculations. IBM Systems Journal, 8(2):118–135, 1969. CODEN IBMSA7. ISSN 0018-8670.

Reisman:1991:PMS

[RC91] S. Reisman and W. A. Carr.

Perspectives on multimedia systems in education. IBM Systems Journal, 30(3):280–295, 1991. CODEN IBMSA7. ISSN 0018-8670.

Rubin:1998:JIS

[RCB98] B. S. Rubin, A. R. Christ, and K. A. Bohrer. Java and the IBM San Francisco project. IBM Systems Journal, 37(3): 365–371, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/373/rubin.html.

Ramanathan:2007:RAR

[RCPR07] J. Ramanathan, R. J. Cohen, E. Plassmann, and K. Ramamoorthy. Role of an auditing and reporting service in compliance management.

IBM Systems Journal, 46(2): 305–??, ???? 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/462/ramanathan.html.

Ritsko:2001:Pa

[RD01] John J. Ritsko and Alfred G. Davis. Preface. IBM Systems Journal, 40(1):2–3, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/401/preface.html.

Ritsko:2002:Pb

[RD02] John J. Ritsko and Alfred G. Davis. Preface. IBM Systems Journal, 41(2):168–177, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/412/preface.html.

Ritsko:2003:P

[RD03] John J. Ritsko and Alfred G. Davis. Preface. IBM Systems Journal, 42(3):403–404, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/423/preface.pdf.

Rechtschaffen:1972:QSU

[Rec72] R. N. Rechtschaffen. Queuing simulation using a random number generator. IBM Systems Journal, 11(3):255–271, 1972. CODEN IBMSA7. ISSN 0018-8670.

Reiser:1976:IMC

[Rei76] Martin Reiser. Interactive modeling of computer systems. IBM Systems Journal, 15(4):309– 327, 1976. CODEN IBMSA7. ISSN 0018-8670.

Reich:1979:FMV

[Rei79] David L. Reich. Forum: Managing VM/CMS systems. IBM Systems Journal, 18(4):604– 605, 1979. CODEN IBMSA7. ISSN 0018-8670. See [DK79b, DK79a].

Reinsch:1988:DDS

[Rei88]  R. Reinsch.  Distributed

database for SAA. IBM Systems Journal, 27(3):362–369, 1988. CODEN IBMSA7. ISSN 0018-8670.

Reilly:1989:DVA

[Rei89] P. Reilly. Data visualization in archaeology. IBM Systems Journal, 28(4):569–579, 1989. CODEN IBMSA7. ISSN 0018-8670.

Resnick:1996:TTW

[Res96] M. Resnick. Things to think with. IBM Systems Journal, 35 (3&4):441–442, 1996. CODEN IBMSA7. ISSN 0018-8670.

Resnick:2000:PEC

[Res00] M. Resnick. Part 3: Everywhere computing: It’s not just information. IBM Systems Journal, 39(3/4):816–817, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/

sj/393/part2/resnick.html;
 http://www.research.ibm.
com/journal/sj/393/part2/
resnick.pdf;   http://www.
research.ibm.com/journal/
sj/393/part2/resnick.txt.

Roberts:1985:DCI

[RG85] M. L. Roberts and P. D. Griffiths. Design considerations for IBM Personal Computer Professional Fortran, an optimizing compiler. IBM Systems Journal, 24(1):49–60, 1985. CO-DEN IBMSA7. ISSN 0018-8670.

Raimondi:1976:LED

[RGH^{+}76] Donald L. Raimondi, Henry M. Gladney, Gerd Hochweller, Robert W. Martin, and Linda L. Spencer. Labs/7 — a distributed real-time operating system. IBM Systems Journal, 15(1):81–101, 1976. CODEN IBMSA7. ISSN 0018-8670.

Raver:1977:ALD

[RH77] N. Raver and G. U. Hubbard. Automated logical data base design: Concepts and applications. IBM Systems Journal, 16(3):287–312, 1977. CODEN IBMSA7. ISSN 0018-8670.

Roth:2006:XMT

[RHC^{+}06] M. Roth, M. A. Hernandez, P. Coulthard, L. Yan, L. Popa, H. C.-T. Ho, and C. C. Salter. XML mapping technology: Making connections in an XML-centric world. IBM Systems Journal, 45(2):

389–409, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/roth.html.

Radice:1985:PPS

[RHMP85] R. A. Radice, J. T. Harding, P. E. Munnis, and R. W. Phillips. A programming process study. IBM Systems Journal, 24(2):91–101, 1985. CO-DEN IBMSA7. ISSN 0018-8670.

Radice:1999:PPS

[RHMP99] R. A. Radice, J. T. Harding, P. E. Munnis, and R. W. Phillips. A programming process study. IBM Systems Journal, 38(2–3):297–307, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/382/radice.pdf. Reprinted

from [RROC85].

Ricke:2000:TNT

[Ric00] D. D. Ricke. Technical note — A three-dimensional framework for information technology solutions. IBM Systems Journal, 39(2):336–359, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/392/ricke.html.

Reinwald:1996:SUO

[RLPG96] B. Reinwald, T. J. Lehman, H. Pirahesh, and V. Gottemukkala. Storing and using objects in a relational database. IBM Systems Journal, 35(2):172–191, 1996. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj35-2.html#five.

Ricour:1967:IDM

[RM67] D. H. Ricour and V. Mei. Internal data management techniques for DOS/360. IBM Systems Journal, 6(1):38–48, 1967. CODEN IBMSA7. ISSN 0018-8670.

Rose:1988:SNM

[RM88] D. B. Rose and J. E. Munn. SNA network management directions. IBM Systems Journal, 27(1):3–14, January 1988. CODEN IBMSA7. ISSN 0018-8670.

Rhodes:2000:JTI

[RM00] B. J. Rhodes and P. Maes. Just-in-time information retrieval agents. IBM Systems Journal, 39(3/4):685–704, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part2/rhodes.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part2/
                                        rhodes.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/rhodes.txt.

Russell:2003:CNM

[RMC03] L. W. Russell, S. P. Morgan, and E. G. Chron. Clockwork: a new movement in autonomic systems. IBM Systems Journal, 42(1):77–84, ????

2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/russell.pdf.

Russell:2003:DGM

[RMDN03] D. M. Russell, P. P. Maglio, R. Dordick, and C. Neti. Dealing with ghosts: Managing the user experience of autonomic computing. IBM Systems Journal, 42(1):177–195, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/maglio.pdf.

Rooney:2008:ITP

[RMH08] William J. Rooney, Gregory E.

McBride, and Tariq Hanif. IBM TotalStorage Productivity Center for Replication for z/OS. IBM Systems Journal, 47(4): 681–694, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/rooney.pdf.

Radhakrishnan:2008:ISM

[RMP08] Rajesh Radhakrishnan, Kari Mark, and Bill Powell. IT service management for high availability. IBM Systems Journal, 47(4):549–561, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/474/radhakrishnan.pdf.

Rose:2006:VXT

[RMS06] K. H. Rose, S. Malaika, and R. J. Schloss. Virtual XML: a toolbox and use cases for the XML world view. IBM Systems Journal, 45(2):411–424, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/452/rose.html.

Resnick:1996:PBT

[RMSS96] M. Resnick, F. Martin, R. Sargent, and B. Silverman. Programmable bricks: Toys to think with. IBM Systems Journal, 35(3&4):443–452, 1996. CODEN IBMSA7. ISSN 0018-8670.

Robson:2001:FSU

[Rob01] B. Robson. Fastfinger: a study into the use of compressed residue pair separation matrices for protein sequence comparison. IBM Systems Journal, 40(2):442–463, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/robson.html;   http:
//www.research.ibm.com/journal/
sj/402/robson.pdf;   http:
//www.research.ibm.com/journal/
sj/402/robson.txt.

Rofrano:1992:DCD

[Rof92] J. J. Rofrano, Jr. Design considerations for distributed applications. IBM Systems Journal, 31(3):564–589, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5487.

Rofrano:2001:IWC

[Rof01] J. J. Rofrano. IBM Web-Sphere Commerce Suite Product Advisor. IBM Systems Journal, 40(1):91–103, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/401/rofrano.html.

Rogers:1975:SPV

[Rog75] John G. Rogers. Structured programming for virtual storage systems. IBM Systems Journal, 14(4):385–406, 1975. CODEN IBMSA7. ISSN 0018-8670.

Rogers:1983:SAC

[Rog83] G. R. Rogers. A simple architecture for consistent application program design. IBM Systems Journal, 23(3):199–213, 1983. CODEN IBMSA7. ISSN 0018-8670.

Rometty:2008:MSV

[Rom08]  Ginni Rometty.  Message

from the Senior Vice President, Global Business Services. IBM Systems Journal, 47(1):??, January/March 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/471/message.html.

Resnick:1996:CCP

[RR96] M. Resnick and N. Rusk. The Computer Clubhouse: Preparing for life in a digital world. IBM Systems Journal, 35 (3&4):431–440, 1996. CODEN IBMSA7. ISSN 0018-8670.

Radice:1985:PPA

[RROC85] R. A. Radice, N. K. Roth, A. C. O’Hara, Jr., and W. A. Ciarfella. A programming process architecture. IBM Systems Journal, 24(2):79–90, 1985. CODEN IBMSA7. ISSN 0018-8670.

Rettus:1972:ACD

[RS72] R. C. Rettus and R. A. Smith. Accounting control of data processing. IBM Systems Journal, 11(1):74–92, 1972. CODEN IBMSA7. ISSN 0018-8670.

Ravin:1973:IGS

[RS73] J. Ravin and M. Schatzoff. An interactive graphics system for analysis of business decisions. IBM Systems Journal, 12(3):238–256, 1973. CODEN IBMSA7. ISSN 0018-8670.

Ryan:1985:CPL

[RS85] R. R. Ryan and H. Spiller. The C programming language and a C compiler. IBM Systems Journal, 24(1):37–48, 1985. CO-DEN IBMSA7. ISSN 0018-8670.

Retallick:1998:EKS

[RS98] R. Retallick and S. Sánchez. Enterprise knowledge sharing, activity management, and a fabric for commitment. IBM Systems Journal, 37(2):189– 199, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/372/retallick.html.

Ritsko:2004:Pb

[RS04a] John J. Ritsko and David I. Seidman. Preface. IBM Systems Journal, 43(2):211–212, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/432/preface.html.

Ritsko:2004:Pc

[RS04b] John J. Ritsko and David I. Seidman. Preface. IBM Systems Journal, 43(3):449–450, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/preface.html;   http:
//www.research.ibm.com/journal/
sj/433/preface.pdf;   http:
//www.research.ibm.com/journal/
sj/433/preface.txt.

Ritsko:2005:Pa

[RS05] John J. Ritsko and David I.

Seidman. Preface. IBM Systems Journal, 44(1):3–4, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/441/preface.html;   http:
//www.research.ibm.com/journal/
sj/441/preface.pdf;   http:
//www.research.ibm.com/journal/
sj/441/preface.txt.

Ritsko:2006:Pb

[RS06] John J. Ritsko and David I. Seidman. Preface. IBM Systems Journal, 45(3):449–450, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/453/preface.html.

Rubsam:1989:MDS

[Rub89] K. G. Rubsam. MVS data services. IBM Systems Journal, 28(1):151–164, 1989. CODEN IBMSA7. ISSN 0018-8670.

Rully:1968:IGD

[Rul68] A. D. Rully. Interactive graphics in data processing: a subroutine package for FOR-TRAN. IBM Systems Journal, 7(3/4):248–256, 1968. CODEN IBMSA7. ISSN 0018-8670.

Ross:2004:PUC

[RW04] J. W. Ross and G. Westerman. Preparing for utility computing: The role of IT architecture and relationship management. IBM Systems Journal, 43(1):5–19, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/431/ross.pdf.

Roth:2002:IIN

[RWKN02] M. A. Roth, D. C. Wolfson, J. C. Kleewein, and C. J. Nelin. Information integration: a new generation of information technology. IBM Systems Journal, 41(4):563–577, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/414/roth.pdf.

Ryder:1970:HAT

[Ryd70]  K. D. Ryder.  A heuris-

tic approach to task dispatching. IBM Systems Journal, 9 (3):189–198, 1970. CODEN IBMSA7. ISSN 0018-8670.

Ryder:1974:OPR

[Ryd74] K. D. Ryder. Optimizing program replacement in virtual systems. IBM Systems Journal, 13(4):292–306, 1974. CODEN IBMSA7. ISSN 0018-8670.

Ryder:1983:EAS

[Ryd83]  K. D. Ryder.  An exper-

imental address space isolation technique for SNA networks. IBM Systems Journal, 23(4):367–386, 1983. CODEN IBMSA7. ISSN 0018-8670.

Ryman:1990:PSI

[Rym90] A. Ryman. Personal systems image application architecture: Lessons learned from the ImagEdit program. IBM Systems Journal, 29(3):408–420, 1990. CODEN IBMSA7. ISSN 0018-8670.

Skinner:1969:ESC

[SA69] C. E. Skinner and J. R. Asher.

Effects of storage contention on system performance. IBM Systems Journal, 8(4):319–333, 1969. CODEN IBMSA7. ISSN 0018-8670.

Senko:1973:DSAa

[SAAF73a] M. E. Senko, E. B. Altman, M. M. Astrahan, and P. L. Fehder. Data structures and accessing in data-base systems, part I: Evolution of information systems. IBM Systems Journal, 12(1):30–44, 1973. CO-DEN IBMSA7. ISSN 0018-

Senko:1973:DSAb

[SAAF73b] M. E. Senko, E. B. Altman, M. M. Astrahan, and P. L. Fehder. Data structures and accessing in data-base systems, part II: Information organization. IBM Systems Journal, 12(1):45–63, 1973. CODEN IBMSA7. ISSN 0018-8670.

Senko:1973:DSAc

[SAAF73c] M. E. Senko, E. B. Altman, M. M. Astrahan, and P. L. Fehder. Data structures and accessing in data-base systems, part III: Data representations and the data independent accessing model. IBM Systems Journal, 12(1):64–93, 1973. CODEN IBMSA7. ISSN 0018-8670.

Sabbah:2006:I

[Sab06]  Daniel Sabbah.  Introduc-

tion. IBM Systems Journal, 45(3):??, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/453/intro.html.

Sagawa:1990:RMT

[Sag90] J. M. Sagawa. Repository Manager technology (specifications manager). IBM Systems Journal, 29(2):209–227, 1990. CO-DEN IBMSA7. ISSN 0018-8670. See erratum [Ano90f].

Sales:1988:CNG

[Sal88] R. Sales. COBOL/2: The next generation in applications programming. IBM Systems Journal, 27(2):158–169, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Samukawa:1988:PSI

[Sam88] H. Samukawa. Programming style on the IBM 3090 Vector Facility considering both performance and flexibility. IBM Systems Journal, 27(4):453– 474, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Sandberg:1981:PRD

[San81] G. Sandberg. A primer on relational data base concepts. IBM Systems Journal, 20(1):23–40, 1981. CODEN IBMSA7. ISSN 0018-8670.

Sanford:2007:MSV

[San07] Linda S. Sanford. Message from the Senior Vice President, Enterprise On Demand Transformation & Information Technology. IBM Systems Journal, 46(4):??, October/ December 2007. CODEN IBMSA7. ISSN 0018-8670.

Sarrna:1977:IEM

[Sar77] K. S. Sarrna. An input-output econometric model. IBM Systems Journal, 16(4):398–420, 1977. CODEN IBMSA7. ISSN 0018-8670.

Sakamoto:1982:SBS

[SB82a] J. G. Sakamoto and F. W.

Ball. Supporting business systems planning studies with the DB/DC data dictionary. IBM Systems Journal, 21(1):54–80, 1982. CODEN IBMSA7. ISSN 0018-8670.

Schick:1982:DIA

[SB82b] Thomas Schick and Robert F. Brockish. The Document Interchange Architecture: a member of a family of architectures in the SNA environment. IBM Systems Journal, 21(2):220–244, 1982. CODEN IBMSA7. ISSN 0018-8670.

Saracelli:1993:PAS

[SB93] K. D. Saracelli and K. F. Bandat. Process automation in software application development. IBM Systems Journal, 32(3):376–396, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5517.

Sultan:1995:API

[SB95] R. A. Sultan and C. Basso. ATM: Paving the information superhighway. IBM Systems Journal, 34(3):375–389, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj34-3.html#four.

Seaman:1997:COS

[SB97] C. B. Seaman and V. R. Basili. Communication and organization in software development: An empirical study. IBM Systems Journal, 36(4):550–563, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

                                        research.ibm.com/journal/
                                        sj/364/seaman.html.

Selker:2000:CAD

[SB00a] T. Selker and W. Burleson.

Context-aware design and interaction in computer systems. IBM Systems Journal, 39(3/4): 880–891, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part3/selker.html;
 http://www.research.ibm.
com/journal/sj/393/part3/
selker.pdf;   http://www.
research.ibm.com/journal/
sj/393/part3/selker.txt.

Smith:2000:JIM

[SB00b] B. K. Smith and E. Blankinship. Justifying imagery: Multimedia support for learning through explanation. IBM Systems Journal, 39(3/4):749– 767, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part2/blankinship.
html;   http://www.research.
ibm.com/journal/sj/393/part2/
blankinship.pdf;   http:/
/www.research.ibm.com/journal/
sj/393/part2/blankinship.
txt.

Sockut:1997:MLR

[SBC97] G. H. Sockut, T. A. Beavin, and C.-C. Chang. A method for on-line reorganization of a database. IBM Systems Journal, 36(3):411–436, 1997. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.

                                        research.ibm.com/journal/
                                        sj/363/sockut.html.

Scadden:2008:RHC

[SBC^{+}08] Russell R. Scadden, Richard J. Bogdany, Jade W. Clifford, Herbert D. (Herbie) Pearthree, and Richard A. Locke. Resilient hosting in a continuously available virtualized environment. IBM Systems Journal, 47(4):535–548, ???? 2008. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/474/scadden.pdf.

Smith:2000:SSJ

[SBE^{+}00] B. K. Smith, W. Bender, I. Endter, J. Driscoll, M. Turpeinen, and D. Quan. Silver Stringers and Junior Journalists: Active information producers. IBM Systems Journal, 39(3/4): 730–748, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/smith.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part2/
                                        smith.pdf;   http://www.
                                        research.ibm.com/journal/
                                        sj/393/part2/smith.txt.

Slonim:1994:CAS

[SBL^{+}94] J. Slonim, M. A. Bauer, P.-Å. Larson, J. Schwarz, C. Butler, E. B. Buss, and D. Sabbah. The Centre for Advanced Studies: a model for applied research and development. IBM Systems Journal, 33(3):382–398, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-3.html#two. G321-5547.

Srinivasan:1997:OPO

[SC97] V. Srinivasan and D. T. Chang.

Object persistence in objectoriented applications. IBM Systems Journal, 36(1):??, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/361/srinivasan.html.

Scherr:1973:FSI

[Sch73] A. L. Scherr. Functional structure of IBM Virtual Storage Operating Systems — 2 OS/ VS2-2 concepts and philosophies. IBM Systems Journal, 12(4):382–400, 1973. CODEN IBMSA7. ISSN 0018-8670.

Scherr:1978:DDP

[Sch78] Allan L. Scherr. Distributed data processing. IBM Systems Journal, 17(4):324–343, 1978. CODEN IBMSA7. ISSN 0018-8670.

Schardt:1980:MTA

[Sch80a] Richard M. Schardt. An MVS tuning approach. IBM Systems Journal, 19(1):102–119, 1980. CODEN IBMSA7. ISSN 0018-8670.

Schiller:1980:SCP

[Sch80b] Donald C. Schiller. System capacity and performance evaluation. IBM Systems Journal, 19(1):46–67, 1980. CODEN IBMSA7. ISSN 0018-8670.

Scherr:1983:PCC

[Sch83] A. L. Scherr. A perspective on communications and computing. IBM Systems Journal, 23(1/2):5–10, 1983. CODEN IBMSA7. ISSN 0018-8670.

Scherr:1987:SNS

[Sch87] Allan L. Scherr. Structures for networks of systems. IBM Systems Journal, 26(1):4–12, January 1987. CODEN IBMSA7. ISSN 0018-8670.

Scherr:1988:SDP

[Sch88] A. L. Scherr. SAA distributed processing. IBM Systems Journal, 27(3):370–383, 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Scherr:1993:NAB

[Sch93] A. L. Scherr. A new approach to business processes. IBM Systems Journal, 32(1):80–98, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5504.

Scherr:1999:DDP

[Sch99] A. L. Scherr. Distributed data processing. IBM Systems Journal, 38(2–3):354–374, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/382/scherr.pdf. Reprinted

from [Sch78].

Schirmer:2003:PKM

[Sch03] A. L. Schirmer. Privacy and knowledge management: Challenges in the design of the Lotus Discovery Server. IBM Systems Journal, 42(3):519–531, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/423/schirmer.pdf.

Somani:2002:BTC

[SCK02] A. Somani, D. Choy, and J. Kleewein. Bringing together content and data management systems: Challenges and opportunities. IBM Systems Journal, 41(4):686–696, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/414/somani.pdf.

Sawin:2003:ATR

[SCO03] D. A. Sawin, J. A. Calcaterra, and K. Olka. Access ThinkPad: The right information at the right time and place. IBM Systems Journal, 42(4):669–685, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/424/sawin.pdf.

Shi:2003:SMF

[SD03] D. Shi and R. L. Daniels. A survey of manufacturing flexibility: Implications for e-business flexibility. IBM Systems Journal, 42(3):414–427, ???? 2003. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/423/shi.pdf.

Shabo:2007:SLC

[SD07] A. Shabo and D. Dotan. The seventh layer of the clinicalgenomics information infrastructure. IBM Systems Journal, 46(1):57–67, ???? 2007. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/461/shabo.html.

Sow:2006:UHY

[SDE^{+}06] D. M. Sow, J. S. Davis, M. R. Ebling, A. Misra, and L. Bergman. Uncovering the to-dos hidden in your in-box. IBM Systems Journal, 45(4): 739–757, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/454/sow.html.

Sparacino:2000:MPI

[SDP00] F. Sparacino, G. Davenport, and A. Pentland. Media in performance: Interactive spaces for dance, theater, circus, and museum exhibits. IBM Systems Journal, 39(3/4):479–510, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/393/part1/sparacino.html;
                                         http://www.research.ibm.
                                        com/journal/sj/393/part1/
                                        sparacino.pdf;   http://
                                        www.research.ibm.com/journal/
                                        sj/393/part1/sparacino.txt.

Sharon:2008:EPN

[SE08] G. Sharon and O. Etzion. Event-processing network model and implementation. IBM Systems Journal, 47(2):321– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/sharon.html.

Seaman:1966:TSDb

[Sea66] P. H. Seaman. On teleprocessing system design, part VI: The role of digital simulation. IBM Systems Journal, 5(3):175–189, 1966. CODEN IBMSA7. ISSN 0018-8670.

Seaman:1980:MCP

[Sea80] Phillip H. Seaman. Modeling considerations for predicting performance of CICS/VS systems. IBM Systems Journal, 19(1):68–80, 1980. CO-DEN IBMSA7. ISSN 0018-8670.

Selinger:1987:DT

[Sel87] Patricia G. Selinger. Database technology. IBM Systems Journal, 26(1):96–106, January 1987. CODEN IBMSA7. ISSN 0018-8670.

Selic:2006:UMD

[Sel06] B. Selic. UML 2: a modeldriven development tool. IBM Systems Journal, 45(3):607– 620, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/453/selic.html.

Selic:2008:APE

[Sel08]  B. Selic.  Accounting for

platform effects in the design of real-time software using model-based methods. IBM Systems Journal, 47(2):309– ??, April/June 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/472/selic.html.

Senko:1977:DSD

[Sen77] M. E. Senko. Data structures and data accessing in data base systems past, present, future. IBM Systems Journal, 16(3):208–257, 1977. CODEN IBMSA7. ISSN 0018-8670.

Seward:1967:AMS

[Sew67] F. K. Seward. An applicationoriented multiprocessing system, part VI: Programs for the intended application. IBM Systems Journal, 6(2):124–132, 1967. CODEN IBMSA7. ISSN 0018-8670.

Siew:2001:CEP

[SF01] N. Siew and D. Fischer. Convergent evolution of protein structure prediction and computer chess tournaments: CASP, Kasparov, and CAFASP. IBM Systems Journal, 40(2): 410–425, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/402/siew.html;   http:
//www.research.ibm.com/journal/
sj/402/siew.pdf;   http:/

/www.research.ibm.com/journal/
sj/402/siew.txt.

Sabbah:2005:I

[SF05] Danny Sabbah and Daniel Frye. Introduction. IBM Systems Journal, 44(2):??, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/442/intro.html;   http:
//www.research.ibm.com/journal/
sj/442/intro.pdf;   http:
//www.research.ibm.com/journal/
sj/442/intro.txt.

Schlatter:1994:BOM

[SFJ^{+}94] M. Schlatter, R. Furegati, F. Jeger, H. Schneider, and H. Streckeisen. The Business Object Management System. IBM Systems Journal, 33(2):239–263, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-2.html#three. G321-

5541.

Sawyer:1998:SDP

[SG98] S. Sawyer and P. J. Guinan.

Software development: Processes and performance. IBM Systems Journal, 37(4):552– 569, ???? 1998. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/374/sawyer.html.

Scalzi:1989:ESA

[SGS89] C. A. Scalzi, A. G. Ganek, and R. J. Schmalz. Enterprise Systems Architecture/370: an architecture for multiple virtual space access and authorization. IBM Systems Journal, 28(1): 15–38, 1989. CODEN IBMSA7. ISSN 0018-8670.

Stanley:1968:SGS

[SH68] W. I. Stanley and H. F. Hertel. Statistics gathering and simulation for the Apollo realtime operating system. IBM Systems Journal, 7(2):85–102, 1968. CODEN IBMSA7. ISSN 0018-8670.

Simpson:1987:IRP

[SH87] R. O. Simpson and P. D. Hester. The IBM RT PC ROMP processor and memory management unit architecture. IBM Systems Journal, 26(4): 346–360, December 1987. CO-DEN IBMSA7. ISSN 0018-8670.

Shapiro:1965:FAO

[Sha65a] S. Shapiro. Fabrication and assembly operations, part VII: Adaptive control in production planning. IBM Systems Journal, 4(3):250–255, 1965. CO-DEN IBMSA7. ISSN 0018-8670.

Shapiro:1965:TCW

[Sha65b] S. Shapiro. A technique to control waiting time in a queue. IBM Systems Journal, 4(1):53– 57, 1965. CODEN IBMSA7. ISSN 0018-8670.

Shah:1979:APE

[Sha79] M. J. Shah. Automatic programming for energy management using sensor based computers. IBM Systems Journal, 18(3):457–469, 1979. CODEN IBMSA7. ISSN 0018-8670.

Shepherd:1980:EPM

[She80] B. J. Shepherd. Experimental page makeup of text with graphics on a raster printer. IBM Systems Journal, 19(3): 345–355, 1980. CODEN IBMSA7. ISSN 0018-8670.

Snir:1995:CSP

[SHFG95] M. Snir, P. Hochschild, D. D.

Frye, and K. J. Gildea. The communication software and parallel environment of the IBM SP2. IBM Systems Journal, 34(2):205–221, 1995. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj34-2.html#five.

Schmidt:2005:ESB

[SHLP05] M.-T. Schmidt, B. Hutchison, P. Lambros, and R. Phippen. The Enterprise Service Bus: Making service-oriented architecture real. IBM Systems Journal, 44(4):781–797, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/schmidt.html.

Stacy:1969:DES

[SHP69] E. W. Stacy, L. E. Hunsinger, and J. F. Price. Determining economic sampling plans. IBM Systems Journal, 8(3):220–244, 1969. CODEN IBMSA7. ISSN 0018-8670.

Shu:1989:VPP

[Shu89] N. C. Shu. Visual programming: Perspectives and approaches. IBM Systems Journal, 28(4):525–547, 1989. CO-DEN IBMSA7. ISSN 0018-8670.

Shu:1999:VPP

[Shu99] N. C. Shu. Visual programming: Perspectives and approaches. IBM Systems Journal, 38(2–3):199–224, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/382/shu.pdf. Reprinted

from [Shu89].

Simon:2005:VOS

[Sim05] K. D. Simon. The value of open standards and opensource software in government environments. IBM Systems Journal, 44(2):227–238, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/442/simon.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/simon.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/442/simon.txt.

Singer:2003:CSI

[Sin03] C. A. Singer. Context-specific intellectual capital—the next link in the knowledge chain. IBM Systems Journal, 42(3): 446–461, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.

research.ibm.com/journal/
sj/423/singer.pdf.

Siwiec:1977:HPD

[Siw77] Joseph E. Siwiec. High performance DB/DC system. IBM Systems Journal, 16(2):169– 195, 1977. CODEN IBMSA7. ISSN 0018-8670.

Su:2004:MIP

[SJL^{+}04] Z. Su, J. Jiang, T. Liu, G. T.

Xie, and Y. Pan. Market Intelligence Portal: An entity-based system for managing market intelligence. IBM Systems Journal, 43(3):534–545, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/su.html;   http://
www.research.ibm.com/journal/
sj/433/su.pdf;   http://
www.research.ibm.com/journal/
sj/433/su.txt.

Singh:1986:IPB

[SKA86] Y. Singh, G. M. King, and J. W. Anderson. IBM 3090 performance: a balanced system approach. IBM Systems Journal, 25(1):20–35, 1986. CO-DEN IBMSA7. ISSN 0018-8670.

Sultan:1987:ISA

[SKG^{+}87] R. A. Sultan, P. Kermani, G. A. Grover, T. P. Barzilai, and A. E. Baratz. Implementing System/36 Advanced Peer-to-Peer Networking. IBM Systems Journal, 26(4):429– 452, December 1987. CODEN IBMSA7. ISSN 0018-8670.

Spangler:2006:MCD

[SKN06] W. S. Spangler, J. T. Kreulen, and J. F. Newswanger. Machines in the conversation: Detecting themes and trends in informal communication streams. IBM Systems Journal, 45(4): 785–799, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/454/spangler.html.

Shedler:1967:ERP

[SL67] G. S. Shedler and M. M. Lehman. Evaluation of redundancy in a parallel algorithm. IBM Systems Journal, 6(3):142–149, 1967. CODEN IBMSA7. ISSN 0018-8670.

Savir:1975:CDU

[SL75] David Savir and George J. Laurer. The characteristics and decodability of the Universal Product Code symbol. IBM Systems Journal, 14(1):16–34, 1975. CODEN IBMSA7. ISSN 0018-8670.

Schoch:1995:RSC

[SL95] D. J. Schoch and P. A. Laplante. A real-time systems context for the framework for information systems architecture. IBM Systems Journal, 34(1):20–38, 1995. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj34-1.html#three.

Sulack:1989:NDR

[SLD89] R. A. Sulack, R. J. Lindner, and D. N. Dietz. A new development rhythm for AS/400 software. IBM Systems Journal, 28(3):386–406, 1989. CODEN IBMSA7. ISSN 0018-8670.

Seaman:1966:TSDa

[SLW66] P. H. Seaman, R. A. Lind, and T. L. Wilson. On teleprocessing system design, part IV: An analysis of auxiliary-storage activity. IBM Systems Journal, 5(3):158–170, 1966. CODEN IBMSA7. ISSN 0018-8670.

Seawright:1979:VSM

[SM79] L. H. Seawright and R. A.

MacKinnon. VM/370—a study of multiplicity and usefulness. IBM Systems Journal, 18(1):4–17, 1979. CO-DEN IBMSA7. ISSN 0018-

8670. URL https://pages.
cs.wisc.edu/~stjones/proj/
vm_reading/ibmsj1801C.pdf.

Szabat:1992:INM

[SM92] M. M. Szabat and G. E.

Meyer. IBM network management strategy. IBM Systems Journal, 31(2):154–160, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5467.

Straatsma:2001:LBM

[SM01] T. P. Straatsma and J. A. Mc-Cammon. Load balancing of molecular dynamics simulation with NWChem. IBM Systems Journal, 40(2):328–341, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/402/straatsma.html;
                                        http://www.research.ibm.
                                        com/journal/sj/402/straatsma.
                                        pdf;   http://www.research.
                                        ibm.com/journal/sj/402/straatsma.
                                        txt.

Small:1996:NLB

[Sma96] D. Small. Navigating large bodies of text. IBM Systems Journal, 35(3&4):514–525, 1996. CODEN IBMSA7. ISSN 0018-8670.

Stevens:1974:SD

[SMC74] W. P. Stevens, G. J. Myers, and L. L. Constantine. Structured design. IBM Systems Journal, 13(2):115–139, 1974. CODEN IBMSA7. ISSN 0018-8670.

Stevens:1999:SD

[SMC99] W. P. Stevens, G. J. Myers, and L. L. Constantive. Structured design. IBM Systems Journal, 38(2–3):231–257, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/382/stevens.pdf. Reprinted

from [SMC74].

Smedley:1978:DSS

[Sme78] Bevan S. Smedley. The development of software systems to aid in physical planning. IBM Systems Journal, 17(4):359–382, 1978. CODEN IBMSA7. ISSN 0018-8670.

Smith:1962:SSE

[Smi62] E. C. Smith, Jr. Simulation in systems engineering. IBM Systems Journal, 1:33–50, September 1962. CODEN IBMSA7. ISSN 0018-8670.

Smith:1963:DCM

[Smi63] E. C. Smith, Jr. A directly coupled multiprocessing system. IBM Systems Journal, 2: 218–229, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Smith:1996:FME

[Smi96] J. R. Smith. Field mice: Extracting hand geometry from electric field measurements. IBM Systems Journal, 35 (3&4):587–608, 1996. CODEN IBMSA7. ISSN 0018-8670.

Smith:2001:TBT

[Smi01] R. A. Smith. Trends in ebusiness technologies. IBM Systems Journal, 40(1):4– 7, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/401/smith.html.

Sairamesh:2002:PBB

[SMK^{+}02] J. Sairamesh, R. Mohan, M. Kumar, L. Hasson, and C. Bender. A platform for business-to-business sell-side, private exchanges and marketplaces. IBM Systems Journal, 41(2):242–252, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.

research.ibm.com/journal/
sj/412/sairamesh.html;
http://www.research.ibm.
com/journal/sj/412/sairamesh.
pdf;   http://www.research.
ibm.com/journal/sj/412/sairamesh.
txt.

Schmandt:2000:EM

[SMM^{+}00] C. Schmandt, N. Marmasse, S. Marti, N. Sawhney, and S. Wheeler. Everywhere messaging. IBM Systems Journal, 39(3/4):660–677, ???? 2000. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/393/part1/schmandt.html;
 http://www.research.ibm.
com/journal/sj/393/part1/
schmandt.pdf;   http://www.
research.ibm.com/journal/
sj/393/part1/schmandt.txt.

Stone:1991:VCS

[SNC91] R. L. Stone, T. S. Nettleship, and J. Curtiss. VM/ESA CMS Shared File System. IBM Systems Journal, 30(1):52–71, 1991. CODEN IBMSA7. ISSN 0018-8670.

Strosnider:2008:MDS

[SNK^{+}08] Jay K. Strosnider, Prabir Nandi, Santhosh Kumaran, Shuvanker Ghosh, and Ali Arsnajani. Model-driven synthesis of SOA solutions. IBM Systems Journal, 47(3):415–432, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/473/strosnider.pdf. See

errata [Ano08a] with a corrected Table 4.

Somerville:1984:UIC

[Som84] Peter J. Somerville. Use of images in commercial and office systems. IBM Systems Journal, 23(3):281–296, 1984. CODEN IBMSA7. ISSN 0018-8670. See erratum [Ano85b].

Sordi:1984:QMF

[Sor84] Joseph J. Sordi. The Query Management Facility. IBM Systems Journal, 23(2):126–150, 1984. CODEN IBMSA7. ISSN 0018-8670.

Suganuma:2000:OIJ

[SOT^{+}00] T. Suganuma, T. Ogasawara, M. Takeuchi, T. Yasue, M. Kawahito, K. Ishizaki, H. Komatsu, and T. Nakatani. Overview of the IBM Java Just-in-Time Compiler. IBM Systems Journal, 39 (1):175–193, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/391/suganuma.html.

Sowa:2002:AIS

[Sow02] J. F. Sowa. Architectures for intelligent systems. IBM Systems Journal, 41(3):331–349, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/413/bigus.html.

Simpson:1978:NJE

[SP78] Richard O. Simpson and Grady H. Phillips. Network job entry facility for JES2. IBM Systems Journal, 17(3):221– 240, 1978. CODEN IBMSA7. ISSN 0018-8670.

Spector:2002:MVP

[Spe02] Alfred Spector. Message from the Vice President, Services and Software, Research Division. IBM Systems Journal, 41(3):??, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/message.html.

Spielberg:1966:CEU

[Spi66] K. Spielberg. Computation of e^{x} with the use of large tables. IBM Systems Journal, 5(2):102–114, 1966. CODEN IBMSA7. ISSN 0018-8670.

Spohrer:1999:IP

[Spo99] J. C. Spohrer. Information in places. IBM Systems Journal, 38(4):602–605, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/spohrer.html.

Sahulka:1991:FCI

[SPS^{+}91] R. J. Sahulka, E. C. Plachy, L. J. Scarborough, R. G. Scarborough, and S. W. White. FORTRAN for clusters of IBM ES/3090 multiprocessors. IBM Systems Journal, 30(3):296– 311, 1991. CODEN IBMSA7. ISSN 0018-8670.

Shedletsky:1993:ARD

[SR93] J. J. Shedletsky and J. J.

Rofrano. Application reference designs for distributed systems. IBM Systems Journal, 32(4):625–646, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5527.

Schmandt:1996:UAS

[SR96] C. Schmandt and D. Roy. Using acoustic structure in a hand-held audio playback device. IBM Systems Journal, 35 (3&4):453–472, 1996. CODEN IBMSA7. ISSN 0018-8670.

Sharp:2006:OGB

[SR06] C. E. Sharp and M. Rowe. Online games and e-business: Architecture for integrating business models and services into online games. IBM Systems Journal, 45(1):161–179, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/451/sharp.html.

Seidman:2008:P

[SR08] David Seidman and John J.

Ritsko. Preface. IBM Systems Journal, 47(3):355–356, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/473/preface.pdf.

Shabo:2006:RIX

[SRCV06] A. Shabo, S. Rabinovici-Cohen, and P. Vortman. Revolutionary impact of XML on biomedical information interoperability. IBM Systems Journal, 45 (2):361–372, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/452/shabo.html.

Schmidt:1998:PDR

[SRM^{+}98] W. J. Schmidt, R. R. Roediger, C. S. Mestad, B. Mendelson, I. Shavit-Lottem, and V. Bortnikov-Sitnitsky. Profiledirected restructuring of operating system code. IBM Systems Journal, 37(2):270– 305, ???? 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/372/schmidt.html.

Saracco:2004:EDE

[SRW04] C. M. Saracco, M. A. Roth, and D. C. Wolfson. Enabling distributed enterprise integration with WebSphere and DB2 Information Integrator. IBM Systems Journal, 43(2): 255–269, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/432/saracco.html;   http:
//www.research.ibm.com/journal/
sj/432/saracco.pdf;   http:
//www.research.ibm.com/journal/
sj/432/saracco.txt.

Seaman:1969:SOS

[SS69] P. H. Seaman and R. C. Soucy. Simulating operating systems. IBM Systems Journal, 8(4):264–279, 1969. CODEN IBMSA7. ISSN 0018-8670.

Slate:1978:ECL

[SS78] Laurel Slate and Kurt Spielberg. The Extended Control Language of MPSX/370 and possible applications. IBM Systems Journal, 17(1):64–81, 1978. CODEN IBMSA7. ISSN 0018-8670.

Singleton:1994:DAW

[SS94] J. P. Singleton and M. M.

Schwartz. Data access within the Information Warehouse framework. IBM Systems Journal, 33(2):300–325, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-2.html#six. G321-5544.

Smith:2001:PSP

[SS01] S. W. Smith and D. Safford. Practical server privacy with secure coprocessors. IBM Systems Journal, 40(3): 683–695, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/smith.html;   http:
//www.research.ibm.com/journal/
sj/403/smith.pdf.

Stunkel:1995:SHP

[SSA^{+}95] C. B. Stunkel, D. G. Shea, B. Abali, M. G. Atkins, C. A. Bender, D. G. Grice, P. Hochschild, D. J. Joseph, B. J. Nathanson, R. A. Swetz, R. F. Stucke, M. Tsao, and P. R. Varker. The SP2 high performance switch. IBM Systems Journal, 34(2):185–204, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-2.html#four.

Shi:2000:EDS

[SSB^{+}00] S. S. B. Shi, E. Stokes, D. Byrne, C. F. Corn, D. Bachmann, and T. Jones. An enterprise directory solution with DB2. IBM Systems Journal, 39(2):360–383, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/392/shi.html.

Sakaki:1988:EUI

[SSH88] M. Sakaki, H. Samukawa, and N. Honjou. Effective utilization of IBM 3090 large virtual storage in the numerically intensive computations of ab initio molecular orbitals. IBM Systems Journal, 27(4): 528–540, November 1988. CO-DEN IBMSA7. ISSN 0018-8670.

Stojanovic:2004:ROA

[SSM^{+}04] L. Stojanovic, J. Schneider, A. Maedche, S. Libischer, R. Studer, Th. Lumpp, A. Abecker, G. Breiter, and J. Dinger. The role of ontologies in autonomic computing systems. IBM Systems Journal, 43(3):598–616, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/stojanovic.html;
http://www.research.ibm.
com/journal/sj/433/stojanovic.
pdf;   http://www.research.
ibm.com/journal/sj/433/stojanovic.
txt.

Schutt:1987:MHS

[SSR87] Thomas E. Schütt, James B.

Staton, III, and Wilhelm F. Racke. Message handling systems based on the CCITT X.400 recommendations. IBM Systems Journal, 26(3):235– 254, 1987. CODEN IBMSA7. ISSN 0018-8670.

Shaikh:2006:DPO

[SSR^{+}06] A. Shaikh, S. Sahu, M.-C.

Rosu, M. Shea, and D. Saha. On demand platform for online games. IBM Systems Journal, 45(1):7–19, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/451/shaikh.html.

Sundstrom:1987:SCR

[SSS^{+}87] Robert J. Sundstrom, James B.

Staton, III, Gary D. Schultz, Matthew L. Hess, George A. Deaton, Jr., Leo J. Cole, and Robert M. Amy. SNA: Current requirements and direction. IBM Systems Journal, 26 (1):13–36, January 1987. CO-DEN IBMSA7. ISSN 0018-8670.

Spector:2004:I

[SSS04] Alfred Spector, John Swainson, and Danny Sabbah. Introduction. IBM Systems Journal, 43(2):??, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/432/intro.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/intro.pdf;   http:
                                        //www.research.ibm.com/journal/
                                        sj/432/intro.txt.

Scharfenberg:1964:CG

[SSV64] K. F. Scharfenberg, P. H. Smith, Jr., and R. D. Villani. A concordance generator. IBM Systems Journal, 3(1):104–111, 1964. CODEN IBMSA7. ISSN 0018-8670.

Sy:1987:OI

[SSY^{+}87] K. K. Sy, M. O. Shiobara, M. Yamaguchi, Y. Kobayashi, S. Shukuya, and T. Tomatsu. OSI-SNA interconnections. IBM Systems Journal, 26(2):157– 173, 1987. CODEN IBMSA7. ISSN 0018-8670.

Schleicher:1989:SOA

[ST89] D. L. Schleicher and R. L. Taylor. System overview of the Application System/400. IBM Systems Journal, 28(3):360– 375, 1989. CODEN IBMSA7. ISSN 0018-8670.

Schleicher:1999:SOA

[ST99] D. L. Schleicher and R. L. Taylor. System overview of the Application System/400. IBM Systems Journal, 38(2– 3):398–413, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.

research.ibm.com/journal/
sj/382/schleicher.pdf. Reprinted

from [ST89].

Stanley:1969:MSO

[Sta69] W. I. Stanley. Measurement of system operational statistics. IBM Systems Journal, 8(4):299–308, 1969. CODEN IBMSA7. ISSN 0018-8670.

Stay:1976:HIP

[Sta76] James F. Stay. HIPO and integrated program design. IBM Systems Journal, 15(2):143– 154, 1976. CODEN IBMSA7. ISSN 0018-8670.

Starner:1996:HWC

[Sta96] T. Starner. Human-powered wearable computing. IBM Systems Journal, 35(3&4):618– 629, 1996. CODEN IBMSA7. ISSN 0018-8670.

Sternad:1963:PC

[Ste63] Nicholas Sternad. Programming considerations for the 7750. IBM Systems Journal, 2: 57–85, March 1963. CODEN IBMSA7. ISSN 0018-8670.

Stevens:1964:SSP

[Ste64] W. Y. Stevens. The structure of SYSTEM/360, part II: System implementations. IBM Systems Journal, 3(2):136–143, 1964. CODEN IBMSA7. ISSN 0018-8670.

Stewart:1979:PAC

[Ste79] H. M. Stewart. Performance analysis of complex communications systems. IBM Systems Journal, 18(3):356–373, 1979. CODEN IBMSA7. ISSN 0018-8670.

Stevens:1982:HDF

[Ste82] Wayne P. Stevens. How data flow can improve application development productivity. IBM Systems Journal, 21(2):162– 178, 1982. CODEN IBMSA7. ISSN 0018-8670.

Stevenson:1992:MMN

[Ste92] J. G. Stevenson. Management of multivendor networks. IBM Systems Journal, 31(2):189– 205, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5469.

Stecher:1993:BBA

[Ste93] P. Stecher. Building business and application systems with the Retail Application Architecture. IBM Systems Journal, 32(2):278–306, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5514.

Siepel:2001:IPH

[STF^{+}01] A. C. Siepel, A. N. Tolopko, A. D. Farmer, P. A. Steadman, F. D. Schilkey, B. D. Perry, and W. D. Beavis. An integration platform for heterogeneous bioinformatics software components. IBM Systems Journal, 40(2):570–591, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/siepel.html;   http:
//www.research.ibm.com/journal/
sj/402/siepel.pdf;   http:

//www.research.ibm.com/journal/
sj/402/siepel.txt.

Stierhoff:1994:SR

[Sti94] George C. Stierhoff. Suggested reading. IBM Systems Journal, 33(4):700–701, 1994. CODEN IBMSA7. ISSN 0018-8670.

Streeter:1972:CES

[Str72] D. N. Streeter. Cost-benefit evaluation of scientific computing services. IBM Systems Journal, 11(3):219–233, 1972. CODEN IBMSA7. ISSN 0018-8670.

Streeter:1973:CDC

[Str73] D. N. Streeter. Centralization or dispersion of computing facilities. IBM Systems Journal, 12(3):283–301, 1973. CODEN IBMSA7. ISSN 0018-8670.

Streeter:1975:PCW

[Str75] Donald N. Streeter. Productivity of computer-dependent workers. IBM Systems Journal, 14(3):292–305, 1975. CODEN IBMSA7. ISSN 0018-8670.

Strickland:1997:VRL

[Str97] J. P. Strickland. VSAM recordlevel data sharing. IBM Systems Journal, 36(2):361–371, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://www.

research.ibm.com/journal/
sj/362/strickland.html.

Smith:2008:AAB

[STTA08] W. Earl Smith, Kishor S.

Trivedi, Lorrie A. Tomek, and Jerry Ackaret. Availability analysis of blade server systems. IBM Systems Journal, 47(4):621–640, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/smith.pdf.

Suda:1967:AMS

[Sud67] R. Suda. An applicationoriented multiprocessing system, part V: The diagnostic monitor. IBM Systems Journal, 6(2):116–123, 1967. CODEN IBMSA7. ISSN 0018-8670.

Sullivan:1983:CNM

[Sul83] Timothy P. Sullivan. Communications network management enhancements for SNA networks: an overview. IBM Systems Journal, 23(1-2):129– 142, 1983. CODEN IBMSA7. ISSN 0018-8670.

Summers:1984:OCS

[Sum84] R. C. Summers. An overview of computer security. IBM Systems Journal, 23(4):309–325, 1984. CODEN IBMSA7. ISSN 0018-8670.

Summers:1989:LDS

[Sum89] R. C. Summers. Local-area distributed systems. IBM Systems Journal, 28(2):227–240, 1989. CODEN IBMSA7. ISSN 0018-8670.

Strickland:1982:IES

[SUW82] Jimmy P. Strickland, Peter P. Uhrowczik, and Vern L. Watts.

IMS/VS: an evolving system. IBM Systems Journal, 21(4):490–510, 1982. CODEN IBMSA7. ISSN 0018-8670.

Sarkar:2003:IPS

[SVM^{+}03] P. Sarkar, K. Voruganti, K. Meth, O. Biran, and J. Satran. Internet Protocol storage area networks. IBM Systems Journal, 42(2): 218–231, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/422/sarkar.pdf.

Swartzlander:1990:CAa

[Swa90] Earl E. Swartzlander, Jr. Computer Arithmetic, volume 1 of IEEE Computer Society Press tutorial. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-8931-5 (hardcover), 0-8186-5931-9 (microfiche). xiii + 378 pp. LCCN QA76.6 .C633 1990.

Sweeney:1965:AFP

[Swe65] D. W. Sweeney. An analysis of floating-point addition. IBM Systems Journal, 4(1):31–42, 1965. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/041/ibmsjIVRID.pdf. Reprinted

in [Swa90, pp. 317–328].

Swope:2001:DCL

[Swo01] W. C. Swope. Deep computing for the life sciences.

IBM Systems Journal, 40(2): 248–263, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/402/swope.html;   http:
//www.research.ibm.com/journal/
sj/402/swope.pdf;   http:
//www.research.ibm.com/journal/
sj/402/swope.txt.

Sinha:2006:MFE

[SWS06] A. Sinha, C. E. Williams, and P. Santhanam. A measurement framework for evaluating model-based test generation tools. IBM Systems Journal, 45(3):501–514, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/453/sinha.html.

Shedler:1971:SMP

[SY71] G. S. Shedler and S. C. Yang. Simulation of a model of paging system performance. IBM Systems Journal, 10(2):113–128, 1971. CODEN IBMSA7. ISSN 0018-8670.

Sykes:1988:RIC

[Syk88] S. C. G. Sykes. The Realtime Interface Co-Processor Multiport/2 adapter. IBM Systems Journal, 27(2):198–205, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Symonds:1968:IGD

[Sym68]  A. J. Symonds.  Interac-

tive graphics in data processing: Auxiliary-storage associative data structure for PL/I. IBM Systems Journal, 7 (3/4):229–245, 1968. CODEN IBMSA7. ISSN 0018-8670.

Symonds:1986:IIK

[Sym86] A. J. Symonds. Introduction to IBM’s knowledge-systems products. IBM Systems Journal, 25(2):134–146, 1986. CO-DEN IBMSA7. ISSN 0018-8670.

So:1988:PLV

[SZ88] K. So and V. Zecca. Program locality of vectorized applications running on the IBM 3090 with Vector Facility. IBM Systems Journal, 27(4):436– 452, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Sowa:1992:EFF

[SZ92] J. F. Sowa and J. A. Zachman.

Extending and formalizing the framework for information systems architecture. IBM Systems Journal, 31(3):590–616, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5488.

Tagg:1990:ITS

[Tag90] B. S. Tagg. Implementing tool support for box structures. IBM Systems Journal, 29(1): 79–89, 1990. CODEN IBMSA7. ISSN 0018-8670.

Talmadge:1963:DIP

[Tal63] R. B. Talmadge. Design of an integrated programming and operating system, part II: The assembly program and its language. IBM Systems Journal, 2:162–181, June 1963. CODEN IBMSA7. ISSN 0018-8670.

Tavera:1985:ASI

[TAR85] M. L. Tavera, M. Alfonseca, and J. Rojas. An APL system for the IBM Personal Computer. IBM Systems Journal, 24(1):61–70, 1985. CODEN IBMSA7. ISSN 0018-8670.

Tetzlaff:1984:VAP

[TB84] W. H. Tetzlaff and W. M. Buco. VM/370, attached processor, and multiprocessor performance study. IBM Systems Journal, 23(4):375–385, 1984. CODEN IBMSA7. ISSN 0018-8670.

Turk:2005:VLS

[TB05] D. Turk and J. Bausch. Virtual Linux servers under z/VM: Security, performance, and administration issues. IBM Systems Journal, 44(2):341– 351, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/turk.html;   http:
//www.research.ibm.com/journal/
sj/442/turk.pdf;   http:/
/www.research.ibm.com/journal/
sj/442/turk.txt.

Tetzlaff:1987:PPV

[TBB^{+}87] W. H. Tetzlaff, T. Beretvas, W. M. Buco, J. Greenberg, D. R. Patterson, and G. A. Spivak. A page-swapping prototype for VM/HPO. IBM Systems Journal, 26(2):215–230, 1987. CODEN IBMSA7. ISSN 0018-8670.

Tang:1969:CEC

[TC69] D. T. Tang and R. T. Chien.

Coding for error control. IBM Systems Journal, 8(1):48–86, 1969. CODEN IBMSA7. ISSN 0018-8670.

Thomas:1981:HFC

[TC81] J. C. Thomas and J. M. Carroll. Human factors in communication. IBM Systems Journal, 20(2):237–263, 1981. CODEN IBMSA7. ISSN 0018-8670.

Temoshenko:1992:MSP

[Tem92] L. Temoshenko. Managing session performance using the NetView Performance Monitor. IBM Systems Journal, 31 (2):286–299, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5474.

Tetzlaff:1979:SSI

[Tet79] W. H. Tetzlaff. State sampling of interactive VM/370 users. IBM Systems Journal, 18(1):164–180, 1979. CODEN IBMSA7. ISSN 0018-8670.

Teng:1984:MID

[TG84] J. Z. Teng and R. A. Gumaer. Managing IBM Database 2 buffers to maximize performance. IBM Systems Journal, 23(2):211–218, 1984. CODEN IBMSA7. ISSN 0018-8670.

Tirso:1993:TFM

[TG93] J. R. Tirso and H. Gregorius.

Technical forum: Management of reuse at IBM. IBM Systems Journal, 32(4):612–615, 1993.

CODEN IBMSA7. ISSN 0018-8670.

Thadhani:1981:IUP

[Tha81] A. J. Thadhani. Interactive user productivity. IBM Systems Journal, 20(4):407–423, 1981. CODEN IBMSA7. ISSN 0018-8670.

Thadhani:1984:FAP

[Tha84] Arvind J. Thadhani. Factors affecting programmer productivity during application development. IBM Systems Journal, 23(1):19–35, 1984. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1977:P

[Thi77] Connie A. Thiel. Preface. IBM Systems Journal, 16(4):322– 323, 1977. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1978:Pc

[Thi78a] Connie Thiel. Preface. IBM Systems Journal, 17(3):204– 205, 1978. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1978:Pd

[Thi78b] Connie Thiel. Preface. IBM Systems Journal, 17(4):322– 323, 1978. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1978:Pa

[Thi78c] Connie A. Thiel. Preface. IBM Systems Journal, 17(1): 2–3, 1978. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1978:Pb

[Thi78d] Connie A. Thiel. Preface. IBM Systems Journal, 17(2):104– 105, 1978. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1979:Pa

[Thi79a] Connie Thiel. Preface. IBM Systems Journal, 18(1):2–3, 1979. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1979:Pb

[Thi79b] Connie Thiel. Preface. IBM Systems Journal, 18(2):186– 187, 1979. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1979:Pc

[Thi79c] Connie Thiel. Preface. IBM Systems Journal, 18(3):354– 355, 1979. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1979:Pd

[Thi79d] Connie Thiel. Preface. IBM Systems Journal, 18(4):482– 483, 1979. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1980:Pa

[Thi80a] Connie Thiel. Preface. IBM Systems Journal, 19(1):2–3, 1980. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1980:Pb

[Thi80b] Connie Thiel. Preface. IBM Systems Journal, 19(2):170, 1980. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1980:Pc

[Thi80c] Connie Thiel. Preface. IBM Systems Journal, 19(3):290– 291, 1980. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1980:Pd

[Thi80d] Connie A. Thiel. Preface. IBM Systems Journal, 19(4):412– 413, 1980. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1981:Pa

[Thi81a] Connie A. Thiel. Preface. IBM Systems Journal, 20(1): 2–3, 1981. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1981:Pb

[Thi81b] Connie A. Thiel. Preface. IBM Systems Journal, 20(2): 122, 1981. CODEN IBMSA7. ISSN 0018-8670.

Thiel:1981:Pc

[Thi81c] Connie A. Thiel. Preface. IBM Systems Journal, 20(3):270– 271, 1981. CODEN IBMSA7. ISSN 0018-8670.

Telford:2003:UDC

[THL^{+}03] R. Telford, R. Horman, S. Lightstone, N. Markov, S. O’Connell, and G. Lohman. Usability and design considerations for an autonomic relational database management system. IBM Systems Journal, 42(4): 568–581, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/424/telford.pdf.

Thompson:1988:VDC

[Tho88] S. Thompson. VGA—design choices for a new video subsystem. IBM Systems Journal, 27(2):185–197, June 1988.

CODEN IBMSA7. ISSN 0018-8670.

Thomson:1991:ASL

[Tho91] N. D. Thomson. APL2 as a specification language for statistics. IBM Systems Journal, 30(4):539–542, 1991. CO-DEN IBMSA7. ISSN 0018-8670.

Turunen:2005:AAS

[THR^{+}05] M. Turunen, J. Hakulinen, K.-J. Räihä, E.-P. Salonen, A. Kainulainen, and P. Prusi. An architecture and applications for speech-based accessibility systems. IBM Systems Journal, 44(3):485–504, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/443/turunen.html.

Tost:2000:UJC

[TJ00] A. Tost and V. M. Johnson. Using JavaBeans components as accessors to Enterprise JavaBeans components. IBM Systems Journal, 39(2): 293–300, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/tost.html.

TenDyke:1989:OP

[TK89] R. P. Ten Dyke and J. C.

Kunz. Object-oriented programming. IBM Systems Journal, 28(3):465–478, 1989. CO-DEN IBMSA7. ISSN 0018-

Thomas:2001:KMP

[TKE01] J. C. Thomas, W. A. Kellogg, and T. Erickson. The knowledge management puzzle: Human and social factors in knowledge management. IBM Systems Journal, 40(4): 863–884, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/404/thomas.html;   http:
//www.research.ibm.com/journal/
sj/404/thomas.pdf.

Tan:1998:IEH

[TLRH98] Y. S. Tan, D. B. Lindquist, T. O. Rowe, and J. R. Hind. IBM eNetwork Host On-Demand: The beginning of a new era for accessing host information in a Web environment. IBM Systems Journal, 37(1):133–151, 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/371/tan.html.

Treinen:2006:FSC

[TMF06] J. J. Treinen and S. L. Miller-Frost. Following the sun: Case studies in global software development. IBM Systems Journal, 45(4):773–783, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/454/treinen.html.

Todd:1976:PRT

[Tod76] Stephen J. P. Todd. The Peterlee Relational Test Vehicle—a system overview. IBM Systems Journal, 15(4):285–308, 1976. CODEN IBMSA7. ISSN 0018-8670.

Tedijanto:1995:NPS

[TOV^{+}95] T. E. Tedijanto, R. O. Onvural, D. C. Verma, L. Gun, and R. A. Guerin. NBBS path selection framework. IBM Systems Journal, 34(4):629–639, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-4.html#six.

Toomey:1988:IPF

[TPS^{+}88] L. J. Toomey, E. C. Plachy, R. G. Scarborough, R. J. Sahulka, J. F. Shaw, and A. W. Shannon. IBM Parallel FOR-TRAN. IBM Systems Journal, 27(4):416–435, November 1988. CODEN IBMSA7. ISSN 0018-8670.

Trimble:1991:SMI

[Tri91] R. Trimble. Storage management in IBM APL systems. IBM Systems Journal, 30(4):456–468, 1991. CODEN IBMSA7. ISSN 0018-8670.

Tian:2008:BFB

[TRL^{+}08] C. H. Tian, B. K. Ray, J. Lee, R. Cao, and W. Ding. BEAM: a framework for business ecosystem analysis and modeling. IBM Systems Journal, 47(1): 101–??, January/March 2008.

CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/471/tian.html.

Tucker:1967:MCS

[Tuc67] S. G. Tucker. Microprogram control for SYSTEM/360. IBM Systems Journal, 6(4):222–241, 1967. CODEN IBMSA7. ISSN 0018-8670.

Tucker:1986:ISO

[Tuc86] S. G. Tucker. The IBM 3090 System: an overview. IBM Systems Journal, 25(1):4–19, 1986. CODEN IBMSA7. ISSN 0018-8670.

Turnburke:1963:SDP

[Tur63] V. P. Turnburke, Jr. Sequential data processing design. IBM Systems Journal, 2:37–48, March 1963. CODEN IBMSA7. ISSN 0018-8670.

Tai:1994:EPB

[TVPL94] K.-C. Tai, M. A. Vouk, A. M. Paradkar, and P. Lu. Evaluation of a predicate-based software testing strategy. IBM Systems Journal, 33(3):445–457, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj33-3.html#five.  G321-

5550.

Tan:2004:SD

[TVX^{+}04] Y.-S. Tan, V. Vellanki, J. Xing, B. Topol, and G. Dudley. Service domains. IBM Systems Journal, 43(4):734–755, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/434/tan.html;   http:/
/www.research.ibm.com/journal/
sj/434/tan.pdf;   http://
www.research.ibm.com/journal/
sj/434/tan.txt.

Uhlir:1988:EUI

[Uhl88] S. Uhlir. Enabling the user interface. IBM Systems Journal, 27(3):306–314, 1988. CODEN IBMSA7. ISSN 0018-8670.

Uhrowczik:1973:DD

[Uhr73] P. P. Uhrowczik. Data Dictionary/Directories. IBM Systems Journal, 12(4):332–350, 1973. CODEN IBMSA7. ISSN 0018-8670.

Ullmer:2000:EFT

[UI00] B. Ullmer and H. Ishii. Emerging frameworks for tangible user interfaces. IBM Systems Journal, 39(3/4):915–931, ???? 2000. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/393/part3/ullmer.html;
http://www.research.ibm.
com/journal/sj/393/part3/
ullmer.pdf;   http://www.
research.ibm.com/journal/
sj/393/part3/ullmer.txt.

Uramoto:2004:TMS

[UMN^{+}04] N. Uramoto, H. Matsuzawa, T. Nagano, A. Murakami, H. Takeuchi, and K. Takeda. A text-mining system for knowledge discovery from biomedical documents. IBM Systems Journal, 43(3):516–533, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/433/uramoto.html;   http:
//www.research.ibm.com/journal/
sj/433/uramoto.pdf;   http:
//www.research.ibm.com/journal/
sj/433/uramoto.txt.

Unland:2008:MVP

[Unl08] Holly D. Unland. Message from the Vice President, Solutions & Global Technical Sales, IBM Global Sales & Distribution. IBM Systems Journal, 47(4):??, ???? 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/474/message.pdf.

Vilalta:2002:PAM

[VAH^{+}02] R. Vilalta, C. V. Apte, J. L. Hellerstein, S. Ma, and S. M. Weiss. Predictive algorithms in the management of computer systems. IBM Systems Journal, 41(3):461–474, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/413/vilalta.html.

VanLeer:1976:TDU

[Van76] P. Van Leer. Top-down development using a program design language. IBM Systems Journal, 15(2):155–170, 1976. CO-DEN IBMSA7. ISSN 0018-8670.

vanderSalm:1998:ISF

[van98] R. L. van der Salm. Introducing shareable frameworks into a procedural development environment. IBM Systems Journal, 37(2):200–214, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/372/vandersalm.html.

vanEmdeBoas-Lubsen:2000:BCP

[van00] H. van Emde Boas-Lubsen.

Business Component Prototyper for SanFrancisco: An experiment in architecture for application development tools. IBM Systems Journal, 39(2): 248–266, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/vanemdeboas.html.

Varadan:2008:IBF

[VCS^{+}08] Radhu Varadan, Kishore Channabasavaiah, Siljan Simpson, Kerrie Holley, and Abdul Allam. Increasing business flexibility and SOA adoption through effective SOA governance. IBM Systems Journal, 47(3):473–488, ???? 2008. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/473/varadan.pdf.

Salm:2000:TNI

[vdS00] R. van der Salm. Technical note — IBS Consist two years later: The LEGO brick dream.

IBM Systems Journal, 39(2): 301–314, ???? 2000. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/392/vandersalm.html.

Velasco:1964:GPD

[Vel64] C. R. Velasco. A general purpose digital simulator and examples of its application, part II: Simulation of a telephone intercept system. IBM Systems Journal, 3(1):35–40, 1964. CODEN IBMSA7. ISSN 0018-8670.

Verplaetse:1996:IPD

[Ver96] C. Verplaetse. Inertial proproceptive devices: Self-motion-sending toys and tools. IBM Systems Journal, 35(3&4):639– 650, 1996. CODEN IBMSA7. ISSN 0018-8670.

Vetter:1995:BCN

[Vet95] R. J. Vetter. Books: “Routing in Communications Networks”, Martha Streenstrup, Editor. IBM Systems Journal, 34(4):751–752, 1995. CODEN IBMSA7. ISSN 0018-8670.

vonKanel:1998:IMF

[vGLS98] J. von Känel, J. S. Givler, B. Leiba, and W. Segmuller. Internet messaging frameworks. IBM Systems Journal, 37(1):4–18, 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/371/vonkanel.html.

Vayghan:2007:IIT

[VGW^{+}07] J. A. Vayghan, S. M. Garfinkle, C. Walenta, D. C. Healy, and Z. Valentin. The internal information transformation of IBM. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-8670.

Viswanathan:2000:JVM

[VL00] D. Viswanathan and S. Liang.

Java Virtual Machine Profiler Interface. IBM Systems Journal, 39(1):82–95, ???? 2000. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/391/viswanathan.html.

Vetere:2005:MSI

[VL05] G. Vetere and M. Lenzerini.

Models for semantic interoperability in service-oriented architectures. IBM Systems Journal, 44(4):887–903, ???? 2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/vetere.html.

VanBuskirk:2003:EP

[VM03] R. Van Buskirk and B. W.

Moroney. Extending prototyping. IBM Systems Journal, 42 (4):613–623, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/424/vanbuskirk.pdf.

Venton:2005:ULH

[VMKB05] T. Venton, M. Miller, R. Kalla, and A. Blanchard. Using Linux for hardware bring up, development, and manufacturing. IBM Systems Journal, 44(2): 319–329, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/venton.html;   http:
//www.research.ibm.com/journal/
sj/442/venton.pdf;   http:
//www.research.ibm.com/journal/
sj/442/venton.txt.

Voss:1992:ADA

[Vos92] F. W. Voss. APPC/MVS distributed application support. IBM Systems Journal, 31(2):381–408, 1992. CODEN IBMSA7. ISSN 0018-8670. G321-5479.

Verma:2008:PCC

[VPH^{+}08] R. Verma, G. R. Plaschka, B. Hanlon, A. Livingston, and K. Kalcher. Predicting customer choice in services using discrete choice analysis. IBM Systems Journal, 47(1):179– ??, January/March 2008. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/471/verma.html.

Voelker:1986:GKE

[VR86] J. A. Voelker and G. B. Ratica. The genesis of a knowledgebased expert system. IBM Systems Journal, 25(2):181–189, 1986. CODEN IBMSA7. ISSN 0018-8670.

Vredenburg:2003:BEU

[Vre03]  Karel Vredenburg.  Build-

ing ease of use into the IBM user experience. IBM Systems Journal, 42(4):517–531, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/424/preface.pdf.

Vredenburg:2003:P

[VRS03] Karel Vredenburg, John J.

Ritsko, and David I. Seidman. Preface. IBM Systems Journal, 42(4):515–516, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/424/preface.pdf.

Verma:2003:SSR

[VSC^{+}03] D. C. Verma, S. Sahu, S. Calo, A. Shaikh, I. Chang, and A. Acharya. SRIRAM: a scalable resilient autonomic mesh. IBM Systems Journal, 42(1): 19–28, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/421/verma.pdf.

Viegas:2006:TFC

[VW06] F. B. Viégas and M. Wattenberg. Technical forum: Communication-minded visualization: A call to action. IBM Systems Journal, 45(4): 801–812, ???? 2006. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/454/viegas.html.

Weiss:2002:AGM

[WA02a] S. M. Weiss and C. V. Apte. Automated generation of model cases for help-desk applications. IBM Systems Journal, 41(3):421–427, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/413/weiss.html.

Weyuker:2002:MPP

[WA02b] E. J. Weyuker and A. Avritzer. A metric for predicting the performance of an application under a growing workload. IBM Systems Journal, 41(1):45–54, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/411/weyuker.html;   http:
                                        //www.research.ibm.com/journal/
                                        sj/411/weyuker.pdf.

Ward:2007:ICC

[WAB^{+}07] C. Ward, V. Aggarwal, M. Buco, E. Olsson, and S. Weinberger. Integrated change and configuration management. IBM Systems Journal, 46(3):459–??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/463/ward.html.

Wagner:1973:IDC

[Wag73] R. E. Wagner. Indexing design considerations. IBM Systems Journal, 12(4):351–367, 1973. CODEN IBMSA7. ISSN 0018-8670. See errata [Ano74k].

Wahi:1972:GMB

[Wah72] P. N. Wahi. A general management business simulation in APL. IBM Systems Journal, 11(2):169–180, 1972. CODEN IBMSA7. ISSN 0018-8670. See addendum [Ano73e].

Walker:2007:IBT

[Wal07] L. Walker. IBM business transformation enabled by serviceoriented architecture. IBM Systems Journal, 46(4):??, October/December 2007. CODEN IBMSA7. ISSN 0018-8670.

Wasmund:1993:ICS

[Was93] M. Wasmund. Implementing Critical Success Factors in software reuse. IBM Systems Journal, 32(4):595–611, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5526.

Waters:1979:DID

[Wat79] F. C. H. Waters. Design of the IBM 8100 Data Base and Transaction Management System—DTMS. IBM Systems Journal, 18(4):565–581, 1979. CODEN IBMSA7. ISSN 0018-8670.

Watson:1988:OQM

[Wat88] S. L. Watson. OS/2 query manager overview and prompted interface. IBM Systems Journal, 27(2):119–133, June 1988. CODEN IBMSA7. ISSN 0018-8670.

Watkins:1994:TNR

[Wat94a] A. J. Watkins. Technical note: On reliability modeling and software quality. IBM Systems Journal, 33(1):220–222, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-1.html#twelve. G321-

5539.

Watson:1994:BVC

[Wat94b] D. Watson. Books: Visual Cues: Practical Data Visualization by Peter R. Keller and Mary M. Keller. IBM Systems Journal, 33(3):530–531, 1994. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj33-3.html.

Wolf:2004:STS

[WAV^{+}04] C. G. Wolf, S. R. Alpert, J. G. Vergo, L. Kozakov, and Y. Doganata. Summarizing technical support documents for search: Expert and user studies. IBM Systems Journal, 43 (3):564–586, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/433/wolf.html;   http:
//www.research.ibm.com/journal/
sj/433/wolf.pdf;   http:/
/www.research.ibm.com/journal/
sj/433/wolf.txt.

White:1969:NAE

[WB69] W. W. White and A. M.

Bomberault. A network algorithm for empty freight car allocation. IBM Systems Journal, 8(2):147–169, 1969. CODEN IBMSA7. ISSN 0018-8670.

WladawskyBerger:1995:PPP

[WB95] I. Wladawsky-Berger. The power and promise of parallel computing. IBM Systems Journal, 34(2):146–151, 1995. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj34-2.html#two.

Wladawsky-Berger:1999:TPI

[WB99] I. Wladawsky-Berger. Turning points in information technology. IBM Systems Journal, 38 (2–3):449–453, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/382/wladawsky.pdf.

Willenborg:2004:DWA

[WBC04] R. Willenborg, K. Brown, and G. Cuomo. Designing Web-Sphere Application Server for performance: An evolutionary approach. IBM Systems Journal, 43(2):327–340, ???? 2004. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/432/willenborg.html;
http://www.research.ibm.
com/journal/sj/432/willenborg.
pdf;   http://www.research.

                                        ibm.com/journal/sj/432/willenborg.
                                        txt.

Waters:1987:AUE

[WBSK87] F. C. H. Waters, R. G. Bias, and P. L. Smith-Kerker. AIX usability enhancements and human factors. IBM Systems Journal, 26(4):383–394, December 1987. CODEN IBMSA7. ISSN 0018-8670.

Weller:1980:SAG

[WCG^{+}80] D. L. Weller, E. D. Carlson, G. M. Giddings, F. P. Palermo, R. Williams, and S. N. Zilles. Software architecture for graphical interaction. IBM Systems Journal, 19(3):314–330, 1980. CODEN IBMSA7. ISSN 0018-8670.

Wade:1969:TAU

[WCV69] R. D. Wade, G. P. Cawsey, and R. A. K. Veber. A teleprocessing approach using standard equipment. IBM Systems Journal, 8(1):28–47, 1969. CODEN IBMSA7. ISSN 0018-8670.

Weingarten:1979:IAC

[Wei79] R. A. Weingarten. An integrated approach to centralized communications network management. IBM Systems Journal, 18(4):484–506, 1979. CO-DEN IBMSA7. ISSN 0018-8670.

Weisser:2005:MVP

[Wei05] Jason Weisser. Message from the Vice President, IBM Asset and Integration Technology, Software Group. IBM Systems Journal, 44(4):??, ????

2005. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/444/message.html.

Welch:1965:RPS

[Wel65] P. D. Welch. On the reliability of polymorphic systems. IBM Systems Journal, 4(1):43– 52, 1965. CODEN IBMSA7. ISSN 0018-8670.

Walston:1977:FAR

[WF77a] C. E. Walston and C. P. Felix. Forum: Authors’ reply. IBM Systems Journal, 16 (4):422–423, 1977. CODEN IBMSA7. ISSN 0018-8670. See [WF77b, Hal77].

Walston:1977:MPM

[WF77b] Claude E. Walston and Charles P.

Felix. A method of programming measurement and estimation. IBM Systems Journal, 16(1):54–73, 1977. CODEN IBMSA7. ISSN 0018-8670. See letters [Hal77, WF77a].

Wood:1980:DBS

[WFS80] Christopher Wood, Eduardo B.

Fernandez, and Rita C. Summers. Data base security: Requirements, policies, and models. IBM Systems Journal, 19 (2):229–252, 1980. CODEN IBMSA7. ISSN 0018-8670.

Wheeler:1988:ISA

[WG88] E. F. Wheeler and A. G. Ganek.

Introduction to Systems Application Architecture. IBM Systems Journal, 27(3):250–263, 1988. CODEN IBMSA7. ISSN 0018-8670.

Weiser:1999:OUC

[WGB99] M. Weiser, R. Gold, and J. S. Brown. The origins of ubiquitous computing research at PARC in the late 1980s. IBM Systems Journal, 38(4):677, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/weiser.html.

Wheeler:1974:OCP

[Whe74] Thomas F. Wheeler, Jr. OS/ VS1 concepts and philosophies. IBM Systems Journal, 13(3):213–229, 1974. CODEN IBMSA7. ISSN 0018-8670.

Wheatley:1991:EDA

[Whe91] M. T. Wheatley. Extending the domain of APL. IBM Systems Journal, 30(4):446–455, 1991. CODEN IBMSA7. ISSN 0018-8670.

Whitmore:2001:MDS

[Whi01] J. J. Whitmore. A method for designing secure solutions. IBM Systems Journal, 40(3): 747–768, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/403/whitmore.html;   http:
//www.research.ibm.com/journal/
sj/403/whitmore.pdf.

Weingarten:1983:LPD

[WI83] R. A. Weingarten and E. E. Iacobucci. Logical problem determination for SNA networks. IBM Systems Journal, 23(4):387–403, 1983. CODEN IBMSA7. ISSN 0018-8670.

Wilson:1986:PAP

[Wil86] W. G. Wilson. Prolog for applications programming. IBM Systems Journal, 25(2):190– 206, 1986. CODEN IBMSA7. ISSN 0018-8670.

Willhoft:1991:PEA

[Wil91] R. G. Willhoft. Parallel expression in the APL2 language. IBM Systems Journal, 30(4):498–512, 1991. CODEN IBMSA7. ISSN 0018-8670.

Wimbrow:1971:LIA

[Wim71] J. H. Wimbrow. A large-scale interactive administrative system. IBM Systems Journal, 10(4):260–282, 1971. CODEN IBMSA7. ISSN 0018-8670.

Witt:1963:DSA

[Wit63] B. I. Witt. Dynamic storage allocation for a real-time system. IBM Systems Journal, 2: 230–239, September/December 1963. CODEN IBMSA7. ISSN 0018-8670.

Witt:1966:FSO

[Wit66] B. I. Witt. The functional structure of OS/360, part II: Job and task management. IBM Systems Journal, 5(1):12– 29, 1966. CODEN IBMSA7. ISSN 0018-8670.

Whisnant:2003:SMD

[WKI03] K. Whisnant, Z. T. Kalbarczyk, and R. K. Iyer. A system model for dynamically reconfigurable software. IBM Systems Journal, 42(1):45–59, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/whisnant.pdf.

Weinstein:2007:IBT

[WLB^{+}07] R. S. Weinstein, A. M. López, G. P. Barker, E. A. Krupinski, M. R. Descour, K. M. Scott, L. C. Richter, S. J. Beinar, M. J. Holcomb, P. H. Bartels, R. A. McNeely, and A. K. Bhattacharyya. The innovative bundling of teleradiology, telepathology, and teleoncology services. IBM Systems Journal, 46(1):69–84, ???? 2007. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/461/weinstein.html.

Wong:1997:ETD

[WLE^{+}97] J. W. Wong, K. A. Lyons, D. Evans, R. J. Velthuys, G. v. Bochmann, E. Dubois, N. D. Georganas, G. Neufeld, M. T. Özsu, J. Brinskelle, A. Hafid, N. Hutchinson, P. Iglinski, B. Kerhervé, L. Lamont, D. Makaroff, and D. Szafron. Enabling technology for distributed multimedia applications. IBM Systems Journal, 36(4):489–507, 1997. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/364/wong.html.

Walrad:1993:MKA

[WM93] C. Walrad and E. Moss. Measurement: the key to application development quality. IBM Systems Journal, 32(3):445– 460, 1993. CODEN IBMSA7. ISSN 0018-8670. G321-5520.

Wyckoff:1998:S

[WMLF98] P. Wyckoff, S. W. McLaughry, T. J. Lehman, and D. A. Ford. T spaces. IBM Systems Journal, 37(3):454–481, ???? 1998. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/373/wyckoff.html.

Wang:2001:NTE

[WMSW01] J. T. L. Wang, Q. Ma, D. Shasha, and C. H. Wu. New techniques for extracting features from protein sequences. IBM Systems Journal, 40(2): 426–441, ???? 2001. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/402/wang.html;   http:
//www.research.ibm.com/journal/
sj/402/wang.pdf;   http:/
/www.research.ibm.com/journal/
sj/402/wang.txt.

Wang:2007:ICD

[WND07] X. S. Wang, L. Nayda, and R. Dettinger. Infrastructure for a clinical-decision-intelligence system. IBM Systems Journal, 46(1):151–169, ???? 2007. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.

                                        research.ibm.com/journal/
                                        sj/461/wang.html.

Wolford:1988:AES

[Wol88] D. E. Wolford. Application enabling in SAA. IBM Systems Journal, 27(3):301–305, 1988. CODEN IBMSA7. ISSN 0018-8670.

Woodrum:1969:ISM

[Woo69]  L. J. Woodrum.  Internal

sorting with minimal comparing. IBM Systems Journal, 8 (3):189–203, 1969. CODEN IBMSA7. ISSN 0018-8670. See errata [Ano70k].

Woodrum:1970:MFB

[Woo70] L. J. Woodrum. A model of floating buffering. IBM Systems Journal, 9(2):118–144, 1970. CODEN IBMSA7. ISSN 0018-8670.

Woodworth:2002:MCD

[Woo02]  Bill Woodworth.  Message

from the Corporate Director, IBM Software Test. IBM Systems Journal, 41(1):??, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/411/execletter.html;
                                        http://www.research.ibm.
                                        com/journal/sj/411/execletter.
                                        pdf.

Wothke:1993:MBA

[Wot93] K. Wothke. Morphologically based automatic phonetic transcription. IBM Systems Journal, 32(3):486–511, 1993. CO-DEN IBMSA7. ISSN 0018-8670. G321-5522.

Waszkowycz:2001:LSV

[WPSL01] B. Waszkowycz, T. D. J.

Perkins, R. A. Sykes, and J. Li. Large-scale virtual screening for discovering leads in the postgenomic era. IBM Systems Journal, 40(2):360–376, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/waszkowycz.html;
http://www.research.ibm.
com/journal/sj/402/waszkowycz.
pdf;   http://www.research.
ibm.com/journal/sj/402/waszkowycz.
txt.

Want:2003:CAP

[WPT03] R. Want, T. Pering, and D. Tennenhouse. Comparing autonomic and proactive computing. IBM Systems Journal, 42(1):129–135, ???? 2003. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/421/want.pdf.

Wilson:1992:IIO

[WR92] D. G. Wilson and B. D. Rudin.

Introduction to the IBM Optimization Subroutine Library. IBM Systems Journal, 31(1):4– 10, 1992. CODEN IBMSA7. ISSN 0018-8670.

Will:2004:WPU

[WRS04] R. Will, S. Ramaswamy, and T. Schaeck. WebSphere Portal: Unified user access to content, applications, and services.

IBM Systems Journal, 43(2): 420–429, ???? 2004. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/432/will.html;   http:
//www.research.ibm.com/journal/
sj/432/will.pdf;   http:/
/www.research.ibm.com/journal/
sj/432/will.txt.

Winslow:2001:MMV

[WSG^{+}01] R. L. Winslow, D. F. Scollan, J. L. Greenstein, C. K. Yung, W. Baumgartner, Jr., G. Bhanot, D. L. Gresh, and B. E. Rogowitz. Mapping, modeling, and visual exploration of structure-function relationships in the heart. IBM Systems Journal, 40(2):342–359, ???? 2001. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/402/winslow.html;   http:
//www.research.ibm.com/journal/
sj/402/winslow.pdf;   http:
//www.research.ibm.com/journal/
sj/402/winslow.txt.

Willuhn:2003:DAS

[WSKW^{+}03] D. Willuhn, C. Schulz, L. Knoth-Weber, S. Feger, and Y. Saillet. Developing accessible software for data visualization. IBM Systems Journal, 42(4): 652–668, ???? 2003. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/424/willuhn.pdf.

Williams:2002:STT

[WSP^{+}02] C. Williams, H. Sluiman, D. Pitcher, M. Slavescu, J. Spratley, M. Brodhun, J. McLean, C. Rankin, and K. Rosengren. The STCL test tools architecture. IBM Systems Journal, 41(1):74–88, ???? 2002. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/411/williams.html;   http:
//www.research.ibm.com/journal/
sj/411/williams.pdf.

White:1964:DIP

[WT64] B. White and J. Trimble. Design of an integrated programming and operating system, part VI: Implementation on the 7040/44 data processing system. IBM Systems Journal, 3(1):79–94, 1964. CODEN IBMSA7. ISSN 0018-8670.

Wu:1998:SWU

[WYB98] K.-L. Wu, P. S. Yu, and A. Ballman. SpeedTracer: a Web usage mining and analysis tool. IBM Systems Journal, 37(1):89–105, 1998. CO-DEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (elec-

tronic). URL http://www.
research.ibm.com/journal/
sj/371/wu.html.

Yagil:1963:GID

[Yag63] S. Yagil. Generation of input data for simulations. IBM Systems Journal, 2:288–296, September/December 1963.

CODEN IBMSA7. ISSN 0018-8670.

Yardeni:1965:ATC

[Yar65] L. A. Yardeni. Algorithms for traffic-signal control. IBM Systems Journal, 4(2):148–161, 1965. CODEN IBMSA7. ISSN 0018-8670.

Ye:2006:SPM

[YC06] M. Ye and L. Cheng. Systemperformance modeling for massively multiplayer online roleplaying games. IBM Systems Journal, 45(1):45–58, ???? 2006. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

                                        research.ibm.com/journal/
                                        sj/451/ye.html.

Yelavich:1985:CIC

[Yel85] B. M. Yelavich. Customer information control system — an evolving system facility. IBM Systems Journal, 24(3-4):264– 278, 1985. CODEN IBMSA7. ISSN 0018-8670.

Yellin:2003:CAD

[Yel03] D. M. Yellin. Competitive algorithms for the dynamic selection of component implementations. IBM Systems Journal, 42(1):85–97, ???? 2003. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/421/yellin.pdf.

Yeoh:2005:BAL

[Yeo05] C. Yeoh. Building applications for the Linux Standard Base. IBM Systems Journal, 44(2):369–380, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/yeoh.html;   http:
//www.research.ibm.com/journal/
sj/442/yeoh.pdf;   http:/
/www.research.ibm.com/journal/
sj/442/yeoh.txt.

Yakhnis:1994:DPU

[YFS94] V. R. Yakhnis, J. A. Farrell, and S. S. Shultz. Deriving programs using generic algorithms. IBM Systems Journal, 33(1):158–181, 1994. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj33-1.html#nine.  G321-

5537.

Yglesias:1993:TFI

[Ygl93] K. P. Yglesias. Technical forum: Information reuse parallels software reuse. IBM Systems Journal, 32(4):615–620, 1993. CODEN IBMSA7. ISSN 0018-8670.

Yan:2005:UOS

[YLG05] N. Yan, D. Leip, and K. Gupta.

The use of open-source software in the IBM corporate portal. IBM Systems Journal, 44 (2):419–425, ???? 2005. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/442/yan.html;   http:/
/www.research.ibm.com/journal/
sj/442/yan.pdf;   http://
www.research.ibm.com/journal/
sj/442/yan.txt.

Youngs:1999:SAD

[YRPSK99] R. Youngs, D. Redmond-Pyle, P. Spaas, and E. Kahan. A standard for architecture description. IBM Systems Journal, 38(1):32–50, ???? 1999. CODEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/381/youngs.html.

Yeh:1991:EIC

[YS91] P. C. Yeh and R. M. Smith, Sr. ESA/390 integrated cryptographic facility: An overview. IBM Systems Journal, 30(2): 192–205, 1991. CODEN IBMSA7. ISSN 0018-8670.

Yuval:1980:SCA

[Yuv80] Asher Yuval. System contention analysis — an alternate approach to system tuning. IBM Systems Journal, 19 (2):208–228, 1980. CODEN IBMSA7. ISSN 0018-8670.

Zatti:1992:NRI

[ZABM92] S. Zatti, J. Ashfield, J. Baker, and E. Miller. Naming and registration for IBM distributed systems. IBM Systems Journal, 31(2):353–380, 1992. CO-DEN IBMSA7. ISSN 0018-8670. G321-5478.

Zachman:1982:BSP

[Zac82] J. A. Zachman. Business Systems Planning and Business Information Control Study: a comparison. IBM Systems Journal, 21(1):31–53, 1982.

CODEN IBMSA7. ISSN 0018-8670.

Zachman:1987:FIS

[Zac87] John A. Zachman. A framework for information systems architecture. IBM Systems Journal, 26(3):276–292, 1987. CODEN IBMSA7. ISSN 0018-8670.

Zachman:1999:FIS

[Zac99] J. A. Zachman. A framework for information systems architecture. IBM Systems Journal, 38(2–3):454–471, ???? 1999. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/382/zachman.pdf. Reprinted

from [Zac87].

Zhao:1998:TNW

[Zha98] Y. Zhao. Technical note — WebEntree: a Web service aggregator. IBM Systems Journal, 37(4):584–595, ???? 1998. CODEN IBMSA7. ISSN 0018-8670. URL http://www.

research.ibm.com/journal/
sj/374/zhao.html.

Ziegler:1979:DIS

[Zie79] K. Ziegler, Jr. A distributed information system study. IBM Systems Journal, 18(3):374– 401, 1979. CODEN IBMSA7. ISSN 0018-8670.

Zimmerman:1996:PAN

[Zim96] T. G. Zimmerman. Personal Area Networks: Near-field intrabody communication. IBM Systems Journal, 35(3&4):609– 617, 1996. CODEN IBMSA7. ISSN 0018-8670.

Zimmerman:1999:WND

[Zim99] T. G. Zimmerman. Wireless networked digital devices: a new paradigm for computing and communication. IBM Systems Journal, 38(4): 566–574, ???? 1999. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/zimmerman.html.

Zeng:2007:MAB

[ZLK^{+}07] L. Zeng, H. Lei, T. Koyanagi, H. Ohsaki, and H. Chang. Model analysis for business event processing. IBM Systems Journal, 46(4):??, October/December 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/464/zeng.html.

Zloof:1977:QEA

[Zlo77] Moshe M. Zloof. Query-by-Example: a data base language. IBM Systems Journal, 16(4):324–343, 1977. CODEN IBMSA7. ISSN 0018-8670.

Zloof:1982:OBL

[Zlo82]  M. M. Zloof.  Office-by-

Example: a business language that unifies data and word processing and electronic mail. IBM Systems Journal, 21(3):272–304, 1982. CODEN IBMSA7. ISSN 0018-8670.

Zhou:2002:AML

[ZMF02] M. X. Zhou, S. Ma, and Y. Feng. Applying machine learning to automated information graphics generation. IBM Systems Journal, 41(3): 504–523, ???? 2002. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/413/zhou.html.

Zollar:2007:I

[Zol07] Al Zollar. Introduction. IBM Systems Journal, 46(3):??, July/September 2007. CO-DEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/463/intro.html.

Zhai:1999:MIE

[ZS99] S. Zhai and B. A. Smith. Multistream input: An experimental study of document scrolling methods. IBM Systems Journal, 38(4):642–651, ???? 1999. CODEN IBMSA7. ISSN 0018-

8670.  URL http://www.
research.ibm.com/journal/
sj/384/zhai.html.

Zhang:2008:SMP

[ZZC^{+}08] Liang-Jie Zhang, Nianjun Zhou, Yi-Min Chee, Ahamed Jalaldeen, Karthikeyan Ponnalagu, Renuka R. Sindhgatta, Ali Arsanjani, and Fausta Bernardini. SOMA-ME: a platform for the model-driven design of SOA solutions. IBM Systems Journal, 47(3):397–

413, ????  2008.  CO-

DEN IBMSA7. ISSN 0018-

                                        8670.  URL http://www.
                                        research.ibm.com/journal/
                                        sj/473/zhang.pdf.